A Selected Bibliography of Publications by, and about, Mario Bunge

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

18 April 2024 Version 1.11

Title word cross-reference

$1.75 [Bun53]. $13.50 [Par81]. $19.95 [BS96]. $2.00 [Bun53]. $21.00 [Lin69]. $23.90 [Bun76a]. $26.00 [Bun99a]. $32.00 [Bun11a]. $50 [Bun07c]. 512 × 814 [Ano59]. $58.00 [Bun86c]. 6 × 9 [Lin69]. $7.50 [Ano59]. $9.50 [Sne68b]. ^{∘} [Den14a].

-. [KPSP74].

. [KPSP74].

0 [Kil04]. 0-8020-8860-0 [Kil04].

1 [Sne68b]. 11-012 [Har77]. 12.50 [Sin81]. 13 [Bun75b]. 13th [SW89]. 14th [SW89]. 16. [Die78]. 16ème [Die78]. 16th [Die78]. 184.£ [Sin81]. 1940s [Ort19]. 1960s [Ort19]. 1962 [Fea64]. 1963 [Mul63]. 1973 [Bru73, Rob74].

1975 [Har77]. 1976 [SA76]. 1978 [BAB^{+}80]. 1986 [HD88]. 1990 [Aga91]. 1992 [Mon92]. 1996 [Asp98b]. 1997 [AL99]. 1999 [Van04].

2000 [AF01]. 21st [KFT89, SW89]. 25 [Bun84a]. 250-year [Wie99]. 250jährigen [Wie99].

3-540-13923-0 [Bun87a]. 38th [Mon92].

5th [OW79].

60s [Din60]. 60th [SD91]. 68 [Bun53].

7.50 [Bun79b]. 7th [OW79].

90th [BAG17]. 95th [Den14a]. 978-3-319-29250-2 [Mat16].

= [Die78, Flø86, Mor81, SW89].

âge [Vac93]. âme [Bun83c].

Aarhus [AF01]. Abd [Obi22]. ability [GBS^{+}05]. Absolute [Bun00a]. abstraccion [Bun85f]. abstraction [Bun85d, Bun85f]. Academia [Bun96c, Bun96d]. Academy [AL99, AF01, Bun78d]. accessibility [QdAY13]. acción [Bun85b]. acciones [Sca14]. According [Bun01-30, Ore22b, Van01, Pie99, Sil13c]. Account [Bun12c]. accounting [BRG04]. Acerca [Roj13]. acessibilidade [QdAY13]. Achievements [Mat19b]. Acid [Bun90h]. Acknowledgements [Mat19d]. Acquisition [Bun86g]. Across [Bun18d]. actes [Cau88]. Action [Vol93, Bun85b]. actions [Sca14]. actual [Bun95g]. actuales [Agu06, DM00]. Actualidad [Bun87i]. actualizacion [Bun97b, Bun09c, Bun80d]. Adam [Bun62b]. Addenda [Bun71b]. Addition [Bun69f, Bun75d]. adición [Bun75d]. Admissible [Bun12u]. Advancement [Ray15]. Advances [SD91, Slo99, Bun19f, AL99]. Advantages [Bun09a]. Affairs [Bun99a]. afraid [BvL91]. after [Bun84b]. against [BB10, OB84]. Agassi [BAG17, JL95]. Agathonism [Bun90h, Kar19a]. Age [Bun55c, Kra80, Vac93]. Agnostic [Bun87b]. Aharonov [Bun15a]. Aires [Fonxx]. Akten [SW89]. al [Bun39, Bun75d, Obi22]. al-Rahman [Obi22]. Albert [Fra18, Bun18c, Fra18]. Alchourrón [GKvWZ97]. Alexander [Bun69a, Pfe69]. Alfonso [Bar11]. algebra [Kon95]. alguna [Bun09d]. Algunos [Qui14]. alternativa [Ama14]. Alternative [Bun01a, Bun01-31, Bun04h, Ama14, Bun88b, Bun00f, WL96]. Alternatives [Bun00g]. Alwin [HD88]. am [SW89]. ambivalent [Bun88a]. Ameghino [Bun45b]. America [Bun89f, FvWK03]. American [Bun73c, NSB10]. among [Bun60f]. amos [Bar14]. Analogy [Bun68a, Bun69b, Bun81a].

Analysis [Bun11d, Hob19, Bla19, Bun62a]. Analytic [Bun88b, Bun01a]. Analytical [Bun07b, zW12]. Analyticity [Bun61a]. Animal [Bun93d, YCVP^{+}19, Yan14]. aniversario [DM00, Den14a]. anlässlich [Wie99]. Anniversary [SD91, Den14a]. Announcements [Ano79a]. Annual [AF01]. Años [BWM^{+}04]. Ansätze [Kla03a, Kla03b, Kla04]. Anthropology [Cri19, TBG^{+}85]. Anti [Bun90a, Bun95d]. Anti-Realism [Bun90a, Bun95d]. Antirealism [Bun01x, Bun85l, Bun93c]. antirrealismo [Bun85l, Bun95d]. Anything [Bun12b]. août [Bun78d]. Apocalypse [Bun88h]. apoiadas [QdAY13]. Apostol [Bun85c]. Appendix [Bun10b, Bun10a]. applicability [LB78b]. Application [YCVP^{+}19, Bel00, Bel01]. applications [Har77]. Applied [Bun66e, Bun01o, Bra82, Bun83j, Bun88c, Bun91a]. Applying [Set96]. Appraisal [Mar12]. Appreciation [zW12]. Approach [Bun55d, Bun64a, Bun80k, Bun99c, Bun01a, Cor19, zW11a, Bos14, Bun88b, Bun90b, BT^{+}03, Bun15e, Bun17g, Ber82, Cau84, Puc81, Par81]. Approaches [Kla03a, dSMdC19, GML14, Har77, Kla03b, Kla04, SA93]. approcci [SA93]. Approximation [Kon95]. April [Har77, HD88]. aproximación [Bos14]. aquellos [Bar14]. Arab [Obi22]. Arabic [Obi11]. Arbeit [Kla03b, Kla04, Obr05, Kla03a]. architectonique [Sil13b]. architectural [Sil13b]. architecture [Mac11]. Archivo [Fonxx]. Argentina [Bun87b, Ort19]. Argentine [BC14, Bun61b, Den19]. argentino [BC14, SA01]. Arguments [Bun18d]. Aris [Bun79b]. Arising [Bas71]. Aristote [Esp92, Sin91]. Aristotle [Esp92]. Arquitetura [Mac11]. arrow [Bun70a]. Art [Aga88a]. Article [Erl67]. Aspectos [Bun75j]. aspects [BG16, Bun75j]. Assaying [Bun61h, Bun66f]. Association [SA76]. asymmetry [Bun70g]. atom [Bun52]. atomic [Bun51d]. atómica [Bun51d]. átomo [Bun52]. Attachment [GBS^{+}05]. Attachment-style [GBS^{+}05]. Auge [Bun44c]. August [Bun78d, SW89]. Augusto [For89a]. ausgewählte [SW89]. Austria [SW89]. authors [Ano14b]. Autobiography [Bun90d]. Autónoma [Fonxx]. Autonomous [Bun90g]. autores [Ano14b]. avances [Bun19f]. Axiología [Bun95f]. Axiology [Bun95f]. Axiomatic [Mat82]. Axiomatics [Bun67f, Kah19, Sal71]. axiomatization [Bun67e]. Axiomatize [Bun17e]. axioms [Var17].

B [Bun10b]. babies [WL96]. Back [Ano10a, Ano12a, Bun10g]. Background [Vel22]. bad [Bun89b, Bun90c]. bankruptcy [Bun79a]. Based [Bun18d, CD19, LSP22]. bases [Cha06]. Basic [Bar81, Bow80, Bun88c, Bun91a, Bun03g, Bun13k, Esp86, McF78, Qui80, Rus86, She76, Rus80]. Basis [Bun01n, VOHPMJ12]. bathwater [WL96]. Baum [Bun74a]. Baunge [vR77]. Bayesian [Pin13, Roj13, Sil13f]. bayesiana [Pin13, Roj13, Sil13f]. Bayesianism [Bun08a]. Be [Bun10h, Bou91, Bun88c, Bun91a, Bun14a, Moe87, Pan90, Bun02b]. Beacon [Bun53]. Become [Bun10h, Bun88j]. Bedeutung [Bun92b]. Begegnungen [Fra18]. Behavior [Bun80a, Bun80e, Pan90]. Behavioral [Bun73j, SB01, Bun73k]. Behind [Bun06f, Bun09e, Bun06d]. Being [Bun76d, Bun12x, Bun12y, Bun76d]. Beiträge [Vol93, SW89]. beliefs [Bun91i]. believer [Mos91]. Bell [Bun89a]. Bemerkungen [Nar75]. Berg [Bun01-30]. Bericht [Mon92]. Berlin [Bun87a]. Berlin-Heidelberg-New [Bun87a]. Bertrand [Bun73a]. Besonderheiten [BK75]. better [Poo96]. Between [Aga12, Bun54, Bun00c, Bun01d, Bun14e, Bun16a, Bun18a, Bli91, Bun81a, Bun94d, Bun95o, Bun01h, Bun01p, Cha06, Jod10, Lac96, Mat82, Per91, Bun16b, Mat16]. Bewusstsein [PB89]. Beyond [Bas71, Bun71a, Bun79l]. Bhaskar [Kai09]. Bibliographie [Sil13a]. Bibliography [Bli90a, Sil13a]. Biblioteca [Fonxx]. Biennial [SA76]. Big [Bun14a]. biography [Ano74]. biología [Bun76j, Bun95h]. Biologic [BL71]. Biological [Mac85, Bun90g]. Biologie [MB00]. Biology [BB89, MB00, Bun76j, Bun87a, Bun87h, Bun95h, LB78a]. biomedical [Bun78a, Bun78g]. Biophilosophy [MB97, Bun79k]. Biopopulations [Bun81b]. biospecies [Bun81b]. Birthday [BAG17, SD91]. Bis [SW89]. Black [Bun63c, Bun85b]. Blackwell [BKR96, BS96]. Blushing [Bun07a]. Bodies [Bun10d]. Body [Ber82, BAB^{+}80, Bun80j, Bun80k, Bun96g, Bun10l, Cau84, Lev83, Puc81, Bun79g, Bun79h, Bun87d, Bun91f, LB78a, Par81]. Bogus [Bun10i, Bun11b]. Bohm [Bun15a, Jun19]. Bohr [Bun79e, Bun85j, Bun88e, Bun92c]. bons [Bun89b, Bun90c]. Book [Aga72, Ano59, Asp98a, Asp98b, Bar11, Bar14, Bar81, Ber82, BS96, Bow80, Bra97, Bru91, Bru73, Bub68a, Bub68b, Bun51a, Bun58a, Bun58b, Bun60b, Bun62b, Bun63a, Bun66a, Bun68b, Bun71a, Bun72a, Bun72b, Bun73b, Bun74a, Bun76a, Bun76b, Bun76c, Bun78a, Bun78b, Bun78c, Bun79b, Ano81, Bun84a, Bun84b, Bun86b, Bun86a, Bun86c, Bun87a, Bun93a, Bun99a, Bun99b, Bun00b, Bun01b, Bun04a, Bun05a, Bun05b, Bun07b, Bun07c, Bun11a, Cau84, Col99, Cor86, Déc10, Dem97, Din60, Dri87, Duc60a, Duc60b, Ech85, Edg69, Esp85, Esp86, Fea64, Fin70, For07, G.73, Haj65, Haj76, Har66, Har75, Hes60, Hin61a, Hin61b, Hin64, Hof84, Kir75a, Kir75b, Lev83, Lin69, Lon62, Lun00, Mat16, McF78, Met73a, Met73b, Mor61, Mul63, Par81, Puc81, Qui74, Qui80, Ric08, Rob74, Rus80, Rus86, Sin81]. Book [Sle96, Sle12a, Sne68a, Sne68c, Sne68b, Tre01, Tri06, Tuo74a, Tuo74b, Vol73a, Vol73b, Vol79, Vol81, Whe80, dP16, Bun78h, Bun78i, Bun78g, Bun93d, Bun06h, Bun06a]. Borges [B^{+}99]. Borgoñoz [Bar11]. Boston [Bow80, Bru73, Bun53, KPSP74, Rob74]. Boston-U.S.A [KPSP74]. Boudon [Bun90a, CH09]. Boundary [Bun92f, Cav12, Mar96]. Bowles [Bun07c]. Box [Bun63c]. boxes [Bun85b]. Boyd [Bun07c]. Bradley [Bun72a, Bun72b]. Brain [BAB^{+}80, Bun10c, Bun10n, Gam14, OW79, Bun99g, Pan90, OW79, PB89]. Brainless [Bun01g, Bun81g, Bun90b]. Branch [Bun82b]. Breves [Yan14]. bridge [Bun95o]. Bridging [Bun12a]. Bridgman [Bun58b]. brief [Ano74, Yan14]. Brookes [Bun84a]. Bruce [Bun13a, CC13]. buds [Bun79f]. Buenos [Fonxx]. Builder [Aga12]. Building [KFT89]. Bulygin [GKvWZ97]. bundles [Bun14a]. BUNGE [G.73, Ano59, Asp98b, Bar11, Bow80, Bru73, Ano81, Cor86, Din60, Ech85, Fea64, KPSP74, Kil04, Kla03a, Mat16, Mul63, Par81, Rob74, Sin81, Sle12a, Sne68b, Whe80, Qui80, Aga72, AC82, Aga12, Aga13, ABA19, Gru05, Ano74, Ano04, Ano09, Ano14a, Ano15, Ano16, Ano23, Apo85, Art19, Asp98a, AC13, Bar20, BRG04, Bar81, BvL91, Bel00, Bel01, Ber82, Bla19, BC14, Bra97, Bru91, Bub68a, Bub68b, Bun83a, Bun14d, Bun17d, Cau84, Cav12, CD90, Col01, Col99, Cor14, Cor12, Cor19, Cri14, Cri19, Cup19, CD19, Déc10, Del12, Del13, Del19, Dem97, DM00, Den14a, Den14b, Den19, Den03, Dri87, Dro15, Dro19, Duc60a, Duc60b, Edg69, Erl67, Esp85, Esp86, Esp92, Fin70, For07, For89a, Gam14, Gil19, Góm14, Haj65, Haj76, Har66]. Bunge [Har75, Hes60, Hin77, Hin61a, Hin61b, Hin64, Hof84, Ing19, Jod10, Kai09, Kan15, Kir75a, Kir75b, KJ10, Kla03a, Kla03b, Kla04, Kon95, Lan89, Lev83, Lon62, LSP22, Lun00, MB16, Mah01, Mah14, Mah21, Man85, Mar72, Mar96, Mar12, Mar19, Mat82, Mat03, Mat12, Mat19a, Mat19c, Mat19b, Mau22a, May90, McF78, Met73a, Met73b, Mil77, Moe87, Mor61, Mos14, Moy15, Nar75, Obi11, Obi19, Obi22, Obr05, OHS02, Ore22a, Ore22b, Ort19, Pan90, Pat90, Pic01, Pic04, Pic07, Pie99, Pin13, Puc81, QdAY13, Qui74, QB79, Qui79, Ric08, Roj13, Roj02, Rom17, Rom18, Rom19, RG02, RW05, Roy90, Rus72, Rus73, Rus80, Rus86, Sal71, Sca14, Sch61, Scr63, SC89, Sil13a, Sil13c, Sil19, Sil13f, Sle12b, Sle19, Sne68a]. Bunge [Sne68c, Sol22, Str69, Tre01, Tri06, Tul95, Tuo74a, Tuo74b, Vac93, Van01, VOHPMJ12, Vol73a, Vol73b, Vol79, Vol81, WW90, WD90, Wet82, WL96, Zie16, Zie22, dF19, dP16, Lin69]. Bungean [Hal19, zW11a, zW12, YCVP^{+}19]. Bunges [Dro15, Obr05, Sad04].

C [Bun86b, SL19]. C. [SB92]. C1428EGA [Fonxx]. cafe [Bun06h, Bun06a, Bun06h, Bun06a]. Cajas [Bun85b]. California [Bun74a]. called [BvL91, Bun70e]. Cambridge [Ano59, Bun53, Bun07c]. Camilo [Slo99]. campus [BD78]. Can [Bun12b, Bun12c, MB16, Obi19, Pan90, Bun88c, Bun91a, KK07]. Canguilhem [Bun76b, Bun76c]. cannot [Bun88j]. caos [Bun95o]. Capitalism [Bun01-31, Bun99a]. Cápsulas [Bun03a]. Capsules [Bun03a]. Care [Bun91f]. caricatura [Bun91l]. caricature [Bun91l, Bun98g]. Carlos [GKvWZ97]. cas [Bel00, Bel01, Jod13]. Case [MB16, Bel00, Bel01, Bli91, Jod13, Mil93]. Cat [Bun01-28]. catalán [Mor16, Mor16]. Causa [BWM^{+}04, Tor14]. Causal [Bun55b, Bun59a, Bun63b, Bun70b, Din60, Duc60a, Duc60b, Hes60, Hin61a, Hin61b, Ing19, Mor61, Riv19, Bun56f, Bun70c, Bun70i, Bun81c, Cha69]. causale [Bun70c, Bun70i]. causales [Cha69]. Causalidad [Bun61i, Bun65b, Bun72h, BR97]. causalità [Bun70c, Bun70i]. causalité [Bun71c, B^{+}71]. Causality [Art19, Bun55b, Bun59a, Bun61c, Bun62c, Bun63b, Bun70b, Bun79c, Bun81c, BR97, Bun01z, Bun09b, Sch61, Bun61i, Bun62g, Bun65b, Bun71c, B^{+}71, Bun72h, Bun81j, Tul95, Bun61i, Bun62g, Bun65b, Bun70c, Bun70i, Bun72h, Bun72d, Bun87c, BR97, Bun73b, Ano59, Din60, Hes60, Hin61a, Hin61b, Mor61]. Causality. [Duc60a, Duc60b]. Causation [Ore22b, Bun68c]. cause [Tor14]. causes [Esp92]. Cellerier [Bun90e]. Centenary [BS96, Mat19a]. centrality [Bun04b]. Centuries [Bun03h]. Century [BS79, KFT89, PS94, Ano81, Sin81, Whe80]. Cerebral [Bun79d]. Challenge [Bun77d, Bun01i, Bun15e]. Challenges [Bun06c, Mad94, Qui14]. Chance [Bun51b, Bun61c, Kar19c]. change [Bun15e, Bun99b]. Changes [Göt95]. chaos [Bun95o]. Charlatanism [Bun96c, Bun96d]. Chasing [Bun06b, Bun14b, Ric08, For07, Tri06]. Chemistry [Bun82b, Vel22]. cherish [Bun91k]. Chicago [BD78]. Chichester [Bun84a]. Choice [Bun95b, Bun95c, Bun01w]. Chosen [Pot70, Pot92]. Christian [Bun79h]. chroniques [Flø86, Kli68]. Ciba [OW79]. Cien [Bun06h]. Ciencia [Bun80s, Bun81i, Bun93e, Bun96a, BR97, Bun97a, Bun98h, Bun13e, Bun14c, Aga14, Agu06, Ano16, Bun51c, Bun57b, Bun58f, Bun60i, Bun60c, Bun61i, Bun63f, Bun65b, Bun65c, Bun68q, Bun72h, Bun73f, Bun80t, Bun81m, Bun85p, Bun86j, Bun87j, Bun91l, Bun98j, Bun13c, DM00, Kan15, dP16]. Ciencias [Bun95d, Fonxx, Bun76k, BO79, Bun95k, Bun95o, Bun98i, Cha06, CCB^{+}90]. Científica [Bun73d, Bos14, Bun69h, Bun76l, Bun83h, Bun89i, Bun00h, Qui14, Roj02]. científicamente [Bun57b]. Cientificismo [BA17]. Científico [Bun17d, B^{+}99, Bun17g, Cha06]. cinemática [Bun60d]. Circle [BD78]. Circumstances [Bun86c, Bun86a]. Ciudad [Fonxx]. Civilizations [Bun04a]. cladism [Bun87h]. Clarifying [Bun04c]. Classical [Bun77d, Bun01i, Bun10d, BK75]. Clifford [Bun86a, Bun86c]. Cliffs [Fea64, Mul63]. Climate [Bun12d, Bun15e]. Clinical [Ahm19]. Close [Bun97g]. cloth [BS96, Sle96]. cloud [Bun93d]. Co [Bun74a, Bun11a, Wie99]. Cognition [Bun10e, Hal19, Bun90b]. Cognitive [Bun01y, GG86, Sle12b, Sle19]. cohesion [BGS76]. Colin [BS96]. College [BS96]. Collins [Bun99b]. Colloque [Bun76b, Bun76c, Bun78d]. Colloquium [Aga91, Bas71, LM68, WZ70]. combining [Bun76f]. come [Bun14a]. command [LB78b]. Comment [Bun83a, Bun85c, Erl67, Tul95]. Commentators [Bun03g, Bun85i, Bun91j]. Comments [Aga88a, TBG^{+}85, Nar75]. Communication [Bun68b]. Community [KFT89]. Cómo [Bun45b, Aga14, Ama14, Bun95o, Bun96i]. Companion [NSB10]. Company [Bow80, Bru73, KPSP74]. compare [Kla03b]. compared [BK75]. Comparison [Kla03a, Kla04]. compartilhamento [Mac11]. compatibility [Woo96]. compatible [Jod13, MB96b]. compatibles [Jod13]. competition [Bun76f]. complementaridad [Bun55g]. Complementarity [Bun55e, Bun55f, Bun55g]. Complex [Ort19]. Complexity [Bun61d, Bun62d]. components [BL10]. comprender [Sca14]. computation [CD90]. Computer [Bun10c]. Computerism [Bun90b]. Computers [Bun56b, Bun56c]. con [SC89]. Concept [Bun10h, Bun74c, Bun79l, Bun93b, Bun04g, LB78b]. Conceptología [Bun76d]. Conceptology [Bun76d]. Concepts [Bun01-29, Bun12p, WW90, Bun68e, Bun72f, Bun81f, Bun19e, Bun58a].

Conceptual [Bun81d, Bun13g, MM19, Bun75e]. concerning [Bun70j]. Conclusion [Bun80b, Bun12e]. Concretas [Bun97h]. Concrete [Bun85k, Bun97h, Bun85o]. concreto [Bun85k]. conditions [Sil13e]. conference [Van04]. conferencia [Ano09]. Confirmation [FM66]. Confusing [Bun73j, Bun73k]. Congrès [Die78, Cau88]. Congreso [Gru05]. Congress [Die78, Mon92]. conhecimento [Mac11]. Conjoining [Mah21]. Conjonction [Bun71c]. Conjunction [Bun68c, Bun71c]. Conjuntos [Bun85h]. Connection [Bun99k, Bun13i, Sad04, Tre01]. connections [Bun60f]. conscious [Bun79d]. Consciousness [Bun80c, Bun10e, MB19, PB89]. conservatism [Bun03b]. conservatisme [Bun03b]. consideraciones [Yan14]. Considérations [Bun03b, Yan14, Bun03b]. considered [Bou91]. Construction [Bun61h, Bun66f]. Constructive [Bun20a, Sol22]. Constructivism [Col01]. Constructivist [zW11a]. Constructs [BL71, RG02]. contemporain [Sil13d]. contemporaine [Flø86, Kli68]. contemporaines [Bun74d, Bun75b]. contemporánea [Bun85l]. Contemporary [Bun94b, Bun01k, Flø86, FvWK03, Kli68, Mac85, Sle12b, Sle19, Bun74d, Bun85l, LB78a]. Content [Bun67j, Bun13j]. Context [Bun01-33, SW09, WS06]. Contingency [Ray19]. Contra [OB84, ABA19]. Contrasting [Del19]. Contribute [MB16]. Contributions [Vol93]. Contributors [Ano96a]. control [Bun96i]. Controversial [Hob19, Cha06]. Controversias [Bun82c, Bun83b]. Controversies [Ray15, Bun87h, Bun82c, Bun83b]. Convergence [Bun03c, Kil04, Pic07]. Conversations [Bun71a]. Cooper [Bun74a]. cooperation [Bun76f]. Copenhagen [AF01, Góm14]. Copenhague [Góm14]. corporate [BC14]. corps [Bun87d]. Correct [Bun69c, Cup19]. Corrections [Bun69c, Str69]. correlates [Bun79d, GBS^{+}05]. correspondence [Bun12f]. Corrigenda [Bun71b]. cosmologia [SA93]. cosmological [Bun93d]. Cosmology [Bun62e, SA93]. Counter [Bun94b]. Counter-Enlightenment [Bun94b]. coupled [SBC^{+}08]. Course [Bun97b, Bun09c, Bun80d]. Covariant [BK69]. Creativity [Bun01f]. credo [Mos91]. Crime [SW09, WS06, Ama14, Bun06g, Bun09g]. crise [Bun04e]. Crises [Bun12c]. Crisis [Bun99a, Bun01s, Bun04e, Bun09e]. Criteria [Bun61f]. Criterion [Bun12e]. Crítica [Ano16, Bun75i, Bun55g, Par83, dP16]. Critical [AJ87, Apo85, Bun64a, Bun75a, Bun91b, Bun92a, Bun95c, Bun99c, Bun01w, JL95, Par83, Vol93, Mar96, SL19, Bun55g, Gam14]. Criticism [Bun86a, Bun86c, Bun20a, Bun75i]. criticisms [Bun77j]. critics [MB96a]. Critique [Bun66b, Bun92d, Sol22, Bun56a, Ano16]. Crusade [Bun19b]. Crystal [Vel22]. cuántica [Vuc14]. cuatro [B^{+}99, Bun04f]. cuestión [Cha06]. cuestionamientos [Moy15]. Cultural [TBG^{+}85]. Culture [Bun86d, Bun00b, Pic19, Cau88, Mar81, Bun86d, Bun94a, HBB^{+}94, Cau88]. Cultures [Pic19]. Curiosidad [Bun98a]. Curiosity [Bun98a]. Current [Bun68m, Bun86g, Bun87i, Ing19, Agu06, Bun95g]. Curriculum [Ano14a]. Curso [Bun97b, Bun09c, Bun80d].

D [Bow80, Bru73, Bun84a, Bun84b, KPSP74]. Dagum [Slo99]. dangerous [Tho91]. dangers [Bun19f, LB78b]. dans [Bel00, Bel01, Bun78d]. Dark [Bun19a]. Darwin [BB89]. Data [Bun12a]. David [Jun19]. Davies [Sle96]. Dawson [Sin81]. Dead [Bun01-28]. Debate [Bun86g, Bun98f, Col99, Lun00, Bun79e]. decaying [BK83a, BK84]. December [OW79]. Decision [EB88, Bun73c]. dedicated [Mor81]. Deeper [Bla92]. defence [Bun86e]. defensa [Roj13, Sil13f]. Defense [Bun61b, Mad94, Mue68, Bun15b, Roj13, Sil13f]. Definicões [QdAY13]. definite [Bun71e]. Definition [Cav12]. definitions [QdAY13]. Degrees [Bun85d, Bun85f]. dejarse [Bar14]. del [Ama14, Bar11, Bar14, Bun52, Bun55g, Bun55c, Bun60d, Bun70c, Bun70i, Bun95o, BA17, Cha06, Ech85, Mar72, Par83, SA01]. Delaware [Bun67a, Bun13b, Aga72, Sne68a, Sne68b]. delito [Ama14]. Demarcating [Bun83d]. demarcation [Bli91]. Dematerialization [Bun19b]. Democracias [Bun17a]. Democracies [Bun17a]. Democracy [Kar19b]. Denken [Pfe77]. Denmark [AF01]. Deomocracy [Bun05a]. dependent [Bun90g]. Derecho [Bun96i, BC14]. deren [Kla03a]. derrière [Bun06d]. Desarrollo [Bun97a, Bun14c, Bun80s, Bun98h]. Desarrollos [Agu06]. descriptions [Bun71e, BG16]. Desiderata [Bun87f]. designs [Sca14]. Destructive [Bun20a]. determination [Bun68c, Bun71c, Bun71c]. Determinism [Bun12h, Bun12i, Bun13d]. déterminisme [Bun13d]. Deutschen [Mon92]. deux [Bun16b]. Developing [RG02]. Development [Bun81e, Bun97a, Bun12y, Bun14c, SW09, WS06, Bel00, Bel01, Bun74i, Bun80s, Bun98h, Bun12x]. Developments [LK74, Agu06]. développement [Bel00, Bel01]. d’explication [Cha69]. Dez [HD88]. d’histoire [Bun76b, Bun76c]. diagnosis [Sei06]. diagrams [Ano59]. Dialéctica [Eng47, Moy15]. Dialectical [Del19, Moy15]. Dialectics [Bun75a, Bun01k, Apo85, Eng47]. Dialectiques [Per75]. Dialogue [Bun75b]. Dialogues [Bun54, Bun01p]. Dictionary [BKR96, Bun99d, Bun03e, Bun20b]. Dictionnaire [Bun20b]. Did [Bun07d]. Diemer [HD88]. difference [Per91]. differences [GBS^{+}05, Mat82]. Differentiation [BGS76]. Digital [Bun12q]. dignidad [CFF^{+}14]. dignity [CFF^{+}14]. Dimensional [Bun58c]. Dimensions [AD04, TBG^{+}85, Bun78b, Bun78h]. Dinge [BM04]. Dirac [Bun52, Hin77]. direction [Bun76b, Bun76c]. disabilities [Mil93]. Disarmament [LP88]. disbeliefs [Bun91i]. discipline [Bun90g, Roy90]. Discourse [Mau22a]. discovery [Bun59c]. discusión [Bun98i]. discuss [Bun06h, Bun06a]. discussion [Bun98i]. Discussions [Bas71, WZ70]. discutir [Bun06h, Bun06a]. disease [BT^{+}03]. diseños [Sca14]. dispensable [Bun04g]. Dissecting [Bun07b]. Dissident [Jun19]. dissimulation [Van91]. Dissolution [Bun69f]. distancia [Bun85b]. distant [Ano15]. distinción [Cha06]. distinction [Cha06]. distinguish [Bou91, Bun91j]. Distribution [Slo99]. Do [Bun56b, Bun56c, Bun01k, BK83a, BK84, Bun02b, Kla03b, Kla04, Mac11].

Doctors [Lab14, AC13]. doctrines [Bun76b, Bun76c]. Documents [Bla92]. Does [Bun04d, Bun12g, Bun12h, Bun12i, Bun15a, Bun91j, BGD13, Bun13d]. dogma [Bun79h]. dogmatism [Bun00a]. Doing [Bun17b]. Don’t [Bun17f]. Dordrecht [KPSP74, Mat16]. Dordrecht-Holland [KPSP74]. Double [Bun12k]. Doubts [Bun97c]. Dr [HD88, KPSP74]. Dr. [Mar72]. Dual [Kah19]. Dualism [Bun10c, Bun79a, Bun89g]. Dualismus [Bun89g]. Ducret [Bun90e]. d’un [Bun03b, Bun16b]. d’une [Sil13e]. durable [Bel00, Bel01]. Duties [Bun01-27, Bun91h, Bun94e]. dynamics [BP94, Bun76i]. dynamique [BP94].

Early [Bun04a]. Earth [BS96]. easy [Fey91]. ecología [BM98, GML14]. Ecology [CD19, dSMdC19, BM98, GML14]. Econometrics [Slo99]. Economía [Bun82a, OB84, Cor86, Dri87]. Economic [Bun09e, Bun12c, Bun12y, Bun12x, Den03]. economics [Bun03b]. économique [Bun03b]. economy [OB84, Bun82a]. ecuaciones [Bun43c]. Ed [Sne68b, BS96, KPSP74]. Edad [Bun55c]. edición [Cor86]. Edited [Sin81]. Editorial [Ano90, Ano96b, Mat96, Mau20, Moe91, Mun93]. Eds [Bun87a, Bun07c]. Eduardo [Bar14]. educación [BdCFG^{+}96, Bun43b, BSO^{+}96, OSV96]. Education [Bun43b, Del12, Hal19, Mat12, Mat19c, OSV96, BdCFG^{+}96, BSO^{+}96, MB96b, Poo96, Set96, Woo96, WL96]. Effect [Bun15a]. eighth [Har77]. Einführung [Dro15]. Einsatz [Vol93]. Einstein [Bun79e, Bun51a]. Electron [Bun50, BK69, Bun55a, Bun60d, Hin77]. Elementos [Par83, Ech85]. Elements [Sen93, Par83]. elle [Bun13d]. Ellis [Bun84a]. Elogio [Bun98a, BA17, Den14a, Den14b]. Emergence [Bun77a, Bun01c, Bun01-30, Bun03c, Bun10f, Kai09, Kal19, Kil04, Jod13, Pic07]. Emergent [Bli90b]. Emergentist [Del12, Del19, zW11b, Obr05]. emergentistische [Obr05]. Empirical [VOHPMJ12, Zie22, Zie16]. encarar [Bun57b]. Encounters [Bun71a, Fra18, SC89]. Encouraging [BAG17]. Encuentros [SC89]. Encyclopedia [Kal13, SB01]. End [Bun99e]. Endangered [Bun99a]. Energy [Bun00c, Bun01d, Bun12p, Bun03i]. enfoques [GML14]. Engage [Obi19]. Engels [Bun11a]. Engelsian [Del19]. engenharia [Mac11]. engineering [Bel00, Bel01, Mac11]. Englewood [Fea64, Mul63]. engullir [Bar14]. Enhancing [RW05]. Enlightened [Bun06c]. Enlightenment [Bun94b, Mat19c, Mad94]. Ensayos [Den14a, dF78, Bun89j, Bun12z]. entero [Bun58g]. Enterprise [AD04]. entities [Bun60a]. Entre [Bun14e, Bun16b, Bun94d, Bun95o, BG01, Cha06, Jod10]. Entretien [Bun87e, Lan89]. Entretiens [Vac93]. Entrevista [Ano15, Cor14, QB79]. Entrevistas [Bun87i, Bun97h]. entropie [Jod13]. entropy [Jod13]. Environment [BS96]. Epiphenomenon [MB19]. Epistemologia [Bun97b, Bun09c, Bun77b, Bun80d, Bun44d]. epistemological [BG16, Har91, Sca14]. epistemológicas [Sca14]. Épistémologie [Bun83e]. Epistemology [Bun77b, Bun85a, Bun83f, Bun83g, Bun83e, Bun85e, Bun01e, Bun13k, Bun15d, CD19, Esp86, MB16, Pie99, Res19, Rus86, Bun44d, Bun69e, Rus73, Bun80d, Bun97b, Bun09c]. equalization [Bun01-35]. equals [Bun00a]. equations [Bun43c]. Era [Ray19]. Ernst [Bla92]. Erratum [Bun75b]. Erwin [Bun53]. escolástica [OB84]. espace [Bun91d]. esprit [Bun87d]. Essays [AC82, BAG17, Bas71, Bun64a, Bun86a, Bun86c, Bun99c, EB88, GPA91, JL95, Mah01, MC95, Mor81, SD91, Slo99, Bun89j, Bun12z, CH09, Coh81, dF78, Den14a, Bun60b]. estadística [Pin13]. Estado [Bun95g]. estatuto [Cha06]. estrategia [Bun69h]. estudio [Bun39, Cha06]. estudios [B^{+}99, Bun04f]. états [Jod13]. Ethical [AD04, Kar19a, Kar19b]. Ethics [Bun61e, Bun89c, Bun99f, Kar19a, Kar19b, Kra80, Obi19, Bun78c, Bun78i, Bun95f, Bun07e, KK07, WZ70, Bun60c, Bun96a]. Ethics. [Bru91]. ethnography [Cri14]. Etica [Bun60c, Bun96a, Bun95f]. etnografía [Cri14]. Eugenio [GKvWZ97]. EUR [Mat16]. European [BCD03]. Evaluando [Bun14f]. evaluate [Bun04g]. Evaluating [Bun12j, Bun17c, Ray19]. Evaluation [Bun78a, Bun78g, OHS02]. events [Bun77k]. everyday [Per91]. Everything [Bun03g]. Evidence [Bun01-30, Bun18d]. evolución [Lab14]. Evolution [Bli90b, Bun97g, BRG04, Bun78e, Lab14]. Evolutionary [BB89, Bun79g, Bun87h]. evolve [Bun81b]. Exact [Bun73e, Var17, Dro15, KPSP74]. Exactas [Fonxx]. exactness [Bun91k]. exakte [Dro15]. examen [Sil13e]. Examination [Apo85, Bun75a, Bun91b, Bun92a]. Examining [Sil13e]. Exercise [Bun69d, Bun80e]. Existence [Bun81d, Bun12k, MM19, Bun85d, Bun85f, Man85]. existencia [Bun85f]. Existió [Bun09d]. exorcize [Fey91]. expanding [SL19]. Expeditions [CC13]. Experience [Bun70f, Bun79d]. Explain [Bun12b]. Explaining [Bun01f]. Explains [Bun03g]. Explanation [Bun97d, SW09, WS06, dSMdC19, Bun78a, Bun78g, Bun85n, Bun98d, Bun13f, Cha69, GML14, BM98]. Explanations [Riv19]. Explanatory [Bun04d]. explicación [BM98, GML14]. Exploring [Bun83f, GBS^{+}05]. Exposición [Bun55g]. exposition [Bun55g]. Expressiveness [RW05].

F [Ech85]. Faces [Bun01-36, Bun88i]. Fact [Bun09f, Bun88g]. Facts [Bun04f, Bun75e]. Facultad [Fonxx]. Failed [Bun09e]. failure [Bun44c]. Far [Bun97g]. Fare [Bun01k]. Features [Vel22]. fécond [Bun91e]. Fehr [Bun07c]. fenomenología [Bun51c]. Fernando [BdCFG^{+}96, OSV96]. Ferrater [Coh81]. fertile [Bun91e]. Festschrift [GKvWZ97, Mat19a, Wie99, FJMT88]. Fiat [DB14]. Fiction [Bun09f, Mar19, Bun88g]. Fictionism [Bun01m, Bun97e]. Fictitious [Bun70h]. Fields [Bun10d, Bun18d]. Filosofar [Bun57b]. filosofi [Bun66a]. Filosofia [Bun67g, AC13, BdCFG^{+}96, Bun67h, Bun95h, Bun95k, Bun95i, Bun95j, Bun95e, BSO^{+}96, Bun08b, Bun12z, OSV96, BA88, Bun12l, Aga14, Agu06, BC14, Bun44c, Bun60i, Bun63f, Bun68q, Bun69h, Bun73f, Bun75j, Bun76k, Bun80t, Bun81m, Bun82a, Bun83i, Bun85p, Bun85l, Bun86j, Bun87j, Bun95g, Bun99j, Bun07f, Bun13c, Cor86, DM00, Kan15, Mah14, Moy15, Roj02, BG01, Dri87, Sil13f]. filosofías [Bun14f]. filosófica [Bun98i, SB92]. filosóficamente [Bun57b]. filosófico [Bun95g]. filosóficos [Moy15, Qui14]. filósofo [SA01]. Finding [Bun96b, Asp98b, Asp98a, Bra97, Dem97]. fines [Sca14]. Física [Bun82c, Bun58g, Bun83b, Bun95i]. Físicas [Fonxx]. físico [Bun43a, Bun43c]. Five [Bah81, Bun03h, Bun79f]. Flight [GLL96]. Florentino [Bun45b]. Folkestone [Sin81]. fonction [Sil13b]. Forces [Bun44a, Bun44e]. forecast [Bun73p]. foresight [Bos14]. Foreword [Bou99, Bun72c, Bun19c, Mun93, Par14]. formais [QdAY13]. Formal [Esp86, WW90, Ech85, Par83, QdAY13]. Formalization [Man85]. formes [BP94]. forms [BP94]. Formulation [Bun57a]. Foundation [OW79, WW90]. Foundations [Aga72, Bun67a, Bun67d, Bun71h, Bun95c, Bun01w, Bun13b, HH76, LSP22, MB97, MB00, Met73a, Met73b, Sne68a, Sne68b, Har91, Sei06, Bun69c, Haj76, Str69]. Four [Bun69d, Bun81f, Bun84b, Bun04f, B^{+}99, Esp92]. fourth [Bun70e]. fracaso [Bun44c]. framework [Mat82]. Francisco [Bun74a, Bun79b, Bun01-38, SB92, SA01]. Frank [Bun53, Bun51a]. Free [Bun01-30, Bun10e, Bun67e, Bun76e]. Free-Wheeling [Bun01-30]. Freedom [Bun97c]. Freeman [Bun74a]. Fregean [Bun75i]. fregeana [Bun75i]. French [Bel00, Bel01, BP94, Bun68e, Bun68l, Bun68d, Bun71c, B^{+}71, Bun74d, Bun78d, Bun83c, Bun83e, Bun86d, Bun87d, Bun87e, Bun89b, Bun90c, Bun91d, Bun91e, Bun91c, Bun94d, Bun94c, Bun96f, Bun98g, Bun03b, Bun04e, Bun04i, Bun06e, Bun06d, Bun08c, Bun13d, Bun16b, Bun19e, Bun20b, Cha69, Del13, Esp92, Esp94, Jod10, Jod13, Lan89, Mah13, Moe08, Rus73, Sil13a, Sil13b, Sil13c, Sil13e, Vac93]. frequentist [Bun56a]. Freudenthal [Sal71]. Friedrich [Bun11a]. Front [Ano10b, Ano10c, Ano12b, Ano12c]. fruitful [Woo96]. fuerzas [Bun44e]. Fugitive [Bun86a, Bun86c]. Function [MB01, Sil13b]. Functionalism [MB01]. Functionalist [Bun12b]. Functions [Bun80f]. Furniture [Bar81, Bun77g, Vol79, Rus80]. Future [SK09, Bun09d, BGD13]. fuzzy [Sei06].

Galilean [MB16]. Game [Bun89e]. Games [Bun68m]. Ganze [Bun71a]. Gap [Bun01h, Bun94d]. García [G.73]. Gedenkschr [HD88]. gegenüber [BK75]. Gegenwart [Wie99]. Gehirn [PB89]. General [Bun63c, Bun77c, Bun10h, Bun11a, Mau22a, Ore22a]. Gentle [Aga88a]. Genuine [Bun10i, Bun11b]. Geometry [GSB81, SBC^{+}08]. George [Bun68b, Bun76b, Bun76c, Bun99a, Har77]. German [Bun00b, Mon92, Bun69a, Bun73i, BK75, Bun87c, Bun89g, BA90, Bun92b, BM04, Dro15, Fra18, HG75, Kla03a, Kla03b, Kla04, MB00, Mon92, Nar75, Obr05, Pfe69, Pfe77, PB89, Vol93, Wie99]. Germany [AL99]. Geschichte [Bun87c, Wie99]. Gesellschaft [Mon92]. gewidmet [Mor81]. Gezeitenwechsel [Dro15]. ghost [Bun67e]. ghosts [Fey91]. Gifford [BS96]. Gintis [Bun07c]. Glasgow [BS96]. Global [Bun99a, Bun99b, Bun06c, dCPM94]. globalización [Bun01-35, Bar14]. globalization [Bun01-35]. Gnoseología [Bun95l]. Gnoseology [Bun95l].

Go [Bun14g]. Goals [Bun73e, Sca14, KPSP74]. God [BS96, Sle96, BS96, Bun84e]. González [Bar11]. Gonzalo [Bar14]. Good [Bru91, Bun89c, Bun89b, Bun90c]. Gould [BB89]. Grados [Bun85f]. Gradualism [BB89]. Graham [Ore22b]. grandes [Bun39]. Gravitational [Bun18b, Rom18, Rom17]. great [Bun39]. Growing [Bun87b]. Growth [Bun12y, Bun12x]. Grundlagen [MB00]. Gruyter [Wie99]. GST [Bun77d, Bun01i]. Guide [Bun05b]. guilty [Bun88c, Bun91a]. Güiraldes [Fonxx].

H [Bun87a, Bun07c]. Habermas [Bun74b]. hacer [Bun02b]. Hacking [Col01]. Half [Bun81h]. Hall [Fea64, Mul63]. Handbook [DB93]. Hans [Bun60b, Fra18, Bun18c, Fra18]. Hao [Bun76a]. hardback [Bun11a, Mat16]. Hardware [Bun10c]. Hardware/Software [Bun10c]. Harman [Ore22a, Ore22b]. Harper [Bun71a]. Harvard [Ano59, Din60]. Haven [Asp98b]. Hay [Bun75j]. headless [Bun83c]. hechos [Bun75e, Bun04f]. Hedström [Bun07b]. hegemonic [Bos14]. hegemónica [Bos14]. Heidelberg [Bun87a]. Heisenberg [Bun71a, Bun77e, Pfe77]. Held [BAB^{+}80, OW79, Har77, Van04]. help [KK07]. hence [Bun14a]. Henry [Bun11a]. Hermann [Bun76b, Bun76c]. Hermeneutics [Cup19]. Hidden [Bun84c, Bun95a, Bun01j]. hidrógeno [Bun52]. highlights [BG16]. Hispanic [Bun86i]. historia [BdCFG^{+}96, OSV96]. Historic [Tor14]. histórica [Tor14]. historical [Bun43a, Bun43c, BG16, Ray15]. histórico [Bun43a, Bun43c]. History [BD78, Bun58a, Bun87c, KM70, MB16, MC95, OSV96, Sin81, Wie99, BdCFG^{+}96, Bun88f]. Hitler [Bun00b]. Holism [Bun77c, Bun00b, Bun01a, Bun04h, Bun79l, Bun88b, Bun00f]. Holland [KPSP74]. Holt [Bun11a]. Homage [Fra18, Den14a]. homenaje [BdCFG^{+}96, DM00, Den14a, OSV96]. Homenaxe [Gru05]. Hommage [Fra18]. Honor [AC82, EB88, MC95, Slo99, Coh81]. Honoris [BWM^{+}04, BWM^{+}04]. Honour [Bun64a, Bun99c, GPA91, SD91, CH09]. Horwood [Bun84a]. Hudson [Har77]. Hugues [MC95]. Human [Bun69d, CC13, GPA91, dCPM94, AK98, Bun99g, CFF^{+}14]. humana [CFF^{+}14]. humanidades [Bun95o]. Humanism [Bun53, Bun04e, KFT89, SB92]. humanisme [Bun04e]. humanismo [SB92]. Humanist [Bun97c]. Humanistic [Obi19, Bun90g]. Humanities [Bun76a, Bun95o]. Humanity [BS96, BS96]. Humboldt [Bun69a, Pfe69, Bun69a, Pfe69]. hundred [Bun06h]. Hunt [Bun11a]. hydrogen [Bun52]. Hylas [Bun54, Bun01p]. Hylorealistic [Vel22]. hypothesis [Bun03d].

Ian [SL19]. Iatrofilosofía [Bun78j]. Iatrophilosophy [Bun78j]. Ideals [Obi22]. Ideas [Bun06a, Bun06h]. Identitas [Gam14]. Identity [Gam14, VOHPMJ12, Bun80n]. Ideología [BO79, Bun13h, Bun85m, Bun86k, Bun14h]. ideological [WL96]. Ideology [Bla19, Bun86i, Bun88d, Bun01-30, Bun85m, Bun86k, Bun13h, Bun14h, Per91, BO79]. Idiot [Bun86a, Bun86c]. Ignore [Bun00e]. igualamiento [Bun01-35]. II [Bow80, Bun00b, Edg69, Esp86, McF78, Qui80, Bun55f, Bun56c, Bun79j, Bun83g, Bun13k, Kar19b]. III [Esp86, LM68, Bun85e, HH76]. Illinois [BD78]. illusion [Tho91]. im [Kla03a, Vol93]. image [Bun58g]. imagen [Bun58g]. impact [SL19]. impediment [Tur96]. implications [Obi11]. Imply [Bun01-27, Bun91h]. Impopulares [Bun87i]. Importance [Bun92b, Alc91]. Imprecision [Bun01-30]. including [Har77]. Income [Slo99]. incompatibility [MB96a]. Inconsistent [Bun68p]. incorrect [Bun69c]. inculture [Bun86d]. independence [Mor16]. independentismo [Mor16]. indeterminacy [Bun70e]. Indicator [Bun75h]. Indicators [Bun81e, Bun74i]. Indigenous [dCPM94]. Individual [Bun85k, Bun85o]. Individualism [Bun00g, Bun01a, Bun04h, Bun79l, Bun88b, Bun00f]. Individuals [Bun66c, Bun85g, Del12, Bun81b, Bun85h]. individuo [Bun85k]. Individuos [Bun85h]. Indonesian [Gam14]. inducción [Bun60j]. Induction [Bun60g, Bun01l, Bun60j, WZ70]. Inductive [Bun12g, FM66]. Inequalities [Bun77e, Bun89a]. Inexhaustible [Bun50]. Inference [HH76]. influence [Pfe77]. informal [Ano15, Ano15]. Informatics [Bun12m]. Information [Bun97c, LSP22, PW80, WW90, Bun79h]. informes [Bun76b, Bun76c]. Ingasei [Bun72d]. ingénierie [Bel00, Bel01]. innocent [Bun88c, Bun91a]. inobservables [Jod13]. inputs [Bun78f]. Inquire [Bun97c]. Inquiry [Bun10k, Res87, Sle12a]. Insight [Bun68a]. Insights [Art19]. Inspired [CC13]. Instant [Bun90d]. Institut [Bun76b, Bun76c]. Instruments [Bun10r]. Integral [Bun12y, Kar19b, Bun58g, Bun12x]. Integration [Bun89g, Bun01-37, Bun11d]. Intellectual [Bun99b, Dro19, Jod10]. intellectuel [Jod10]. intelligibilité [BP94]. intelligibility [BP94]. Intendente [Fonxx]. inter [Bun70j]. inter-theory [Bun70j]. Interdependence [dCPM94]. Interests [Bun07c]. International [AL99, AF01, DB93, LM68, SW89, SB01, Pfe69, Bun78d]. internationale [Bun78d]. Internationalen [SW89]. Interpretación [Bun52, Bun56f, Góm14]. Interpretando [Bar20]. Interpretation [Bun57a, Bun74g, Bun77e, Bun03d, McF78, Bun56f, Góm14, Bun52]. Interpretations [Bun56e]. Interpreting [Bar20]. intervals [BK83a, BK84]. Intervention [zW11a]. Interview [Bun87e, QB79, VOHPMJ12, Ano15, Moe87, Cor14, Lan89]. Interviews [Bun87i, Bun97h, Qui79, Vac93]. Intolerance [Bun96c, Bun96d]. Introducción [Bun39, Aga14]. Introduction [Bun80g, Bun80h, Bun02a, Mat12, Mat19b, Aga14, Bun39, Dro15, Bun62b]. Introductory [Ano04]. Intuición [Bun65c, Bun86f, Bun96e]. Intuition [Bun62f, Bun67b, Bun75c, Bun96f, Scr63, Bun65c, Bun86f, Bun96f, Bun96e, Fea64, Hin64, Lon62]. Inverse [Bun19d]. investigación [Ama14, Bun69h, Bun76l, Bun83h, Bun89i, Bun00h]. investigation [Bun76l, Bun83h, Bun89i]. irrationalists [Bun84b]. irreversability [Bun70g]. irritantes [Bun89j]. irritating [Bun89j]. ISBN [Bun87a, Kil04, Mat16]. Islamic [Obi11, Obi19, Obi22]. Issue [Ano04].

Issues [Bun43b, Bun13g, Hob19, Mac85, She76]. Italian [Bun70c, Bun70i, SA93]. IV [Qui80].

J [Bun72a, Bun72b, Bun74a, Bun04g, G.73]. J. [Coh81]. Jammer [Bun58a]. Japanese [Bun72d]. Jarvie [SL19]. Jean [Bar14]. Jersey [Mul63]. John [Bun84a]. Joseph [BAG17, JL95]. judgments [KK07]. June [AF01]. Jung [McG94]. Jurgen [Bun74b]. Just [Bun12t]. justification [Bun98d].

Kálnay [Hin77]. Karl [Bun64a, Bun99c]. Karlsruhe [AL99]. Kausalität [Bun87c]. Kegan [Bun76a]. Key [Art19]. kind [Bun90g]. Kinds [Bun61f, BS77]. kinematics [Bun60d]. Kirchberg [SW89]. Kitcher [Bun05a]. klassischen [BK75]. Klibansky [GPA91]. know [Bun02b]. Knowing [Bun80r]. Knowledge [Bun03c, Bun10i, Bun11b, Bun12n, Kil04, Mac11, Per91, Pic07]. Kongress [Mon92]. Kritis [Gam14].

L [G.73]. Labourer [Aga12]. l’Académie [Bun78d]. lack [Bun86d]. ladder [Bun93d]. Lagrangian [Bun57a]. Language [Bun86g, CCB^{+}90, Wie99]. Languages [Sil19, Wie99]. latest [Bun91l]. Latin [FvWK03, NSB10]. Law [Bun61c, Bun12t, Obi19, BC14, Bun96i]. Laws [Bun61f, Bun61g, Sad04, Bun68o]. Learning [Bun80i, Mil93, Bun87a]. Leaving [Kar19c]. Leblanc [MC95]. L’écart [Bun94d]. lecture [Ano09, Nar75]. Lecturers [Ano75]. Lectures [BS96]. Legacy [Jun19, Bun88a, Jod10, SL19]. Legal [Bun01n, GKvWZ97]. Legitimacy [Bun01n]. Leibnizian [Art19]. Leinfellner [EB88]. leisten [Kla03b, Kla04]. Leite [FJMT88]. lejano [Ano15]. L’émergence [Jod13]. lenguaje [CCB^{+}90]. L’épistémologie [Rus73]. less [Cup19]. Let [Moe87]. Level [Bli90b]. levelmanship [Bun80e]. Levels [Bun60e, Bun77f, Bun10f, Bun60f, Bun69e, Bun89g]. L’héritage [Jod10]. liberalism [Bun03b]. libéralisme [Bun03b]. Liberty [For89b]. libro [Bun06h, Bun06a]. lieu [Bun91d]. Life [Bun51a, Bun75h, Bun11a, Esp86, Mat16, Mat19b, CH09]. Light [Bun17b, BG16, LB78a]. Limits [Bun01v, Göt95, Sad04, Van04, Bun91g, Bun09a]. line [Bli91]. Linguistic [Gil19, Zie16, Zie22]. Lingüística [Bun83i]. Linguistics [Bun84d, Bun86h, Bun99h, Tul95, Bun83i]. Little [Kar19c]. Logic [Aga91, Bun74e, Bun12d, Bun12g, FM66, Iva19, MC95, PS94, Bun59c, Bun77i, CCB^{+}90, Kon95, Qui58, BM92, SB92]. Lógica [BM92, SB92, CCB^{+}90, Qui58]. logical [Bra82]. London [BS96, Bun53, Bun76a, Bun79b, Din60, LM68, OW79, Sle96]. Long [Mat16]. longevidad [Den14b]. longevity [Den14b]. longitudinal [BL10]. Look [Bla92, Bun95c, Bun97f, Bun01q, Bun01w, Bun71e]. Looks [Bun86g]. Lopes [FJMT88]. López [Bar11]. lost [Mor16]. Ltd [Bun84a]. lucha [Bar14]. Luhmann [Kla03a, Kla03b, Kla04, Kla03a, Kla03b, Kla04]. Luhmannian [zW11a]. Luna [Ech85].

M [Bun84a, Pic07, Rob74, Vol79, Vol81]. M. [Bun68i, Ano81]. MA [Bun07c, Rob74]. macchine [Bun66a]. Mach [Bla92, Bun51d, Bun66b, Bun68i, Bun92d, Mue68, dF19, Bun72a, Bun72b]. machine [Bun93d]. macro [Moe08, Déc10]. maestro [Bun75d]. Magic [Bun62e]. Mahner [Del12, WL96]. Mahner/Bunge [WL96]. make [KK07]. Making [Bun17c]. Man [Göt95, Bou91]. management [BRG04]. Manuel [BWM^{+}04]. March [Van04]. Margáin [Bun75d]. Maria [Bun60b]. Mario [Aga72, Ano59, Ano14a, Asp98a, Bar11, Bar81, Ber82, Bow80, Bra97, Bru91, Bru73, Bub68a, Bub68b, Cau84, Col99, Cor86, Déc10, Dem97, Din60, Dri87, Dro15, Duc60a, Duc60b, Ech85, Edg69, Esp85, Esp86, Fea64, Fin70, For07, G.73, Haj65, Haj76, Har66, Har75, Hes60, Hin61a, Hin61b, Hin64, Hof84, Kil04, Kir75a, Kir75b, Kla03a, Kla03b, Kla04, Lev83, Lin69, Lon62, Lun00, McF78, Met73a, Met73b, Mor61, Mul63, Obr05, Par81, Puc81, Qui74, Qui80, Ric08, Rus80, Rus86, Sin81, Sne68a, Sne68c, Sne68b, Tre01, Tri06, Tuo74a, Tuo74b, Vol73a, Vol73b, Whe80, dP16, AC82, Gru05, Ano74, Ano09, Ano15, Ano16, Ano23, Apo85, Art19, Asp98b, AC13, Bel00, Bel01, BC14, Bun83a, Fonxx, Bun14d, Bun17d, Cav12, Cor14, Cor12, Cor19]. Mario [Cri14, CD19, Del13, DM00, Den14a, Den14b, Den19, Den03, Dro15, Dro19, Erl67, For89a, Gam14, Góm14, Ing19, Jod10, KPSP74, Kan15, Kla03a, Kla03b, Kla04, Lan89, MB16, Mah01, Mah14, Mah21, Man85, Mar12, Mat82, Mat03, Mat12, Mat16, Mat19a, Mat19c, Mat19b, Moe87, Mos14, Moy15, Obi11, Obi22, Obr05, Ore22b, Ort19, Pat90, Pic01, Pic04, Pie99, Pin13, QdAY13, QB79, Qui79, Roj13, Roj02, Rom18, Rom19, Rus73, Sad04, Sca14, Sch61, Scr63, SC89, Sil13a, Sil13c, Sil19, Sil13f, Sle12a, Sle12b, Sle19, Sol22, Vac93, VOHPMJ12, WW90, WD90, Wet82, Zie16, Zie22]. Marler [Bun87a]. Marx [Bun11a]. Marxism [Cup19]. Marxist [Bun12o]. Masks [Bun01-36, Bun88i]. Mass [Ano59, Bru73, Bun68i, Bun44b]. mass-system [Bun44b]. Massachusetts [Din60]. Master [Aga12]. Master-Builder [Aga12]. Material [Bun07c]. Materialism [Bun81k, Bun92e, Bun01k, BM04, Bun10j, Bun12h, Bun12i, Bun12v, Sil13e, Bun04i, Bun08c, Bun13d, Del13, Sil13b, Sil13c, Bun81i, Bun04e, Bun06e, Bun13e, Esp85, Hof84]. Matérialisme [Bun04e, Bun06e, Bun04i, Bun08c, Bun13d, Del13, Sil13b, Sil13d, Sil13c, Sil13e]. Materialismo [Bun81i, Bun13e]. Materialisms [Del19]. Materialismus [BM04]. Materialist [Sle12b, Moy15]. materialista [Moy15]. Matériaux [Sil13d]. Mathematical [Bun69d, Bun01m, Mar19, Bun97e, Bun79b, Bun84a]. Mathematics [AD97, Bun01h, MC95, Mar12, PS94, Bun74a, Bun76a, Bun79b, Bun94d]. mathématiques [Bun94d]. mathématisation [Bun76b, Bun76c]. Matrix [Bun12r, Bun13a]. Matter [Ano10a, Ano10b, Ano10c, Ano12a, Ano12b, Ano12c, Bun73g, Bun10d, Bun10h, Bun10n, Bun10k, Bun10q, Bun12p, Bun91j, Sle12a, Har75, Qui74, Tuo74a, Tuo74b, Vol73a]. Maturation [Bun68f]. mauvais [Bun89b, Bun90c]. Max [Bun58a, McF07]. Maxwell [Bun43a, Bun43c]. May [Aga91, AL99, AF01, BD78]. Mayer [Bun90e]. McGill [BAB^{+}80, Van04]. McNally [Bun04g]. Meaning [Bun75f, Qui58, Bun83l, Sle96]. Measure [Bun73j, Bun73k]. Measurement [Bun73j, Bun73k, Bun77h]. measurements [BK83a, BK84]. Measuring [Bun10r]. mecánica [Bun56f, Vuc14]. mecanísmicos [GML14]. Mechanical [Bun57a, Bun13f]. Mechanics [Bun56e, Bun66b, Bun92d, Bun56f, Bun67e, Bun77h, Bun79e, Bun03i, Vuc14]. Mechanism [Bun97d, Bun13f, Joh19]. Mechanismic [dSMdC19]. Mechanisms [Ahm19, Bun04c, Bun04d, Iva19, Pic04, SW09, WS06]. Mechanistic [CD19, GML14]. med [HD88]. médecine [Bun19e]. Medical [Ahm19, Bun13g, Bun78b, Bun78c, Bun78h, Bun78i, Sei06]. Medicine [Bun13g, Bun78a, Bun78b, Bun78c, Bun78h, Bun78i, Bun78g, Bun19e]. Médicos [Lab14, AC13, Bun12l, Sil13f]. Meeting [AL99, AF01, Bun18c, SA76]. Meets [Bun70f]. Melbourne [Bun79b]. mémoires [Bun16b]. Memoirs [Bun18a, Mat16, Bun16a]. Memorias [Bun14e, Kan15]. memories [Bun16b, Kan15, Bun14e]. Memory [Bun80i, Bun01-38]. ménage [Bun91e]. Mentalism [Mil93]. mentation [Bun80e]. Mente [Bun89j, Mah14]. mesto [Bun62g]. meta [RG02]. Metafísica [Bun73d]. Metamodel [KJ10]. Metaphors [Bun01-34, LB78b]. Metaphysical [Bun12p, Bun74d, Mah13]. Metaphysics [Bun71d, Bun00c, Bun01d, For89a, For89b, JL95, Bun68l, Bun69e, Esp94, Bun71b]. métaphysique [Bun68l, Esp94, Mah13]. métaphysiques [Bun74d, Bun75b]. Metascience [Mau22a]. Metascientific [Bun59b, Mau22b, Mau22c]. Method [Aga12, Bun71f, Bun73g, Bun01r, Bun60i, Bun68q, Bun76j, Bun80t, Bun81m, Bun85p, Bun86j, Bun87j, Bun94c, Bun19e, Dah91, Har75, Qui74, Tuo74a, Tuo74b, Vol73a]. méthode [Bun94c, Bun19e]. Methodological [Bun73h, Alc91, Cha06]. methodológicas [Cha06]. Methodology [Bun73j, Bun85a, Bun83f, Bun83g, Bun85e, Bun13k, Den03, Esp86, Slo99, Sne68b, Bun69e, Bun73k, Bun74i, LK74, Mat82, CCB^{+}90, Rus86]. Methods [Bun86a, Bun86c, Bun72f, Bun73f, Bun13c]. método [Bun60i, Bun68q, Bun73f, Bun76j, Bun80t, Bun81m, Bun85p, Bun86j, Bun87j, Bun13c]. Metodología [CCB^{+}90]. MeV [Bun45a]. micro [Déc10, Moe08]. microphysics [Bun71g]. Migration [Bun69d]. milenio [Bun07f]. Miller [Bun68b]. Mind [Ber82, BAB^{+}80, Bun77a, Bun80b, Bun80j, Bun80k, Bun91f, Bun96g, Bun01c, Bun10l, Bun10k, Bun10m, Cau84, Gam14, KM70, Lev83, OW79, Puc81, Sle96, Sle12a, Sle12b, Bun79g, Bun79h, Bun87d, Bun87g, Bun89d, Bun07a, LB78a, Mah14, Pan90, Par81, Bun89j]. Mind-Body [Ber82, Bun80j, Bun80k, Cau84, Lev83, Puc81]. Minding [Bun10n]. Mindless [Bun01g, Bun81g]. Miro [SB92, SB92]. Misunderstandings [Bun04c]. mito [Góm14]. Mitos [Bun04f, Bun01-35]. mobilidade [QdAY13]. mobility [QdAY13]. Model [Bun73g, Har75, OHS02, Qui74, Tuo74a, Tuo74b, Vol73a, Bun73c, Bun76f, Bun78e, Bun79i, RG02, SBC^{+}08, Bun68e]. modèle [Bun68e]. Modelisation [Pie99]. Modelling [MB16, Bun79b]. Models [Bun68g, Bun69d, GG86, Joh19, Pie99, Bun84a]. Moderate [Bun97e, Bun01m]. Modern [Bun55b, Bun59a, Bun63b, Bun70b, Bun79c, BR97, Bun09b, Din60, Duc60a, Duc60b, Hes60, Hin61a, Hin61b, Mor61, Bun61i, Bun62g, Bun65b, Bun70c, Bun70i, Bun72h, Bun81c, Bun84b, Bun60b]. Moderna [BR97, Bun61i, Bun65b, Bun70c, Bun70i, Bun72h]. modernen [HD88]. Modernity [Obi22]. modernization [Obi11]. Modes [Bun00g]. Modules [GG86]. Moessinger [Déc10]. Molina [G.73]. mondes [Bun16b]. mondial [Cau88, Die78]. monism [Bun89g]. Monismus [Bun89g]. Montreal [BAB^{+}80, Cau88]. Mora [Coh81]. Moral [Bun97f, Bun01q, GKvWZ97, Obi19, BdCFG^{+}96, BSO^{+}96, KK07, OSV96, BdCFG^{+}96, BSO^{+}96, Bun07c, OSV96]. Morality [Bun01n, Bun89h]. morally [Bun76e]. morphogenesis [Bun84a]. Morris [Bun56d]. Mosterín [Bun85h, Bun85g]. Motivation [Bun80a]. movement [Mor16]. Much [Bun03g]. Mueller [Bun68i]. Multi [Bun58c]. Multi-Dimensional [Bun58c]. multiagent [Mac11]. multiagentes [Mac11]. Multiple [Bun12m]. München [Bun71a]. mundo [Bar14, Bun45b]. Mundos [Bun14e]. mundus [DB14]. my [Bun77j, Bun91j, Bun03g, Mat82]. Myth [BM63, Bun63d, Haj65, Har66, Mul63, dF19, Góm14]. myths [Bun01-35, Bun04f].

Nacional [Fonxx, Bun14d]. Natur [BM04]. natural [Cha06, Lau91, SA93, SW89]. Naturales [Fonxx, Cha06]. Naturaleza [CFF^{+}14, Moy15, Bun44c, Eng47]. Naturalism [Bun10o, SK09, Bun09a, Mah13]. naturalisme [Mah13]. Naturalization [Mau22a]. Nature [Bun01o, BM04, Dah91, Eng47, Mar81, Bun44c, Bun68h, Bun78c, Bun78i, Bun83j, Bun90e, Cha06, CFF^{+}14, Moy15]. naturleza [Cha06]. Naturphilosophie [SA93]. naturwissenschaften [SW89]. nauke [Bun62g]. nec [DB14]. Necessary [Joh19]. Need [Bun01s]. negative [GBS^{+}05]. negras [Bun85b]. Neil [BS96]. nella [Bun70c, Bun70i]. Neo [Gam14, Bun03b, Bun03b]. néo- [Bun03b]. neo-conservatism [Bun03b]. neo-liberalism [Bun03b]. néo-libéralisme [Bun03b]. Neo-Sadrian [Gam14]. Nervous [PW80]. neural [GBS^{+}05, VOHPMJ12]. Neurath [SPT11]. neuro [Bun78b, Bun78h]. neuro-medical [Bun78b, Bun78h]. Neurological [Gil19]. neurologie [Bun83c]. neurology [Bun83c]. neuron [Bun80e, Bun89d]. neuropsychologie [Bun83c]. Neuropsychology [BKR96, Bun01g, Bun81g, Bun83c]. Neuroscience [BAB^{+}80, Bun01g, Bun81g]. neurosciences [LB78b]. Neuroscientists [Bun91f]. neutral [Bun76e]. Neutron [Bun45a]. Nevertheless [Aga13]. Newtonian [Bun66b, Bun92d]. Niels [Bun85j, Bun88e, Bun92c]. Niklas [Kla03a, Kla03b, Kla04, Kla03a, Kla03b, Kla04]. Ninety [BWM^{+}04]. Niveaus [Bun89g]. NJ [Fea64]. No [Bun76d, Mil77, Bos14, Mor16]. no-hegemónica [Bos14]. noción [Bun75i, Tor14]. non [Bos14, WL96]. non-hegemonic [Bos14]. non-ideological [WL96]. Nonconcept [Bun68i]. None [Bun00g]. nonlinear [SBC^{+}08]. nonscience [Bli91, Bou91].

Nonsense [Bun68a]. nonvanishing [BK83a, BK84]. Normative [GKvWZ97, Bun73i]. Normen [Bun73i]. norms [Bun73i]. North [Bun89f]. Nota [Tor14]. Notable [Jun19]. Notas [Bun76d]. Note [Ano59, Ano04, Mar96, Tor14]. Notes [Bun76d, BS96, Bun79b]. Notion [GPA91, Bun75i, Tor14]. nouveau [Bun04i]. nouvelle [Bun98g]. nouvelles [Flø86]. Novel [Pot92]. Novelty [Bun03c, Kil04, Pic07]. Noventa [BWM^{+}04]. novísima [Bun91l]. Nuclear [Bun44a, Bun44e]. nucleares [Bun44e]. nuestra [Qui58]. Nuestro [Bun97h, Bun01-35]. nueva [Ano16, Bun44e, dP16]. nuevo [Bun95g, Bun07f]. nuevos [Bar14]. Null [Bun66c]. nuovi [SA93].

o [Bun85f, Den14b, Mac11, Bun17d]. Object [KJ10, Cha06]. Object-Oriented [KJ10]. Objective [Bun68k]. Objectivity [Bun07d, McF07]. Objects [Bun10a, Ore22a]. objeto [Cha06]. obra [Roj13, SB92]. Observations [Aga12]. obstacle [Bun71g]. Occasion [BAG17, SD91, Wie99]. Occur [Bun15a]. ohne [Bun73i]. Olgierd [Bun62b]. once [Bun09d]. ondas [Vuc14]. ondulatoria [Bun56f]. One [Bun12m, Bun06h]. Ontología [Bun95m, Mah14, Mos14]. Ontological [OHS02, RG02, Vuc14]. ontológicos [Vuc14]. Ontology [Bar81, Bow80, Bun74e, CD19, KJ10, Mau22b, Mau22c, RW05, Rus80, WW90, YCVP^{+}19, Bun77i, Lau91, Mah14, Mos14, Bun77g, Bun79j, Bun95m, Qui80]. Open [Bun99a]. openmindedness [Set96]. Openness [BAG17]. operationism [Bun88a]. Operator [BK69, Hin77]. operators [Bun03i]. Opinionated [Bun05b]. Opiniones [Bun87i]. Opinions [Bun87i]. orden [Bun95o]. order [Bun95o]. Organ [Bun80l]. orientación [Par83]. Orientation [Par83]. Oriented [KJ10]. Österreich [SW89]. Otak [Gam14]. other [Bun12z]. otros [Bun12z]. Otto [SPT11]. Our [Bun53, Bun97h, Bun01-35, MB96a, Qui58]. Ours [Bun00e]. Outlined [Zie22, Zie16]. outputs [Bun78f]. Overcome [Bun04e]. own [Mat82]. Oxford [BS96, Bun84b, Din60, Par81].

P [Bun58b, Bun73b, Bun87a, Bun76a]. Pabellón [Fonxx]. pacientes [Lab14]. pages [Ano59, Bun99a, Lin69]. Panksepp [Bun90e]. Paolo [Bun66a]. Paper [Bun85c, Sle96]. papers [Van04, Die78, SW89]. Para [Bun76d, AC13, Bun06h, Bun06a, Bun07f, Bun12l, Ech85, Mac11, Par83, Sca14, Sil13f]. Paradigm [Zie22, Obr05, Zie16]. Paradigma [Obr05]. paradigms [AK98, Bun93d]. paradoja [Bun75d, Mar72]. Paradox [Bun69f, Bun75d, Mar72]. paradoxes [BK83b]. Parallel [Bun12q]. parapsychology [Bli91, Bun88j]. Parasitism [YCVP^{+}19, Yan14]. parasitismo [Yan14]. Paris [Bun76b, Bun76c]. Parra [Ech85]. Parsimony [dF19]. Part [Bun91b, Bun92a, Esp86, MC95]. Partial [Bun68p, Mil77]. participation [BGS76]. Particles [Bun70h, Bun44b, Bun58g, Vuc14]. partículas [Bun58g, Vuc14]. Passion [BP94, BP94]. patients [Lab14]. Paul [Bun76a, Mor81, SD91, Sle96]. Peace [LP88]. peculiarities [BK75]. peligros [Bun19f]. Penguin [Sle96]. pensadores [Bun39]. pensamiento [SA01]. pensar [Bun06h, Bun06a]. pensée [Bun04e]. Penser [Sin91]. Peoples [dCPM94]. Perception [BL10, Bun80o, Bun89f]. perdido [Mor16]. pereat [DB14]. Pergamon [Bun84b, Par81]. Personal [Bun03f]. Personality [Bun80c]. persons [Mar81]. perspectiva [Bun98i]. Perspective [BAB^{+}80, Bun91f, Bun98f, Gam14, Hal19, MB01, Pic19, Ray15, Res87, Bun79g, Bun98i, Bun06g, Bun09g, Col99, Lun00]. Perspectives [Bla92]. Perspektif [Gam14]. Pervasive [Kra80]. Peter [Bun07b]. Phenomenological [Bun64b, Bun99i]. Phenomenology [Bun51c]. phil [HD88]. Philip [Bun05a]. Philipp [Bun51a, Bun53]. Philonous [Bun54, Bun01p]. Philosophandi [Bun73a]. philosophe [Bun03b, Bun16b]. Philosopher [Bun86g, Bun18a, Mat03, Mat16, Bun03b, Bun16a, Bun16b]. Philosophers [Bun17f]. Philosophical [Aga88a, Bun76g, Bun78b, Bun78c, Bun78f, Bun78h, Bun78i, Bun84d, Bun86h, Bun91f, Bun98f, Bun03e, Bun10k, Bun12r, Bun13a, Col99, Lun00, Res87, Sle12a, Bun71g, Bun95g, Bun98i, Bun98b, Coh81, Moy15, Qui14, SB92, Bun20b, MB00]. philosophically [Bun57b]. Philosophie [BB75, Bun87e, Bun92b, Bun19e, Die78, HD88, SW89, Bun78d, Bun89b, Bun90c, Bun91e, Bun92c, Bun94c, Bun06d, Cau88, Flø86, Jod10, Kli68, Vac93, Bun69a, BA90, Cau88, Dro15, HG75, Mor81]. Philosophies [Bun77d, Bun99b, Bun01i, Bun12j, Bun12s, Bun19g, Bun87g]. philosophique [Bun20b]. philosophiques [Sil13d]. Philosophische [MB00]. Philosophize [Bun57b]. Philosophy [AC82, Aga91, AD97, AL99, AF01, Ano04, Asp98a, Asp98b, Bah81, Bar81, Bow80, Bra97, BD78, Bun55d, Bun60b, Bun63d, Bun64a, Bun67h, Bun68m, Bun68j, Bun69a, Bun70d, Bun71f, Bun72a, Bun72b, Bun73e, Bun73l, Bun73m, Bun78h, Bun78i, Bun78g, Bun78d, Bun82a, Bun83i, Bun85e, Bun85j, BA87, Bun91f, Bun92b, Bun95e, Bun96b, Bun98c, Bun98d, Bun99c, Bun99d, Bun99h, Bun99k, Bun00e, Bun00d, Bun01a, Bun01r, Bun01s, Bun01t, Bun01-30, Bun01-32, Bun01-33, Bun03f, Bun06f, Bun09f, Bun10p, Bun12o, Bun13g, Bun13i, Bun13k, BA14, Bun17b, Cau88, Dem97, Esp86, Flø86, FM66, HH76, Kal13, KPSP74, LM68, LSP22, Mah21, MC95, Mar12, Mat12, McF78, NSB10, Obi22, OSV96, PS94, Pic04, Qui80, Rom19, Rus86, Sad04, SD91, She76, Sne68b, SA76, Tre01, Vac93, Vol93, Aga14, Agu06, BdCFG^{+}96, BC14]. philosophy [Bun44c, Bun60i, Bun63f, Bun68q, Bun69h, Bun73f, Bun75j, Bun75g, Bun80t, Bun81m, Bun85p, Bun85l, Bun86j, Bun87j, Bun88b, Bun88e, Bun89b, Bun90c, Bun91e, Bun94c, Bun95g, BSO^{+}96, Bun96h, Bun99j, Bun06d, Bun07a, Bun10g, Bun13c, Die78, Dro15, FvWK03, Jod10, Kan15, Kli68, Mah14, Mor81, Moy15, Obi11, Pic01, Rus80, SA93, WZ70, Bun76k, Bun87e, BA90, Bun95h, Bun95k, Bun95i, Bun95j, Bun08b, Bun12z, Bun19e, HG75, SW89, Bun76a, AC13, Bru73, Bun74a, Bun78a, Bun78b, Bun78c, Kir75a, Kir75b, Rob74, Sil13f, Vol73b]. Phobosophies [Bun12s]. photon [BG16]. Physica [SA93]. Physical [Bun61g, Bun63a, Bun67f, Bun67j, Bun68k, BG76, Bun01t, Bun12p, Bun13j, Esp86, HH76, Sal71, Bun58g, Bun69g, Bun43a, Bun43c]. Physicians [Bun00e, Sil13f]. Physicist [Bun70d, FJMT88, Mat03, Bun58b]. Physicists [Bun84e]. Physics [Aga72, Bun53, Bun65a, Bun66d, Bun67a, Bun67d, Bun71h, Bun73l, Bun73m, BS79, Ano81, Bun82b, Bun82c, Bun83b, Bun00c, Bun01d, Bun03h, Bun12h, Bun12i, Bun12q, Bun13b, MB16, MC95, Riv19, Rom19, Sin81, Whe80, Bun68j, BK75, Bun95i, Bun10g, Bun13d, HG75, WZ70, Bun68l, Bun71a, SA93, Bru73, Bun58a, Haj76, Kir75a, Kir75b, Met73a, Met73b, Rob74, Sne68a, Sne68b, Vol73b, Bun69c, Str69]. Physik [BK75, HG75]. Physique [Bun68l, BB75, Bun13d]. piano [BL10]. picture [Bun55a]. Pierre [Déc10]. Pikiran [Gam14]. pillars [Bun96h]. Pioneer [FJMT88]. Piper [Bun71a]. Pitman [Bun79b]. Place [Bun55b, Bun59a, Bun60g, Bun63b, Bun70b, Din60, Duc60a, Duc60b, Hes60, Hin61a, Hin61b, Mor61, Wet82, Bun62g, Bun70c, Bun70i, Bun81c, Bun90e, Bun91d]. planning [Bun73p, Bun90f]. plans [Sen93]. planteamiento [Bun17g]. Plastic [Bun10n]. Plata [Bun14d]. Plato [Bun74a]. Plattsburgh [Har77]. Plausibility [Gil19]. Polémica [Cha06]. Polemics [Aga88a]. policy [Den03, Qui14]. política [Qui14]. Political [Bun01n, Bun09f, Bun19g, Kar19a, Kar19b, Bun89e]. Politics [Kar19a, Kar19b, Bun87e]. politique [Bun87e]. Pool [Ano75]. Pope [Bun97g]. Popper [ABA19, Bun59c, Bun64a, Bun96h, Bun99c, Bun01u, Bun84b]. Popperian [SL19]. popular [Bun43b, Bun89f, Bun43b]. por [Bar14]. Portuguese [Mac11, QdAY13]. porvenir [Bun09d, BGD13]. Posible [Bun73d, Ama14]. Position [BK69, Hin77]. Positivism [Cup19]. Possibility [Bun76h, Iva19]. Possible [Bun71d, Kar19c, Riv19, Ama14, Moe87]. Post [Ray19]. Post-Truth [Ray19]. posto [Bun70c, Bun70i]. pour [Bun04e, Sil13d, Vac93]. Poverty [Bun95b]. Power [Bun01v, Bun91g]. Pp [Bun53, Bun07c, Din60, Mul63, Rob74, Sin81, BS96, Bun71a, Bun79b, Bun84a, Bun86c, Bun87a, Bun11a, Fea64, KPSP74, Mat16, Par81, Sle96, Sne68b, Bun76b]. pp. [Bow80, Bun74a, Bun76c]. Practice [Ahm19, Bun04g, Bun74b]. practices [BG16]. Practise [Bun07d]. Praise [Bun96c, Bun96d, Den14a, Den14b, Bun98a, BA17]. praxiology [Bun99f]. Preached [Bun07d]. Prebisch [Cor86]. Precedents [Art19]. precise [Moe87]. predicado [Bun75i]. predicate [Bun75i, Man85]. Preface [Ano79b, Bun80m, Bun15c, Mat19d, Bun08d]. Preliminary [Bun60e]. Prentice [Fea64, Mul63]. Prentice-Hall [Fea64, Mul63]. prescription [Den03]. Present [Wie99, Bun74a]. Presentation [Mau22b]. presented [Van04]. Press [Ano59, Asp98b, BS96, Bun53, Bun76a, Bun84b, Bun07c, Din60, Kil04, Par81, Pic07, Ric08]. présupposés [Bun74d, Bun75b]. Presuppositions [Bun74d]. prevent [BK83a, BK84]. Price [Ano59, Din60, Lin69]. Prichinnost [Bun62g]. prichinnosti [Bun62g]. Principio [BR97, Bun55g, Bun61i, Bun65b, Bun70c, Bun70i, Bun72h]. Principle [Bun55b, Bun59a, Bun63b, Bun70b, BR97, Din60, Duc60a, Duc60b, Hes60, Hin61a, Hin61b, Mor61, dF19, Bun55g, Bun61i, Bun62g, Bun65b, Bun70c, Bun70i, Bun72h, Bun81c]. Principles [Bun07b]. printsipa [Bun62g]. Probabilidad [Bun60j]. Probabilistic [Bun73b]. Probabilities [Bun12u]. Probability [Aga88b, Bun01-36, FM66, HH76, Pat90, Bun56a, Bun76h, Bun81f, Bun88i, Tul95, Bun60j]. Problem [Aga91, AF01, Ber82, BAB^{+}80, Bun64d, Bun80j, Bun80k, Bun91f, Bun96g, Bun10l, Cau84, Puc81, Ama14, Bun79g, Bun79h, Bun87d, Bun98c, LB78a, Par81]. Problem. [Lev83]. Problemas [Bun87i, Bun97h, Vuc14]. problemática [Ama14]. Problematizing [zW11a]. problème [Bun87d, Bun87c]. Problems [Bun63d, Bun68m, Bun70j, Bun71h, Bun73e, Bun84d, Bun86h, Bun87c, Bun87i, Bun97h, Bun19d, KPSP74, LM68, Van04, Bun79k, Vuc14, Met73a, Met73b]. procedures [Bun72f]. Proceedings [AL99, AF01, SA76, Har77, SW89, WZ70, BAB^{+}80, LM68]. Processes [Bun70h, Mac85, Bun76f]. Processing [PW80]. Productive [Mat16]. products [Bun74d]. produits [Bun74d, Bun75b]. Prof [HD88, Nar75]. Prof. [Nar75]. Profesor [Ano14a]. professional [Bun04g]. Professor [BvL91]. Program [Bun72e, Bun73n]. Progress [Bun12r, Bun19a]. Project [Mat19c]. Projects [Bun17c, Bun04g]. Prólogo [Par14]. Promesas [Bun19f]. Promise [Bun12o, Van04]. Promises [Bun19f]. properties [BS77]. proposal [Mar96]. Proposals [Bun97h, Bun75j, Sca14, WL96]. proposiciones [Bun75j]. propósito [Sil13f]. Propuestas [Bun97h, Sca14]. Prospectiva [Bos14]. Proton [Bun45a]. Proved [Bun68p]. Provocaciones [Bun11c]. Provocations [Bun11c]. PSA [SA76]. Pseudociencia [Bun13h]. Pseudociencias [Bun14g, Bar11]. Pseudoscience [Bun06f, Bun08a, Bun83d, Bun84f, Bun91j, Bun04g, Bun06d, Bun85m, Bun86k, Bun13h, Bun14h]. pseudosciences [Van91, Bun14g]. psi [BvL91]. psicología [Bun95j, BA88]. psychiatry [Var17]. Psycho [VOHPMJ12]. Psycho-neural [VOHPMJ12]. psychoanalysis [Bli91, BB10]. Psychobiological [Ber82, Bun80k, Cau84, Puc81, Par81]. Psychological [Mon92, Bun85n]. Psychologie [Mon92, Bun83c, BA90, Bun92b]. Psychologists [Bun91f]. Psychology [Bun68b, BA87, BA90, Bun92b, Bun01g, Bun12b, BA14, CD90, Mac85, VOHPMJ12, Bun81g, Bun83c, Bun85i, Bun90e, Bun90g, Bun95j, May90, Roy90]. Psychoneural [Bun80n, Bun79a]. Public [Bun99a, Dro19]. Publications [Scr63, Sil19, Sin81]. Publishing [Bow80, Bru73, KPSP74, Wie99]. puente [Bun95o]. Pyramids [Bun12n]. Pythagoras [Bun03h].

quadratic [SBC^{+}08]. Quaint [Bun03g]. Qualitative [BCD03, Bun03c, Kil04, Pic07, BP94]. Quality [Bun75h]. Quanta [Bun67g, Bun67h, Bun67h]. quantal [GSB81]. Quantenphysik [BK75]. Quantifiers [MM19]. quantique [Bun13d]. Quantitative [Res19]. Quantons [Bun03g]. Quantum [Aga72, Bas71, Bun55d, Bun56e, Bun67i, Bun68a, Bun77h, Bun03g, Bun03h, Bun10q, Bun12h, Bun12i, Hob19, Jun19, Bun67e, BK75, Bun79e, BK83a, BK83b, BK84, Bun88g, Bun03i, Bun13d, Vuc14, Bub68a, Bub68b, Fin70, Sne68c]. quatre [Esp92]. que [Bar14, Bun85k, Bun44d, Bun58f]. Québec [MC95]. Queries [Bun59b].

Quesada [SB92]. question [Cha06]. Questions [Roy90, Bou91, Bun14a, Moy15]. Quine [Bun75j].

R [Bun71a, Bun79b, Ano81, Bun07c, Sin81, Whe80]. rôle [Mah13]. racionales [Sca14]. Racionalidad [Bun85q, CCB^{+}90, Roj13, Sil13f]. Rafael [Bar11]. Rahman [Obi22]. raison [Bun96f]. raíz [AC13]. Rand [Bun84a]. Randall [Bun99b]. Rational [Bun95b, Bun95c, Bun01w, Bou91, Sca14, Bun86b]. Rationalism [JL95, SL19]. Rationality [AJ87, Bun87f, Bun01-30, Van04, CCB^{+}90, Roj13, Sil13f, Bun85q]. Raúl [Cor86]. Raymond [CH09, GPA91]. Razón [Bun86f, Bun96e, Mor16, SB92]. razones [Bun04f]. Re [Bla19, zW11a]. Re- [zW11a]. Re-analysis [Bla19]. reactions [Mos91]. Reading [Bun10r]. Readings [FM66]. Real [Bun70h, BK83a, Bun03g, Bun10q, BK84]. realidad [Bun72g, G.73]. Realism [Bun90a, Bun93c, Bun95a, Bun97f, Bun01k, Bun01j, Bun01q, Bun01x, Bun03h, Bun06b, Bun12h, Bun12i, Bun14b, Cor12, Cor19, Del19, Ing19, Mah01, Bun84c, Bun85q, Bun86e, Bun95d, Bun13d, Bun17d, Esp94, Bun85l, Bun95d, Ric08, For07, Tri06]. réalisme [Bun13d, Esp94]. Realismo [Bun85l, Bun95d, Bun17d, Bun85q]. Realist [Hob19]. realista [Bun07f]. réalité [Jod13]. Reality [Bli90b, Bun65a, Bun66d, Bun67i, Bun01h, Bun06b, Bun12o, Bun14b, For07, LSP22, Pat90, Pie99, Ric08, Rom18, Tri06, Bun72g, Bun94d, Jod13, Aga72, Bub68a, Bub68b, Fin70, Sne68c]. Reason [Bun86f, Bun96e, GLL96, Mad94, Bun96f, Mor16, SB92]. reasoning [Sei06]. Reasons [Bun04f]. Reception [Kla03a]. Reconceptualizing [Pic19]. Reconsidered [McF07, Har91]. Reconstruction [Bun01s, CD19]. Recuerdo [Bun01-38]. Redefined [Bun61a]. Reduction [Bun77f, Bun01v, Bun01-37, Bun11d, Bun91g, Bun89g]. Reductionism [Aga91]. Reduktion [Bun89g]. réel [Bun94d]. Reenchanted [Bun00b]. Reference [Bun74f]. reflection [BRG04]. Reflections [Mos91, AC13, Bun58b]. Reflexiones [AC13]. reform [Bun96i]. reforma [Bun96i]. Reframing [zW11b]. Refute [Bun12h, Bun12i, Bun13d, Bun13d]. réfute-t-elle [Bun13d]. reglas [Sca14]. Regulae [Bun73a]. Reichenbach [Bun60b, Bun60b]. Reidel [Bow80, Bru73, KPSP74, Rob74]. Rejoinder [Bun62c]. relación [BG01, Lab14]. relating [WL96]. relation [Bun70e]. Relational [Bun68k, BG76]. Relations [Bun74e, Bun70j, Bun77i, Bun03i, Cha69, Lac96]. relationship [Lab14]. relationships [Cha69]. Relativism [Bun01y]. relativista [Bun60d]. Relativistic [BK69, Bun03i, Bun60d]. Relativity [Bun01t, Bun53]. relevance [Bun75g]. Religion [Tur96, Lac96, MB96a, Set96, WL96]. religious [MB96b, Poo96, Woo96]. remote [Bun85b]. René [BP94, Bun84a, BP94]. Réplica [Bun85h]. Reply [Bun68i, Bun77j, Bun85h, Bun85i, Bun03g, Bun04g, TBG^{+}85, Bun85g, Bun90b, Bun90e, Bun91j, MB96a]. Report [Bun03f, Mon92]. representación [Bun44e, Bun75e]. Representation [Bun44a, Bun69b, Bun44e, Bun75e, Mat82]. required [Bun04g]. Requirement [Gil19]. requirements [Mac11]. requisitos [Mac11]. Research [Ahm19, BCD03, Bun67c, Bun13a, Bun17c, Edg69, VOHPMJ12, YCVP^{+}19, Zie22, Ama14, BRG04, Bun69h, Bun79b, Bun85i, Bun00h, Bun04g, Lin69, Zie16]. resisten [Bar14]. Respect [Bun17f]. response [Bun75d, CD90, Den03, May90, Pan90, Roy90]. respuesta [Bun75d]. Restricted [LB78b]. résumés [Die78]. rethinking [Pfe77]. retos [Qui14]. retrato [SA01]. reversal [Bun70g]. Review [Aga72, Asp98a, Bar11, Bar14, Bar81, Ber82, Bow80, Bra97, Bru91, Bru73, Bub68a, Bub68b, Bun51a, Bun53, Bun58a, Bun58b, Bun59c, Bun60b, Bun62b, Bun63a, Bun66a, Bun68b, Bun71a, Bun72a, Bun72b, Bun73b, Bun74a, Bun76a, Bun76b, Bun76c, Bun76i, Bun78a, Bun78b, Bun78c, Bun79b, Ano81, Bun84a, Bun84b, Bun86b, Bun86a, Bun86c, Bun87a, Bun99a, Bun99b, Bun00b, Bun01b, Bun05a, Bun05b, Bun07b, Bun07c, Bun11a, Cau84, Col99, Cor86, Déc10, Dem97, Din60, Dri87, Duc60a, Duc60b, Ech85, Edg69, Esp85, Esp86, Fea64, Fin70, For07, G.73, Haj65, Haj76, Har66, Har75, Hes60, Hin61a, Hin61b, Hin64, Hof84, Kil04, Kir75a, Kir75b, Lev83, Lin69, Lon62, Lun00, Mat16, McF78, Met73a, Met73b, Mor61, Mul63, Par81, Pic07, Puc81, Qui74, Qui80, Ric08, Rob74, Rus80, Rus86, Sin81]. Review [Sle96, Sle12a, Sne68a, Sne68c, Sne68b, Tre01, Tri06, Tuo74a, Tuo74b, Vol73a, Vol73b, Vol79, Vol81, Whe80, dP16, Bun74b, Bun93a]. Reviews [Aga88a, Asp98b, BKB75, Bun58d, Bun58e, Bun60h, Bun63e, Bun64c, Bun68n, Bun73o, BB74, Bun84e, Bun04a, KPSP74, MVR^{+}73, VBH^{+}75, WPG^{+}74]. revisited [Mil93]. Revival [Bun01z, Ing19, Bun81j]. Revolution [Bun97c]. Revolutionary [Bun11a, Göt95]. Rezeption [Kla03a]. Richard [Bun63a, Bun04g]. Richer [Bun53]. Richness [Bun76g]. riders [Bun88h]. Right [AD04, Bun89c, Bun79e, Bru91]. Rights [Bun91h, Bun01-27, GPA91, dCPM94, Bun94e]. rise [Bun71g, Bun44c]. Risk [Bun00e]. river [Bun93d]. Rixensart [Bun78d]. Robert [Bun74a]. role [Bun73p, Mah13]. Romero [Bun01-38, SA01]. Ron [Bun85b]. Rootselaar [Bun77j]. Rosettes [Bun12n]. Rossi [Bun66a]. Routledge [Bun76a]. Row [Bun71a]. Royce [Bun90e]. Rules [Bun05b, Bun12t, Bun68o, Sca14]. Russell [BS96, Bun73a]. Russian [Bun62g]. Rutherford [BS79, Ano81, Sin81, Whe80].

S [Bun84b, Bun87a, Bun07c, Sil13e]. sa [Bun94c]. saber [Bun02b]. sabiduría [Aga14, Den14a, Den14b]. Sadosky [BWM^{+}04]. Sadrian [Gam14]. Salmerón [OSV96]. Saltationism [BB89, Bun87h]. Salzburg [WZ70]. San [Bun74a, Bun79b]. Sánchez [Bun85b]. sanity [Mor16]. sans [Bun83c]. saviour [Tur96]. saw [Bun45b]. scam [Bun14g]. Scattering [Bun45a]. Schaff [Bun62b]. Schlegel [Bun63a]. scholastic [OB84]. Schrödinger [Bun53, Bun01-28]. Schuster [Sle96]. Science [Aga91, AL99, AF01, AD04, Ano04, Bla19, Bun55b, Bun59a, Bun60b, Bun60g, Bun61e, Bun62f, Bun63b, Bun64a, Bun66e, Bun68f, Bun68m, Bun70b, Bun71f, Bun72e, Bun73j, Bun73h, Bun73n, Bun75c, Bun75f, Bun77d, Bun78d, Bun79c, Bun81i, Bun85e, Bun86a, Bun86c, Bun86i, Bun91b, Bun92a, Bun93e, Bun96a, Bun96b, BR97, Bun98c, Bun98d, Bun98f, Bun99c, Bun99e, Bun00b, Bun01a, Bun01i, Bun01k, Bun01l, Bun01r, Bun01o, Bun01x, Bun01-30, Bun01-33, Bun01-37, Bun03f, BM04, Bun05b, Bun08a, Bun09b, Bun12u, Bun13e, Bun17b, Bun17c, Col99, Del12, DB93, Esp86, GG86, GLL96, HH76, JL95, KM70, Mad94, KK07, LM68, Lun00, MB16, Mah21, MC95, Mat12, Mat19c, PS94, Pic04, Pie99, Ray15, She76, Sil13e, Sin81, Sle96, Sle12b, Sle19, Sne68b, SA76, SPT11, Vol93, Aga14, Agu06, Ano16, Bli91, Bou91, Bra82]. science [Bun51c, Bun57b, Bun58f, Bun60c, Bun61i, Bun62g, Bun63f, Bun65b, Bun65c, Bun67b, Bun68g, Bun68h, Bun70c, Bun70i, Bun72h, Bun73i, Bun74d, Bun75b, Bun75g, Bun76e, Bun81c, Bun83d, Bun83j, Bun87e, Bun88b, Bun88c, Bun88d, Bun88j, Bun89f, Bun91a, Bun91l, Bun91e, Bun91j, Bun93c, Bun94c, Bun98g, Bun98j, Bun99g, Bun06e, Bun07e, Har91, Jod10, Kan15, Lac96, LK74, MB96a, MB96b, Mah13, Obr05, Pic01, Set96, Tur96, Vac93, Woo96, WL96, Asp98b, Bun53, Bun60i, Bun68q, Bun73f, Bun80s, Bun80t, Bun81m, Bun85p, Bun86j, Bun87j, Bun94c, Bun97a, Bun98h, Bun05a, Bun13c, Bun14c, LM68, Vol93, WZ70, Asp98a, Bra97, Bun72a, Bun72b, Din60, Duc60a, Duc60b, Fea64, Hes60, Hin61a, Hin61b, Hin64, Lon62, Mor61, Scr63]. Sciences [Aga88b, Ano04, Bun95d, Esp86, HH76, Kal13, SB01, Van01, Van04, Wie99, AK98, Bun73k, Bun76b, Bun76k, Bun76c, Bun78h, Bun78g, Bun78d, BO79, Bun95k, Bun95o, Bun98i, Bun06e, Cha06, CCB^{+}90, Lau91, SW89, Dem97, Bun78a, Bun78b]. scientia [DB14]. Scientific [AC82, Bun61f, Bun61h, Bun63d, Bun66f, Bun67c, Bun68o, Bun71d, Bun72f, Bun81k, Bun12r, Bun13a, Bun17c, Bun18d, Cor12, Cor19, Cri19, For89a, Mah01, Mah21, Obi19, Obi22, Ray15, Res87, SD91, Slo99, Vel22, AD04, Bun59c, Bun68d, Bun87g, Bun88f, Bun90g, Bun17g, Bun17d, Cha06, Dah91, Del13, Mor81, Per91, Qui14, Set96, Bos14, Bun69h, Bun76l, Bun83h, Bun89i, Bun00h, Bun08c, Esp94, Bun71b, Edg69, Esp85, Hof84, Lin69]. scientifically [Bun57b]. scientifique [Bun08c, Bun16b, Del13, Esp94, Mau17]. scientifiques [Bun68d, Sil13d]. Scientism [BA17, Bun86e, Bun15b]. Scientist [Bun18a, Bun89e, B^{+}99, Bun16a, Bun16b, Mat16]. Scientists [Bun17f, LP88]. scienza [Bun70c, Bun70i]. Scopes [Göt95]. se [Bar14]. Search [Bun67c, Bun04d, Edg69, Pie99, Sle96, Bun87g, Lin69, May90]. Second [Ano04]. secrecy [Bun79i]. section [Die78]. sections [Die78]. See [Moe08]. Sektions [Die78]. Sektions-Vorträge [Die78]. selbstkritischen [Vol93]. Selected [Aga88a, Bun60b, Mah01, SW89]. Self [Vol93]. Self-Critical [Vol93]. selon [Sil13c]. Semántica [Bun95n]. Semantical [Bun60e]. Semantics [Bun72e, Bun73n, Bun74e, Bun74g, Bun74f, McF78, Bun77i, Bun78d, Bun95n, Bun62b]. sémantique [Bun78d]. Seminar [Aga72, Bun67a, Bun13b, Sne68a, Sne68b]. semiophysics [BP94]. sémiophysique [BP94]. Semiotic [Bun98e, Pic19]. Sempere [G.73]. Seni [Bun90f]. sens [Bun01b]. Sensation [Bun80o]. Sense [Bun74f, Bun76d]. Sentido [Bun76d, Qui58]. Sentiments [Bun07c]. seny [Mor16].

Separability [Bun95a, Bun01j, Bun84c]. September [Bun78d]. septembre [Bun78d]. Ser [Bun76d, Bun02b]. series [Bun78a, Bun78b, Bun78c, Bun78h, Bun78i, Bun78g]. Sets [Bun85h, Bun85g]. Seudociencia [Bun85m, Bun86k, Bun14h]. Seven [Bun87f, Bun93d, Bun96h]. sharing [Mac11]. Shea [Ano81, Sin81, Whe80]. Should [Bou91, Bun91f, Bun14a]. Side [Bun19a]. siglo [SA01]. Significado [Bun43a, Bun43c, Bun83l]. significance [Bun43a, Bun43c, Bun78i, Bun78c]. Similarities [Mat82]. Simon [Sle96]. Simplicity [Bun61d, Bun61h, Bun62d, BM63, Bun63d, Bun66f, FM66, dF19, Haj65, Har66, Mul63]. Simulation [Bun69b]. Single [Bun12k]. sistema [Bun95g, Ech85, Par83]. Sistemas [Bun85h, Bun95e, Bun99j, Mac11, QdAY13]. Sistemismo [Ama14]. Six [Hob19, Var17]. skeptic [Bun91i]. Skepticism [Kre91, Alc91, Bun00a]. Sloman [Bun90b]. so-called [Bun70e]. Sobre [Ano14b, Bun58g, Bun87i, Bun97h, Tor14, Yan14]. Social [Ano04, Asp98a, Asp98b, BRG04, Bra97, Bun90a, Bun94b, Bun95d, Bun96b, Bun98f, Bun01a, Bun01x, Bun01-30, Bun01-33, Bun01-37, Bun04c, Bun04f, Bun07b, Dem97, Esp86, Iva19, Kai09, Kal13, Kla03a, Pic04, Sad04, SB01, Van01, Van04, zW11b, Bel00, Bel01, Bra82, Bun74c, Bun75g, Bun76k, BO79, Bun88b, Bun93b, Bun93c, Bun95k, Bun96i, Bun96h, Bun03d, Cha06, CCB^{+}90, Kla03b, Kla04, LK74, Obr05, Par83, Pic01, Bun95e, Bun98i, Bun99j, Ech85, Col99, Lun00]. sociale [Bel00, Bel01]. Sociales [Bun95d, Bun95e, Bun76k, BO79, Bun95k, Bun98i, Bun99j, Bun04f, Cha06, CCB^{+}90]. Socialism [Bun56d, Bun01-31, Kar19b, Bun09d, BGD13]. socialismo [Bun09d, BGD13]. Socialismos [Bun17a]. Socialisms [Bun17a]. Sociality [Bun80p]. sociedad [Bun89j, Bun08b]. societario [BC14]. société [Déc10, Moe08]. Society [Aga91, Bun99a, Bun01a, Bun10m, Göt95, Mon92, Bun79l, Bun88b, Bun89j, Bun08b, Moe08]. Socio [Ray15]. Socio-historical [Ray15]. Sociología [Bun93e, Bun98j, Ano16, Bun91l, dP16, BG01]. sociological [Bun90g, Moy15]. sociológicos [Moy15]. sociologie [Bun98g, Cha69]. Sociology [Bun69d, Bun91b, Bun92a, Bun99b, Bun99k, Bun01e, Bun13i, Bun15d, Sad04, Tre01, zW12, Ano16, Bun91l, Bun98g, Cha69, CH09, Bun93e, Bun98j, Bun07b]. Sociology-Philosophy [Tre01]. sociotechnological [Mac11]. sociotecnológicos [Mac11]. Software [Bun10c]. Solar [Bar11]. Solution [BK83b]. Solutions [Bun06c]. Some [Art19, Bun79k, Bun04c, Vel22, Bun72f, Qui14]. son [Bel00, Bel01]. Soros [Bun99a]. Soul [Bun97g]. soulless [Bun83c]. sound [Bun83k]. sous [Bun76b, Bun76c]. sovremennoi [Bun62g]. Sozialarbeitswissenschaft [Obr05]. Sozialen [Kla03a, Kla04, Obr05, Kla03b]. Space [Bun55d, Bun58a, BG76, Bun91d]. Spacetime [Bun18b, Rom18]. Spanish [Aga14, Agu06, Ama14, Ano09, Ano14b, Ano15, Ano16, AC13, Bar20, BdCFG^{+}96, Bos14, BC14, Bun39, Bun43a, Bun43c, Bun43b, Bun44c, Bun44e, Bun44d, Bun45b, Bun51c, Bun51d, Bun52, Bun55g, Bun55c, Bun56f, Bun57b, Bun58g, Bun58f, Bun60d, Bun60i, Bun60c, Bun60j, Bun61i, Bun63f, Bun65b, Bun65c, Bun67h, Bun68q, Bun69h, Bun72h, Bun72g, Bun73f, Bun75i, Bun75j, Bun75d, Bun75e, Bun76k, Bun76l, Bun76j, Bun76d, Bun77b, Bun78j, BO79, Bun80s, Bun80t, Bun80d, Bun81m, Bun81i, Bun82c, Bun82a, Bun83b, Bun83h, Bun83i, Bun83l, Bun85p, Bun85b, Bun85h, Bun85l, Bun85q, Bun85m, Bun85f, Bun85k, Bun86j, Bun86f, Bun86k, Bun87j, Bun87i, Bun89i, Bun89j, Bun91l, BM92, Bun93e, Bun95f, Bun95g, Bun95h, Bun95k, Bun95i, Bun95j, Bun95l, Bun95m, Bun95d, Bun95e, Bun95n, Bun95o, Bun96i, Bun96a, BSO^{+}96, Bun96e, BR97, Bun97a, Bun97b, Bun97h]. Spanish [Bun98i, Bun98h, Bun98a, BM98, Bun98j, B^{+}99, Bun99j, Bun00h, Bun01-38, Bun01-35, Bun02b, Bun03a, BWM^{+}04, Bun04f, Bun06h, Bun06a, Bun08b, Bun09c, Bun09d, Bun11c, Bun12z, Bun13c, BGD13, Bun13e, Bun13h, Bun14c, Bun14e, Bun14g, Bun14h, Bun17a, BA17, Bun17g, Bun17d, Bun19f, CFF^{+}14, Cha06, Cor14, CCB^{+}90, Cri14, Den14a, Den14b, Eng47, Góm14, GML14, Kan15, Lab14, Mah14, Mar72, Mor16, Mos14, Moy15, OB84, OSV96, Par14, Par83, Pin13, Qui58, QB79, Qui14, Roj13, Sca14, SC89, Sil13f, SB92, Tor14, Vuc14, Yan14, dF78]. Species [Del12]. Speculation [Bun83k]. spin [Bun44b, Bun58g]. sponsor [Bun87g]. Sprache [Wie99]. Sprachen [Wie99]. Springer [Bun86c, Bun87a, Lin69, Mat16, Sne68b]. Springer-Verlag [Bun86c, Bun87a, Lin69, Sne68b]. SPSA [Obr05, Obr05]. Spurway [BS96]. Stance [Bun18d]. Standard [Bun12c]. standing [Pfe69]. state [Bun95g]. States [Bun77k, Bun88g, Jod13]. Stating [Bun64d]. Statistical [HH76]. statistics [Pin13]. Status [Bun01-29, Bun88f, Cha06]. Stegmüller [Bun76i]. Stove [Bun84b]. Strategies [Bun01-37, Bun11d, Van91, Bun85i]. strategy [Bun69h]. strawman [BvL91]. Strife [Bun55e, Bun55f, Bun06b, Bun14b, For07, Ric08, Tri06]. string [BL10]. strong [Kon95]. Structuralism [Mar19]. Structure [Bli90b, Bun67j, Bun13j, Bra82, Bun74c, Bun76i]. Studi [Gam14]. Studies [Bun90a, Bun94b, Bun04f, GG86, MB16, MC95, Sne68b, WD90, B^{+}99, Bun03d]. Study [Gam14, Bun39, Cha06]. stuff [Sza98]. style [GBS^{+}05]. subject [Sil13f]. Subjective [Bun12u]. Subjectivism [Bun03h]. succession [Bun68c, Bun71c]. successive [BK83a, BK84]. Such [Mil77]. Superposition [Bun88g]. Suppes [Bun73b]. supported [QdAY13]. suppress [GBS^{+}05]. surmonter [Bun04e]. Survey [Bun56e, Flø86, FvWK03, Kli68]. Survival [Bun94e, Bun89h]. Sus [BWM^{+}04]. sustainable [Bel00, Bel01]. sustained [MB96a]. Swiss [Aga91]. Symbolic [TBG^{+}85]. Symposium [BAB^{+}80, BD78, Bun78d, Har77, OW79, Pic04, SW89]. Symposiums [SW89]. synonymie [Sil13e]. Synonymy [Sil13e]. Synthesis [Bun11d]. Synthetic [MB01]. System [Bun67c, Bun92f, Cav12, Mar96, Par83, PW80, Bel00, Bel01, Bun44b, Bun91e, Bun93b, Bun93d, Bun95g, GSB81, Kla03b, Lin69, Kla03a]. Systematic [Mat12]. systematically [Bun14a]. Systeme [Bun89g, Bel00, Bel01, Bun91e]. systèmes [Cha69]. Systemic [Aga19, Bun01a, CD19, Hal19, Pic19, YCVP^{+}19, Bun88b, BT^{+}03, Bun06g, Bun09g, Bun15e, Kla03b, Kla04, Sil13c]. Systemics [Bun12v]. systemischen [Kla03b, Kla04]. Systemism [Bun00f, Bun04h, Sad04, zW11b, Obr05, Ama14]. Systemismus [Obr05]. Systemist [Del12, LSP22]. systémiste [Sil13c]. Systems [AK98, Bah81, Bun77c, Bun79j, Bun81l, Bun85h, Bun95e, Bun01-30, Bun04c, DB93, GKvWZ97, Kal19, Pic04, Pic19, Qui80, zW11a, WW90, Bun79l, Bun81a, BK83a, BK83b, BK84, Bun85g, Bun89g, Bun93d, Bun98e, Bun99j, Cha69, Kla04, Mac11, Mat82, Obr05, QdAY13, Har77, Bow80, Vol81]. Systemtheoretische [Obr05, Kla03a]. Systemtheorien [Kla03b, Kla04].

t [Bun13d]. tête [Bun83c]. Table [Ray19]. tackle [Bun57b]. tackled [Bun14a]. Taha [Obi22]. take [BK83a, BK84]. teacher [Bun75d]. Teaching [MB16, Pie99, BG16]. technical [Bun91e]. technical-science-philosophy [Bun91e]. technique [Bun74d, Bun75b, Bun91e, Bun96i]. technique-science-philosophie [Bun91e]. techniques [Bun79b]. Techno [AD04]. Techno-scientific [AD04]. Technoethics [Bun77l, Bun80q]. Technoholodemocracy [Bun01-31]. Technological [Bun19a, Bos14, Bun19f]. technologie [Bel00, Bel01]. Technologies [Bun01-32, Bun98b, Bun99f]. Technology [AL99, BD78, Bun66e, Bun76g, Bun85e, Bun86i, Bun96a, Bun97a, Bun01o, Bun01-33, Bun03f, Bun14c, Esp86, Kra80, LSP22, Bun74d, Bun78f, Bun83j, Bun88c, Bun91a, Bun98h, Bun12z, Sca14, Bel00, Bel01, Bun63f]. Technophilosophy [Kal19, Bun79f]. Technoscience [Bun12w]. Técnica [Bun96a, Bun97a, Bun14c, Bun96i, Bun98h]. Tecnología [Bun63f, Bun12z, Sca14]. tecnológica [Bos14]. tecnológicos [Bun19f]. Teil [Bun71a]. Temas [Bun43b]. temps [Bun68l]. Ten [Bun87g, Bun00g]. tenu [Bun76b, Bun76c]. Teorema [QB79, Qui79]. Teori [Gam14]. Teoría [Bun72g, Bun95o, QdAY13, Bun51d, Bun52, Ech85, Par83, Bun43a, G.73]. Terhadap [Gam14]. Terrace [Bun87a]. Test [Bun90h]. Testing [Bun19g]. text [AC13]. texto [AC13]. Their [Art19, Bun00e, Bun00g, Bun04c, Kla03a]. them [BK83a, BK84]. Themes [Sad04]. Theorema [QB79]. theoretic [Bun73c]. Theoretical [Kla03a, Mac85, Bun68g]. Theories [Bun58a, Bun61h, Bun66f, Bun12a, HH76, Ray19, Bun64b, Bun68d, Bun69g, Bun71g, Bun99i, Har77, Kla03b, Kla04, Bun68d, B^{+}71, B^{+}71, Bun76i]. Theorization [VOHPMJ12]. Theory [Aga72, Bas71, Bub68a, Bub68b, Bun55d, Bun63c, Bun67i, Bun67j, Bun68a, Bun68k, Bun68p, Bun70f, Bun73b, BG76, Bun95b, Bun95c, Bun99b, Bun01w, Bun03g, Bun09e, Bun12c, Bun13j, CD19, EB88, Fin70, Gam14, GKvWZ97, Gil19, HH76, Mil77, Ore22a, Par83, Sad04, Sle12b, Sne68c, Sol22, Vol93, zW11b, BRG04, Bun51d, Bun52, Bun56a, Bun70j, Bun77j, BS77, Bun79h, Bun80n, BK83b, Bun89e, Bun90f, Bun98c, Bun12f, Man85, Obr05, QdAY13, Sen93, vR77, Bun72g, Bun95o, Bun43a, Bun74b]. there [Bun60a, Bun75j, Bun09d]. Thermodynamics [Bun86b]. Thing [Mil77]. Things [Bun74h, BM04, Bun77j, vR77]. Think [Bun56b, Bun56c, Bun06h, Bun06a]. Thinker [Den19]. thinkers [Bun39]. Thinking [Aga19, Bun80r, Bun01-34, Pfe77]. Third [FJMT88]. thirteen [Bun88h]. Thom [BP94, Bun84a, BP94]. Thought [Bun04e, Bun72f, Den03, Sei06]. thoughts [GBS^{+}05]. Three [Bun01-36, Bun88i, Bun01-35]. threesome [Bun91e]. Tide [Bun67k, Dro15]. Tiempo [Bun97h, Bun01-35]. Tiene [Bun76d, Bun09d, Mor16, BGD13]. Time [Bun55d, Bun58c, Bun68k, Bun70g, Bun97h, Rus72, Bun68l, Bun70a, BK83a, BK84, Bun01-35, Bun03i, Bun53, Bun63a]. time-energy [Bun03i]. timeless [Bun60a]. Times [Bun51a]. timo [Bun14g, Bar11]. tipos [Bun44e]. Today [AC82, AD97, Bun92e]. Tokyo [Bun87a]. Tolerance [GPA91]. tool [Bun89e]. Tools [Bun73e, KPSP74]. topical [Bun79k]. Tópicos [DM00]. Toronto [Kil04, Pic07, Ric08]. total [Bun44b]. Tradition [BCD03, Wie99]. Training [Bun86a, Bun86c]. Translated [Bun71a, Bun84a]. translucent [Bun85b]. translúcidas [Bun85b]. Transparencies [Coh81]. transport [SBC^{+}08]. Treatise [Bun13k, Pat90, Qui80, Bar81, Bow80, Esp86, McF78, Rus80, Rus86, WD90]. Trees [Iva19]. Tres [Bun01-35]. triade [Mau17]. Triangle [Bun01-33]. Tribute [OSV96, BdCFG^{+}96]. Trier [Mon92]. Trigger [Bun13a, CC13]. Tristram [Bun11a]. trois [Bun91e]. tronquée [Esp94]. true [Mos91]. Truesdell [Bun86b, Bun86a, Bun86c]. truncated [Esp94]. truncheon [Lau91]. Truth [Bun67c, Bun68p, Bun74g, Bun05a, Mil77, Ray19, Sol22, Bun83l, BM92, Bun04b, Bun12f, Kre91, Bun91c, Bun53, Edg69]. Truth. [McF78]. Truths [Bun81h, Bun10b, Joh19]. Turn [Bun67k, BS79, Ano81, Dro15, Sin81, Whe80]. Twentieth [PS94]. Twenty [Bun03h]. Twenty-Five [Bun03h]. Two [Bun87h, Bun88i, Bun01-36, Bun01-37, Bun11d, Bun14e, Bun18a, Mat16, Bou91, Bun44b, BK83b, Bun16a, Bun16b, SBC^{+}08, GML14]. type [Cha69, Man85]. Types [Bah81, Bun44a, Bun85n, Bun44e].

u [Vuc14]. U.S [Bun76a]. U.S.A [KPSP74]. Ultimate [Sle96]. Umdenken [Pfe77]. UML [OHS02]. Under-Labourer [Aga12]. underlying [Bun72f]. Understanding [Bun80b, Bun83g, Bun04a, Bun13k, Göt95, Pan90, Sca14]. understood [Pan90]. Unification [Bun01-37, Bun11d]. Units [Bun17c]. Unity [AF01, Bun73h, Bun03c, Kil04, Pic07, SPT11]. Universal [Den19]. Universe [Bun55c]. Universes [Bun12q]. Universidad [Fonxx, Bun14d]. Universitaria [Fonxx]. Université [Bun76b, Bun76c]. University [Ano59, Asp98b, BAB^{+}80, BS96, BD78, Bun53, Bun61b, Din60, Kil04, Pic07, Ric08, Van04]. Universo [Bun55c]. Unjust [Bun12t]. unobservable [Jod13]. Unpopular [Bun87i]. unstable [BK83a, BK83b, BK84]. Unworldly [Bun01u]. update [Bun80d, Bun97b, Bun09c]. US$20 [BS96]. US$31 [Bun87a]. usages [Bun89b, Bun90c]. USD10 [Sle96]. Use [Del12]. Useful [Sza98, Bun89e]. uses [Bun89b, Bun90c]. Using [Bun18d, OHS02, BRG04].

V [Bun75j, Bun62g]. vagueness [Sei06]. valeurs [Bun01b]. Valuation [Bun12e]. value [Bun62a, Bun76e]. value-free [Bun76e]. values [Bun73i]. Variables [Bun95a, Bun01j, Bun84c, Cha69]. Vaya [Bar11, Bun14g]. veía [Bun45b]. Velocity [Bun03i]. verdad [Bun83l, BM92]. Vergleich [Kla03a, Kla03b, Kla04]. vérification [Bun68d, Bun68d]. vérité [Bun91c]. Verlag [Bun71a, Bun86c, Bun87a, Lin69, Sne68b]. Verlagstradition [Wie99]. Vers [Bun04i]. vez [Bun09d]. vibrations [BL10]. Vietnam [Bun73c]. View [AJ87, Del12]. Views [Bun87i, Bun97h]. viii [BS96, Sin81]. Virtual [Bun70h]. virtualidad [Bun01-35]. virtuality [Bun01-35]. Vision [Bun09f]. Vistas [Bun87i, Bun97h]. Vitae [Ano14a]. Voir [Moe08, Déc10]. Vol [Bun78a, Bun78b, Bun78c, Bun78h, Bun78i, Bun78g, Qui80, Rus80, Bar81, Bun79b, Esp86, Rus86, Sne68b]. Volume [Bow80, Bun98c, Bun98d, HH76, McF78, Sil13d, LM68, JL95]. Vortrag [Nar75]. Vorträge [Die78]. vs [BB89, Sal71].

W [Bun58b, Ano81, Bun84a]. W. [Bun75j]. Walter [Wie99]. Wand [KJ10, OHS02, RG02, RW05]. Wang [Bun76a]. war [Bun73c]. Wars [Bun94a, HBB^{+}94]. was [Bun79e, Bun09d, Kla03b, Kla04]. Wave [BG16, Bun56f]. waveguides [SBC^{+}08]. Waves [Bun18b, Rom18, Rom17, Vuc14]. way [Bun81l]. Wealth [Bun12x, Bun12y]. Weber [Bun07d, KJ10, McF07, OHS02, RG02, RW05]. Wechsel [SW89]. Weigartner [Mor81]. Weight [Bun61h, Bun66f]. Weingartner [Mor81, SD91]. Weird [Bun10q]. Welches [BK75]. Well [Bun12y, Bun12x]. Well-being [Bun12y, Bun12x]. Welt [HD88]. Weltgeltung [Pfe69]. Weltkongress [Die78]. Werk [Pfe69, Pfe77]. Werner [Bun71a, Pfe77, EB88, Pfe77]. Werten [Bun73i]. Western [Obi22]. Western-Scientific [Obi22]. Wheeling [Bun01-30]. Which [Bun00g]. Who [Bun05b, Bun79e]. Wild [Bun83k]. Wiley [Bun84a]. Wilhelm [Bun00b]. Will [Bun10e]. William [Sin81, Whe80, Bun56d]. Winter [BAB^{+}80]. Wirkung [Pfe77]. wisdom [Aga14, Den14a, Den14b]. wise [KK07]. Wissenschaft [Bun73i, Mor81, BM04]. Wissenschaften [Wie99]. Wissenschaftsphilosophie [Vol93]. Wissenschaftstheorie [Vol93]. without [Bun73i, Pan90]. Wittgenstein [SW89]. Wojtasiewicz [Bun62b]. Wolfgang [Bun76i]. Work [Bun04d, Bun12g, Kla03a, Mat19b, Rom19, Kla03b, Kla04, Obr05, Pfe77, Roj13, SB92, Pfe69]. Works [Bun00g, SL19]. World [Bow80, Bun77g, Bun79j, Bun83f, Bun83g, Bun01u, Bun13k, Die78, FJMT88, KFT89, Qui80, Vol81, Bun45b, Mat82, Bun86i, Bar81, Bun63a, Rus80, Vol79]. Worlds [Bun14e, Bun18a, Mat16, Bun16a, Bun16b]. Worldview [Bun10p, Obi11]. Worldviews [Obi22]. Wrong [AD04, McG94].

x [Bun87a, KPSP74]. xii [Bun07c, Sne68b]. xvi [Bun53]. XVIIe [Cau88]. xviii [Bun86c, Mat16]. xx [Din60, SA01]. xxx [Bun99a].

Yale [Asp98b]. yatrofilosofia [dF78]. Yatrophilosophy [dF78]. year [Wie99]. Years [BWM^{+}04]. York [Bun53, Bun71a, Bun76a, Bun84a, Bun87a, Bun99a, Bun11a, Har77, Lin69, Sin81, Sle96, Sne68b]. York-Tokyo [Bun87a].

Zecha [Bun90h]. Ziegler [Bar14]. Zürich [Aga91].

References

Agassi:2019:BCP

[ABA19] Joseph Agassi and Nimrod Bar-Am. Bunge contra Popper. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 15, pages 263–272. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Agassi:1982:SPT

[AC82] Joseph Agassi and Robert S. Cohen, editors. Scientific Philosophy Today: Essays in Honor of Mario Bunge, volume 67 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982. ISBN 90-277-1262-X, 90-277-1263-8 (paperback), 94-009-8462-6 (e-book). ISSN 0068-0346. x + 513 pp.

LCCN Q174 .B67 vol. 67. URL https://link.springer.com/
book/10.1007/978-94-009-8462-2.

Aycaguer:2013:RRZ

[AC13] Silva Ayçaguer and Luis Carlos. Reflexiones a raíz de filosofía para médicos, un texto de Mario Bunge. (Spanish) [Reflections on Philosophy for Doctors, a text by Mario Bunge]. Salud colectiva, 9(1):103–116, April 1, 2013.

Agazzi:1997:PMT

[AD97] Evandro. Agazzi and György Darvas, editors. Philosophy of Mathematics Today, volume 22 of Episteme. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4343-3. LCCN QA8.6.

Agazzi:2004:RWS

[AD04] Evandro Agazzi and Craig Dilworth. Right, Wrong and Science: the Ethical Dimensions of the Techno-scientific Enterprise, volume 81 of Poznań studies in the philosophy of the sciences and the humanities; Monographs-in-debate. Rodopi, Amsterdam, The Netherlands, 2004. ISBN 90-420-0919-5 (hardcover). 354 pp. LCCN BJ57 .A3513 2004.

Agazzi:2001:PUS

[AF01] Evandro Agazzi and Jan Faye, editors. The Problem of the Unity of Science: Proceedings of the Annual Meeting of the International Academy of the Philosophy of Science, Copenhagen– Aarhus, Denmark, 31 May–3 June 2000. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2001. ISBN 981-02-4791-5, 981-279-959-1 (e-book). LCCN

Q175.32.R47 A33 2000. URL http://ebooks.worldscinet.
com/ISBN/9789812799593/9789812799593.html.

Agassi:1972:BRD

[Aga72] Joseph Agassi. Book review: Delaware Seminar in the Foundations of Physics, and Quantum Theory and Reality by Mario Bunge. Philosophy of Science, 39(2):263–265, June 1972. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.jstor.org/stable/186730.

Agassi:1988:GAP

[Aga88a] Joseph Agassi. The Gentle Art of Philosophical Polemics: Selected Reviews and Comments. Open Court, La Salle, IL, USA, 1988. ISBN 0-8126-9036-2 (paperback), 0-8126-9063-2 (paperback), 0-912050-63-2 (hardcover). 521 pp. LCCN B73 .A43 1988.

Agazzi:1988:PS

[Aga88b] Evandro Agazzi, editor. Probability in the Sciences, volume 201 of Synthese library. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2808-9. x + 269 pp. LCCN Q174 .P76 1988.

Agazzi:1991:PRS

[Aga91] Evandro Agazzi, editor. The Problem of Reductionism in Science: Colloquium of the Swiss Society of Logic and Philosophy of Science, Zürich, May 18–19, 1990, volume 18 of Episteme. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1406-9 (hardcover), 94-010-5544-0, 94-011-3492-8 (e-book). LCCN Q174 .P82 1991.

Agassi:2012:BUL

[Aga12] Joseph Agassi. Between the under-labourer and the masterbuilder: Observations on Bunge’s method. Science & Education (Springer), 21(10):1405–1418, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Agassi:2013:BN

[Aga13] Joseph Agassi. Bunge nevertheless. Philosophy of the Social Sciences = Philosophie des sciences sociales, 43(4):542–562, 2013. ISSN 0048-3931 (print), 1552-7441 (electronic).

Agazzi:2014:FCI

[Aga14] Evandro Agazzi. La filosofía de la ciencia como introducción a la sabiduría. (Spanish) [The philosophy of science as an introduction to wisdom]. In Denegri [Den14a], pages 27–38. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Agazzi:2019:ST

[Aga19] Evandro Agazzi. Systemic thinking. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 13, pages 219–240. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

AguirreGarcia:2006:DAF

[Agu06] Juan Carlos Aguirre Garcia. 10. desarrollos actuales en filosofía de la ciencia. (Spanish) [10. Current developments in philosophy of science]. Cuadernos de epistemología, ??(1):41–54, 2006.

URL https://issuu.com/juanaguirre/docs/cuadernos_de_
epistemolog_a__n_mero_1.

Ahmad:2019:MCR

[Ahm19] Omar Ahmad. Mechanisms in clinical research and medical practice. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 39, pages 735–749. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Agassi:1987:RCV

[AJ87] Joseph Agassi and I. C. (Ian Charles) Jarvie, editors. Rationality: the Critical View, volume 23 of Nijhoff international philosophy series. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1987. ISBN 90-247-3275-1, 90-247-3455-X. xi + 477 pp. LCCN

B833 .R344 1986. URL http://www.springer.com/gb/.

Altmann:1998:SNP

[AK98] Gabriel Altmann and Walter A. Koch, editors. Systems: new paradigms for the human sciences. Walter de Gruyter, New York, NY, USA, 1998. ISBN 3-11-015619-9, 3-11-080119-1 (e-book). xiv + 781 pp. LCCN H61 .S993 1998.

Agazzi:1999:APT

[AL99] Evandro Agazzi and Hans Lenk, editors. Advances in the Philosophy of Technology: Proceedings of a Meeting of the International Academy of the Philosophy of Science, Karlsruhe, Germany, May 1997. Society for Philosophy and Technology, Newark, DE, USA, 1999.

Alcock:1991:IMS

[Alc91] James E. Alcock. On the importance of methodological skepticism. New Ideas in Psychology, 9(2):151–155, 1991. ISSN 0732-118X (print), 1873-3522 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0732118X9190018H.

Special issue on Mario Bunge and Martin Mahner.

AmadoYannarella:2014:SAI

[Ama14] Ana María Amado Yannarella. El Sistemismo como una alternativa de investigación posible a la problemática del delito. (Spanish) [Systemism as a possible research alternative to the problem of crime]. In Denegri [Den14a], pages 39–72. ISBN 950-23-2339-4.

LCCN BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Anonymous:1959:BNC

[Ano59] Anonymous. Book note: Causality: by Mario Bunge. 380 pages, diagrams, 512 × 814 in. Cambridge, Mass. Harvard University Press, 1959. Price, $7.50. Journal of The Franklin Institute, 268(2):129, 1959. CODEN JFINAB. ISSN 0016-0032 (print),

1879-2693 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0016003259904648.

Anonymous:1974:MBB

[Ano74] Anonymous. [Mario Bunge: brief biography]. International Journal of General Systems, 1(4):281, 1974. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Anonymous:1975:PL

[Ano75] Anonymous. Pool of lecturers. International Journal of General Systems, 2(2):73–74, January 1975. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Anonymous:1979:A

[Ano79a] Anonymous. Announcements. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 11(1):141–142, March 1979. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Anonymous:1979:P

[Ano79b] Anonymous. Preface. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 10(1–4):1–6, January 1979. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Bunge:1981:BRR

[Ano81] Anonymous. Book review: M. Bunge and W. R. Shea, Rutherford and Physics at the Turn of the Century. Journal for the History of Astronomy, 12:75, January 1981. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1981JHA....12...75S.

Anonymous:1990:E

[Ano90] Anonymous. Editorial. New Ideas in Psychology, 8(2–3):????,

1990.

Anonymous:1996:C

[Ano96a] Anonymous. Contributors. Science & Education (Springer), 5 (2):219–220, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Anonymous:1996:E

[Ano96b] Anonymous. Editorial. Science & Education (Springer), 5(2): 91–99, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Anonymous:2004:INS

[Ano04] Anonymous. Introductory note to the second issue on Bunge’s philosophy of social science philosophy of the social sciences. Philosophy of the Social Sciences = Philosophie des sciences sociales, 34(3):325–327, 2004. ISSN 0048-3931 (print), 1552-7441 (electronic).

Anonymous:2009:CMB

[Ano09] Anonymous. Una conferencia de Mario Bunge. (Spanish) [A lecture by Mario Bunge]. Mientras Tanto, ??(112):23, Otoño 1, 2009. CODEN ???? ISSN 0210-8259 (print), 2014-7317 (elec-

tronic). URL https://www.jstor.org/stable/27821351.

Anonymous:2010:BMf

[Ano10a] Anonymous. Back matter. In Matter and Mind: a Philosophical Inquiry [Bun10k], pages 287–319. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https://link.springer.

com/book/10.1007/978-90-481-9225-0.

Anonymous:2010:FMq

[Ano10b] Anonymous. Front matter. In Matter and Mind: a Philosophical Inquiry [Bun10k], page 141. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664

2010; Q175 .B73 v.287. URL https://link.springer.com/
book/10.1007/978-90-481-9225-0.

Anonymous:2010:FMr

[Ano10c] Anonymous. Front matter. In Matter and Mind: a Philosophical Inquiry [Bun10k], page 265. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664

2010; Q175 .B73 v.287. URL https://link.springer.com/
book/10.1007/978-90-481-9225-0.

Anonymous:2012:BMj

[Ano12a] Anonymous. Back matter. In Evaluating Philosophies [Bun12j], pages 185–200. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL http://www.loc.gov/

catdir/enhancements/fy1506/2012941366-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1506/2012941366-
t.html;   https://link.springer.com/book/10.1007/978-
94-007-4408-0.

Anonymous:2012:FMaa

[Ano12b] Anonymous. Front matter. In Evaluating Philosophies [Bun12j], page 103. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL http://www.loc.gov/

catdir/enhancements/fy1506/2012941366-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1506/2012941366-
t.html;   https://link.springer.com/book/10.1007/978-
94-007-4408-0.

Anonymous:2012:FMz

[Ano12c] Anonymous. Front matter. In Evaluating Philosophies [Bun12j], page 49. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL http://www.loc.gov/

catdir/enhancements/fy1506/2012941366-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1506/2012941366-
t.html;   https://link.springer.com/book/10.1007/978-
94-007-4408-0.

Anonymous:2014:PMB

[Ano14a] Anonymous. Profesor Mario Bunge. Curriculum vitae. In Denegri [Den14a], pages 337–382. ISBN 950-23-2339-4. LCCN

BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Anonymous:2014:ASA

[Ano14b] Anonymous. Sobre los autores. (Spanish) [About the authors]. In Denegri [Den14a], pages 383–389. ISBN 950-23-2339-4.

LCCN BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Anonymous:2015:EIM

[Ano15] Anonymous. Entrevista informal a un Mario Bunge lejano. (Spanish) [Informal interview with a distant Mario Bunge]. Revista española de física, 29(2):53–55, ???? 2015.

Anonymous:2016:CTN

[Ano16] Anonymous. Crítica de la nueva sociología de la ciencia, de Mario Bunge. (Spanish) [Critique of the new sociology of science, by Mario Bunge]. Teorema: Revista Internacional de Filosofía, 35 (2):209–214, ???? 2016. ISSN 0210-1602.

Anonymous:2023:MB

[Ano23] Anonymous. Mario Bunge. Web site, November 28, 2023. URL

https://en.wikipedia.org/wiki/Mario_Bunge.

Apostol:1985:AMB

[Apo85] Ravel Apostol. About Mario Bunge’s “A Critical Examination of Dialectics”. Studies in Soviet Thought, 29(2):89–136, February 1985. CODEN SSVTBD. ISSN 0039-3797 (print), 2214-7136 (electronic).

Arthur:2019:MBC

[Art19] Richard T. W. Arthur. Mario Bunge on causality: Some key insights and their Leibnizian precedents. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 11, pages 185–204. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Aspers:1998:BRF

[Asp98a] Patrik Aspers. Book review: Finding Philosophy in Social Science by Mario Bunge. Acta Sociologica, 41(2):184–185, ???? 1998. CODEN ASOGAC. ISSN 0001-6993 (print), 1502-3869

(electronic). URL https://www.jstor.org/stable/4201077.

Aspers:1998:BRM

[Asp98b] Patrik Aspers. Book reviews: Mario Bunge: Finding Philosophy in Social Science: New Haven: Yale University Press, 1996. Acta Sociologica, 41(2–3):184–185, 1998. CODEN ASOGAC. ISSN 0001-6993 (print), 1502-3869 (electronic).

Bunge:1971:TC

[B^{+}71] Mario Bunge et al. Les théories de la causalité. (French) [Theories of causality], volume 25 of Études d’épistémologie génétique; Bibliothèque scientifique internationale. Presses universitaires de France, Paris, France, 1971. 208 pp. LCCN QC6 .T475.

Bunge:1999:BCC

[B^{+}99] Mario Bunge et al. Borges científico: cuatro estudios. (Spanish) [Borges scientist: four studies]. Colección Fin del milenio. Ediciones Biblioteca Nacional, Buenos Aires, Argentina, 1999. ISBN 987-503-197-6. 79 pp. LCCN PQ7797.B635 Z63446 1999.

Bunge:1987:PP

[BA87] Mario Bunge and Rubén Ardila. Philosophy of Psychology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96442-8, 1-4612-9118-6,1-4612-4696-2 (e-book). xii + 327 pp. LCCN BF38 .B85 1987.

Bunge:1988:FP

[BA88] Mario Augusto Bunge and Rubén Ardila. Filosofía de la psicología. Ariel psicología. Ariel, Barcelona, Spain, 1988. ISBN 84-344-0853-8 (paperback). 435 pp. LCCN BF38 .B8618 1988.

Bunge:1990:PP

[BA90] Mario Bunge and Rubén Ardila. Philosophie der Psychologie. (German) [Philosophy of Psychology], volume Band 65 of Die Einheit der Gesellschaftswissenschaften. Mohr, Tübingen, Germany, 1990. ISBN 3-16-345550-6. ISSN 0424-6985. 469 pp. LCCN IN PROCESS (WBC).

Bunge:2014:PP

[BA14] Mario Bunge and Rubén Ardila. Philosophy of Psychology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 1-4612-9118-6, 1-4612-4696-2 (ebook). LCCN ????

Bunge:2017:EDC

[BA17] Mario Bunge and Gabriel (Andrade Campo-Redondo) Andrade. Elogio del Cientificismo. (Spanish) [Praise of Scientism], volume 10 of Biblioteca Bunge. Laetoli, Pamplona, Spain, 2017. ISBN 84-946742-3-4. 255 pp. LCCN ????

Bindra:1980:BMN

[BAB^{+}80] Dalbir Bindra, James A. Anderson, Mario Bunge, John P. Flynn, Peter M. Milner, and Theodore L. Sourkes, editors. The Brain’s Mind: a Neuroscience Perspective on the Mind–Body Problem. Proceedings of a Symposium Held at McGill University, Montreal in the Winter of 1978. Gardner Press, New York, NY, USA, 1980. ISBN 0-89876-003-8. LCCN BF161 .B5.

Bar-Am:2017:EOE

[BAG17] Nimrod Bar-Am and Stefano Gattei. Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday, volume 325 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). xiv + 584 pp. LCCN B833.E53 2017. URL https://

link.springer.com/book/10.1007/978-3-319-57669-5.

Bahm:1981:FTS

[Bah81] Archie J. Bahm. Five types of systems philosophy. International Journal of General Systems, 6(4):233–237, March 1981. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Barreau:1981:BRT

[Bar81] H. Barreau. Book review: Treatise on Basic Philosophy. vol. 3, Ontology I: The Furniture of the World by Mario Bunge. Revue de métaphysique et de morale, ??(1):112–117, January 1981. ISSN 0035-1571 (print), 2102-5177 (electronic). URL https://www.

jstor.org/stable/40901788.

Barcelo:2011:BRP

[Bar11] Alfons Barceló. Book review: Las pseudociencias. ‘!Vaya timo! by Mario Bunge, Alfonso López Borgoñoz, Rafael González del Solar. Mientras Tanto, ??(116):119–123, ???? 2011. CODEN ???? ISSN 0210-8259 (print), 2014-7317 (electronic). URL

https://www.jstor.org/stable/41703219.

Barcelo:2014:BRN

[Bar14] Alfons Barceló. Book review: Los nuevos amos del mundo Y la lucha de aquellos que se resisten a dejarse engullir por la globalización by Jean Ziegler, Eduardo Gonzalo. Mientras Tanto, ?? (122–123):209–222, ???? 2014. CODEN ???? ISSN 0210-8259 (print), 2014-7317 (electronic). URL https://www.jstor.org/

stable/43491774.

Barcelo:2020:IBS

[Bar20] Alfons Barceló. Interpretando a Bunge. (Spanish) [Interpreting Bunge], volume 33 of Las dos culturas. Editorial Laetoli, Madrid and Pamplona, Spain, 2020. ISBN 84-12-18564-1, 84-949717-3-5. 153 pp. LCCN B1034.B864 A65 2020.

Bastin:1971:QTB

[Bas71] Ted Bastin, editor. Quantum Theory and Beyond: Essays and Discussions Arising from a Colloquium. Cambridge University Press, Cambridge, UK, 1971. ISBN 0-521-07956-X. LCCN QC174.1 .Q37. URL http://www.loc.gov/

catdir/enhancements/fy0907/77127237-d.html;   http://
www.loc.gov/catdir/enhancements/fy0907/77127237-t.html.

Bunge:1974:R

[BB74] Mario Bunge and Radu J. Bogdan. Reviews. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 5(2):243–247, August 1974. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Bunge:1975:PP

[BB75] Mario Bunge and Françoise Balibar. Philosophie de la Physique. Éditions du Seuil, Paris, France, 1975. ISBN 2-02-002834-4. 299 pp. LCCN ????

Blitz:1989:GVS

[BB89] D. Blitz and Mario Bunge. Gradualism vs. saltationism in evolutionary biology from Darwin to Gould. In Schurz and Weingartner [SW89], pages 297–301. ISBN 3-209-00861-2 (paperback). LCCN B3376.W564; Q174 .I586 1988.

Bud:2010:AP

[BB10] Robert Bud and Mario Bunge. For and against psychoanalysis. New Scientist, 208(2780):22–23, 2010. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0262407910624001.

Botteri:2014:DSA

[BC14] José David Botteri and Diego Coste. El derecho societario argentino y la filosofía de Mario Bunge. (Spanish) [Argentine corporate law and the philosophy of Mario Bunge]. In Denegri [Den14a], pages 99–120. ISBN 950-23-2339-4. LCCN

BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Boudon:2003:ETQ

[BCD03] Raymond Boudon, Mohamed Cherkaoui, and Pierre Demeulenaere, editors. The European Tradition in Qualitative Research, volume 4 of Sage benchmarks in social research methods. Sage, London, UK, 2003. ISBN 0-7619-7438-5 (hardcover), 1-4462-6360-6 (e-book). 426 pp. LCCN H62 .E82 2003. URL http:

//sk.sagepub.com/navigator/the-european-tradition-in-
qualitative-research.

Bugliarello:1978:HPT

[BD78] George Bugliarello and Dean B. Doner, editors. The History and Philosophy of Technology: [Symposium on the History and Philosophy of Technology, May 14–16, 1973, at the Chicago Circle campus of the University of Illinois]. University of Illinois Press, Urbana, IL, USA, 1978. ISBN 0-252-00462-0. LCCN T15 .S96 1973.

BarraganVargas:1996:FME

[BdCFG^{+}96] Miguel Barragán Vargas, María del Carmen Freyssinier Vera, Raúl Gutiérrez Romero, León Olivé, Fernando Salmerón Roiz, and Luis Villro, editors. Filosofía moral, educación e historia: homenaje a Fernando Salmerón. (Spanish) [Moral philosophy, education, and history: tribute to Fernando Salmerón]. Facultad de Filosofía y Letras, Instituto de Investigaciones Filosóficas, Universidad Nacional Autónoma de México, México, DF, México, 1996. ISBN 968-36-5425-8. 776 pp. LCCN BJ424.S253 F5 1996.

Belley:2000:TSM

[Bel00] Philippe Belley. La technologie dans le système de Mario Bunge et son application à l’ingénierie sociale: Le cas du développement durable. (French) [Technology in Mario Bunge’s system and its application to social engineering: The case of sustainable development]. Maître ès arts (M.A.), Faculté de philosophie, Université Laval, Québec, QC, Canada, December 2000. vii

+ 122 pp. URL https://search.proquest.com/pqdtglobal/
docview/304657210/.

Belley:2001:TSM

[Bel01] P. Belley. La technologie dans le système de Mario Bunge et son application a l’ingéniérie sociale: Le cas du développement durable. (French) [Technology in Mario Bunge’s system and its application to social engineering: the case of sustainable development]. ????, ??(??):??, ???? 2001.

Berger:1982:BRM

[Ber82] George Berger. Book review: The Mind-Body Problem, a Psychobiological Approach by Mario Bunge. Erkenntnis (1975–), 17 (3):399–403, May 1982. CODEN ERKEDQ. ISSN 0165-0106 (print), 1572-8420 (electronic). URL https://www.jstor.org/

stable/20010785.

Bunge:1976:RTP

[BG76] Mario Bunge and Adalberto García Máynez. A relational theory of physical space. International Journal of Theoretical Physics, 15(12):961–972, December 1976. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1976IJTP...15..961B.

Bunge:2001:RES

[BG01] Mario Augusto Bunge and María Dolores González Rodríguez. La relación entre la sociología y la filosofía. Edaf, Madrid, Spain, second edition, 2001. ISBN 84-414-0767-3 (paperback). 359 pp. LCCN HM585 .B8518 2001.

Buonaura:2016:WPD

[BG16] Biagio Buonaura and Giuseppe Giuliani. Wave and photon descriptions of light: historical highlights, epistemological aspects and teaching practices. European Journal of Physics, 37(5): 055303:1–055303:35, July 2016. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

Bunge:2013:EPS

[BGD13] Mario Bunge, Carlos Gabetta, and Antoni Domènech. ¿Tiene porvenir el socialismo? (Spanish) [Does socialism have a future?]. Lectores. Eudeba, Ciudad de Buenos Aires, Argentina, 2013. ISBN 950-23-2197-9. 197 pp. LCCN HX73 .T53 2013.

Bunge:1976:DPC

[BGS76] Mario Bunge and Máximo García-Sucre. Differentiation, participation and cohesion. Quality & Quantity, 10(??):171–178, ???? 1976. CODEN QQEJAV. ISSN 0033-5177 (print), 1573-7845 (electronic).

Bunge:1969:CPO

[BK69] Mario Bunge and Andrés J. Kálnay. A covariant position operator for the relativistic electron. Progress of Theoretical Physics, 42(6):1445–1459, December 1, 1969. CODEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1969PThPh..42.1445B.

Bunge:1975:WBQ

[BK75] Mario Bunge and Andrés J. Kálnay. Welches sind die Besonderheiten der Quantenphysik gegenüber der klassischen Physik?.

(German) [What are the peculiarities of quantum physics compared to classical physics?]. In Philosophie und Physik. (German) [Philosophy and physics] [HG75], pages 25–38. ISBN 3-322-86073-6, 3-528-08335-2 (print). LCCN B1-5802. URL https:/

/link.springer.com/chapter/10.1007/978-3-322-86073-6_
4.

Bunge:1983:RSM

[BK83a] Mario Bunge and Andrés J. Kálnay. Real successive measurements on unstable quantum systems take nonvanishing time intervals and do not prevent them from decaying. Il Nuovo Cimento B (11), 77(1):10–18, September 1983. CODEN NCIBAW. ISSN 0369-3554 (print), 1826-9877 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/1983NCimB..77...10B.

Bunge:1983:STP

[BK83b] Mario Bunge and Andrés J. Kálnay. Solution to two paradoxes in the quantum theory of unstable systems. Il Nuovo Cimento B (11), 77(1):1–9, September 1983. CODEN NCIBAW. ISSN 0369-3554 (print), 1826-9877 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1983NCimB..77....1B.

Bunge:1984:RSM

[BK84] Mario Bunge and Andrés J. Kálnay. Real successive measurements on unstable quantum systems take nonvanishing time intervals and do not prevent them from decaying. Nuovo Cimento B

Serie, 80:143, March 1984. URL https://ui.adsabs.harvard.
edu/#abs/1984NCimB..80..143B.

Berghel:1975:R

[BKB75] H. L. Berghel, Hermann Kotthoff, and Mario Bunge. Reviews. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 6(2):239–247, May 1975. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Beaumont:1996:MBP

[BKR96] J. Graham Beaumont, Pamela M. Kenealy, and Marcus Rogers, editors. The Blackwell Dictionary of Neuropsychology. Blackwell Publishers, Cambridge, MA, USA, 1996. ISBN 0-631-17896-1 (hardcover), 0-631-21435-6 (paperback). xix + 788 pp. LCCN

QP360 .B577 1996. URL http://www.tandfonline.com/toc/
rwhi20/.

Bunge:1971:BC

[BL71] Mario Bunge and Ralph W. Lewis. Biologic constructs. The American Biology Teacher, 33(6):363–364, September 1971. ISSN 0002-7685 (print), 0002-8444 (electronic), 1938-4211, 1938-422X.

URL http://www.jstor.org/stable/4443540.

Bank:2010:PLC

[BL10] Balaźs Bank and Heidi-Maria Lehtonen. Perception of longitudinal components in piano string vibrations. Journal of the Acoustical Society of America, 128(3):EL117, 2010. CODEN JASMAN. ISSN 0001-4966.

Blackmore:1992:EMD

[Bla92] John Blackmore. Ernst Mach — a Deeper Look: Documents and New Perspectives, volume 143 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. xix + 463 pp.

LCCN C5.53; QC6.9. URL https://link.springer.com/book/
10.1007/978-94-011-2771-4.

Blackford:2019:BSI

[Bla19] Russell Blackford. Bunge on science and ideology: A re-analysis. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 26, pages 439–463. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Blitz:1990:B

[Bli90a] David Blitz. Bibliography. In Weingartner and Dorn [WD90], pages 685–709. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990.

Blitz:1990:EEL

[Bli90b] David Blitz. Emergent evolution and the level structure of reality. In Weingartner and Dorn [WD90], pages 153–169. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990.

Blitz:1991:LDB

[Bli91] David Blitz. The line of demarcation between science and nonscience: The case of psychoanalysis and parapsychology. New Ideas in Psychology, 9(2):163–170, 1991. ISSN 0732-118X (print),

1873-3522 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0732118X9190020M. Special issue on

Mario Bunge and Martin Mahner.

Bunge:1963:MS

[BM63] Mario Bunge and Albert A. Mullin. The myth of simplicity. American Journal of Physics, 31:891, November 1963. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL https://ui.adsabs.harvard.edu/#abs/1963AmJPh..31.
.891B.

Bunge:1992:LVS

[BM92] Mario Bunge and Jean-Pierre Marquis. Lógica y verdad. (Spanish) [Logic and truth]. In Sobrevilla and Balaúnde [SB92], pages 359–369.

Bunge:1998:EE

[BM98] Mario Bunge and Luis Marone. La explicación en ecología. (Spanish) [Explanation in ecology]. Boletín de la Asociación Argentina de Ecología, 7(2):35–37, ???? 1998.

Bunge:2004:NDM

[BM04] Mario Bunge and Martin Mahner. Über die Natur der Dinge: Materialismus und Wissenschaft. (German) [On the Nature of Things: Materialism and Science]. S. Hirzel Verlag, Stuttgart, Germany, 2004. ISBN 3-7776-1321-5. xi + 273 pp. LCCN B825 .B96 2004.

Bunge:1979:ICS

[BO79] Mario Bunge and Mario H. Otero. Ideología y ciencias sociales. (Spanish) [Ideology and social sciences]. Universidad nacional autonoma de México, México D.F., México, 1979. 231 pp. LCCN H83 .I3.

Bosch:2014:PCF

[Bos14] Marcelo Bosch. Prospectiva científica y tecnológica: una aproximación no-hegemónica. (Spanish) [Scientific and technological foresight: a non-hegemonic approach]. In Denegri [Den14a], pages 73–98. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL

http://core.cambeiro.com.ar/0-100933-7.pdf;   http://
www.loc.gov/catdir/toc/gc01/9789502323398.pdf.

Boudon:1991:TQS

[Bou91] Raymond Boudon. On two questions: Should man be considered as rational? How to distinguish science from nonscience? New Ideas in Psychology, 9(2):171–179, 1991. ISSN 0732-118X (print),

1873-3522 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0732118X9190021D. Special issue on

Mario Bunge and Martin Mahner.

Boudon:1999:F

[Bou99] Raymond Boudon. Foreword. In The Sociology–Philosophy Connection [Bun99k], pages xi–xv. ISBN 1-56000-416-9. LCCN HM585 .B85 1999.

Bowler:1980:BRT

[Bow80] T. Downing Bowler. Book review: Treatise on Basic Philosophy, Volume 4, Ontology II: a World of Systems, Mario Bunge, D. Reidel Publishing Company, Boston, 1979. 291 pp. International Journal of General Systems, 6(1):51–53, March 1980. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Bompard-Porte:1994:PFD

[BP94] Michèle Bompard-Porte. Passion des formes: dynamique qualitative, sémiophysique et intelligibilité: à René Thom (French) [Passion of forms: qualitative dynamics, semiophysics and intelligibility: to René Thom]. Theoria. ENS éditions, Fontenayaux-Roses, France, 1994. ISBN 2-902126-09-3 (paperback). 821 (two volumes) pp. LCCN Q175 .P34238 1994. URL http://

catalogue.bnf.fr/ark:/12148/cb35763316f.

Bunge:1997:CPC

[BR97] Mario Bunge and Hernán Rodríguez. La Causalidad: el Principio de Causalidad en la Ciencia Moderna. (Spanish) [Causality: the Principle of Causality in Modern Science]. Biblioteca Mario Bunge. Editorial Sudamericana, Buenos Aires, Argentina, revised edition, 1997. ISBN 950-07-1288-1. 507 pp. LCCN BD541 .B8518 1997.

Braun:1982:LSA

[Bra82] Günther E. Braun. The logical structure of applied social science. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 14(1):1–18, March 1982. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Branaman:1997:BRF

[Bra97] Ann Branaman. Book review: Finding Philosophy in Social Science by Mario Bunge. Contemporary Sociology — a Journal of Reviews, 26(6):783–784, November 1997. ISSN 0094-3061 (print),

1939-8638 (electronic). URL https://www.jstor.org/stable/
2654688.

Barrachina:2004:SRE

[BRG04] Mercedes Barrachina, Vicente Ripoll, and Susana Gago. Social research evolution in management accounting: reflection using Bunge’s theory. Critical Perspectives on Accounting, 15(4): 701–713, 2004. ISSN 1045-2354 (print), 1095-9955 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S104523540300056X. A Critical Response to Managerialism in the Academy.

Brush:1973:BRP

[Bru73] Stephen G. Brush. Book review: Philosophy of Physics, by Mario Bunge. D. Reidel Publishing Company, Boston, Mass. 1973. Physics Today, 26(9):61, September 1973. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/1973PhT....26i..61B.

Bruck:1991:BRE

[Bru91] Andreas Bruck. Book review: Ethics. The Good and the Right by Mario Bunge. Anthropos, ??(??):234–236, ???? 1991. URL

https://www.jstor.org/stable/40462417.

Bunge:1977:TPK

[BS77] Mario Bunge and Arturo A. L. Sangalli. A theory of properties and kinds. International Journal of General Systems, 3(3):183– 190, January 1977. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Bunge:1979:RPT

[BS79] Mario Bunge and William R. Shea, editors. Rutherford and Physics at the Turn of the Century. Dawson, Kent, UK, 1979. ISBN 0-88202-184-2 (New York), 0-7129-0918-4 (Cannon House Folkestone). 184 pp. LCCN QC7.5 .R87 1979.

Birch:1996:BNC

[BS96] Charles Birch and John Stenhouse. Book notes: [Colin A. Russell, The Earth, Humanity and God, London, University College Press, 1994. pp. viii + 193, US$20. Neil Spurway (ed.), Humanity, Environment and God: Glasgow Centenary Gifford Lectures, Oxford, Blackwell, 1993. pp. viii + 232, $19.95 (cloth)]. Science & Education (Springer), 5(2):213–218, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Bunge:1996:FME

[BSO^{+}96] Mario Bunge, Ernesto Sora, Leon Olive, et al. Filosofía moral y educación. (Spanish) [Moral philosophy and education]. In Olivé et al. [OSV96], page ?? ISBN 968-36-5425-8. LCCN BJ1142.

Bunge:2003:TSA

[BT^{+}03] Mario Bunge, G. Thurler, et al. Toward a systemic approach to disease. ComPlexUs, 1(??):117–122, ???? 2003. ISSN 1424-8492 (print), 1424-8506 (electronic).

Bub:1968:BRB

[Bub68a] Jeffrey Bub. Book review: Quantum Theory and Reality by Mario Bunge. Philosophy of Science, 35(4):425–429, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/186257.

Bub:1968:BRQ

[Bub68b] Jeffrey Bub. Book review: Quantum Theory and Reality by Mario Bunge. Philosophy of Science, 35(4):425–429, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL https://www.jstor.org/stable/186257.

Bunge:1939:IAE

[Bun39] Mario Augusto Bunge. Introducción al estudio de los grandes pensadores. (Spanish) [Introduction to the study of great thinkers]. In ????, editor, Conferencias (Buenos Aires) III, pages 105–109, 124–126. ????, ????, 1939.

Bunge:1943:SFHa

[Bun43a] Mario Bunge. Significado físico e histórico de la Teoría de Maxwell. (Spanish) [Physical and historical significance of the Maxwell’s theory]. Tall. Gráficos Suárez y Dini., Buenos Aires, Argentina, 1943. 16 pp.

Bunge:1943:TEP

[Bun43b] Mario Bunge. Temas de educación popular. (Spanish) [Popular education issues]. El Ateneo (Buenos Aires), ??(??):??, 1943.

Bunge:1943:SFHb

[Bun43c] Mario Augusto Bunge. Significado físico e histórico de las ecuaciones de Maxwell. (Spanish) [Physical and historical significance of the Maxwell equations]. Universidad Obrera Argentina, Buenos Aires, Argentina, 1943. ???? pp.

Bunge:1944:NRTb

[Bun44a] Mario Bunge. A new representation of types of nuclear forces. Physical Review, 65(7–8):249–250, April 1944. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL https://ui.adsabs.harvard.edu/#abs/1944PhRv...65.
.249B.

Bunge:1944:TSM

[Bun44b] Mario Bunge. The total spin of a mass-system of two particles. Revista Union Mat. Argentina, 10:13–14, 1944. ISSN 0041-6932 (print), 1669-9637 (electronic).

Bunge:1944:AFF

[Bun44c] Mario Augusto Bunge. Auge y fracaso de la filosofía de la naturaleza. (Spanish) [Rise and failure of the philosophy of nature]. Minerva (Buenos Aires), 1(??):213–235, ???? 1944.

Bunge:1944:QES

[Bun44d] Mario Augusto Bunge. ‘?Qué es la epistemología?. (Spanish) [What is epistemology?]. Minerva (Buenos Aires), 1(??):27–43, ???? 1944.

Bunge:1944:NRTa

[Bun44e] Mario Augusto Bunge. Una nueva representación de los tipos de fuerzas nucleares. (Spanish) [A new representation of the types of nuclear forces]. Revista de la Facultad de Ciencias Físicomatemáticas (La Plata, Argentina), ??(??):221–

239, 1944. URL http://catalogosuba.sisbi.uba.ar/vufind/
Record/201603080901575367.

Bunge:1945:NPS

[Bun45a] Mario Bunge. Neutron–proton scattering at 8.8 and 13 MeV. Nature, 156(3958):301, September 8, 1945. CODEN NATUAS.

ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/1945Natur.156..301B.

Bunge:1945:CVM

[Bun45b] Mario Augusto Bunge. Cómo veía el mundo Florentino Ameghino. (Spanish) [How Florentino Ameghino saw the world]. Minerva (Buenos Aires), 2(??):184–185, ???? 1945.

Bunge:1950:IE

[Bun50] Mario Bunge. The inexhaustible electron. Science & Society, 14(2):115–121, Spring 1950. ISSN 0036-8237 (print), 1943-2801

(electronic). URL https://www.jstor.org/stable/40399998.

Bunge:1951:BRE

[Bun51a] Mario Bunge. Book review: Einstein, His Life and Times by Philipp Frank. Science & Society, 15(1):89–93, Winter 1951. ISSN 0036-8237 (print), 1943-2801 (electronic). URL https:/

/www.jstor.org/stable/40400053.

Bunge:1951:WC

[Bun51b] Mario Bunge. What is chance? Science & Society, 15(3):209–231, Summer 1951. ISSN 0036-8237 (print), 1943-2801 (electronic).

URL https://www.jstor.org/stable/40400071.

Bunge:1951:FCS

[Bun51c] Mario Augusto Bunge. La fenomenología y la ciencia. (Spanish) [Phenomenology and science]. Cuadernos Americanos, ??(4):108– 122, ???? 1951.

Bunge:1951:MTA

[Bun51d] Mario Augusto Bunge. Mach y la teoría atómica. (Spanish) [Mach and atomic theory]. Boletin del Químico Peruano, 3(16):12–17, ???? 1951.

Bunge:1952:IAH

[Bun52] Mario Augusto Bunge. Interpretación del átomo de hidrógeno en la teoría de Dirac. (Spanish) [Interpretation of the hydrogen atom in the Dirac theory]. Doctor en ciencias físicas, Facultad de Ciencias Exactas y Naturales, Universidad de Buenos Aires, Buenos Aires, Argentina, 1952. 121 pp. URL http://catalogosuba.

sisbi.uba.ar/vufind/Record/20160331012339142.

Bunge:1953:RRR

[Bun53] Mario Bunge. Review: Relativity — a Richer Truth, by Philipp Frank. Boston: The Beacon Press, 1950. $2.00. Pp. xvi + 142; Science and Humanism. Physics in Our Time, by Erwin Schrödinger, London and New York: Cambridge University Press, 1951. $1.75. Pp. 68. Science & Society, 17(3):266–269, Summer 1953. CODEN ???? ISSN 0036-8237 (print), 1943-2801

(electronic). URL http://www.jstor.org/stable/40400200.

Bunge:1954:NDB

[Bun54] Mario Bunge. New dialogues between Hylas and Philonous. Philosophy and Phenomenological Research, 15(2):192–199, December 1954. CODEN PPHRAI. ISSN 0031-8205 (print), 1933-1592

(electronic). URL https://www.jstor.org/stable/2103573.

Bunge:1955:PE

[Bun55a] M. Bunge. A picture of the electron. Il Nuovo Cimento (10), 1(6):977–985, June 1955. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1955NCimS...1..977B.

Bunge:1955:CPC

[Bun55b] Mario Bunge. Causality: the Place of the Causal Principle in Modern Science. Dean Funes, ????, 1955. xx + 380 pp. LCCN BD541 .B85 1955.

Bunge:1955:EDU

[Bun55c] Mario Bunge. La Edad del Universo. (Spanish) [The Age of the Universe]. Editorial U.M.S.A., La Paz, Bolivia, 1955. 94 pp.

LCCN QB981 .B83. URL https://ui.adsabs.harvard.edu/#
abs/1955edun.book.....B.

Bunge:1955:PST

[Bun55d] Mario Bunge. The philosophy of the space–time approach to the quantum theory. Methodos, 7(??):295–308, 1955.

Bunge:1955:SACa

[Bun55e] Mario Bunge. Strife about complementarity (I). British Journal for the Philosophy of Science, VI(21):1–12, May 1955. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://bjps.oxfordjournals.org/content/
VI/21/1.full.pdf+html;   http://www.jstor.org/stable/
685570.

Bunge:1955:SACb

[Bun55f] Mario Bunge. Strife about complementarity (II). British Journal for the Philosophy of Science, VI(22):141–154, August 1955. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/VI/22/141.
full.pdf+html;   http://www.jstor.org/stable/685522.

Bunge:1955:ECP

[Bun55g] Mario Augusto Bunge. Exposición y crítica del principio de complementaridad. (Spanish) [Critical exposition of the principle of complementarity]. In Notas del Curso Iteramericano de Física Moderna, pages 27–36. Laboratorio de Física Cósmica, La Paz, 1955.

Bunge:1956:CFT

[Bun56a] Mario Bunge. A critique of the frequentist theory of probability. In ????, editor, Congresso International de Filosofia, vol. III, page ?? Sociedade Brasileira de Filosofia, São Paulo, Brazil, 1956.

Bunge:1956:DCTa

[Bun56b] Mario Bunge. Do computers think? (I). British Journal for the Philosophy of Science, VII(26):139–148, August 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/VII/26/139.full.pdf+html;   http://www.jstor.
org/stable/685697.

Bunge:1956:DCTb

[Bun56c] Mario Bunge. Do computers think? (II). British Journal for the Philosophy of Science, VII(27):212–219, November 1956. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/VII/27/212.full.pdf+html;   http://www.jstor.
org/stable/685873.

Bunge:1956:WMS

[Bun56d] Mario Bunge. On William Morris’ socialism. Science & Society, 20(2):142–144, Spring 1956. ISSN 0036-8237 (print), 1943-2801

(electronic). URL https://www.jstor.org/stable/40400398.

Bunge:1956:SIQ

[Bun56e] Mario Bunge. Survey of the interpretations of quantum mechanics. American Journal of Physics, 24(4):272–286, April 1956. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL https://ui.adsabs.harvard.edu/#abs/1956AmJPh..24.
.272B.

Bunge:1956:ICM

[Bun56f] Mario Augusto Bunge. La interpretación causal de la mecánica ondulatoria. (Spanish) [The causal interpretation of wave mechanics]. Ciencia e Investigación, 12(??):448–457, ???? 1956. CODEN CIBAAH. ISSN 0009-6733 (print), 2314-3134 (electronic).

Bunge:1957:LFM

[Bun57a] Mario Bunge. Lagrangian formulation and mechanical interpretation. American Journal of Physics, 25(4):211–218, April 1957. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1957AmJPh..25..211B.

Bunge:1957:FCE

[Bun57b] Mario Augusto Bunge. Filosofar científicamente y encarar la ciencia filosóficamente. (Spanish) [Philosophize scientifically and tackle science philosophically]. Ciencia e Investigación, 13(??): 244–257, ???? 1957. CODEN CIBAAH. ISSN 0009-6733 (print), 2314-3134 (electronic).

Bunge:1958:BRC

[Bun58a] Mario Bunge. Book review: Concepts of Space: The History of Theories of Space in Physics by Max Jammer. British Journal for the Philosophy of Science, 9(33):77–79, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL

https://www.jstor.org/stable/686111.

Bunge:1958:BRR

[Bun58b] Mario Bunge. Book review: Reflections of a Physicist by P. W. Bridgman. British Journal for the Philosophy of Science, 9(33): 74–77, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://www.jstor.org/stable/
686110.

Bunge:1958:MDT

[Bun58c] Mario Bunge. On multi-dimensional time. British Journal for the Philosophy of Science, IX(33):39, May 1958. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/IX/33/39.
full.pdf+html;   http://www.jstor.org/stable/686099.

Bunge:1958:Ra

[Bun58d] Mario Bunge. Reviews. British Journal for the Philosophy of Science, IX(33):74–77, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/IX/33/74.full.pdf+html;
http://www.jstor.org/stable/686110.

Bunge:1958:Rb

[Bun58e] Mario Bunge. Reviews. British Journal for the Philosophy of Science, IX(33):77–79, May 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/IX/33/77.full.pdf+html;
http://www.jstor.org/stable/686111.

Bunge:1958:QCS

[Bun58f] Mario Augusto Bunge. ‘?Qué es la ciencia?. (Spanish) [What is science?]. Universidad de Buenos Aires, Buenos Aires, Argentina, 1958. ???? pp.

Bunge:1958:IFP

[Bun58g] Mario Augusto Bunge. Sobre la imagen física de la partículas de spin entero. (Spanish) [On the physical image of integral spin particles]. Ciencia e Investigación, 14(??):311–315, ???? 1958. CO-DEN CIBAAH. ISSN 0009-6733 (print), 2314-3134 (electronic).

Bunge:1959:CPC

[Bun59a] Mario Bunge. Causality: the Place of the Causal Principle in Modern Science. Harvard University Press, Cambridge, MA, USA, 1959. xx + 380 pp. LCCN BD541 .B85 1959.

Bunge:1959:MQ

[Bun59b] Mario Bunge. Metascientific Queries, volume 341 of American Lecture Series. Charles C. Thomas Publisher, Springfield, IL, 1959. ISSN 0099-5371. xiv + 313 pp. LCCN Q175 .B825.

Bunge:1959:RPL

[Bun59c] Mario Augusto Bunge. Review of Popper’s The logic of scientific discovery. Ciencia e Investigación, 15(??):216, ???? 1959. CO-DEN CIBAAH. ISSN 0009-6733 (print), 2314-3134 (electronic).

Bunge:1960:TTE

[Bun60a] Mario Bunge. Are there timeless entities? Miscelanea de Estudos a Joaquim de Carvalho. Figueira da Foz (Portugal), 3(??):290– 292, ???? 1960.

Bunge:1960:BRM

[Bun60b] Mario Bunge. Book review: Modern Philosophy of Science: Selected Essays by Hans Reichenbach, Maria Reichenbach. British Journal for the Philosophy of Science, 11(42):153–156, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://www.jstor.org/stable/685591.

Bunge:1960:EC

[Bun60c] Mario Bunge. Etica y ciencia. (Spanish) [Ethics and science], volume 35 of Colección Panorama. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1960. 77 pp. LCCN BJ1411 .B8.

Bunge:1960:CER

[Bun60d] Mario Bunge. La cinemática del electron relativista. (Spanish) [Relativistic electron kinematics]. Facultad dee Ciencias Exactas y Tecnología, Universidad Nacional de Tucumán, San Miguel de Tucumán, Argentina, 1960. 254 pp.

URL http://catalogosuba.sisbi.uba.ar/vufind/Record/
201603080902035037.

Bunge:1960:LSP

[Bun60e] Mario Bunge. Levels: a semantical preliminary. Review of Metaphysics, 13(3):396–406, March 1960. CODEN ???? ISSN 1549-4853 (print), 0034-6632 (electronic). URL http://www.jstor.

org/stable/20123770.

Bunge:1960:CAL

[Bun60f] Mario Bunge. On the connections among levels. In ????, editor, Proceedings of the XIIth International Congress of Philosophy VI, pages 63–70. Sansoni, Firenze, Italy, 1960.

Bunge:1960:PISb

[Bun60g] Mario Bunge. The place of induction in science. Philosophy of Science, 27(3):262–270, July 1960. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185969.

Bunge:1960:R

[Bun60h] Mario Bunge. Reviews. British Journal for the Philosophy of Science, XI(42):153–156, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/XI/42/153.
full.pdf+html;   http://www.jstor.org/stable/685591.

Bunge:1960:CMF

[Bun60i] Mario Augusto Bunge. La ciencia: su método y su filosofía. (Spanish) [Science: its method and its philosophy]. Siglo Veinte, Buenos Aires, Argentina, 1960. ???? pp.

Bunge:1960:PISa

[Bun60j] Mario Augusto Bunge. Probabilidad e inducción. (Spanish) [Probability and induction]. Ciencia y Técnica (Buenos Aires), 129(??):240, ???? 1960.

Bunge:1961:AR

[Bun61a] Mario Bunge. Analyticity redefined. MIND: A Quarterly Review of Pyschology and Philosophy, 70(278):239–245, April 1961. CO-DEN MNDDAB. ISSN 0026-4423 (print), 1460-2113 (electronic).

URL http://www.jstor.org/stable/2251754.

Bunge:1961:AUD

[Bun61b] Mario Bunge. The Argentine university: a defense. AAUP Bulletin, 47(1):53–55, March 1961. ISSN 0277-3945 (print),

1930-322X (electronic). URL http://www.jstor.org/stable/
40222624.

Bunge:1961:CCL

[Bun61c] Mario Bunge. Causality, chance, and law. American Scientist, 49(4):432–448, December 1961. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.

org/stable/27827906.

Bunge:1961:CS

[Bun61d] Mario Bunge. The complexity of simplicity. The Journal of Philosophy, LIX(5):113–135, March 1, 1961. ISSN 0022-362X (print),

1939-8549 (electronic). URL http://www.jstor.org/stable/
2023305.

Bunge:1961:ES

[Bun61e] Mario Bunge. Ethics as a science. Philosophy and Phenomenological Research, XXII(2):139–152, December 1961. CODEN PPHRAI. ISSN 0031-8205 (print), 1933-1592 (electronic). URL

http://www.jstor.org/stable/2104835.

Bunge:1961:KCS

[Bun61f] Mario Bunge. Kinds and criteria of scientific laws. Philosophy of Science, 28(3):260–281, July 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185111.

Bunge:1961:LPL

[Bun61g] Mario Bunge. Laws of physical laws. American Journal of Physics, 29(8):518–529, August 1961. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1961AmJPh..29..518B.

Bunge:1961:WSC

[Bun61h] Mario Bunge. The weight of simplicity in the construction and assaying of scientific theories. Philosophy of Science, 28(2):120– 149, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/
185804.

Bunge:1961:CPC

[Bun61i] Mario Augusto Bunge. Causalidad: el principio de causalidad en la ciencia moderna. (Spanish) [Causality: the principle of causality in modern science]. Temas de Eudeba. Filosofia. Eudeba, Buenos Aires, Argentina, 1961. 403 pp.

Bunge:1962:AV

[Bun62a] Mario Bunge. An analysis of value. Mathematicae Notae, 18(??): 95–108, ???? 1962.

Bunge:1962:BRI

[Bun62b] Mario Bunge. Book review: Introduction to Semantics by Adam Schaff, Olgierd Wojtasiewicz. The Journal of Philosophy, 59(25): 805–808, December 1962. ISSN 0022-362X (print), 1939-8549

(electronic). URL http://www.jstor.org/stable/2023140.

Bunge:1962:CR

[Bun62c] Mario Bunge. Causality: a rejoinder. Philosophy of Science, 29(3):306–317, July 1962. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186050.

Bunge:1962:CS

[Bun62d] Mario Bunge. The complexity of simplicity. The Journal of Philosophy, 59(5):113–135, March 1962. ISSN 0022-362X (print),

1939-8549 (electronic). URL http://www.jstor.org/stable/
2023305.

Bunge:1962:CM

[Bun62e] Mario Bunge. Cosmology and magic. The Monist, 47(1):116–141, Fall 1962. ISSN 0026-9662 (print), 2153-3601 (electronic). URL

http://www.jstor.org/stable/27901496.

Bunge:1962:IS

[Bun62f] Mario Bunge. Intuition and Science. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1962. x + 142 pp.

Bunge:1962:PMP

[Bun62g] Mario Bunge. Prichinnost’: mesto printsipa prichinnosti v sovremennoi nauke. (Russian) [Causality: the place of the principle of causality in modern science]. Izdatelstvo inostranno, Moscow, USSR, 1962. 511 pp.

Bunge:1963:BRT

[Bun63a] Mario Bunge. Book review: Time and the Physical World by Richard Schlegel. British Journal for the Philosophy of Science, 14(53):65–69, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.org/

stable/685355.

Bunge:1963:CPC

[Bun63b] Mario Bunge. Causality: the Place of the Causal Principle in Modern Science. World Publishing Company [Meridian Books], Cleveland, OH, USA, second edition, 1963. xx + 394 pp. LCCN BD541 .B85 1959.

Bunge:1963:GBB

[Bun63c] Mario Bunge. A general black box theory. Philosophy of Science, 30(4):346–358, October 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186066.

Bunge:1963:MSP

[Bun63d] Mario Bunge. The Myth of Simplicity: Problems of Scientific Philosophy. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1963. xii + 239 pp. LCCN BD241 .B85.

Bunge:1963:R

[Bun63e] Mario Bunge. Reviews. British Journal for the Philosophy of Science, XIV(53):65–69, May 1963. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/XIV/53/65.full.pdf+html;
http://www.jstor.org/stable/685355.

Bunge:1963:TCF

[Bun63f] Mario Augusto Bunge. Tecnología, ciencia y filosofía. (Spanish) [Technology, science and philosophy]. Revista de la Universidad de Chile, 121(126):64–92, ???? 1963.

Bunge:1964:CAS

[Bun64a] Mario Bunge, editor. The Critical Approach to Science and Philosophy: Essays in Honour of Karl Popper. Free Press of Glencoe, New York, NY, USA, 1964. xv + 480 pp. LCCN B29 .B87.

Bunge:1964:PT

[Bun64b] Mario Bunge. Phenomenological theories. In The Critical Approach to Science and Philosophy: Essays in Honour of Karl Popper [Bun64a], pages 234–254. LCCN B29 .B87.

Bunge:1964:R

[Bun64c] Mario Bunge. Reviews. British Journal for the Philosophy of Science, XV(58):168–170, August 1964. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/XV/58/168.
full.pdf+html;   http://www.jstor.org/stable/685262.

Bunge:1964:SP

[Bun64d] Mario Bunge. Stating the problem. Science, 143(3613): 1394, March 27, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1713065;   https://ui.adsabs.harvard.edu/#abs/
1964Sci...143.1394B.

Bunge:1965:PR

[Bun65a] Mario Bunge. Physics and reality. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):195–222, September/December 1965. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL http://www.jstor.org/stable/
42964458.

Bunge:1965:CPC

[Bun65b] Mario Augusto Bunge. Causalidad : el principio de causalidad en la ciencia moderna. (Spanish) [Causality: the principle of causality in modern science]. Temas de Eudeba. Filosofía. Eudeba, Buenos Aires, Argentina, second edition, 1965. 402 pp. Spanish translation of [Bun63b].

Bunge:1965:ICS

[Bun65c] Mario Augusto Bunge. Intuición y ciencia. (Spanish) [Intuition and science]. Colección ensayos. Eudeba, Buenos Aires, Argentina, 1965. 174 pp.

Bunge:1966:BRF

[Bun66a] Mario Bunge. Book review: I filosofi e le macchine (1400–1700) by Paolo Rossi. Technology and Culture, 7(1):73–74, Winter 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL http://www.jstor.org/stable/3101607;
https://muse.jhu.edu/pub/1/article/893858/pdf.

Bunge:1966:MCN

[Bun66b] Mario Bunge. Mach’s critique of Newtonian mechanics. American Journal of Physics, 34(7):585–596, July 1966. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL https:

//ui.adsabs.harvard.edu/#abs/1966AmJPh..34..585B. See

comment [Erl67]. Reprinted in [Bla92, pages 243–261].

Bunge:1966:NI

[Bun66c] Mario Bunge. On null individuals. The Journal of Philosophy, 63(24):776–778, December 1966. ISSN 0022-362X (print),

1939-8549 (electronic). URL http://www.jstor.org/stable/
2023807.

Bunge:1966:PR

[Bun66d] Mario Bunge. Physics and reality. Dialectica: International Review of Philosophy of Knowledge, 20(2):174–195, June 1966. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL

http://www.jstor.org/stable/42964486.

Bunge:1966:TAS

[Bun66e] Mario Bunge. Technology as applied science. Technology and Culture, 7(3):329–347, Summer 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL http://www.

jstor.org/stable/3101932.

Bunge:1966:WSC

[Bun66f] Mario Bunge. The weight of simplicity in the construction and assaying of scientific theories. In Foster and Martin [FM66], page ??

Bunge:1967:DSF

[Bun67a] M. Bunge, editor. Delaware Seminar in the Foundations of Physics, volume 1 of Studies in the Foundations, Methodology and Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. x + 193 pp. LCCN QC71 .D385.

Bunge:1967:IS

[Bun67b] M. Bunge. Intuition and science. Verlag ‘Progress’, Moscow, USSR, 1967. 188 (est.) pp.

Bunge:1967:SRS

[Bun67c] M. Bunge. Scientific Research. 1: The Search for System. 2: The Search for Truth, volume 3/1–3/2 of Studies in the Foundations, Methodology and Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. xii + 536 pp. LCCN Q180.A1 B77.

Bunge:1967:FP

[Bun67d] Mario Bunge. Foundations of Physics, volume 10 of Springer tracts in natural philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-642-49289-4 (print), 3-642-49287-8 (e-book). ISSN 0081-3877. xii + 311 pp. LCCN QC21 .B95.

Bunge:1967:GFA

[Bun67e] Mario Bunge. A ghost free axiomatization of quantum mechanics. In Quantum Theory and Reality [Bun67i], pages 105–117.

ISBN 3-642-88028-2 (print), 3-642-88026-6 (e-book). ISSN 0081-

8577. LCCN QC1-75. URL http://public.eblib.com/choice/
publicfullrecord.aspx?p=3098628.

Bunge:1967:PA

[Bun67f] Mario Bunge. Physical axiomatics. Reviews of Modern Physics, 39(2):463–474, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.463;   http://
rmp.aps.org/abstract/RMP/v39/i2/p463_1;   https://ui.
adsabs.harvard.edu/#abs/1967RvMP...39..463B.

Bunge:1967:QF

[Bun67g] Mario Bunge. Quanta y filosofia. Crítica: Revista Hispanoamericana de Filosofía, 1(3):41–64, September 1967. CO-DEN ???? ISSN 0011-1503. URL http://www.jstor.org/

stable/40103877.

Bunge:1967:QFS

[Bun67h] Mario Bunge. Quanta y filosofía. (Spanish) [Quanta and philosophy]. Crítica: Revista Hispanoamericana de Filosofía, 1(3): 41–64, September 1967. CODEN ???? ISSN 0011-1503. URL

http://www.jstor.org/stable/40103877.

Bunge:1967:QTR

[Bun67i] Mario Bunge, editor. Quantum Theory and Reality, volume 2 of Studies in the Foundations, Methodology and Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-642-88028-2 (print), 3-642-88026-6 (e-book). ISSN 0081-8577. iv + 117

pp. LCCN QC1-75. URL http://public.eblib.com/choice/
publicfullrecord.aspx?p=3098628.

Bunge:1967:SCP

[Bun67j] Mario Bunge. The structure and content of a physical theory. In Bunge [Bun67a], pages 15–27. LCCN QC71 .D385.

Bunge:1967:TT

[Bun67k] Mario Bunge. The turn of the tide. In Quantum Theory and Reality [Bun67i], pages 1–6. ISBN 3-642-88028-2 (print), 3-642-88026-6 (e-book). ISSN 0081-8577. LCCN QC1-75.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=3098628.

Bunge:1968:AQT

[Bun68a] Mario Bunge. Analogy in quantum theory: From insight to nonsense. British Journal for the Philosophy of Science, 18 (4):265–286, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.

oxfordjournals.org/content/18/4/265.full.pdf+html;
http://www.jstor.org/stable/686357.

Bunge:1968:BRP

[Bun68b] Mario Bunge. Book review: The Psychology of Communication by George A. Miller. British Journal for the Philosophy of Science, 18(4):350–352, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/686370.

Bunge:1968:CSD

[Bun68c] Mario Bunge. Conjunction, succession, determination and causation. International Journal of Theoretical Physics, 1(3):299–315, October 1968. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-

9575 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1968IJTP....1..299B.

Bunge:1968:VTS

[Bun68d] Mario Bunge. La vérification des théories scientifiques. (French) [Verification of scientific theories]. Logique et Analyse, 11 (41–42):145–179, 1968. ISSN 0024-5836 (print), 2295-5836

(electronic). URL http://virthost.vub.ac.be/lnaweb/ojs/
index.php/LogiqueEtAnalyse/article/view/299.

Bunge:1968:CMFb

[Bun68e] Mario Bunge. Les concepts de modèle. (French) [Model concepts]. L’âge de la science, I(??):165–180, 1968.

Bunge:1968:MS

[Bun68f] Mario Bunge. The maturation of science. In Lakatos and Musgrave [LM68], pages 120–147. ISBN 0-444-53413-X. ISSN 0049-237X. LCCN Q175 .I518 1965. URL http://www.

sciencedirect.com/science/article/pii/S0049237X08705007.

Bunge:1968:MTS

[Bun68g] Mario Bunge. Models in theoretical science. In ????, editor, Proceedings of the XIVth International Congress of Philosophy = Actes du XIVème Congrès International de Philosophie: Vienne, du 2 au 9 septembre 1968, pages 208–217. Herder, Vienna, Austria, 1968.

Bunge:1968:NS

[Bun68h] Mario Bunge. The nature of science. In Klibansky [Kli68], pages 3–15.

Bunge:1968:MNM

[Bun68i] Mario Bunge. On Mach’s nonconcept of mass — a reply to M. Mueller. American Journal of Physics, 36(2):167–168, February 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1968AmJPh..36..167B. See [Mue68].

Bunge:1968:PP

[Bun68j] Mario Bunge. Philosophy and physics. In Klibansky [Kli68], pages 167–199.

Bunge:1968:PTO

[Bun68k] Mario Bunge. Physical time: The objective and relational theory. Philosophy of Science, 35(4):355–388, December 1968. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186253. See note [Rus72].

Bunge:1968:PMD

[Bun68l] Mario Bunge. Physique et métaphysique du temps. (French) [Physics and metaphysics of time]. In ????, editor, Proceedings of the XIVth International Congress of Philosophy = Actes du XIVème Congrès International de Philosophie: Vienne, du 2 au 9 septembre 1968, pages 623–629. Herder, Vienna, Austria, 1968.

Bunge:1968:PGC

[Bun68m] Mario Bunge. Problems and games in the current philosophy of science. In ????, editor, Proceedings of the XIVth International Congress of Philosophy = Actes du XIVème Congrès International de Philosophie: Vienne, du 2 au 9 septembre 1968, volume I, pages 623–629. Herder, Vienna, Austria, 1968.

Bunge:1968:R

[Bun68n] Mario Bunge. Reviews. British Journal for the Philosophy of Science, 18(4):350–352, February 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/18/4/350.
full.pdf+html;   http://www.jstor.org/stable/686370.

Bunge:1968:SLR

[Bun68o] Mario Bunge. Scientific laws and rules. In Klibansky [Kli68], pages 128–140.

Bunge:1968:TPT

[Bun68p] Mario Bunge. Theory of partial truth: Not proved inconsistent. Philosophy and Phenomenological Research, 29(2):297–298, December 1968. CODEN PPHRAI. ISSN 0031-8205 (print),

1933-1592 (electronic). URL http://www.jstor.org/stable/
2105863.

Bunge:1968:CMFa

[Bun68q] Mario Augusto Bunge. La ciencia : su método y su filosofía. (Spanish) [Science: its method and its philosophy]. Editorial Siglo Veinte, Buenos Aires, Argentina, 1968. 159 pp.

Bunge:1969:AHP

[Bun69a] Mario Bunge. Alexander von Humboldt und die Philosophie. (German) [Alexander von Humboldt and philosophy]. In Pfeiffer [Pfe69], pages 17–30. LCCN Q143.H9 A68.

Bunge:1969:ASR

[Bun69b] Mario Bunge. Analogy, simulation, representation. Revue Internationale de Philosophie, 23(87):16–33, ???? 1969. URL http:/

/www.jstor.org/stable/23940647.

Bunge:1969:CFP

[Bun69c] Mario Bunge. Corrections to Foundations of Physics: Correct and incorrect. Synthese, 19(3–4):443–452, April 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485661;
 http://link.springer.com/content/pdf/10.1007/BF00485661.
pdf;   http://www.jstor.org/stable/20114653.

Bunge:1969:FMH

[Bun69d] Mario Bunge. Four models of human migration: an exercise in mathematical sociology. ARSP: Archiv für Rechts- und Sozialphilosophie / Archives for Philosophy of Law and Social Philosophy, 55(4):451–466, ???? 1969. ISSN 0001-2343. URL http://

www.jstor.org/stable/23678432.

Bunge:1969:MEM

[Bun69e] Mario Bunge. The metaphysics, epistemology and methodology of levels. In L. L. Whyte, A. G. Wilson, and D. Wilson, editors, Hierarchical Levels, pages 17–28. American Elsevier, New York, NY, USA, 1969.

Bunge:1969:PAD

[Bun69f] Mario Bunge. The paradox of addition and its dissolution. Crítica: Revista Hispanoamericana de Filosofía, 3(9):27–31, September 1969. CODEN ???? ISSN 0011-1503. URL http:

//www.jstor.org/stable/40103932.

Bunge:1969:WPT

[Bun69g] Mario Bunge. What are physical theories about? In N. Rescher, editor, Studies in the Philosophy of Science, volume 3 of American Philosophical Quarterly Monograph, pages 61–99. Blackwell, Oxford, UK, 1969.

Bunge:1969:ICE

[Bun69h] Mario Augusto Bunge. La investigación científica : su estrategia y su filosofía. (Spanish) [Scientific research: its strategy and its philosophy]. Ariel, Barcelona, Spain, 1969. 955 pp.

Bunge:1970:AT

[Bun70a] Mario Bunge. The arrow of time. International Journal of Theoretical Physics, 3(1):77–79, January 1970. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/1970IJTP....3...77B.

Bunge:1970:CPC

[Bun70b] Mario Bunge. Causality: the Place of the Causal Principle in Modern Science. Harvard University Press, Cambridge, MA, USA, second [??check??] edition, 1970. xx + 394 pp. LCCN BD541 .B85 1959.

Bunge:1970:CPD

[Bun70c] Mario Bunge. La causalità: il posto del principio causale nella scienza moderna. (Italian) [Causality: the place of the causal principle in modern science], volume 47–48 of Universale Scientifica Boringhieri. P. Boringhieri, Torino, Italy, 1970. 486 pp.

Bunge:1970:PP

[Bun70d] Mario Bunge. The physicist and philosophy. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):196–208, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170181.

Bunge:1970:CFI

[Bun70e] Mario Bunge. The so-called fourth indeterminacy relation. Canadian Journal of Physics = Journal canadien de physique, 48 (11):1410–1411, June 1970. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1970CaJPh..48.1410B.

Bunge:1970:TME

[Bun70f] Mario Bunge. Theory meets experience. In Kiefer and Munitz [KM70], pages 138–165. ISBN 0-87395-052-6. LCCN B29

.M556 1970. URL http://catalogue.bnf.fr/ark:/12148/
cb353806733.

Bunge:1970:TAT

[Bun70g] Mario Bunge. Time asymmetry, time reversal and irreversability. Studium Generale, 23(??):562–570, ???? 1970.

Bunge:1970:VPV

[Bun70h] Mario Bunge. Virtual processes and virtual particles: Real or fictitious? International Journal of Theoretical Physics, 3:507–508, December 1970. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/1970IJTP....3..507B.

Bunge:1970:CPP

[Bun70i] Mario Augusto Bunge. La causalità: il posto del principio causale nella scienza moderna. (Italian) [Causality: the place of the causal principle in modern science.], volume 47–48 of Universale scientifica. Boringhieri, Torino, Italy, 1970. 486 pp.

Bunge:1970:PCI

[Bun70j] Mario Augusto Bunge. Problems concerning inter-theory relations. In Weingartner and Zecha [WZ70], pages 285–315. ISBN 90-277-0158-X. LCCN Q175 .S235 1968.

Bunge:1971:BRT

[Bun71a] Mario Bunge. Book review: Der Teil Und Das Ganze, by Werner Heisenberg, 334 pp. R. Piper Verlag, München, 1969. [Translated as Physics and Beyond: Encounters and Conversations, 272 pp. Harper and Row, New York. 1971]. Physics Today, 24(1):63–64, January 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v24/i1/p63/s1.

Bunge:1971:SMA

[Bun71b] Mario Bunge. “Scientific Metaphysics”: Addenda et corrigenda. The Journal of Philosophy, 68(23):876, December 1971. ISSN 0022-362X (print), 1939-8549 (electronic). URL http://www.

jstor.org/stable/2024982.

Bunge:1971:CSD

[Bun71c] Mario Bunge. Conjonction, succession, détermination, causalité. (French) [Conjunction, succession, determination, causality]. In Les théories de la causalité, pages 112–132. Presses universitaires de France, Paris, France, 1971.

Bunge:1971:SMP

[Bun71d] Mario Bunge. Is scientific metaphysics possible? The Journal of Philosophy, 68(17):507–520, September 1971. ISSN 0022-362X (print), 1939-8549 (electronic). URL http://www.jstor.org/

stable/2024691.

Bunge:1971:NLD

[Bun71e] Mario Bunge. A new look at definite descriptions. Philosophy of Science (Japan), 4(??):131–146, ???? 1971.

Bunge:1971:MPS

[Bun71f] Mario Bunge. On method in the philosophy of science. Archives de Philosophie, 34(4):551–574, October/December 1971. CO-DEN ???? ISSN 0003-9632 (print), 1769-681X (electronic). URL

http://www.jstor.org/stable/43033380.

Bunge:1971:POR

[Bun71g] Mario Bunge. A philosophical obstacle to the rise of new theories in microphysics. In Bastin [Bas71], page ?? ISBN 0-521-07956-X. LCCN QC174.1 .Q37. URL http://www.loc.gov/

catdir/enhancements/fy0907/77127237-d.html;   http://
www.loc.gov/catdir/enhancements/fy0907/77127237-t.html.

Bunge:1971:PFP

[Bun71h] Mario Bunge, editor. Problems in the Foundations of Physics, volume 4 of Studies in the Foundations, Methodology, and Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05490-1 (New York), 3-642-80626-0 (print), 3-642-80624-4 (e-book). 162 (est.)

pp. LCCN QC20 .B84. URL https://ui.adsabs.harvard.edu/
#abs/1971pfp..book.....B.

Bunge:1972:BRB

[Bun72a] Mario Bunge. Book review: Mach’s Philosophy of Science by J. Bradley. Philosophy of Science, 39(2):266–267, June 1972. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186731.

Bunge:1972:BRM

[Bun72b] Mario Bunge. Book review: Mach’s Philosophy of Science by J. Bradley. Philosophy of Science, 39(2):266–267, June 1972. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186731.

Bunge:1972:F

[Bun72c] Mario Bunge. Foreword. Journal of Philosophical Logic, 1(3– 4):263–265, August 1972. CODEN JPLGA7. ISSN 0022-3611 (print), 1573-0433 (electronic). URL http://www.jstor.org/

stable/30226042.

Bunge:1972:I

[Bun72d] Mario Bunge. Ingasei. (Japanese) [Causality]. ????, ????, Japan, 1972. 401 pp. LCCN BD541 .B8516.

Bunge:1972:PSS

[Bun72e] Mario Bunge. A program for the semantics of science. Journal of Philosophical Logic, 1(3–4):317–328, August 1972. CODEN JPLGA7. ISSN 0022-3611 (print), 1573-0433 (electronic). URL

http://www.jstor.org/stable/30226046.

Bunge:1972:STS

[Bun72f] Mario Bunge. Scientific thought. Some underlying concepts, methods and procedures, volume 9 of New Babylon. Mouton, The Hague, The Netherlands, 1972. ???? pp. LCCN Q160.2 .S35.

Bunge:1972:TR

[Bun72g] Mario Bunge. Teoría y realidad. (Spanish) [Theory and reality]. Editorial Ariel, Barcelona, Spain, 1972. ???? pp.

Bunge:1972:CPC

[Bun72h] Mario Augusto Bunge. Causalidad : el principio de causalidad en la ciencia moderna. (Spanish) [Causality: the principle of causality in modern science]. Eudeba, Buenos Aires, Argentina, fourth edition, 1972. 403 pp.

Bunge:1973:BRR

[Bun73a] Mario Bunge. Bertrand Russell’s Regulae Philosophandi. In The Methodological Unity of Science [Bun73h], pages 3–12. ISBN 90-277-0354-X, 90-277-0404-X (paperback), 94-010-2667-

X. LCCN Q174 .B47 1973. URL http://www.springerlink.
com/content/978-94-010-2667-3.

Bunge:1973:BRP

[Bun73b] Mario Bunge. Book review: A Probabilistic Theory of Causality by P. Suppes. British Journal for the Philosophy of Science, 24 (4):409–410, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.org/

stable/686623.

Bunge:1973:DTM

[Bun73c] Mario Bunge. A decision theoretic model of the American war in Vietnam. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 3(4):323–338, June 1973. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Bunge:1973:TPM

[Bun73d] Mario Bunge. ‘?Es posible una metafísica científica? Teorema: Revista Internacional de Filosofía, 3(4):435–454, 609–610, ???? 1973. CODEN ???? ISSN 0210-1602. URL http://www.jstor.

org/stable/43047219.

Bunge:1973:EPP

[Bun73e] Mario Bunge, editor. Exact Philosophy. Problems, Tools, and Goals, Synthese library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. ISBN 90-277-0251-9. LCCN B105.E75 S95 1971.

Bunge:1973:CSM

[Bun73f] Mario Bunge. La ciencia, su método y su filosofía. (Spanish) [Science, its methods and its philosophy]. Editorial Siglo Veinte, Buenos Aires, Argentina, 1973. ???? pp.

Bunge:1973:MMM

[Bun73g] Mario Bunge. Method, Model, and Matter. Synthese library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0252-7 (paperback). vi + 198 pp. LCCN Q175 .B826.

Bunge:1973:MUS

[Bun73h] Mario Bunge, editor. The Methodological Unity of Science, volume 3 of Theory and decision library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0354-X, 90-277-0404-X (paperback), 94-010-2667-X. LCCN Q174 .B47 1973. URL http://www.

springerlink.com/content/978-94-010-2667-3.

Bunge:1973:NWN

[Bun73i] Mario Bunge. Normative Wissenschaft ohne Normen—aber mit Werten. (German) [Normative science without norms — but with values]. Conceptus, 7(21-22):57–64, 1973. ISSN 0010-5155 (print), 2196-9523 (electronic).

Bunge:1973:CMMb

[Bun73j] Mario Bunge. On confusing ‘measure’ with ‘measurement’ in the methodology of behavioral science. In The Methodological Unity of Science [Bun73h], pages 105–122. ISBN 90-277-0354-X, 90-277-0404-X (paperback), 94-010-2667-X. LCCN Q174 .B47

1973. URL http://www.springerlink.com/content/978-94-
010-2667-3.

Bunge:1973:CMMa

[Bun73k] Mario Bunge. On confusing ‘measurement’ with ‘measure’ in the methodology of the behavioral sciences. In Exact Philosophy. Problems, Tools, and Goals [Bun73e], pages 105–122. ISBN 90-277-0251-9. LCCN B105.E75 S95 1971.

Bunge:1973:PPa

[Bun73l] Mario Bunge. Philosophy of Physics. Synthese library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0253-5. vii + 248 pp. LCCN QC6 .B87 1973.

Bunge:1973:PPb

[Bun73m] Mario Bunge. Philosophy of Physics, volume 45 of Synthese Library, Monographs on Epistemology, Logic, Methodology, Philosophy of Science, Sociology of Science and of Knowledge, and on the Mathematical Methods of Social and Behavioral Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. ISBN 94-010-2522-3 (e-book), 94-010-2524-

X. ix + 251 pp. LCCN ???? URL http://public.eblib.com/
choice/PublicFullRecord.aspx?p=3109392.

Bunge:1973:PSS

[Bun73n] Mario Bunge. A program for the semantics of science. In Exact Philosophy. Problems, Tools, and Goals [Bun73e], pages 51–62. ISBN 90-277-0251-9. LCCN B105.E75 S95 1971.

Bunge:1973:R

[Bun73o] Mario Bunge. Reviews. British Journal for the Philosophy of Science, 24(4):409–410, December 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL http://bjps.oxfordjournals.org/content/24/4/409.
full.pdf+html;   http://www.jstor.org/stable/686623.

Bunge:1973:RFP

[Bun73p] Mario Bunge. The role of forecast in planning. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 3(3):207–221, March 1973. CODEN THD-CBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Bunge:1974:BRP

[Bun74a] Mario Bunge. Book review: Philosophy and mathematics: From Plato to the present: By Robert J. Baum. San Francisco, California (Freeman, Cooper & Co.), 1973. 332 pp. Historia Mathematica, 1(3):350, August 1974. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
031508607490086X.

Bunge:1974:TPJ

[Bun74b] Mario Bunge. Theory and Practice by Jurgen Habermas (review). Technology and Culture, 15(3):539, July 1974. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/892773/pdf.

Bunge:1974:CSS

[Bun74c] Mario Bunge. The concept of social structure. In Leinfellner and Köhler [LK74], pages 175–215. ISBN 90-277-0493-7 (hardcover), 90-277-0539-9 (paperback). LCCN H61 .D45. URL http://

swbplus.bsz-bw.de/bsz011135883cov.htm.

Bunge:1974:PPM

[Bun74d] Mario Bunge. Les présupposés et les produits métaphysiques de la science et de la technique contemporaines. (French) [Presuppositions and metaphysical products of contemporary science and technology]. Dialogue: Canadian Philosophical Review / Revue canadienne de philosophie, 13:443–453, 1974. CODEN ???? ISSN 0012-2173 (print), 1759-0949 (electronic). See errata [Bun75b].

Bunge:1974:RLS

[Bun74e] Mario Bunge. The relations of logic and semantics to ontology. Journal of Philosophical Logic, 3(3):195–209, July 1974. CODEN JPLGA7. ISSN 0022-3611 (print), 1573-0433 (electronic). URL

http://www.jstor.org/stable/30226978.

Bunge:1974:SSR

[Bun74f] Mario Bunge. Semantics 1: Sense and Reference, volume 1 of His Treatise on basic philosophy; v. 1-2. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0534-8. xii + 185 pp. LCCN B840 .B77.

Bunge:1974:SIT

[Bun74g] Mario Bunge. Semantics 2: Interpretation and Truth, volume 2 of Treatise on Basic Philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0535-6. xii + 210 pp. LCCN B840 .B77 vol. 2.

Bunge:1974:T

[Bun74h] Mario Bunge. Things. International Journal of General Systems, 1(4):229–236, 1974. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Bunge:1974:MDI

[Bun74i] Mario Augusto Bunge. The methodology of development indicators. UNESCO, Methods and Analysis Division, Dept. of Social Sciences, New York, NY, USA, 1974. 20 pp.

Bunge:1975:CED

[Bun75a] Mario Bunge. A critical examination of dialectics. In Perelman [Per75], pages 63–77. ISBN 90-247-1720-5. LCCN B105.D48 I57 1975; B809.7 .E611 1973. See remarks [Nar75].

Bunge:1975:EPP

[Bun75b] Mario Bunge. Erratum: “Les présupposés et les produits métaphysiques de la science et de la technique contemporaines” (Dialogue 13 (1974), 443–453). Dialogue: Canadian Philosophical Review / Revue canadienne de philosophie, 14(1):i, 1975. ISSN 0012-2173 (print), 1759-0949 (electronic). See [Bun74d].

Bunge:1975:IS

[Bun75c] Mario Bunge. Intuition and Science. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1975. ISBN 0-8371-8066-X. x + 142 pp. LCCN Q175 .B823 1975.

Bunge:1975:PAR

[Bun75d] Mario Bunge. La paradoja de la adición: respuesta al maestro Margáin. (Spanish) [The paradox of addition: response to the teacher Margáin]. Crítica: Revista Hispanoamericana de Filosofía, 7(20):105–107, October 1975. CODEN ???? ISSN

0011-1503. URL http://www.jstor.org/stable/40104010.

See [Mar72].

Bunge:1975:RCH

[Bun75e] Mario Bunge. La representación conceptual de los hechos. (Spanish) [The conceptual representation of facts]. Teorema: Revista Internacional de Filosofía, 5(3–4):317–360, ???? 1975. CODEN

???? ISSN 0210-1602. URL http://www.jstor.org/stable/
43045872.

Bunge:1975:MS

[Bun75f] Mario Bunge. Meaning in science. Poznań Studies in the Philosophy of the Sciences and Humanities, 1(??):56–64, ???? 1975.

Bunge:1975:RPS

[Bun75g] Mario Bunge. The relevance of philosophy to social science. In Shea [She76], pages 136–155. ISBN 0-88202-160-5. LCCN Q175.3

.B37. URL http://catalog.hathitrust.org/api/volumes/
oclc/2331476.html.

Bunge:1975:WQL

[Bun75h] Mario Bunge. What is a quality of life indicator? Social Indicators Research, 2(1):65–79, June 1975. CODEN SINRDZ. ISSN 0303-8300 (print), 1573-0921 (electronic). URL http://

www.jstor.org/stable/27521741.

Bunge:1975:CNF

[Bun75i] Mario Augusto Bunge. Crítica de la noción fregeana de predicado. (Spanish) [Criticism of the Fregean notion of predicate]. Revista Latinoamericana de Filosofía, 1(??):5–8, ???? 1975.

Bunge:1975:HPA

[Bun75j] Mario Augusto Bunge. ‘?Hay proposiciones? aspectos de la filosofía de W. V. Quine. (Spanish) [‘Are there proposals? Aspects of the philosophy of W. V. Quine]. Teorema: Revista Internacional de Filosofía, ??(??):53–68, ???? 1975. ISSN 0210-1602.

Bunge:1976:BRMa

[Bun76a] Mario Bunge. Book review: From mathematics to philosophy: By Hao Wang. London (Routledge and Kegan Paul), New York (Humanities Press) 1974. 442 p. U.S. $23.90. Historia Mathematica, 3(1):100–102, February 1976. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0315086076900240.

Bunge:1976:BRMb

[Bun76b] Mario Bunge. Book review: La mathématisation des doctrines informes: Colloque tenu à l’Institut d’histoire des sciences de l’Université de Paris sous la direction de George Canguilhem. Paris (Hermann). 1972. 237 pp. Historia Mathematica, 3(2): 235, May 1976. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086076900574.

Bunge:1976:MDI

[Bun76c] Mario Bunge. Book review: La mathématisation des doctrines informes: Colloque tenu à l’Institut d’histoire des sciences de l’Université de Paris sous la direction de George Canguilhem. Paris (Hermann). 1972. 237 pp. Historia Mathematica, 3(2): 235, May 1976. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086076900574.

Bunge:1976:SNT

[Bun76d] Mario Bunge. El ser no tiene sentido y el sentido no tiene ser: Notas para una conceptología. (Spanish) [Being has no sense and sense has no being: Notes for a conceptology]. Teorema: Revista Internacional de Filosofía, 6(2):201–212, ???? 1976. CODEN

???? ISSN 0210-1602. URL http://www.jstor.org/stable/
43045900.

Bunge:1976:SVF

[Bun76e] Mario Bunge. Is science value-free and morally neutral? Philosophy and Social Action II, 4(??):5–18, ???? 1976.

Bunge:1976:MPC

[Bun76f] Mario Bunge. A model for processes combining competition with cooperation. Applied Mathematical Modelling, 1(1):21–23, 1976. CODEN AMMODL. ISSN 0307-904x (print), 1872-8480

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0307904X76900196.

Bunge:1976:PRT

[Bun76g] Mario Bunge. The philosophical richness of technology. In Suppe and Asquith [SA76], pages 153–172. ISBN 0-917586-02-6 (hardcover), 0-917586-01-8 (paperback). LCCN Q174.B67.

URL https://philpapers.org/rec/BUNTPR;   https://www.
jstor.org/stable/192379.

Bunge:1976:PP

[Bun76h] Mario Bunge. Possibility and probability. In Harper and Hooker [HH76], pages 17–33. ISBN 90-277-0621-2, 94-010-1438-8 (ebook). ISSN 1566-659X. LCCN B67. URL https://link.

springer.com/chapter/10.1007/978-94-010-1438-0_2.

Bunge:1976:RWS

[Bun76i] Mario Bunge. Review of Wolfgang Stegmüller’s The structure and dynamics of theories. Mathematical Reviews, 55(2480):333, ???? 1976. CODEN MAREAR. ISSN 0025-5629.

Bunge:1976:MBS

[Bun76j] Mario Augusto Bunge. El método en la biología. (Spanish) [The method in biology]. Naturaleza (México), 7(??):70–81, ???? 1976.

Bunge:1976:FCS

[Bun76k] Mario Augusto Bunge. La filosofía y las ciencias sociales. (Spanish) [Philosophy and the social sciences]. Grijalbo, México, DF, México, 1976. 43–69 pp.

Bunge:1976:ICF

[Bun76l] Mario Augusto Bunge. La investigación científica. (Spanish) [Scientific investigation], volume 8 of Convivium. Editorial Ariel, Barcelona, Spain, fifth edition, 1976. 955 pp.

Bunge:1977:EM

[Bun77a] Mario Bunge. Emergence and the mind. Neuroscience, 2(4):501– 509, August 1977. CODEN NRSCDN. ISSN 0306-4522 (print),

1873-7544 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0306452277900471.

Bunge:1977:ESE

[Bun77b] Mario Bunge. Epistemología. (Spanish) [Epistemology]. Siglo XXI, Buenos Aires, Argentina, 1977. ???? pp.

Bunge:1977:GSH

[Bun77c] Mario Bunge. General systems and holism. General Systems, 12 (??):87–90, ???? 1977. ISSN 0072-0798.

Bunge:1977:GCC

[Bun77d] Mario Bunge. The GST challenge to the classical philosophies of science. International Journal of General Systems, 4(1):29– 37, January 1977. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Bunge:1977:IHI

[Bun77e] Mario Bunge. The interpretation of Heisenberg’s inequalities. In Pfeiffer [Pfe77], pages 146–156. ISBN 3-492-02275-8. LCCN QC16.H35 D46. DM28.00.

Bunge:1977:LR

[Bun77f] Mario Bunge. Levels and reduction. American Journal of Physiology, 233(3):R75–R82, September 1977. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Bunge:1977:OFW

[Bun77g] Mario Bunge. Ontology I: The Furniture of the World, volume 3 of Treatise on Basic Philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1977. ISBN 90-277-0780-4 (hardcover), 90-277-0785-5 (paperback). xvi + 352 pp. LCCN BD311 .B84 vol. 1 BD331.

Bunge:1977:QMM

[Bun77h] Mario Bunge. Quantum mechanics and measurement. International Journal of Quantum Chemistry, 12(S1):1–13, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bunge:1977:RLS

[Bun77i] Mario Bunge. The relations of logic and semantics to ontology. In Problems in logic and ontology (Colloq., Inst. Wissenschaftstheorie, Salzburg, 1973), pages 29–43. Akad. Druck- und Verlagsanstalt, Graz, Austria, 1977.

Bunge:1977:RVR

[Bun77j] Mario Bunge. Reply to van Rootselaar’s criticisms of my theory of things. International Journal of General Systems, 3(3):181– 182, January 1977. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic). See [vR77].

Bunge:1977:SE

[Bun77k] Mario Bunge. States and events. In Hartnett [Har77], pages 71–95. ISBN 90-277-0822-3. LCCN Q295 .G46 1975.

Bunge:1977:TT

[Bun77l] Mario Bunge. Towards a technoethics. The Monist, 60(1):96–107, January 1977. ISSN 0026-9662 (print), 2153-3601 (electronic).

URL http://www.jstor.org/stable/27902461.

Bunge:1978:BRPa

[Bun78a] Mario Bunge. Book review: Philosophy of medicine: a new book series. Vol. 1: Evaluation and explanation in the biomedical sciences. Neuroscience, 3(7):652–653, 1978. CODEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0306452278900106.

Bunge:1978:BRPb

[Bun78b] Mario Bunge. Book review: Philosophy of medicine: a new book series. Vol. 2: Philosophical dimensions of the neuromedical sciences. Neuroscience, 3(7):652–653, July 1978. CO-DEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0306452278900118.

Bunge:1978:BRPc

[Bun78c] Mario Bunge. Book review: Philosophy of medicine: a new book series. Vol. 3: Philosophical medical ethics: Its nature and significance. Neuroscience, 3(7):652–653, 1978. CODEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030645227890012X.

Bunge:1978:SSC

[Bun78d] Mario Bunge, editor. La sémantique dans les sciences: colloque de l’Académie internationale de philosophie des sciences: [30 août au 3 septembre 1974, Rixensart]. (French) [Semantics in Science: Symposium of the International Academy of Philosophy of Science: [August 30 to September 3, 1974, Rixensart]], volume 26 of Bibliothèque des archives de philosophie: Nouvelle série. Beauchesne, Paris, France, 1978. ISSN 0768-3030. 221 pp. LCCN Q174. .A3 1978.

Bunge:1978:ME

[Bun78e] Mario Bunge. A model of evolution. Applied Mathematical Modelling, 2(3):201–204, September 1978. CO-DEN AMMODL. ISSN 0307-904X (print), 1872-8480 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/0307904X78900082.

Bunge:1978:PIO

[Bun78f] Mario Bunge. Philosophical inputs and outputs of technology. In Bugliarello and Doner [BD78], pages 262–281. ISBN 0-252-00462-0. LCCN T15 .S96 1973.

Bunge:1978:PMNc

[Bun78g] Mario Bunge. Philosophy of medicine: a new book series. vol. 1: Evaluation and explanation in the biomedical sciences. Neuroscience, 3(7):652–653, 1978. CODEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0306452278900106.

Bunge:1978:PMNa

[Bun78h] Mario Bunge. Philosophy of medicine: a new book series. vol. 2: Philosophical dimensions of the neuro-medical sciences. Neuroscience, 3(7):652–653, 1978. CODEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0306452278900118.

Bunge:1978:PMNb

[Bun78i] Mario Bunge. Philosophy of medicine: a new book series. vol. 3: Philosophical medical ethics: Its nature and significance. Neuroscience, 3(7):652–653, 1978. CODEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
030645227890012X.

Bunge:1978:IS

[Bun78j] Mario Augusto Bunge. Iatrofilosofía. (Spanish) [Iatrophilosophy]. In de Florida [dF78], pages 3–5.

Bunge:1979:BPD

[Bun79a] Mario Bunge. The bankruptcy of psychoneural dualism. Filosofskie Nauki, ??(2):77–87, ???? 1979.

Bunge:1979:BRM

[Bun79b] Mario Bunge. Book review: Mathematical modelling techniques: research notes in mathematics, vol. 24: R. Aris Pitman, London–San Francisco–Melbourne, 1978, 191 pp., £7.50. Applied Mathematical Modelling, 3(1):79, 1979. CO-DEN AMMODL. ISSN 0307-904x (print), 1872-8480 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0307904X79900787.

Bunge:1979:CMS

[Bun79c] Mario Bunge. Causality and Modern Science. Dover Publications, Inc., New York, NY, USA, third revised edition, 1979.

ISBN 0-486-23728-1. xxx + 394 pp. LCCN BD541 .B85 1979.

US$6.50. URL http://www.loc.gov/catdir/enhancements/
fy1304/78074117-d.html;   http://www.loc.gov/catdir/
enhancements/fy1318/78074117-t.html.

Bunge:1979:CCC

[Bun79d] Mario Bunge. Cerebral correlates of conscious experience. Neuroscience, 4(4):563–564, April 1979. CODEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0306452279901325.

Bunge:1979:EBD

[Bun79e] Mario Bunge. The Einstein–Bohr debate over quantum mechanics: who was right about what? In H. Nelkowski, A. Hermann, H. Poser, R. Schrader, and R. Seiler, editors, Einstein Symposion, Berlin (1979), volume 100 of Lecture Notes in Physics, pages 204– 219. Springer-Verlag, Berlin, Germany / Heidelberg, Germany /

London, UK / etc., 1979. URL https://ui.adsabs.harvard.
edu/#abs/1979LNP...100..204B.

Bunge:1979:FBT

[Bun79f] Mario Bunge. The five buds of technophilosophy. Technology in Society, 1(1):67–74, Spring 1979. CODEN ???? ISSN 0160-791x (print), 1879-3274 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0160791X79900095.

Bunge:1979:MBPa

[Bun79g] Mario Bunge. The mind–body problem in an evolutionary perspective. In O’Connor and Wolstenholme [OW79], pages 53–63. ISBN 0-444-90095-0 (Elsevier/NorthHolland), 0-470-71820-X (ebook), 90-219-4075-2. LCCN QP376 .S878 1978. URL http://

onlinelibrary.wiley.com/book/10.1002/9780470720523.

Bunge:1979:MBPb

[Bun79h] Mario Bunge. The mind–body problem, information theory and Christian dogma. Neuroscience, 4(3):453–454, March 1979. CO-DEN NRSCDN. ISSN 0306-4522 (print), 1873-7544 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0306452279901088.

Bunge:1979:MS

[Bun79i] Mario Bunge. A model of secrecy. Journal of Irreproducible Results, 25(??):25–26, ???? 1979. CODEN JIRRAL. ISSN 0022-

Bunge:1979:OIW

[Bun79j] Mario Bunge. Ontology II: a World of Systems, volume 4 of His Treatise on basic philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0944-0, 90-277-0945-9 (paperback). xvi + 314 pp. LCCN BD311 .B84 vol. 2 Q175.

Bunge:1979:STP

[Bun79k] Mario Bunge. Some topical problems in biophilosophy. Journal of Social and Biological Structures, 2(2):155–172, 1979. ISSN 0140-1750 (print), 1878-2787 (electronic). URL http://www.

sciencedirect.com/science/article/pii/014017507990006X.

Bunge:1979:SCS

[Bun79l] Mario Bunge. A systems concept of society: Beyond individualism and holism. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 10 (1–4):13–30, January 1979. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic). Methodological developments in social sciences.

Bunge:1980:BM

[Bun80a] Mario Bunge. Behavior and motivation. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 5, pages 120–134. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN BF161 .B78 1980. URL http://www.

sciencedirect.com/science/article/pii/B9780080247205500104.

Bunge:1980:CTU

[Bun80b] Mario Bunge. Conclusion: Towards understanding mind. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 10, pages 209–219. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN BF161 .B78 1980.

URL http://www.sciencedirect.com/science/article/pii/
B9780080247205500153.

Bunge:1980:CP

[Bun80c] Mario Bunge. Consciousness and personality. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 8, pages 174–190. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN BF161 .B78 1980. URL http://www.

sciencedirect.com/science/article/pii/B978008024720550013X.

Bunge:1980:ECA

[Bun80d] Mario Bunge. Epistemología: curso de actualización. (Spanish) [Epistemology: update course], volume 4 of Ciencia de la ciencia. Editorial Ariel, Barcelona, Spain, 1980. ISBN 84-344-8004-2. 275 pp. LCCN Q175 .B8227 1980.

Bunge:1980:NBM

[Bun80e] Mario Bunge. From neuron to behavior and mentation: an exercise in levelmanship. In Pinsker and Willis [PW80], pages 1–16. ISBN 0-89004-422-8. LCCN QP364.5 I53.

Bunge:1980:F

[Bun80f] Mario Bunge. The functions. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 3, pages 65– 95. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback).

LCCN BF161 .B78 1980. URL http://www.sciencedirect.
com/science/article/pii/B9780080247205500086.

Bunge:1980:Ia

[Bun80g] Mario Bunge. Introduction. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], pages xiii–xvi. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN BF161

.B78 1980. URL http://www.sciencedirect.com/science/
article/pii/B9780080247205500050.

Bunge:1980:Ib

[Bun80h] Mario Bunge. Introduction. In Bindra et al. [BAB^{+}80], pages 1–5. ISBN 0-89876-003-8. LCCN BF161 .B5.

Bunge:1980:MLb

[Bun80i] Mario Bunge. Memory and learning. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 6, pages 135–154. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback).

LCCN BF161 .B78 1980. URL http://www.sciencedirect.
com/science/article/pii/B9780080247205500116.

Bunge:1980:MBPa

[Bun80j] Mario Bunge. The mind-body problem. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 1, pages 1– 31. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback).

LCCN BF161 .B78 1980. URL http://www.sciencedirect.
com/science/article/pii/B9780080247205500062.

Bunge:1980:MBPb

[Bun80k] Mario Bunge. The Mind-Body Problem: a Psychobiological Approach. Foundations and philosophy of science and technology; Pergamon international library of science, technology, engineering, and social studies. Pergamon, New York, NY, USA, 1980. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). xvi + 250 pp. LCCN BF161 .B78 1980. URL http://

www.sciencedirect.com/science/book/9780080247205.

Bunge:1980:O

[Bun80l] Mario Bunge. The organ. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 2, pages 32–64. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN

BF161 .B78 1980. URL http://www.sciencedirect.com/
science/article/pii/B9780080247205500074.

Bunge:1980:P

[Bun80m] Mario Bunge. Preface. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], pages ix–xi. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN BF161

.B78 1980. URL http://www.sciencedirect.com/science/
article/pii/B9780080247205500049.

Bunge:1980:PIT

[Bun80n] Mario Bunge. The psychoneural identity theory. In Bindra et al. [BAB^{+}80], pages 89–108 (??). ISBN 0-89876-003-8. LCCN BF161 .B5.

Bunge:1980:SP

[Bun80o] Mario Bunge. Sensation and perception. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 4, pages 96–119. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN BF161 .B78 1980. URL http://www.

sciencedirect.com/science/article/pii/B9780080247205500098.

Bunge:1980:S

[Bun80p] Mario Bunge. Sociality. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 9, pages 191–208. ISBN 0-08-024720-2 (hardcover), 0-08-024719-9 (paperback). LCCN

BF161 .B78 1980. URL http://www.sciencedirect.com/
science/article/pii/B9780080247205500141.

Bunge:1980:T

[Bun80q] Mario Bunge. Technoethics. In Kranzberg [Kra80], pages 139– 142. ISBN 0-89158-686-5. LCCN ????

Bunge:1980:TK

[Bun80r] Mario Bunge. Thinking and knowing. In The Mind-Body Problem: a Psychobiological Approach [Bun80k], chapter 7, pages 155–173. ISBN 0-08-024720-2. LCCN BF161 .B78 1980.

URL http://www.sciencedirect.com/science/article/pii/
B9780080247205500128.

Bunge:1980:CDS

[Bun80s] Mario Augusto Bunge. Ciencia y desarrollo. (Spanish) [Science and development]. Siglo Veinte, Buenos Aires, Argentina, 1980. ???? pp.

Bunge:1980:CMS

[Bun80t] Mario Augusto Bunge. La ciencia : su método y su filosofía. (Spanish) [Science: its method and its philosophy]. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1980. 109 pp.

Bunge:1981:ABS

[Bun81a] Mario Bunge. Analogy between systems. International Journal of General Systems, 7(4):221–223, 1981. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Bunge:1981:BBI

[Bun81b] Mario Bunge. Biopopulations, not biospecies, are individuals and evolve. Behavioral and Brain Sciences, 4(2):284–285, June 1981. CODEN BBSCDH. ISSN 0140-525X (print), 1469-1825 (elec-

tronic). URL https://www.cambridge.org/core/journals/
behavioral-and-brain-sciences/article/biopopulations-
not-biospecies-are-individuals-and-evolve/.

Bunge:1981:CPC

[Bun81c] Mario Bunge. Causality: the place of the causal principle in modern science. University Microfilms International, Ann Arbor, MI, USA, 1981. LCCN BD541.

Bunge:1981:CE

[Bun81d] Mario Bunge. Conceptual existence. In Cohn [Coh81], pages 5–14. ISBN 0-391-02361-6. LCCN B29 .T75.

Bunge:1981:DI

[Bun81e] Mario Bunge. Development indicators. Social Indicators Research, 9(3):369–385, September 1981. CODEN SINRDZ. ISSN 0303-8300 (print), 1573-0921 (electronic). URL http://www.

jstor.org/stable/27520981.

Bunge:1981:FCP

[Bun81f] Mario Bunge. Four concepts of probability. Applied Mathematical Modelling, 5(5):306–312, 1981. CODEN AMMODL. ISSN 0307-904x (print), 1872-8480 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0307904X81800510.

Bunge:1981:MNB

[Bun81g] Mario Bunge. From mindless neuroscience and brainless psychology to neuropsychology. Annals of Theoretical Psychology, 3: 115–133, ???? 1981. ISSN 0747-5241. Comments by de M. C. Corballis and P.C. Dodwell.

Bunge:1981:HT

[Bun81h] Mario Bunge. Half truths. In Philosophie als Wissenschaft: Paul Weingartner gewidmet = Essays in scientific philosophy: dedicated to Paul Weigartner [Mor81], pages 87–92. ISBN 3-88820-000-8. LCCN B790 .P48 1981.

Bunge:1981:MC

[Bun81i] Mario Bunge. Materialismo y Ciencia. (Spanish) [Materialism and Science], volume 164 of Ariel quincenal. Editorial Ariel, Barcelona, Spain, 1981. ISBN 84-344-0828-7. 235 pp. LCCN Q175 .B824 1981.

Bunge:1981:RC

[Bun81j] Mario Bunge. The revival of causality. In Fløistad [Flø86], pages 133–155. ISBN 90-247-2518-6 (hardcover), 90-247-3298-0 (paperback). LCCN ????

Bunge:1981:SM

[Bun81k] Mario Bunge. Scientific Materialism, volume 9 of Episteme. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1981. ISBN 90-277-1304-9 (hardcover), 90-277-1305-7 (paperback). xiv + 219 pp. LCCN Q175 .B827.

Bunge:1981:SAW

[Bun81l] Mario Bunge. Systems all the way. Nature and System: Philosophical Studies of Natural and Artificial Systems, 3(??):37–47, ???? 1981. ISSN 0191-2941.

Bunge:1981:CMS

[Bun81m] Mario Augusto Bunge. La ciencia : su método y su filosofía. (Spanish) [Science: its method and its philosophy]. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1981. 110 pp.

Bunge:1982:EF

[Bun82a] Mario Bunge. Economía y filosofía. (Spanish) [Economy and Philosophy]. Cuadernos de filosofía y ensayo. Tecnos, Madrid, Spain, 1982. ISBN 84-309-0922-2. 125 pp. LCCN MLCS 90/11839 (H).

Bunge:1982:CBP

[Bun82b] Mario Bunge. Is chemistry a branch of physics? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):209–223, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170620.

Bunge:1982:CFS

[Bun82c] Mario Augusto Bunge. Controversias en Física. (Spanish) [Controversies in Physics]. Cuadernos de Filosofia y Ensayo. Tecnos, Madrid, Spain, 1982. ISBN 84-309-0981-8. 252 pp. LCCN QC6.2 .B86 1983.

Bunge:1983:CMB

[Bun83a] Mario Bunge. Comment by Mario Bunge. Journal of Social and Biological Structures, 6(2):159–160, 1983. ISSN 0140-1750 (print), 1878-2787 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0140175083800141.

Bunge:1983:CF

[Bun83b] Mario Bunge. Controversias en física. (Spanish) [Controversies in Physics]. Cuadernos de filosofía y ensayo. Tecnos, Madrid, Spain, 1983. ISBN 84-309-0981-8. 252 pp. LCCN QC6.2 .B86 1983.

Bunge:1983:NSA

[Bun83c] Mario Bunge. De la neurologie sans âme et de la psychologie sans tête à la neuropsychologie. (French) [From soulless neurology and headless psychology to neuropsychology]. La petite revue de philosophie, 5(??):1–45, 1983. ISSN 0709-4469.

Bunge:1983:DSP

[Bun83d] Mario Bunge. Demarcating science from pseudoscience. Fundamenta Scientiae, 3(??):369–388, ???? 1983. CODEN FUSCDO. ISSN 0160-7847.

Bunge:1983:PFE

[Bun83e] Mario Bunge. Épistémologie. (French) [Epistemology]. Maloine, Paris, France, 1983. ???? pp.

Bunge:1983:EME

[Bun83f] Mario Bunge. Epistemology and Methodology I: Exploring the World, volume 5 of Treatise on basic philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1511-4, 90-277-1523-8 (paperback). xv + 404 pp. LCCN BD161 .B86 1983 no. 1.

Bunge:1983:EMI

[Bun83g] Mario Bunge. Epistemology and Methodology II: Understanding the World, volume 6 of Treatise on basic philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1634-X, 90-277-1635-8 (paperback), 94-015-6921-5 (e-book). xii + 296 pp. LCCN BD161

.B86 1983 no. 2. URL http://link.springer.com/10.1007/
978-94-015-6921-7.

Bunge:1983:ICF

[Bun83h] Mario Bunge. La investigación científica. (Spanish) [Scientific investigation]. Editorial Ariel, Barcelona, Spain, 1983. ???? pp.

Bunge:1983:LF

[Bun83i] Mario Bunge. Lingüística y filosofía. (Spanish) [Linguistics and Philosophy]. Ariel quincenal. Editorial Ariel, Barcelona, Spain, 1983. ISBN 84-344-0816-3. 139 pp. LCCN P123 .B79 1983.

Bunge:1983:NAS

[Bun83j] Mario Bunge. The nature of applied science and technology. In Cauchy [Cau88], pages 599–604. ISBN 2-920449-10-9 (vol. 1), 2-9801018-1-8 (vol. 2), 2-9801018-2-6 (vol. 3), 2-9801018-3-4 (vol. 4), 2-9801018-4-2 (vol. 5). LCCN B20 .W63 1983; HM101 .W88

Bunge:1983:SWS

[Bun83k] Mario Bunge. Speculation: Wild and sound. New Ideas in Psychology, 1(1):3–6, 1983. ISSN 0732-118x (print), 1873-3522

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0732118X83900247.

Bunge:1983:SVS

[Bun83l] Mario Augusto Bunge. Significado y verdad. (Spanish) [Meaning and truth]. Editorial Ariel, Barcelona, Spain, 1983. 347 pp.

Bunge:1984:BRM

[Bun84a] Mario Bunge. Book review: Mathematical models of morphogenesis: René Thom (translated by W. M. Brookes and D. Rand) Ellis Horwood Ltd, Chichester (John Wiley, New York), 1983, 305 pp., £25. Applied Mathematical Modelling, 8(3):218, 1984. CO-DEN AMMODL. ISSN 0307-904x (print), 1872-8480 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0307904X84900945.

Bunge:1984:BRP

[Bun84b] Mario Bunge. Book review: Popper and after: Four modern irrationalists: by D. S. Stove, Pergamon Press, Oxford (1982). New Ideas in Psychology, 2(1):81–82, 1984. ISSN 0732-118x (print),

1873-3522 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0732118X84900369.

Bunge:1984:HVS

[Bun84c] Mario Bunge. Hidden variables, separability, and realism. Revista Brasileira de Física, ??(??):150–168, ???? 1984. CODEN RBFSA3. ISSN 0374-4922. Special volume in homage of Mário Schenberg.

Bunge:1984:PPL

[Bun84d] Mario Bunge. Philosophical problems in linguistics. Erkenntnis, 21(2):107–173, July 1984. ISSN 0165-0106 (print),

1572-8420 (electronic). URL https://link.springer.com/
article/10.1007/BF00166382;   https://www.jstor.org/
stable/20010903.

Bunge:1984:RGP

[Bun84e] Mario Bunge. Reviews: God and the physicists. Free Inquiry, 4(2):??, Spring 1984. CODEN FRINFM. ISSN 0272-

0701. URL https://secularhumanism.org/1984/04/god-and-
the-physicists/.

Bunge:1984:WP

[Bun84f] Mario Bunge. What is pseudoscience? The Skeptical Inquirer, IX(??):36–46, ???? 1984. ISSN 0194-6730.

Bunge:1983:EM

[Bun85a] Mario Bunge. Epistemology and Methodology, volume 5–7 of Treatise on basic philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983– 1985. ISBN 90-277-1511-4 (vol. 1), 90-277-1523-8 (paperback: v. 1). ???? pp. LCCN BD161 .B86 1983.

Bunge:1985:CNT

[Bun85b] Mario Bunge. Cajas negras y translúcidas y acción a distancia: Sánchez Ron. (Spanish) [Black and translucent boxes and remote action: Sánchez Ron]. Teorema: Revista Internacional de Filosofía, 15(1–2):271–274, ???? 1985. CODEN ???? ISSN

0210-1602. URL http://www.jstor.org/stable/43046172.

Bunge:1985:CAP

[Bun85c] Mario Bunge. Comment on Apostol’s paper. Studies in Soviet Thought, 29(2):137–138, February 1985. CODEN SSVTBD. ISSN 0039-3797 (print), 2214-7136 (electronic). URL http://

www.jstor.org/stable/20099997.

Bunge:1985:DEA

[Bun85d] Mario Bunge. Degrees of existence or of abstraction? Theoria: An International Journal for Theory, History and Foundations of Science, 1(2):547–549, 1985. ISSN 0495-4548 (print), 2171-679X (electronic).

Bunge:1985:EMI

[Bun85e] Mario Bunge. Epistemology and Methodology III: Philosophy of Science and Technology, volume 7 of Epistemology and methodology; Treatise on basic philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-277-1903-9 (part 1), 90-277-1904-7 (paperback: part 1). xi + 263 pp. LCCN BD161 .B86 1983 no. 3 Q175.

Bunge:1985:TGE

[Bun85f] Mario Bunge. ‘?Grados de existencia o de abstraccion?. (Spanish) [Degrees of existence or abstraction?]. Theoria: An International Journal for Theory, History and Foundations of Science, 1(2): 547–549, ???? 1985. CODEN ???? ISSN 0495-4548 (print),

2171-679X (electronic). URL http://www.jstor.org/stable/
23916146.

Bunge:1985:ISS

[Bun85g] Mario Bunge. Individuals, sets and systems: reply to Mosterín. Theoria: An International Journal for Theory, History and Foundations of Science, 1(2):555–560, 1985. ISSN 0495-4548 (print), 2171-679X (electronic).

Bunge:1985:ICS

[Bun85h] Mario Bunge. Individuos, conjuntos y sistemas: Réplica a Mosterín. (Spanish) [Individuals, sets and systems: Reply to Mosterín]. Theoria: An International Journal for Theory, History and Foundations of Science, 1(2):555–560, ???? 1985. ISSN 0495-4548 (print), 2171-679X (electronic). URL http://www.

jstor.org/stable/23916148.

Bunge:1985:RSP

[Bun85i] Mario Bunge. On research strategies in psychology. reply to commentators. Annals of Theoretical Psychology, 3(??):151–156, ???? 1985. ISSN 0747-5241.

Bunge:1985:PNB

[Bun85j] Mario Bunge, editor. The Philosophy of Niels Bohr, volume 85/03 of Cahiers d’épistémologie. Université du Québec à Montréal, Montréal, QC, Canada, 1985. ISBN 2-920250-23-X. ISSN 0228-7080. LCCN QC16.B63 B86 1985.

Bunge:1985:TQI

[Bun85k] Mario Bunge. ‘?Que es un individuo concreto?. (Spanish) [What is a concrete individual?]. Theoria: An International Journal for Theory, History and Foundations of Science, 1(1):121–128, ???? 1985. CODEN ???? ISSN 0495-4548 (print), 2171-679X

(electronic). URL http://www.jstor.org/stable/23912568.

Bunge:1985:RAF

[Bun85l] Mario Bunge. Realismo y antirrealismo en la filosofía contemporánea. (Spanish) [Realism and antirealism in contemporary philosophy]. Arbor, 473(??):13–40, ???? 1985. ISSN ????

Bunge:1985:SIS

[Bun85m] Mario Bunge. Seudociencia e ideología. (Spanish) [Pseudoscience and ideology], volume 440 of Alianza Universidad. Alianza Editorial, Madrid, Spain, 1985. ISBN 84-206-2440-3. 253 pp. LCCN ????

Bunge:1985:TPE

[Bun85n] Mario Bunge. Types of psychological explanation. In MacGaugh [Mac85], pages 489–501. ISBN 0-444-87781-9 (vol. 1), 0-444-87782-7 (vol. 2), 0-444-87783-5 (vol. 3), 0-444-87784-3 (vol. 4), 0-444-87785-1 (vol. 5), 0-444-87786-X (vol. 6), 0-444-87787-8 (vol. 7), 0-444-87788-6 (vol. 8), 0-444-87789-4 (vol. 9), 0-444-87790-8 (set). LCCN BF121 .I564 1984 vol. 1.

Bunge:1985:WCI

[Bun85o] Mario Bunge. What is a concrete individual? Theoria: An International Journal for Theory, History and Foundations of Science, 1(1):121–128, 1985. ISSN 0495-4548 (print), 2171-679X (electronic).

Bunge:1985:CMF

[Bun85p] Mario Augusto Bunge. La ciencia : su método y su filosofía. (Spanish) [Science: its method and its philosophy]. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1985. ???? pp.

Bunge:1985:RRS

[Bun85q] Mario Augusto Bunge. Racionalidad y realismo. (Spanish) [Rationality and realism], volume 445 of Alianza universidad. Alianza Editorial, Madrid, Spain, 1985. ISBN 84-206-2445-4 (paperback). 191 pp. LCCN B833 .B8 1985; BD115 .B86

1985; BD181 .B86 1985. URL http://catalogue.bnf.fr/ark:
/12148/cb36184595w.

Bunge:1986:BRI

[Bun86a] Mario Bunge. Book review: An Idiot’s Fugitive Essays on Science. Methods, Criticism, Training, Circumstances by Clifford Truesdell. British Journal for the Philosophy of Science, 37(4):

520–523, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.org/

stable/687379.

Bunge:1986:BRB

[Bun86b] Mario Bunge. Book review: Rational Thermodynamics by C. Truesdell. Philosophy of Science, 53(2):305–306, June 1986. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187703.

Bunge:1986:BRT

[Bun86c] Mario Bunge. Book review: Truesdell, Clifford [1984]: An Idiot’s Fugitive Essays on Science. Methods, Criticism, Training, Circumstances. Springer-Verlag. xviii + 654 pp. $58.00. British Journal for the Philosophy of Science, 37(4):520–523, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic). URL http://www.jstor.org/stable/687379.

Bunge:1986:CIF

[Bun86d] Mario Bunge. Culture et inculture. (French) [Culture and lack of culture]. Philosophiques: revue de la Société de philosophie du Québec, 13(2):347–351, 1986. ISSN 0316-2923 (print), 1492-

1391 (electronic). URL https://laspq.org/philosophiques/
volume-13-numero-2-automne-1986.

Bunge:1986:DRS

[Bun86e] Mario Bunge. In defence of realism and scientism. Annals of Theoretical Psychology, 4(??):23–26, ???? 1986. ISSN 0747-5241.

Bunge:1986:IRS

[Bun86f] Mario Bunge. Intuición y Razón. (Spanish) [Intuition and Reason]. Cuadernos de filosofía y ensayo. Tecnos, Madrid, Spain, 1986. ISBN 84-309-1335-1. 180 pp. LCCN BD181 .B87 1986.

Bunge:1986:PLC

[Bun86g] Mario Bunge. A philosopher looks at the current debate on language acquisition. In Gopnik and Gopnik [GG86], pages 229–239. ISBN 0-89391-355-3. LCCN BF311 .F75 1986.

Bunge:1986:PPL

[Bun86h] Mario Bunge. Philosophical Problems in Linguistics. Seishin-Shobo, Tokyo, Japan, 1986. ???? pp.

Bunge:1986:STI

[Bun86i] Mario Bunge. Science, technology and ideology in the Hispanic World. Free Inquiry, 6(3):36–49, Summer 1986. CODEN

FRINFM. ISSN 0272-0701. URL https://secularhumanism.
org/1986/07/science-technology-and-ideology-in-the-hispanic-
world/.

Bunge:1986:CMF

[Bun86j] Mario Augusto Bunge. La ciencia : su método y su filosofía. (Spanish) [Science: its method and its philosophy]. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1986. 110 pp.

Bunge:1986:SIS

[Bun86k] Mario Augusto Bunge. Seudociencia e ideología. (Spanish) [Pseudoscience and ideology]. Alianza, México, DF, México, second edition, 1986. ISBN 968-6001-55-7. 253 pp. LCCN Q175 B86.

Bunge:1987:BRB

[Bun87a] Mario Bunge. Book review: The biology of learning: P. Marler and H. S. Terrace, Eds., Berlin-Heidelberg-New York-Tokyo: Springer-Verlag, 1984. pp. x + 738. ISBN: 3-540-13923-0. US$31. Behavioural Processes, 14(2):241–244, 1987. ISSN 0376-6357.

URL http://www.sciencedirect.com/science/article/pii/
0376635787900581.

Bunge:1987:GAA

[Bun87b] Mario Bunge. Growing up agnostic in Argentina. Free Inquiry, 7(3):??, Summer 1987. CODEN FRINFM. ISSN 0272-

0701. URL https://secularhumanism.org/1987/07/growing-
up-agnostic-in-argentina/.

Bunge:1987:KGP

[Bun87c] Mario Bunge. Kausalität, Geschichte und Probleme. (German) [Causality, History, and Problems], volume 51 of Die Einheit der Gesellschaftswissenschaften [The Unity of Social Sciences]. J. C. B. Mohr (Paul Siebeck), Tübingen, Germany, 1987. ISBN 3-16-944806-4. xxiv + 450 pp. Translated from the English by Herbert Spengler.

Bunge:1987:PCE

[Bun87d] Mario Bunge. Le problème corps–esprit. (French) [The mind– body problem]. Médecine Psychosomatique, 15(??):85–94, ????

Bunge:1987:PSP

[Bun87e] Mario Bunge. Philosophie, science, politique (entretien). (French) [Philosophy, science, politics (interview)]. Concordia: revista internacional de filosofia : internationale Zeitschrift für Philosophie, 10(??):51–67, 1987. ISSN 0179-0846.

Bunge:1987:SDR

[Bun87f] Mario Bunge. Seven desiderata for rationality. In Agassi and Jarvie [AJ87], pages 5–15. ISBN 90-247-3275-1, 90-247-3455-X.

LCCN B833 .R344 1986. URL http://www.springer.com/gb/.

Bunge:1987:TPM

[Bun87g] Mario Bunge. Ten philosophies of mind in search of a scientific sponsor. In Logik, Wissenschaftstheorie und Erkenntnistheorie (Kirchberg am Wechsel, 1986) [Proceedings of the 11th International Wittgenstein Symposium], volume 13 of Schriftenreihe Wittgenstein-Gesellschaft, pages 285–293. Hölder–Pichler– Tempsky, Vienna, Austria, 1987.

Bunge:1987:TCE

[Bun87h] Mario Bunge. Two controversies in evolutionary biology: saltationism and cladism. In Rescher [Res87], pages 129–145. ISBN 0-8191-5798-8, 0-8191-5799-6 (paperback). LCCN Q175 .S4235 1987.

Bunge:1987:VEO

[Bun87i] Mario Bunge. Vistas y Entrevistas: Opiniones Impopulares Sobre Problemas de Actualidad. (Spanish) [Views and Interviews: Unpopular Opinions on Current Problems]. Problemas de nuestro tiempo. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1987. ISBN 950-516-469-6. 239 pp. LCCN F1414 .B93 1987.

Bunge:1987:CMF

[Bun87j] Mario Augusto Bunge. La ciencia : su método y su filosofía. (Spanish) [Science: its method and its philosophy]. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1987. 110 pp.

Bunge:1988:ALO

[Bun88a] Mario Bunge. The ambivalent legacy of operationism. Philosophia Naturalis, 25(3–4):337–345, 1988. ISSN 0031-8027.

Bunge:1988:APS

[Bun88b] Mario Bunge. Analytic philosophy of society and social science: The systemic approach as an alternative to holism and individualism. Revue internationale de Systémique, 2(??):1–13, ???? 1988.

Bunge:1988:BSI

[Bun88c] Mario Bunge. Basic science is innocent, applied science and technology can be guilty. In Lemarchand and Pedace [LP88], pages 245–261. ISBN 9971-5-0761-7. LCCN ????

Bunge:1988:IS

[Bun88d] Mario Bunge. Ideology and science. In Eberlein and Berghel [EB88], pages 79–89. ISBN 90-277-2519-5 (hardcover). LCCN H61 .T466 1988.

Bunge:1988:NBP

[Bun88e] Mario Bunge. Niels Bohr’s philosophy. Philosophia Naturalis, 25 (3–4):399–415, 1988. ISSN 0031-8027.

Bunge:1988:SSH

[Bun88f] Mario Bunge. The scientific status of history. In Hinke-Dörnemann [HD88], pages 593–602. ISBN 3-8204-7349-1 (set).

Bunge:1988:SQS

[Bun88g] Mario Bunge. Superposition of quantum states: fact or fiction? In Fleury et al. [FJMT88], pages 135–142. ISBN 9971-5-0693-9. LCCN QC170 .L447 1988.

Bunge:1988:TRA

[Bun88h] Mario Bunge. The thirteen riders of the Apocalypse. Free Inquiry, 8(3):9, Summer 1988. CODEN FRINFM. ISSN 0272-0701.

Bunge:1988:TFT

[Bun88i] Mario Bunge. Two faces and three masks of probability. In Agazzi [Aga88b], pages 27–49. ISBN 90-277-2808-9. LCCN Q174 .P76 1988.

Bunge:1988:WPC

[Bun88j] Mario Bunge. Why parapsychology cannot become a science. Behavioral and Brain Sciences, 10(4):576–577, December 1988. CODEN BBSCDH. ISSN 0140-525X (print),

1469-1825 (electronic). URL https://www.cambridge.org/
core/journals/behavioral-and-brain-sciences/article/
why-parapsychology-cannot-become-a-science/.

Bunge:1989:BIA

[Bun89a] Mario Bunge. The Bell inequalities and all that. Philosophia Naturalis, 26(1):121–134, 1989. ISSN 0031-8027.

Bunge:1989:BMU

[Bun89b] Mario Bunge. Des bons et des mauvais usages de la philosophie. (French) [Good and bad uses of philosophy]. L’Enseignement philosophique: revue de l’Association des Professeurs de Philosophie de l’Enseignement Public, 40(3):97–110, 1989. ISSN 0035-1393.

Bunge:1989:EGR

[Bun89c] Mario Bunge. Ethics: the Good and the Right, volume 8 of Treatise on basic philosophy. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1989. ISBN 90-277-2839-9 (hardcover), 90-277-2840-2 (paperback), 94-009-2601-4 (e-book). xv + 427 pp. LCCN BJ1012 .B75 1989.

Bunge:1989:NM

[Bun89d] Mario Bunge. From neuron to mind. News in Physiological Sciences, 4(5):206–210, October 1989. CODEN NEPSEY. ISSN 0886-1714 (print), 1522-161X (electronic). URL https://www.

physiology.org/doi/abs/10.1152/physiologyonline.1989.
4.5.206.

Bunge:1989:GTU

[Bun89e] Mario Bunge. Game theory is not a useful tool for the political scientist. Epistemologia (Genoa, Italy), 12(1):195–212, ???? 1989. CODEN ???? ISSN 0392-9760 (print), 1825-652X (electronic).

Bunge:1989:PPS

[Bun89f] Mario Bunge. The popular perception of science in North America. Transactions of the Royal Society of Canada, V(??):269–289, ???? 1989. CODEN TRSCAI. ISSN 0035-9122.

Bunge:1989:RIS

[Bun89g] Mario Bunge. Reduktion und Integration, Systeme und Niveaus, Monismus und Dualismus. (German) [Reduction and integration, systems and levels, monism and dualism]. In Pöppel and Beckermann [PB89], pages 87–104. ISBN 3-527-27901-6. LCCN

BF311 .G42 1989. URL http://digitool.hbz-nrw.de:1801/
webclient/DeliveryManager?pid=3277305.

Bunge:1989:TSM

[Bun89h] Mario Bunge. Toward a survival morality. In Kurtz et al. [KFT89], pages 36–41. ISBN 0-87975-538-5. LCCN BL2747.6 .B85 1989.

Bunge:1989:ICS

[Bun89i] Mario Augusto Bunge. La investigación científica. (Spanish) [Scientific investigation]. Editorial Ariel, Barcelona, Spain, second corrected edition, 1989. 942 pp.

Bunge:1989:MSE

[Bun89j] Mario Augusto Bunge. Mente y sociedad: ensayos irritantes. (Spanish) [Mind and society: irritating essays], volume 612 of Alianza universidad Filosofía. Alianza, Madrid, Spain, 1989. ISBN 84-206-2612-0. 205 pp. LCCN B63 .B86 1989; H61 .B8518 1989.

Bunge:1990:BAR

[Bun90a] Mario Bunge. Boudon on anti-realism in social studies. In Weingartner and Dorn [WD90], pages 613–616. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990.

Bunge:1990:CBA

[Bun90b] Mario Bunge. Computerism: a brainless approach to cognition: A reply to Sloman. New Ideas in Psychology, 8(3): 377–379, 1990. ISSN 0732-118x (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0732118X9490023X.

Bunge:1990:BMU

[Bun90c] Mario Bunge. Des bons et mauvais usages de la philosophie. (French) [Good and bad uses of philosophy]. L’enseignement de la Philosophie, 40(2):97–110, ???? 1990.

Bunge:1990:IA

[Bun90d] Mario Bunge. Instant autobiography. In Weingartner and Dorn [WD90], pages 677–681. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990.

Bunge:1990:NPP

[Bun90e] Mario Bunge. The nature and place of psychology: a reply to Panksepp, Mayer, Royce, and Cellerier and Ducret. New Ideas in Psychology, 8(2):177–188, 1990. ISSN 0732-118X (print),

1873-3522 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0732118X9090007O. See [Bun90g, CD90,

May90, Pan90, Roy90].

Bunge:1990:SPT

[Bun90f] Mario Bunge. Seni on planning theory. In Weingartner and Dorn [WD90], page ?? ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990. See [Sen93].

Bunge:1990:WKD

[Bun90g] Mario Bunge. What kind of discipline is psychology: Autonomous or dependent, humanistic or scientific, biological or sociological? New Ideas in Psychology, 8(2):121– 137, 1990. ISSN 0732-118x (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0732118X9090002J. See comments [CD90, May90, Pan90, Roy90] and reply [Bun90e].

Bunge:1990:ZAT

[Bun90h] Mario Bunge. Zecha’s acid test of agathonism. In Weingartner and Dorn [WD90], pages 660–663. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990.

Bunge:1991:BSI

[Bun91a] Mario Bunge. Basic science is innocent: applied science and technology can be guilty. In Dahlstrom [Dah91], page ?? ISBN 0-8132-0723-1. LCCN B21 .S78 vol. 22 Q175.3. URL http://

www.jstor.org/stable/10.2307/j.ctv176d3.

Bunge:1991:CEN

[Bun91b] Mario Bunge. A critical examination of the new sociology of science: Part 1. Philosophy of the Social Sciences = Philosophie des sciences sociales, 21(4):524–560, December 1991. CODEN ???? ISSN 0048-3931 (print), 1552-7441 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/004839319102100406.

Bunge:1991:VFT

[Bun91c] Mario Bunge. La vérité. (French) [Truth]. In Sinaceur [Sin91], pages 453–457. ISBN 2-86586-191-0 (paperback). LCCN B485 .P39 1991.

Bunge:1991:LEF

[Bun91d] Mario Bunge. Le lieu et l’espace. (French) [The place and the space]. In Sinaceur [Sin91], pages 483–488. ISBN 2-86586-191-0 (paperback). LCCN B485 .P39 1991.

Bunge:1991:STS

[Bun91e]  Mario Bunge.  Le système technique-science-philosophie:

un ménage à trois fécond. (French) [The technical-science-philosophy system: a fertile threesome]. Revue internationale de systémique, 5:171–180, 1991. ISSN 0980-1472. URL

http://www.res-systemica.org/ris/vol-05/vol05-num-02/
ris-vol05-num02-p171-180.pdf.

Bunge:1991:PPM

[Bun91f] Mario Bunge. A philosophical perspective on the mind–body problem or, why neuroscientists and psychologists should care about philosophy. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 135(4): 513–523, December 1991. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/

stable/986813.

Bunge:1991:PLR

[Bun91g] Mario Bunge. The power and limits of reduction. In Agazzi [Aga91], pages 31–49. ISBN 0-7923-1406-9 (hardcover), 94-010-5544-0, 94-011-3492-8 (e-book). LCCN Q174 .P82 1991.

Bunge:1991:RID

[Bun91h] Mario Bunge. Rights imply duties. In Grofier et al. [GPA91], pages 47–54. ISBN 0-88629-164-X (hardcover), 0-88629-163-1 (paperback), 0-7735-8220-7 (e-book). LCCN HM1271

.N685 1991. URL http://www.jstor.org/stable/10.2307/j.
ctt7zt3sb.

Bunge:1991:SBD

[Bun91i] Mario Bunge. A skeptic’s beliefs and disbeliefs. New Ideas in Psychology, 9(2):131–149, 1991. ISSN 0732-118x (print), 1873-3522

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0732118X9190017G. Special issue on Mario Bunge

and Martin Mahner.

Bunge:1991:WSD

[Bun91j] Mario Bunge. What is science? Does it matter to distinguish it from pseudoscience? A reply to my commentators. New Ideas in Psychology, 9(2):245–283, 1991. ISSN 0732-118x (print),

1873-3522 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0732118X9190030P. Special issue on

Mario Bunge and Martin Mahner.

Bunge:1991:WWC

[Bun91k] Mario Bunge. Why we cherish exactness. In Schurz and Dorn [SD91], pages 591–598. ISBN 90-5183-317-2. LCCN B29 .A28 1991; Q175.3A48 1991.

Bunge:1991:CCN

[Bun91l] Mario Augusto Bunge. Una caricatura de la ciencia: la novísima sociología de la ciencia. (Spanish) [A caricature of science: the latest sociology of science]. Interciencia, 16(??):69–77, ???? 1991.

Bunge:1992:CEN

[Bun92a] Mario Bunge. A critical examination of the new sociology of science: Part 2. Philosophy of the Social Sciences = Philosophie des sciences sociales, 22(1):46–76, March 1992. ISSN 0048-3931 (print), 1552-7441 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/004839319202200103.

Bunge:1992:BPF

[Bun92b] Mario Bunge. Die Bedeutung der Philosophie für die Psychologie. (German) [The importance of philosophy for psychology]. In Montada [Mon92], pages 51–63. ISBN 3-8017-0667-2. LCCN ????

Bunge:1992:PNB

[Bun92c] Mario Bunge. La philosophie de Niels Bohr. Horizons philosophiques, 2(2):27–50, 1992. ISSN 1181-9227 (print), 1920-

2954 (electronic). URL https://id.erudit.org/iderudit/
800894ar.

Bunge:1992:MCN

[Bun92d] Mario Bunge. Mach’s critique of Newtonian mechanics. In Ernst Mach — a Deeper Look: Documents and New Perspectives [Bla92], chapter 13, pages 243–261. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. LCCN C5.53; QC6.9.

URL https://link.springer.com/chapter/10.1007/978-94-
011-2771-4_13.

Bunge:1992:MT

[Bun92e] Mario Bunge. Materialism today. Theoria Hist. Sci., 2:7–15 (1993), 1992. ISSN 0867-4159.

Bunge:1992:SB

[Bun92f] Mario Bunge. System boundary. International Journal of General Systems, 20(3):215–219, March 1992. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Bunge:1993:BR

[Bun93a] Mario Bunge. Book review. Foundations of Physics, 23:149–151, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://ui.adsabs.harvard.edu/#
abs/1993FoPh...23..149B.

Bunge:1993:CSS

[Bun93b] Mario Bunge. The concept of a social system. In Delgado and Banathy [DB93], pages 210–221. ISBN 84-604-6236-6 (paperback). LCCN Q295 .I58 1993; TA168 .I57x 1993.

Bunge:1993:RAS

[Bun93c] Mario Bunge. Realism and antirealism in social science. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 35(3):207–235, November 1993. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Bunge:1993:SCP

[Bun93d] Mario Bunge. Seven cosmological paradigms: Animal, ladder, river, cloud, machine, book, and system of systems. In Sánchez Sorondo and Antiseri [SA93], pages 115–131.

Bunge:1993:SCS

[Bun93e] Mario Augusto Bunge. Sociología de la Ciencia. (Spanish) [Sociology of Science]. Ediciones Siglo Veinte, Buenos Aires, Argentina, 1993. ???? pp.

Bunge:1994:CW

[Bun94a] M. Bunge. “Culture wars”. Science, 265(5174):854, August 12, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bunge:1994:CEC

[Bun94b] Mario Bunge. Counter-enlightenment in contemporary social studies. In Madigan [Mad94], pages 25–42. ISBN 0-87975-869-4. LCCN ????

Bunge:1994:SSM

[Bun94c] Mario Bunge. La science, sa méthode et sa philosophie. (French) [Science, its method and philosophy]. Éditions Vigdor, Paris, France, 1994. ISBN 2-910243-01-X. ???? pp.

Bunge:1994:LMR

[Bun94d] Mario Bunge. L’écart entre les mathématiques et le réel. (French) [The gap between mathematics and reality]. In Passion des formes: dynamique qualitative, sémiophysique et intelligibilité: à René Thom (French) [Passion of forms: qualitative dynamics, semiophysics and intelligibility: to René Thom] [BP94], pages 165–173. ISBN 2-902126-09-3 (paperback). LCCN Q175

.P34238 1994. URL http://catalogue.bnf.fr/ark:/12148/
cb35763316f.

Bunge:1994:SRD

[Bun94e] Mario Bunge. Survival, rights, and duties. In del Carmen Patricia) Morales [dCPM94], pages 109–114. ISBN 90-802139-1-8. LCCN K3240.4 .I52 1994; K3247 I54.1994.

Bunge:1995:HVS

[Bun95a] Mario Bunge. Hidden variables, separability, and realism. In Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc [MC95], chapter 12, pages 217–227. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346. LCCN ????

URL https://link.springer.com/chapter/10.1007/978-94-
009-1575-6_12.

Bunge:1995:PRC

[Bun95b] Mario Bunge. The poverty of rational choice theory. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95], chapter 10, pages 149–168. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-011-0471-5_10.

Bunge:1995:RCT

[Bun95c] Mario Bunge. Rational choice theory: a critical look at its foundations. In Götschl [Göt95], pages 211–228. ISBN 0-7923-3627-5 (hardcover). ISSN 0921-3384. LCCN H61 .R474 1995.

Bunge:1995:RAC

[Bun95d] Mario Bunge. Realismo y antirrealismo en las ciencias sociales. (Spanish) [Realism and anti-realism in the social sciences]. Mientras Tanto, ??(61):21–47, Primavera 1995. CODEN ???? ISSN 0210-8259 (print), 2014-7317 (electronic). URL http://www.

jstor.org/stable/27820198.

Bunge:1995:SSF

[Bun95e] Mario Bunge. Sistemas Sociales y Filosofía. (Spanish) [Social Systems and Philosophy]. Editorial Sudamericana, Buenos Aires, Argentina, 1995. ISBN 950-07-1074-9. 197 pp. LCCN HM131 .B76 1995.

Bunge:1995:AES

[Bun95f] Mario Augusto Bunge. Axiología y ética. (Spanish) [Axiology and ethics]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament Contemporani., 1995. URL http://hdl.handle.

net/10256.1/601.

Bunge:1995:EAF

[Bun95g] Mario Augusto Bunge. Estado actual de la filosofía: un nuevo sistema filosófico. (Spanish) [Current state of philosophy: a new philosophical system]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament Contemporani., 1995. URL http://

hdl.handle.net/10256.1/593.

Bunge:1995:FBS

[Bun95h] Mario Augusto Bunge. Filosofía de la biología. (Spanish) [Philosophy of biology]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament Contemporani., 1995. URL http://hdl.

handle.net/10256.1/598.

Bunge:1995:FFS

[Bun95i] Mario Augusto Bunge. Filosofía de la física. (Spanish) [Philosophy of physics]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament Contemporani., 1995. URL http://hdl.

handle.net/10256.1/598.

Bunge:1995:FPS

[Bun95j] Mario Augusto Bunge. Filosofía de la psicología. (Spanish) [Philosophy of psychology]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament Contemporani., 1995. URL http:

//hdl.handle.net/10256.1/599.

Bunge:1995:FCS

[Bun95k] Mario Augusto Bunge. Filosofía de las ciencias sociales. (Spanish) [Philosophy of social sciences]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament Contemporani., 1995. URL

http://hdl.handle.net/10256.1/600.

Bunge:1995:GSG

[Bun95l] Mario Augusto Bunge. Gnoseología. (Spanish) [Gnoseology]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament Contemporani., 1995. URL http://hdl.handle.net/

10256.1/596.

Bunge:1995:OSO

[Bun95m] Mario Augusto Bunge. Ontología. (Spanish) [Ontology]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament

Contemporani., 1995. URL http://hdl.handle.net/10256.1/
594.

Bunge:1995:SSS

[Bun95n] Mario Augusto Bunge. Semántica. (Spanish) [Semantics]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament

Contemporani., 1995. URL http://hdl.handle.net/10256.1/
595.

Bunge:1995:TOC

[Bun95o] Mario Augusto Bunge. Teoría del orden y el caos como puente entre las ciencias y las humanidades. (Spanish) [Theory of order and chaos as a bridge between the sciences and the humanities]. Universitat de Girona, Spain. Càtedra Ferrater Mora de Pensament

Contemporani., 1995. URL http://hdl.handle.net/10256.1/

Bunge:1996:ECT

[Bun96a] Mario Bunge. Etica, Ciencia y Técnica. (Spanish) [Ethics, Science and Technology]. Editorial Sudamericana, Buenos Aires, Argentina, 1996. ISBN 950-07-1131-1. 181 pp. LCCN Q175.35 .B86 1996.

Bunge:1996:FPS

[Bun96b] Mario Bunge. Finding Philosophy in Social Science. Yale University Press, New Haven, CT, USA, 1996. ISBN 0-300-06606-6 (hardcover), 0-300-14623-X (e-book). xii + 432 pp. LCCN H61

.B793 1996. URL http://www.jstor.org/stable/10.2307/j.
ctt32br4n.

Bunge:1996:PICa

[Bun96c] Mario Bunge. In praise of intolerance to charlatanism in academia. Annals of the New York Academy of Sciences, 775: 96–115, June 1996. CODEN ANYAA9. ISSN 0077-8923 (print),

1749-6632 (electronic). URL https://ui.adsabs.harvard.
edu/#abs/1996NYASA.775...96B.

Bunge:1996:PICb

[Bun96d] Mario Bunge. In praise of intolerance to charlatanism in academia. In Gross et al. [GLL96], pages 96–115. ISBN 1-57331-002-6. LCCN Q175 .F58 1996. URL http://onlinelibrary.

wiley.com/doi/10.1111/nyas.1995.775.issue-1/issuetoc;
 https://nyaspubs.onlinelibrary.wiley.com/toc/17496632/

775/1. This volume represents the proceedings of a conference entitled The Flight from Science and Reason, which was sponsored by the New York Academy of Sciences and held in New York, New York on May 31–June 2, 1995.

Bunge:1996:IRS

[Bun96e] Mario Bunge. Intuición y Razón. (Spanish) [Intuition and Reason]. Biblioteca Mario Bunge. Editorial Sudamericana, Buenos Aires, Argentina, 1996. ISBN 950-07-1171-0. 227 pp. LCCN MLCS 2001/02130 (B).

Bunge:1996:IRF

[Bun96f] Mario Bunge. Intuition et raison. (French) [Intuition and reason]. Éditions Vigdor, Paris, France, 1996. ???? pp.

Bunge:1996:MBP

[Bun96g] Mario Bunge. Mind–body problem. In Beaumont et al. [BKR96], pages 488–492. ISBN 0-631-17896-1 (hardcover), 0-631-21435-6 (paperback). LCCN QP360 .B577 1996. URL http://www.

tandfonline.com/toc/rwhi20/.

Bunge:1996:SPP

[Bun96h] Mario Bunge. The seven pillars of Popper’s social philosophy. Philosophy of the Social Sciences = Philosophie des sciences sociales, 26(4):528–556, December 1996. ISSN 0048-3931 (print),

1552-7441 (electronic). URL https://journals.sagepub.com/
doi/pdf/10.1177/004839319602600405.

Bunge:1996:DTS

[Bun96i] Mario Augusto Bunge. Derecho como tecnica social de control y reforma. (Spanish) [Law as a social control and reform technique]. Fondo Editorial de la Facultad de Derecho de la Universidad de San Martin de Porres, Lima, Peru, 1996. 73 pp.

Bunge:1997:CTD

[Bun97a] Mario Bunge. Ciencia, Técnica y Desarrollo. (Spanish) [Science, Technology and Development]. Biblioteca Mario Bunge. Editorial Sudamericana, Buenos Aires, Argentina, expanded and corrected edition, 1997. ISBN 950-07-1240-7. 229 pp. LCCN Q180.A1 B76 1997.

Bunge:1997:ECA

[Bun97b] Mario Bunge. Epistemologia: Curso de actualizacion. (Spanish) [Epistemology: Course of update]. Siglo XXI, Coyoacan, México, DF, México, second edition, 1997. 252 pp.

Bunge:1997:FIH

[Bun97c] Mario Bunge. The freedom to inquire: a humanist’s doubts about the information revolution. Free Inquiry, 17(2):??, Spring 1997. CODEN FRINFM. ISSN 0272-0701. URL https://

secularhumanism.org/1997/03/a-humanists-doubts-about-
the-information-revolution/.

Bunge:1997:ME

[Bun97d] Mario Bunge. Mechanism and explanation. Philosophy of the Social Sciences = Philosophie des sciences sociales, 27(4):410– 465, ???? 1997. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:1997:MMF

[Bun97e] Mario Bunge. Moderate mathematical fictionism. In Agazzi and Darvas [AD97], pages 51–71. ISBN 0-7923-4343-3. LCCN QA8.6.

Bunge:1997:NLM

[Bun97f] Mario Bunge. A new look at moral realism. In Garzón Valdés et al. [GKvWZ97], pages 17–26. ISBN 3-428-08691-0 (hardcover). LCCN K235 .G27 1997.

Bunge:1997:CYF

[Bun97g] Mario Bunge. So close and yet so far: The Pope, evolution, and the soul. Free Inquiry, 17(1):??, Winter 1997. CODEN FRINFM.

ISSN 0272-0701. URL https://secularhumanism.org/1997/
01/the-pope-evolution-and-the-soul/.

Bunge:1997:VEP

[Bun97h] Mario Bunge. Vistas y Entrevistas: Propuestas Concretas sobre Problemas de Nuestro Tiempo. (Spanish) [Views and Interviews: Concrete Proposals on Problems of Our Time]. Biblioteca Mario Bunge. Editorial Sudamericana, Buenos Aires, Argentina, 1997. ISBN 950-07-1311-X. 306 pp. LCCN F1414 .B932 1997.

Bunge:1998:ECS

[Bun98a] Mario Bunge. Elogio de la Curiosidad. (Spanish) [Praise of Curiosity]. Biblioteca Mario Bunge. Editorial Sudamericana, Buenos Aires, Argentina, 1998. ISBN 950-07-1412-4. 253 pp. LCCN MLCS 2006/15583 (B).

Bunge:1998:PT

[Bun98b] Mario Bunge. The philosophical technologies. Technology in Society, 20(4):377–383, November 1998. CODEN ???? ISSN 0160-791x (print), 1879-3274 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0160791X98000232.

Bunge:1998:PSVa

[Bun98c] Mario Bunge. Philosophy of Science: Volume 1: From problem to theory, volume 1 of Science and technology studies. Transaction Publishers, New Brunswick, NJ, USA, revised edition, 1998. ISBN 0-7658-0413-1 (vol. 1: paperback), 0-7658-0415-8 (set: paperback). xv + 605 (vol. 1) pp. LCCN Q180.55.M4 B64 1998.

Bunge:1998:PSVb

[Bun98d] Mario Bunge. Philosophy of Science. Volume 2: From explanation to justification, volume 2 of Science and technology studies. Transaction Publishers, New Brunswick, NJ, USA, revised edition, 1998. ISBN 0-7658-0414-X (vol. 2: paperback), 0-7658-0415-8 (set: paperback). xiii + 423 (vol. 2) pp. LCCN Q180.55.M4 B64 1998.

Bunge:1998:SS

[Bun98e] Mario Bunge. Semiotic systems. In Altmann and Koch [AK98], pages 337–349. ISBN 3-11-015619-9, 3-11-080119-1 (e-book). LCCN H61 .S993 1998.

Bunge:1998:SSU

[Bun98f] Mario Bunge. Social Science under Debate: a Philosophical Perspective. University of Toronto Press, Toronto, ON, Canada, 1998. ISBN 0-8020-4298-8 (hardcover), 0-8020-8357-9 (paperback), 1-4426-8003-2 (e-book). xvi + 538 pp. LCCN H61

.B794 1998. URL http://www.jstor.org/stable/10.3138/
9781442680036.

Bunge:1998:CSN

[Bun98g] Mario Bunge. Une caricature de la science: la nouvelle sociologie de la science. (French) [A caricature of science: the new sociology of science]. Éditions Vigdor, Paris, France, 1998. ???? pp. URL

http://www.vigdor.com/.

Bunge:1998:CTD

[Bun98h] Mario Augusto Bunge. Ciencia, tecnica y desarrollo. (Spanish) [Science, technology and development]. Editorial Sudamericana, Buenos Aires, Argentina, 1998. ???? pp.

Bunge:1998:CSD

[Bun98i] Mario Augusto Bunge. Las ciencias sociales en discusión: una perspectiva filosófica. (Spanish) [Social sciences in discussion: a philosophical perspective]. Sudamericana, Buenos Aires, Argentina, 1998. ISBN 950-07-1566-X (paperback). 573 pp. LCCN H62 .B8518 1999.

Bunge:1998:SCS

[Bun98j] Mario Augusto Bunge. Sociología de la ciencia. (Spanish) [Sociology of science]. Editorial Sudamericana, Buenos Aires, Argentina, 1998. 171 pp.

Bunge:1999:BRC

[Bun99a] Mario Bunge. Book review: The Crisis of Global Capitalism [Open Society Endangered]: George Soros; New York, Public Affairs, 1998, xxx + 247 pages, (US $26.00). Journal of Socio-Economics, 28(4):533–536, 1999. CODEN JSECFK. ISSN 1053-5357 (print), 1879-1239 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1053535799000438.

Bunge:1999:BRS

[Bun99b] Mario Bunge. Book review: The Sociology of Philosophies: a Global Theory of Intellectual Change by Randall Collins. Contemporary Sociology — a Journal of Reviews, 28(3):280–281, May 1999. ISSN 0094-3061 (print), 1939-8638 (electronic). URL http:

//www.jstor.org/stable/2654141.

Bunge:1999:CAS

[Bun99c] Mario Bunge, editor. Critical Approach to Science and Philosophy: Essays in Honour of Karl Popper. Science and technology studies. Transaction Publishers, New Brunswick, NJ, USA, 1999. ISBN 0-7658-0427-1 (paperback), 1-351-31306-1 (e-book). xvii + 480 pp. LCCN B29 .C695 1999.

Bunge:1999:DP

[Bun99d] Mario Bunge. Dictionary of Philosophy. Prometheus Books, Amherst, NY, USA, 1999. ISBN 1-57392-257-9. 316 pp. LCCN B791 .B76 1999.

Bunge:1999:ES

[Bun99e]  Mario Bunge. The end of science?  In Slottje [Slo99],

pages 293–300. ISBN 3-7908-1226-9 (hardcover), 3-642-93643-1 (print), 3-642-93641-5 (e-book). LCCN HB523 .A3 19994.

URL https://link.springer.com/chapter/10.1007/978-3-
642-93641-8_14.

Bunge:1999:EPT

[Bun99f] Mario Bunge. Ethics and praxiology as technologies. In Agazzi and Lenk [AL99], pages 375–379.

Bunge:1999:HBS

[Bun99g] Mario Bunge. The human brain and science. Free Inquiry, 19(2): 17, ???? 1999. CODEN FRINFM. ISSN 0272-0701.

Bunge:1999:LP

[Bun99h] Mario Bunge. Linguistics and philosophy. In Wiegand [Wie99], pages 269–293. ISBN 3-11-015579-6 (hardcover), 3-11-080132-9, 3-11-174876-6. LCCN Q226. URL http://www.degruyter.com/

doi/book/10.1515/9783110801323.

Bunge:1999:PT

[Bun99i] Mario Bunge. Phenomenological theories. In Critical Approach to Science and Philosophy: Essays in Honour of Karl Popper [Bun99c], pages 234–254. ISBN 0-7658-0427-1 (paperback), 1-351-31306-1 (e-book). LCCN B29 .C695 1999.

Bunge:1999:SSF

[Bun99j] Mario Bunge. Sistemas sociales y filosofía. (Spanish) [Social systems and philosophy]. Editorial Sudamericana, Buenos Aires, Argentina, second edition, 1999. ISBN 950-07-1074-9 (paperback). 197 pp. LCCN HM131 .B76 1995.

Bunge:1999:SPC

[Bun99k] Mario Bunge. The Sociology–Philosophy Connection. Science and technology studies. Transaction, New Brunswick, NJ, USA, 1999. ISBN 1-56000-416-9. xix + 244 pp. LCCN HM585 .B85 1999.

Bunge:2000:ASE

[Bun00a] Mario Bunge. Absolute skepticism equals dogmatism. Free Inquiry, 24(4):34–36, ???? 2000. CODEN FRINFM. ISSN 0272-0701.

Bunge:2000:BRR

[Bun00b] Mario Bunge. Book review: Reenchanted Science: Holism in German Culture from Wilhelm II to Hitler. Philosophy of the Social Sciences = Philosophie des sciences sociales, 30(1):124– 126, 2000. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2000:EBP

[Bun00c] Mario Bunge. Energy: Between physics and metaphysics. Science & Education (Springer), 9(5):457–461, September 2000. CO-DEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

URL https://ui.adsabs.harvard.edu/#abs/2000Sc&Ed...9.
.459B.

Bunge:2000:PO

[Bun00d] Mario Bunge. Philosophy from the outside. Philosophy of the Social Sciences = Philosophie des sciences sociales, 30(2):227– 245, 2000. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2000:PIP

[Bun00e] Mario Bunge. Physicians ignore philosophy at their risk — and ours. Facta Philosophica, 2(??):149–160, ???? 2000. ISSN 1424-0602 (print), 2296-4231 (electronic).

Bunge:2000:SAI

[Bun00f] Mario Bunge. Systemism: the alternative to individualism and holism. Journal of Socio-Economics, 29(2):147–157, 2000. CO-DEN JSECFK. ISSN 1053-5357 (print), 1879-1239 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1053535700000585.

Bunge:2000:TMI

[Bun00g] Mario Bunge. Ten modes of individualism — none of which works — and their alternatives. Philosophy of the Social Sciences = Philosophie des sciences sociales, 30(3):384–406, 2000. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2000:ICS

[Bun00h] Mario Augusto Bunge. La investigación científica. (Spanish) [Scientific research]. Siglo xxi, México, DF, México, 2000. ???? pp.

Bunge:2001:APS

[Bun01a] Mario Bunge. Analytic philosophy of society and social science: The systemic approach as an alternative to holism and individualism (1988). In Mahner [Mah01], chapter 21, pages 291–302. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:BRS

[Bun01b] Mario Bunge. Book review: Le sens des valeurs. Philosophy of the Social Sciences = Philosophie des sciences sociales, 31(1): 122–125, 2001. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2001:EM

[Bun01c] Mario Bunge. Emergence and the mind (1977). In Mahner [Mah01], chapter 5, pages 75–91. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:EBP

[Bun01d] Mario Bunge. Energy: Between physics and metaphysics (2000). In Mahner [Mah01], chapter 3, pages 51–56. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:ES

[Bun01e] Mario Bunge. Epistemology of sociology. In Smelser and Baltes [SB01], pages 14569–14574. ISBN 0-08-043076-7.

Bunge:2001:EC

[Bun01f] Mario Bunge. Explaining creativity (1993). In Mahner [Mah01], chapter 20, pages 281–288. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:MNB

[Bun01g] Mario Bunge. From mindless neuroscience and brainless psychology to neuropsychology (1985). In Mahner [Mah01], chapter 19, pages 265–280. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:GBM

[Bun01h] Mario Bunge. The gap between mathematics and reality (1994). In Mahner [Mah01], chapter 14, pages 204–210. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:GCC

[Bun01i] Mario Bunge. The GST challenge to the classical philosophies of science (1977). In Mahner [Mah01], chapter 10, pages 153–166. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:HVS

[Bun01j] Mario Bunge. Hidden variables, separability, and realism (1995). In Mahner [Mah01], chapter 17, pages 244–255. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:HDR

[Bun01k] Mario Bunge. How do realism, materialism, and dialectics fare in contemporary science? (1973). In Mahner [Mah01], chapter 1, pages 27–41. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:IS

[Bun01l] Mario Bunge. Induction in science (1963). In Mahner [Mah01], chapter 9, pages 142–152. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:MMF

[Bun01m] Mario Bunge. Moderate mathematical fictionism (1997). In Mahner [Mah01], chapter 13, pages 187–203. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:MBL

[Bun01n] Mario Bunge. Morality is the basis of legal and political legitimacy (1992). In Mahner [Mah01], chapter 29, pages 397–405. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:NAS

[Bun01o] Mario Bunge. The nature of applied science and technology (1988). In Mahner [Mah01], chapter 24, pages 345–351. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:NDB

[Bun01p] Mario Bunge. New dialogues between Hylas and Philonous (1954). In Mahner [Mah01], chapter 2, pages 42–50. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:NLM

[Bun01q] Mario Bunge. A new look at moral realism (1993). In Mahner [Mah01], chapter 27, pages 377–385. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:MPS

[Bun01r] Mario Bunge. On method in the philosophy of science (1973). In Mahner [Mah01], chapter 8, pages 121–141. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:PCN

[Bun01s] Mario Bunge. Philosophy in Crisis: the Need for Reconstruction. Prometheus lecture series. Prometheus Books, Amherst, NY, USA, 2001. ISBN 1-57392-843-7 (hardcover). 245 pp. LCCN B1034.B83 P45 2001.

Bunge:2001:PRP

[Bun01t] Mario Bunge. Physical relativity and philosophy (1979). In Mahner [Mah01], chapter 16, pages 229–243. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:PUW

[Bun01u] Mario Bunge. Popper’s unworldly world 3 (1981). In Mahner [Mah01], chapter 7, pages 103–119. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:PLR

[Bun01v] Mario Bunge. The power and limits of reduction (1991). In Mahner [Mah01], chapter 11, pages 167–180. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:RCT

[Bun01w] Mario Bunge. Rational choice theory: a critical look at its foundations (1995). In Mahner [Mah01], chapter 22, pages 303–319. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:RAS

[Bun01x] Mario Bunge. Realism and antirealism in social science (1993). In Mahner [Mah01], chapter 23, pages 320–342. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:RCb

[Bun01y] Mario Bunge. Relativism: Cognitive. In Smelser and Baltes [SB01], pages 13009–13012. ISBN 0-08-043076-7.

Bunge:2001:RCa

[Bun01z] Mario Bunge. The revival of causality (1982). In Mahner [Mah01], chapter 4, pages 57–74. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:RID

[Bun01-27] Mario Bunge. Rights imply duties (1999). In Mahner [Mah01], chapter 28, pages 386–394. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:SCD

[Bun01-28] Mario Bunge. Schrödinger’s cat is dead (1999). In Mahner [Mah01], chapter 18, pages 256–262. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:SC

[Bun01-29] Mario Bunge. The status of concepts (1981). In Mahner [Mah01], chapter 6, pages 92–102. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:SER

[Bun01-30] Mario Bunge. Systems and emergence, rationality and imprecision, free-wheeling and evidence, science and ideology: Social science and its philosophy according to van den Berg. Philosophy of the Social Sciences = Philosophie des sciences sociales, 31(3): 404–423, 2001. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2001:TAC

[Bun01-31] Mario Bunge. Technoholodemocracy: an alternative to capitalism and socialism (1994). In Mahner [Mah01], chapter 30, pages 406–412. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:TP

[Bun01-32] Mario Bunge. The technologies in philosophy (1999). In Mahner [Mah01], chapter 26, pages 365–374. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:TSP

[Bun01-33] Mario Bunge. The technology–science–philosophy triangle in its social context (1999). In Mahner [Mah01], chapter 25, pages 352–364. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:TM

[Bun01-34] Mario Bunge. Thinking in metaphors (1999). In Mahner [Mah01], chapter 12, pages 181–184. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:TMN

[Bun01-35] Mario Bunge. Tres mitos de nuestro tiempo: virtualidad, globalización, igualamiento. (Spanish) [Three myths of our time: virtuality, globalization, equalization]. Colección Honoris causa. Universidad Nacional del Litoral, Santa Fe, Argentina, 2001. ISBN 987-508-152-3. 57 pp. LCCN HN17.5 .B845 2001.

Bunge:2001:TFT

[Bun01-36] Mario Bunge. Two faces and three masks of probability (1988). In Mahner [Mah01], chapter 15, pages 211–226. ISBN 1-57392-892-5. LCCN Q175.32 .R42.

Bunge:2001:TUS

[Bun01-37] Mario Bunge. Two unification strategies in social science: Reduction and integration. In Agazzi and Faye [AF01], pages 141–152. ISBN 981-02-4791-5, 981-279-959-1 (e-book). LCCN

Q175.32.R47 A33 2000. URL http://ebooks.worldscinet.
com/ISBN/9789812799593/9789812799593.html.

Bunge:2001:RFR

[Bun01-38] Mario Augusto Bunge. Recuerdo de Francisco Romero. (Spanish) [Memory of Francisco Romero]. In Speroni and Alonso [SA01], pages 175–182. ISBN 987-97861-2-2. LCCN B1034.R64 S64 2001; D107 .L4.

Bunge:2002:I

[Bun02a] Mario Bunge. Introduction. In Robert K. Merton, editor, Teoría y estructura sociales. (Spanish) [Social theory and structure], pages 1–8. Fondo de Cultura Económica, México, DF, México, 2002.

Bunge:2002:SSH

[Bun02b] Mario Bunge. Ser, saber, hacer. (Spanish) [Be, know, do]. Paidos-Universidad Nacional Autonoma de México, México, México, 2002. ISBN 968-853-484-6. 156 pp. LCCN B67 B85 2002.

Bunge:2003:C

[Bun03a] Mario Bunge. Cápsulas. (Spanish) [Capsules]. Colección Libertad y cambio. Editorial Gedisa, Barcelona, Spain, 2003. ISBN 84-7432-952-3. 255 pp. LCCN MLCM 2007/42056 (B).

Bunge:2003:CPE

[Bun03b] Mario Bunge. Considérations d’un philosophe sur l’économique du néo- conservatisme (néo-libéralisme). (French) [Considerations of a philosopher on the economics of neo-conservatism (neoliberalism)]. J.-M. Tremblay, Chicoutimi, QC, Canada, 2003. ISBN 1-55442-572-7. ???? pp.

Bunge:2003:ECQ

[Bun03c] Mario Bunge. Emergence and Convergence: Qualitative Novelty and the Unity of Knowledge. Toronto studies in philosophy. University of Toronto Press, Toronto, ON, Canada, 2003. ISBN 0-8020-8860-0 (hardcover), 1-281-99234-8, 1-4426-2821-9 (paperback), 1-4426-7435-0 (e-book). xiii + 330 pp. LCCN Q175

.B8224 2003. URL http://www.jstor.org/stable/10.3138/
9781442674356.

Bunge:2003:IHS

[Bun03d] Mario Bunge. Interpretation and hypothesis in social studies. In Boudon et al. [BCD03], pages 20–40. ISBN 0-7619-7438-5 (hardcover), 1-4462-6360-6 (e-book). LCCN H62

.E82 2003. URL http://sk.sagepub.com/navigator/the-
european-tradition-in-qualitative-research.

Bunge:2003:PD

[Bun03e] Mario Bunge. Philosophical Dictionary. Prometheus Books, Amherst, NY, USA, 2003. ISBN 1-59102-037-9. 315 pp. LCCN B791 .B765 2003.

Bunge:2003:PST

[Bun03f] Mario Bunge. Philosophy of science and technology: a personal report. In Fløistad et al. [FvWK03], pages 245–272. ISBN 1-4020-1476-7 (hardcover), 90-247-2436-8 (set). LCCN B804 .C573 vol. 8.

Bunge:2003:QQB

[Bun03g] Mario Bunge. Quantons are quaint but basic and real, and the quantum theory explains much but not everything: Reply to my commentators. Science & Education (Springer), 12 (5–6):587–597, August 2003. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/2003Sc&Ed..12..587B. See [Bun03h].

Bunge:2003:TFC

[Bun03h] Mario Bunge. Twenty-five centuries of quantum physics: From Pythagoras to us, and from subjectivism to realism. Science & Education (Springer), 12(5–6):445–466, August 2003. CO-DEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

URL https://ui.adsabs.harvard.edu/#abs/2003Sc&Ed..12.

.445B. See reply [Bun03g].

Bunge:2003:VOT

[Bun03i] Mario Bunge. Velocity operators and time-energy relations in relativistic quantum mechanics. International Journal of Theoretical Physics, 42(1):135–142, 2003. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Bunge:2004:BRU

[Bun04a] Mario Bunge. Book reviews: Understanding early civilizations. Philosophy of the Social Sciences = Philosophie des sciences sociales, 34(4):588–590, 2004. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2004:CT

[Bun04b] Mario Bunge. The centrality of truth. In Right, Wrong and Science: the Ethical Dimensions of the Techno-scientific Enterprise [AD04], pages 233–241. ISBN 90-420-0919-5 (hardcover). LCCN BJ57 .A3513 2004.

Bunge:2004:CSM

[Bun04c] Mario Bunge. Clarifying some misunderstandings about social systems and their mechanisms. Philosophy of the Social Sciences = Philosophie des sciences sociales, 34(3):371–381, 2004. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2004:HDI

[Bun04d] Mario Bunge. How does it work?: The search for explanatory mechanisms. Philosophy of the Social Sciences = Philosophie des sciences sociales, 34(2):182–210, 2004. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2004:MHP

[Bun04e] Mario Bunge. Matérialisme et humanisme: pour surmonter la crise de la pensée. (French) [Materialism and Humanism: to Overcome the Crisis of Thought]. Liber, Montréal, QC, Canada, 2004. ISBN 2-89578-041-2 (paperback). 290 pp. LCCN B1034.B83 P4514 2004.

Bunge:2004:MHR

[Bun04f] Mario Bunge. Mitos, hechos y razones: cuatro estudios sociales. (Spanish) [Myths, Facts and Reasons: Four Social Studies]. Biblioteca Mario Bunge. Editorial Sudamericana, Buenos Aires, Argentina, 2004. ISBN 950-07-2503-7. 156 pp. LCCN H61 .B7935 2004.

Bunge:2004:PCD

[Bun04g] Mario Bunge. The pseudoscience concept, dispensable in professional practice, is required to evaluate research projects: A reply to Richard J. McNally. Scientific Review of Mental Health Practice, 2(2):111–114, Fall/Winter 2004. CODEN ???? ISSN

1538-4985. URL https://www.srmhp.org/0202/concept.html.

Bunge:2004:SAI

[Bun04h] Mario Bunge. Systemism: The alternative to individualism and holism. In Van den Berg [Van04], pages 109–123. ISBN 0-7658-0232-5. LCCN HM495 .S63 2004.

Bunge:2004:VNM

[Bun04i] Mario Bunge. Vers un nouveau matérialisme. (French) [Towards a new materialism]. In Jean Dubessy, Guillaume Lecointre, and Marc Silberstein, editors, Les matérialismes et leurs détracteurs. (French) [Materialism and their detractors], pages 75–80. Éditions Syllepse, Paris, France, 2004. ISBN 2-84797-092-4 (paperback). LCCN B825 .M3352 2004.

Bunge:2005:BRP

[Bun05a] Mario Bunge. Book review: Philip Kitcher: Science, Truth, and Deomocracy. Isis, 96(4):679–680, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/501414.

Bunge:2005:BRW

[Bun05b] Mario Bunge. Book review: Who rules in science? an opinionated guide to the. Philosophy of the Social Sciences = Philosophie des sciences sociales, 35(2):250–253, 2005. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2006:ILP

[Bun06a] Mario Bunge. 100 Ideas: el libro para pensar y discutir en el café. (Spanish) [100 Ideas: the book to think and discuss in a cafe]. Editorial Sudamericana, Buenos Aires, Argentina, 2006. ISBN 950-07-2762-5 (paperback). 255 pp. LCCN PN6145.S6 B86 2006.

Bunge:2006:CRS

[Bun06b] Mario Bunge. Chasing Reality: Strife over Realism. Toronto studies in philosophy. University of Toronto Press, Toronto, ON, Canada, 2006. ISBN 0-8020-9075-3 (hardcover), 1-4426-2822-7 (paperback), 1-281-99189-9, 1-4426-7285-4 (PDF). xiv + 342 pp.

LCCN B835 .B87 2006. URL ttp://www.jstor.org/stable/
10.3138/9781442672857.

Bunge:2006:ESG

[Bun06c] Mario Bunge. Enlightened solutions for global challenges. Free Inquiry, 26(2):??, February/March 2006. CODEN FRINFM.

ISSN 0272-0701. URL https://secularhumanism.org/2006/
02/enlightened-solutions-for-global-challenges/.

Bunge:2006:PDP

[Bun06d] Mario Bunge. La philosophie derrière la pseudoscience. (French) [The philosophy behind pseudoscience]. Science … et pseudosciences: cahiers bimestriels de l’Association française pour l’information scientifique, 275:36–42, April 22, 2006. ISSN

0982-4022.  URL https://www.pseudo-sciences.org/La-
philosophie-derriere-la-pseudoscience.

Bunge:2006:MS

[Bun06e] Mario Bunge. Matérialisme et sciences. (French) [Materialism and science]. Matière première (Syllepse), 1(??):251–262, 2006.

Bunge:2006:PBP

[Bun06f] Mario Bunge. The philosophy behind pseudoscience. The Skeptical Inquirer, 30(4):29–37, July/August 2006. ISSN 0194-

6730. URL https://skepticalinquirer.org/2006/07/the-
philosophy-behind-pseudoscience/.

Bunge:2006:SPC

[Bun06g] Mario Bunge. A systemic perspective on crime. In Wikström and Sampson [WS06], pages 8–30. ISBN 0-521-85707-4 (hardcover). LCCN HV6025 .E86 2006.

Bunge:2006:CIL

[Bun06h] Mario Augusto Bunge. Cien ideas : el libro para pensar y discutir en el café. (Spanish) [One hundred ideas: the book to think about and discuss in the cafe]. Editorial Sudamericana, Buenos Aires, Argentina, 2006. 255 pp.

Bunge:2007:BPM

[Bun07a] Mario Bunge. Blushing and the philosophy of mind. Journal of Physiology — Paris, 101(4):247–256, 2007. ISSN 0928-4257.

URL http://www.sciencedirect.com/science/article/pii/

S092842570700054X. The Evolution of Human Cognition and Neuroscience: a Dialogue between Scientists and Humanists.

Bunge:2007:BRD

[Bun07b] Mario Bunge. Book review: Dissecting the Social: On the Principles of Analytical Sociology by Peter Hedström. American journal of sociology, 113(1):258–260, July 2007. CODEN AJSOAR.

ISSN 0002-9602 (print), 1537-5390 (electronic). URL http://

www.jstor.org/stable/10.1086/520888.

Bunge:2007:BRG

[Bun07c] Mario Bunge. Book review: Gintis, H., Bowles, S., Boyd, R., & Fehr, E., eds. (2005). Moral Sentiments and Material Interests. Cambridge, MA: MIT Press. Pp. xii + 404. US $50. Philosophy of the Social Sciences = Philosophie des sciences sociales, 37(4): 543–547, 2007. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2007:DWP

[Bun07d] Mario Bunge. Did Weber practise the objectivity he preached? In McFalls [McF07], chapter 4, pages 117–134. ISBN 0-8020-9224-1.

LCCN H61 .M4294 2007. URL http://www.jstor.org/stable/
10.3138/j.ctt2tv3d6.8.;   https://www.degruyter.com/
view/books/9781442684553/9781442684553-006/9781442684553-
006.xml.

Bunge:2007:ESS

[Bun07e] Mario Bunge. The ethics of science and the science of ethics. In Kurtz and Koepsell [KK07], chapter 2, pages 27–40. ISBN 1-59102-537-0. LCCN Q175.35 .S346 2007.

Bunge:2007:FRP

[Bun07f] Mario Augusto Bunge. Una filosofía realista para el nuevo milenio. Obras escogidas. Filosofía. Universidad Inca Garcilazo de la Vega, Fondo Editorial, Lima, Perú, second edition, 2007. ISBN 9972-888-68-1. 295 pp. LCCN B805 B86.2007.

Bunge:2008:BSP

[Bun08a] Mario Bunge. Bayesianism: Science or pseudoscience? International Review of Victimology, 15(2):165–178, September 2008. CODEN IRVIE2. ISSN 0269-7580 (print), 2047-9433

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/026975800801500206.

Bunge:2008:FS

[Bun08b] Mario Bunge. Filosofía y sociedad. (Spanish) [Philosophy and society]. Siglo XXI, México, México, 2008. ISBN 968-23-2729-6. 189 pp. LCCN H61.15 .B86 2008.

Bunge:2008:MSF

[Bun08c] Mario Bunge. Le matérialisme scientifique. (French) [Scientific materialism]. Éditions Syllepse, Paris, France, 2008. ISBN 2-84950-047-X, 2-84950-145-X (paperback). 220 pp. LCCN Q175 .B82714 2008.

Bunge:2008:P

[Bun08d] Mario Bunge. Préface. In Pierre Moessinger, editor, Voir la société: le micro et le macro. (French) [See society: the microphone and the macro], Société et pensées. (French) [Society and thoughts], pages 11–14. Hermann, Paris, France, 2008. ISBN 2-7056-6674-5. LCCN HM588 .M64 2008.

Bunge:2009:ALN

[Bun09a] Mario Bunge. Advantages and limits of naturalism. In Shook and Kurtz [SK09], chapter 3, pages 43–64. ISBN 1-59102-731-4 (hardcover). LCCN B828.2 .F88 2009.

Bunge:2009:CMS

[Bun09b] Mario Bunge. Causality and Modern Science. Transaction Publishers, New Brunswick, NJ, USA, fourth revised edition, 2009. ISBN 1-4128-0824-3 (paperback), 1-4128-1215-1 (e-book). xxxiv + 417 pp. LCCN BD541 .B85 2009.

Bunge:2009:ECA

[Bun09c] Mario Bunge. Epistemologia: Curso de actualizacion. (Spanish) [Epistemology: Course of update]. Siglo XXI, Coyoacan, México, DF, México, 2009. 252 pp.

Bunge:2009:TES

[Bun09d] Mario Bunge. ‘?Existió el socialismo alguna vez, y tiene porvenir?. (Spanish) [Was there socialism once, and has it a future?]. Mientras Tanto, ??(112):27–46, Otoño 2009. CODEN ???? ISSN 0210-8259 (print), 2014-7317 (electronic). URL http:

//www.jstor.org/stable/27821353.

Bunge:2009:FTB

[Bun09e] Mario Bunge. The failed theory behind the 2008 economic crisis. In Cherkaoui and Hamilton [CH09], pages 127–142. ISBN 1-905622-18-X (hardcover 4-volume set). LCCN HM585 .R395 2009.

Bunge:2009:PPF

[Bun09f] Mario Bunge. Political Philosophy: Fact, Fiction and Vision. Transaction Publishers, New Brunswick, NJ, USA, 2009. ISBN 1-4128-0828-6. x + 439 pp. LCCN JA71 .B834 2009. URL http:

//www.loc.gov/catdir/toc/ecip0820/2008025320.html.

Bunge:2009:SPC

[Bun09g] Mario Bunge. A systemic perspective on crime. In Sampson and Wikström [SW09], chapter 1, pages 8–30. ISBN 0-521-11905-7 (paperback). LCCN HV6117 .E96 2009. URL http://www.gbv.

de/dms/bowker/toc/9780521119054.pdf.

Bunge:2010:AO

[Bun10a] Mario Bunge. Appendix a: Objects. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 14, pages 267–275. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_14.

Bunge:2010:ABT

[Bun10b] Mario Bunge. Appendix b: Truths. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 15, pages 277–285. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_15.

Bunge:2010:BCH

[Bun10c] Mario Bunge. Brain and computer: The hardware/software dualism. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 12, pages 227–237. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664

2010; Q175 .B73 v.287. URL https://link.springer.com/
chapter/10.1007/978-90-481-9225-0_12.

Bunge:2010:CMB

[Bun10d] Mario Bunge. Classical matter: Bodies and fields. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 2, pages 23–40. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_2.

Bunge:2010:CCF

[Bun10e] Mario Bunge. Cognition, consciousness, and free will. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 11, pages 201–225. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175

.B73 v.287. URL https://link.springer.com/chapter/10.
1007/978-90-481-9225-0_11.

Bunge:2010:EL

[Bun10f] Mario Bunge. Emergence and levels. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 5, pages 73–92. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_5.

Bunge:2010:PPB

[Bun10g] Mario Bunge. From philosophy to physics and back. In Nuccetelli et al. [NSB10], pages 525–539. ISBN 1-4051-7979-1 (hardcover). LCCN B1001 .C65 2010.

Bunge:2010:GCM

[Bun10h] Mario Bunge. General concept of matter: To be is to become. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 4, pages 61–72. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (ebook). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73

v.287. URL https://link.springer.com/chapter/10.1007/
978-90-481-9225-0_4.

Bunge:2010:KGB

[Bun10i] Mario Bunge. Knowledge: Genuine and bogus. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 13, pages 239– 264. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_13.

Bunge:2010:M

[Bun10j] Mario Bunge. Materialism. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 7, pages 121–140. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https://link.

springer.com/chapter/10.1007/978-90-481-9225-0_7.

Bunge:2010:MMPb

[Bun10k] Mario Bunge. Matter and Mind: a Philosophical Inquiry, volume 287 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. xviii + 319 pp. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https://link.springer.

com/book/10.1007/978-90-481-9225-0.

Bunge:2010:MBP

[Bun10l] Mario Bunge. The mind–body problem. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 8, pages 143–157. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_8.

Bunge:2010:MS

[Bun10m] Mario Bunge. Mind and society. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 10, pages 181–199. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_10.

Bunge:2010:MMPa

[Bun10n] Mario Bunge. Minding matter: The plastic brain. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 9, pages 159– 180. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_9.

Bunge:2010:N

[Bun10o] Mario Bunge. Naturalism. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 6, pages 93–119. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287. URL https://link.

springer.com/chapter/10.1007/978-90-481-9225-0_6.

Bunge:2010:PW

[Bun10p] Mario Bunge. Philosophy as worldview. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 1, pages 3– 22. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_1.

Bunge:2010:QMW

[Bun10q] Mario Bunge. Quantum matter: Weird but real. In Matter and Mind: a Philosophical Inquiry [Bun10k], chapter 3, pages 41–59. ISBN 90-481-9224-2 (hardcover), 90-481-9225-0 (e-book). ISSN 0068-0346. LCCN BD418.3 .B7664 2010; Q175 .B73 v.287.

URL https://link.springer.com/chapter/10.1007/978-90-
481-9225-0_3.

Bunge:2010:RMI

[Bun10r] Mario Bunge. Reading measuring instruments. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):85–93, ???? 2010. CODEN ???? ISSN

1913-0465. URL https://spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/article/
view/11725.

Bunge:2011:BRT

[Bun11a] Mario Bunge. Book review: Tristram Hunt, Marx’s General: The Revolutionary Life of Friedrich Engels, New York: Henry Holt and Co., 2009. 430 pp. $32.00 (hardback). Philosophy of the Social Sciences = Philosophie des sciences sociales, 41(3): 446–449, 2011. ISSN 0048-3931 (print), 1552-7441 (electronic).

Bunge:2011:KGB

[Bun11b] Mario Bunge. Knowledge: Genuine and bogus. Science & Education (Springer), 20(5–6):411–438, May 2011. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/2011Sc&Ed..20..411B.

Bunge:2011:P

[Bun11c] Mario Bunge. Provocaciones. (Spanish) [Provocations]. Edhasa, Buenos Aires, Argentina, 2011. ISBN 987-628-145-3. 165 pp.

LCCN B1034 .B83 2011. URL http://core.cambeiro.com.ar/
0-165494-0.pdf.

Bunge:2011:TUS

[Bun11d] Mario Bunge. Two unification strategies: Analysis or reduction, and synthesis or integration. In Symons et al. [SPT11], pages 145– 157. ISBN 94-007-0142-X (hardcover), 94-007-0143-8 (e-book). ISSN 4020-5034. LCCN B3309.N394 O886 2011; Q175 .O88 2011.

URL https://link.springer.com/chapter/10.1007/978-94-
007-0143-4_11.

Bunge:2012:BTD

[Bun12a] Mario Bunge. Bridging theories to data. In Evaluating Philosophies [Bun12j], chapter 13, pages 119–130. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https://link.

springer.com/chapter/10.1007/978-94-007-4408-0_13.

Bunge:2012:CFP

[Bun12b] Mario Bunge. Can functionalist psychology explain anything? In Evaluating Philosophies [Bun12j], chapter 17, pages 159– 165. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/chapter/10.1007/978-94-
007-4408-0_17.

Bunge:2012:CSE

[Bun12c] Mario Bunge. Can standard economic theory account for crises? In Evaluating Philosophies [Bun12j], chapter 8, pages 77–82. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/chapter/10.1007/978-94-
007-4408-0_8.

Bunge:2012:CL

[Bun12d] Mario Bunge. Climate and logic. In Evaluating Philosophies [Bun12j], chapter 5, pages 57–59. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https://link.springer.

com/chapter/10.1007/978-94-007-4408-0_5.

Bunge:2012:CVC

[Bun12e] Mario Bunge. Conclusion: Valuation criterion. In Evaluating Philosophies [Bun12j], chapter 20, pages 181–183. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.1007/978-94-007-4408-0_
20.

Bunge:2012:CTT

[Bun12f] Mario Bunge. The correspondence theory of truth. Semiotica, 2012(188):65–76, February 2012. CODEN ???? ISSN 0037-1998 (print), 1613-3692 (electronic). URL https://www.degruyter.

com/view/j/sem.2012.2012.issue-188/sem-2012-0004/sem-
2012-0004.xml.

Bunge:2012:DIL

[Bun12g] Mario Bunge. Does inductive logic work? In Evaluating Philosophies [Bun12j], chapter 12, pages 115–118. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https://link.

springer.com/chapter/10.1007/978-94-007-4408-0_12.

Bunge:2012:DQPa

[Bun12h] Mario Bunge. Does quantum physics refute realism, materialism and determinism? In Evaluating Philosophies [Bun12j], chapter 15, pages 139–149. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175

.B8223 2012. URL https://link.springer.com/chapter/10.
1007/978-94-007-4408-0_15.

Bunge:2012:DQPb

[Bun12i] Mario Bunge. Does quantum physics refute realism, materialism and determinism? Science & Education (Springer), 21 (10):1601–1610, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). URL https://ui.adsabs.

harvard.edu/#abs/2012Sc&Ed..21.1601B.

Bunge:2012:EP

[Bun12j] Mario Bunge. Evaluating Philosophies, volume 295 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. xiv + 200 pp. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL http://www.loc.gov/

catdir/enhancements/fy1506/2012941366-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1506/2012941366-
t.html;   https://link.springer.com/book/10.1007/978-
94-007-4408-0.

Bunge:2012:ESD

[Bun12k] Mario Bunge. Existence: Single or double? In Evaluating Philosophies [Bun12j], chapter 19, pages 173–179. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346.

LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.1007/978-94-007-4408-0_
19.

Bunge:2012:FPM

[Bun12l] Mario Bunge. Filosofía para médicos. Editorial Gedisa, Barcelona, Spain, 2012. ISBN 84-9784-740-7. 208 pp. LCCN R723.

Bunge:2012:IOM

[Bun12m] Mario Bunge. Informatics: One or multiple? In Evaluating Philosophies [Bun12j], chapter 6, pages 61–64. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.1007/978-94-007-4408-0_
6.

Bunge:2012:KPR

[Bun12n] Mario Bunge. Knowledge pyramids or rosettes? In Evaluating Philosophies [Bun12j], chapter 18, pages 167–171. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.1007/978-94-007-4408-0_
18.

Bunge:2012:MPP

[Bun12o] Mario Bunge. Marxist philosophy: Promise and reality. In Evaluating Philosophies [Bun12j], chapter 9, pages 83–93. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/chapter/10.1007/978-94-
007-4408-0_9.

Bunge:2012:MEP

[Bun12p] Mario Bunge. Matter and energy: Physical or metaphysical concepts? In Evaluating Philosophies [Bun12j], chapter 14, pages 131–138. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175

.B8223 2012. URL https://link.springer.com/chapter/10.
1007/978-94-007-4408-0_14.

Bunge:2012:PUD

[Bun12q]  Mario Bunge. Parallel universes? Digital physics?  In

Evaluating Philosophies [Bun12j], chapter 16, pages 151–157.

ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/chapter/10.1007/978-94-
007-4408-0_16.

Bunge:2012:PMS

[Bun12r] Mario Bunge. The philosophical matrix of scientific progress. In Evaluating Philosophies [Bun12j], chapter 2, pages 15–34. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/chapter/10.1007/978-94-
007-4408-0_2.

Bunge:2012:PP

[Bun12s] Mario Bunge. Philosophies and phobosophies. In Evaluating Philosophies [Bun12j], chapter 1, pages 3–13. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https://link.

springer.com/chapter/10.1007/978-94-007-4408-0_1.

Bunge:2012:RLJ

[Bun12t] Mario Bunge. Rules of law: Just and unjust. In Evaluating Philosophies [Bun12j], chapter 10, pages 95–102. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https:/

/link.springer.com/chapter/10.1007/978-94-007-4408-0_
10.

Bunge:2012:SPA

[Bun12u] Mario Bunge. Subjective probabilities: Admissible in science? In Evaluating Philosophies [Bun12j], chapter 11, pages 105– 113. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012.

URL https://link.springer.com/chapter/10.1007/978-94-
007-4408-0_11.

Bunge:2012:SM

[Bun12v] Mario Bunge. Systemics and materialism. In Evaluating Philosophies [Bun12j], chapter 3, pages 35–47. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https://link.

springer.com/chapter/10.1007/978-94-007-4408-0_3.

Bunge:2012:T

[Bun12w] Mario Bunge. Technoscience? In Evaluating Philosophies [Bun12j], chapter 4, pages 51–56. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175 .B8223 2012. URL https://link.springer.

com/chapter/10.1007/978-94-007-4408-0_4.

Bunge:2012:WWBa

[Bun12x] Mario Bunge. Wealth and well-being, economic growth, and integral development. International Journal of Health Services, 42(1):65–76, January 2012. CODEN IJUSC3. ISSN 0020-7314 (print), 1541-4469 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.2190/HS.42.1.g.

Bunge:2012:WWBb

[Bun12y] Mario Bunge. Wealth and well-being, economic growth and integral development. In Evaluating Philosophies [Bun12j], chapter 7, pages 65–76. ISBN 94-007-4407-2 (hardcover), 94-007-4408-0 (e-book). ISSN 0068-0346. LCCN Q174 .B67 v. 295; Q175

.B8223 2012. URL https://link.springer.com/chapter/10.
1007/978-94-007-4408-0_7.

Bunge:2012:FTO

[Bun12z] Mario Augusto Bunge. Filosofía de la tecnología y otros ensayos. (Spanish) [Philosophy of technology and other essays]. Universidad Inca Garcilaso de la Vega, Lima, Peru, 2012. ???? pp.

Bunge:2013:BTP

[Bun13a] Mario Bunge. Bruce Trigger and the philosophical matrix of scientific research. In Chrisomalis and Costopoulos [CC13], chapter 11, pages 143–159. ISBN 1-4426-6455-X, 1-4426-6455-X (ebook). LCCN CC115.T74 .H863 2013. URL https://www.

degruyter.com/view/books/9781442664555/9781442664555-
014/9781442664555-014.xml.

Bunge:2013:DSF

[Bun13b] Mario Bunge, editor. Delaware Seminar in the Foundations of Physics, volume 1 of Studies in the Foundations, Methodology and Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-86102-4, 3-642-86104-0. x + 193 pp. LCCN QC71 .D385. URL http:

//www.springerlink.com/content/978-3-642-86102-4.

Bunge:2013:CMF

[Bun13c] Mario Bunge. La ciencia: Su método y su filosofía. (Spanish) [Science: its methods and its philosophy], volume 1 of Biblioteca Bunge. Laetoli, Pamplona, Spain, 2013. ISBN 84-92422-59-9. 138 pp. LCCN MLCM 2016/42410 (Q).

Bunge:2013:PQR

[Bun13d] Mario Bunge. La physique quantique réfute-t-elle le réalisme, le matérialisme et le déterminisme?. (French) [Does quantum physics refute realism, materialism and determinism?]. In Matériaux philosophiques et scientifiques pour un matérialisme contemporain: Volume 1 [Sil13d], pages 417–437. ISBN 2-919694-51-0 (print), 2-919694-20-0 (e-book). LCCN B825 .M384 2016.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=5336199.

Bunge:2013:MC

[Bun13e] Mario Bunge. Materialismo y Ciencia. (Spanish) [Materialism and Science], volume 2 of Biblioteca Bunge. Laetoli, Pamplona, Spain, second edition, 2013. ISBN 84-92422-60-2. 206 pp. LCCN MLCM 2016/43196 (Q).

Bunge:2013:MME

[Bun13f] Mario Bunge. Mechanism and mechanical explanation. In Kaldis [Kal13], pages 590–591. ISBN 1-4129-8689-3 (hardcover: set), 1-4522-7604-8 (e-book), 1-4522-7605-6 (e-book), 1-78402-465-1. LCCN B63 .E53 2013. URL http://sk.

sagepub.com/reference/encyclopedia-of-philosophy-and-
the-social-sciences/n220.xml.

Bunge:2013:MPC

[Bun13g] Mario Bunge. Medical Philosophy: Conceptual Issues in Medicine. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2013. ISBN 981-4508-94-2 (softcover), 981-4508-96-9 (e-book). xvi + 270 pp. LCCN R723 .B83 2013.

Bunge:2013:PI

[Bun13h] Mario Bunge. Pseudociencia e Ideología. (Spanish) [Pseudoscience and ideology], volume 3 of Biblioteca Bunge. Laetoli, Pamplona, Spain, second edition, 2013. ISBN 84-92422-63-7. 313 pp. LCCN Q175.5 .B82 2013.

Bunge:2013:SPC

[Bun13i] Mario Bunge. The Sociology–Philosophy Connection. Transaction Publishers, New Brunswick, NJ, USA, 2013. ISBN 1-4128-4965-9. xix + 244 pp. LCCN HM585 .B85 2013.

Bunge:2013:SCP

[Bun13j] Mario Bunge. The structure and content of a physical theory. In Delaware Seminar in the Foundations of Physics [Bun13b], pages 15–27. ISBN 3-642-86102-4, 3-642-86104-0. LCCN QC71

.D385. URL http://www.springerlink.com/content/978-3-
642-86102-4.

Bunge:2013:TBP

[Bun13k] Mario Bunge. Treatise on Basic Philosophy: Epistemology and Methodology II: Understanding the World, volume 6 of Treatise on Basic Philosophy Series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-015-6921-5. xii + 296 pp. LCCN B67BD143-237. URL http://link.

springer.com/10.1007/978-94-015-6921-7.

Bunge:2014:BQC

[Bun14a] Mario Bunge. Big questions come in bundles, hence they should be tackled systematically. International Journal of Health Services, 44(4):835–844, October 2014. CODEN IJUSC3. ISSN 0020-7314 (print), 1541-4469 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.2190/HS.44.4.i.

Bunge:2014:CRS

[Bun14b] Mario Bunge. Chasing Reality: Strife over Realism. Toronto studies in philosophy. University of Toronto Press, Toronto, ON, Canada, 2014. ISBN 1-4426-2822-7 (paperback). xiv + 342 pp.

LCCN B835 .B87 2014. URL http://www.degruyter.com/doi/
book/10.3138/9781442672857.

Bunge:2014:CTD

[Bun14c] Mario Bunge. Ciencia, Técnica y Desarrollo. (Spanish) [Science, Technology and Development], volume 4 of Biblioteca Bunge. Laetoli, Pamplona, Spain, third revised edition, 2014. ISBN 84-92422-64-5. 154 pp. LCCN Q180.A1 B76 2014.

Bunge:2014:MBU

[Bun14d] Mario Bunge. Mario Bunge en la Universidad Nacional de La Plata. Radio program audio recording, 2014. URL http://hdl.

handle.net/10915/36242.

Bunge:2014:MEM

[Bun14e] Mario Bunge. Memorias: Entre Dos Mundos. (Spanish) [Memories: Between Two Worlds]. Gedisa Editorial, Barcelona, Spain, 2014. ISBN 84-9784-895-0. 398 pp. LCCN B1034.B864 A3 2014.

URL http://core.cambeiro.com.ar/0-176685-0.pdf.

Bunge:2014:EF

[Bun14f] Mario Augusto Bunge. Evaluando filosofías. Gedisa, Barcelona, Spain and Buenos Aires, Argentina, 2014. ???? pp.

Bunge:2014:PVT

[Bun14g] Mario Augusto Bunge. Las Pseudociencias ‘!vaya timo!. (Spanish) [Pseudosciences: Go scam!], volume 5 of Biblioteca Bunge. Laetoli, Pamplona, Spain, second edition, 2014. ISBN 84-92422-70-X. 250 pp. LCCN Q172.5.P77 B86 2014.

Bunge:2014:SIS

[Bun14h] Mario Augusto Bunge. Seudociencia e ideología. (Spanish) [Pseudoscience and ideology]. Laetoli, Pamplona, Spain, 2014. ???? pp.

Bunge:2015:DAB

[Bun15a] Mario Bunge. Does the Aharonov–Bohm effect occur? Foundations of Science, 20(2):129–133, June 2015. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://

link.springer.com/article/10.1007/s10699-014-9354-2.

Bunge:2015:DS

[Bun15b] Mario Bunge. In defense of scientism. Free Inquiry, 35 (1):24–28, January 2015. CODEN FRINFM. ISSN 0272-

0701. URL https://secularhumanism.org/2014/11/cont-in-
defense-of-scientism/.

Bunge:2015:P

[Bun15c] Mario Bunge. Preface. In Scientific Controversies: a Sociohistorical Perspective on the Advancement of Science [Ray15], pages xvii–xxii. ISBN 1-4128-5571-3 (hardcover), 1-315-12903-5 (e-book), 1-351-49178-4, 1-351-49179-2, 1-351-49180-6. LCCN

Q175.5 .R3813 2015. URL http://www.tandfebooks.com/
isbn/9781315129037.

Bunge:2015:SE

[Bun15d] Mario Bunge. Sociology, epistemology of. In James D. Wright, editor, International Encyclopedia of the Social & Behavioral Sciences (Second Edition), pages 984–988. Elsevier, Amsterdam, The Netherlands, second edition, 2015. ISBN 0-08-097087-7.

URL http://www.sciencedirect.com/science/article/pii/
B9780080970868321468.

Bunge:2015:SAC

[Bun15e] Mario Augusto Bunge. A systemic approach to the climate change challenge. Internal document, Academia Argentina de Ciencias Exactas, Físicas y Naturales, Buenos Aires, Argentina, 2015.

Bunge:2016:BTW

[Bun16a] Mario Bunge. Between two worlds — memoirs of a philosopher– scientist. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2016. ISBN 3-319-29250-1 (hardcover), 3-319-29251-X (e-book). xviii + 496 pp. LCCN B1034.B864

B48 2016. URL http://www.springerlink.com/content/978-

3-319-29251-9. With an Appendix by Marta Bunge: My Life with Mario.

Bunge:2016:DMM

[Bun16b] Mario Bunge. Entre deux mondes: mémoires d’un philosophe– scientifique. (French) [Between two worlds: memories of a philosopher–scientist]. Éditions Matériologiques, Paris, France, 2016. ISBN 2-37361-073-6. 656 pp. LCCN ????

Bunge:2017:DSS

[Bun17a] Mario Bunge. Democracias y Socialismos. (Spanish) [Democracies and Socialisms], volume 11 of Biblioteca Bunge. Laetoli, Pamplona, Spain, 2017. ISBN 84-946742-5-0. 80 pp. LCCN JC423 .B8648 2017.

Bunge:2017:DSL

[Bun17b] Mario Bunge. Doing Science: in the Light of Philosophy. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2017. ISBN 981-320-276-9 (hardcover), 981-320-277-7 (paperback), 981-320-278-5 (e-book). xv + 225 pp. LCCN Q175 .B82228 2017.

Bunge:2017:ESR

[Bun17c] Mario Bunge. Evaluating scientific research projects: The units of science in the making. Foundations of Science, 22(3):455– 469, September 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-015-9474-3.

Bunge:2017:RCM

[Bun17d] Mario Bunge. O realismo científico de Mario Bunge. (Spanish) [On Mario Bunge’s scientific realism]. Revista de Filosofia Aurora, 29(46):353, April 17, 2017.

Bunge:2017:WA

[Bun17e] Mario Bunge. Why axiomatize? Foundations of Science, 22(4): 695–707, December 2017. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-016-9493-8.

Bunge:2017:WDS

[Bun17f] Mario Bunge. Why don’t scientists respect philosophers? In Encouraging Openness: Essays for Joseph Agassi on the Occasion of His 90th Birthday [BAG17], chapter 1, pages 3– 12. ISBN 3-319-57668-2 (hardcover), 3-319-57669-0 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN B833.E53

2017. URL https://link.springer.com/chapter/10.1007/
978-3-319-57669-5_1.

Bunge:2017:PCS

[Bun17g] Mario Augusto Bunge. El planteamiento científico. (Spanish) [The scientific approach]. Revista Cubana de Salud Pública., 43(3):1–29, September 2017. ISSN 0864-3466 (print), 1561-

3127 (electronic). URL http://scielo.sld.cu/scielo.php?
script=sci_arttext&pid=S0864-34662017000300016.

Bunge:2018:BTW

[Bun18a] Mario Bunge. Between Two Worlds: Memoirs of a Philosopher– Scientist. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-80522-3 (hardcover), 3-319-80522-3 (paperback), 3-319-29251-X (e-book). xviii + 496 pp. LCCN ????

Bunge:2018:GWS

[Bun18b] Mario Bunge. Gravitational waves and spacetime. Foundations of Science, 23(2):399–403, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http:/

/link.springer.com/article/10.1007/s10699-017-9526-y.

See comments [Rom18].

Bunge:2018:MHA

[Bun18c] Mario Bunge. Meeting Hans Albert. In Begegnungen Mit Hans Albert: Eine Hommage. (German) [Encounters With Hans Albert: a Homage] [Fra18], pages 161–164. ISBN 3-658-22689-7 (hardcover), 3-658-22690-0. LCCN B833 .B44 2019. URL http:

//www.springerlink.com/content/978-3-658-22690-9.

Bunge:2018:SSU

[Bun18d] Mario Bunge. The Scientific Stance: Using Evidence Based Arguments Across Fields. Hamilton Books, Lanham, MD, USA, 2018. ISBN 0-7618-6979-4 (paperback). ???? pp. LCCN ????

Bunge:2019:DST

[Bun19a] Mario Bunge. The dark side of technological progress. In Sassower and Laor [SL19], pages 109–113. ISBN 3-319-90825-1 (hardcover), 3-319-90826-X (PDF e-book), 3-030-08108-7.

LCCN B833 .I473 2019. URL https://philpapers.org/rec/
BUNTDS;https://link.springer.com/chapter/10.1007/978-
3-319-90826-7_10.

Bunge:2019:DC

[Bun19b] Mario Bunge. The dematerialization crusade. The Skeptical Inquirer, 43(2):??, March/April 2019. ISSN 0194-

6730. URL https://skepticalinquirer.org/2019/03/the-
dematerialization-crusade/.

Bunge:2019:F

[Bun19c] Mario Bunge. Foreword. In Mario Bunge: a Centenary Festschrift [Mat19a], pages ix–xviii. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Bunge:2019:IP

[Bun19d] Mario Bunge. Inverse problems. Foundations of Science, 24(3): 483–525, September 2019. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-018-09577-1.

Bunge:2019:PMC

[Bun19e] Mario Bunge. Philosophie de la médecine: concepts et méthode. (French) [Philosophy of medicine: concepts and method]. Éditions Matériologiques, Paris, France, 2019. ISBN 2-37361-202-X. 272 pp.

Bunge:2019:PPA

[Bun19f] Mario Bunge. Promesas y peligros de los avances tecnológicos. (Spanish) [Promises and dangers of technological advances]. Trilogía Ciencia Tecnología Sociedad, 11(21):7–10, ???? 2019.

ISSN 2145-7778. URL https://www.redalyc.org/journal/
5343/534367788006/534367788006.pdf.

Bunge:2019:TPP

[Bun19g] Mario Bunge. Testing political philosophies. Philosophy of the Social Sciences = Philosophie des sciences sociales, 49(3): 232–239, June 2019. ISSN 0048-3931 (print), 1552-7441 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.
1177/0048393119835159.

Bunge:2020:CDC

[Bun20a] Mario Bunge. Criticism: Destructive and constructive. Metascience: Scientific General Discourse, 1(??):161–164, 2020. CO-

DEN ???? ISSN ???? URL https://philpapers.org/rec/
BUNCDA.

Bunge:2020:DPF

[Bun20b] Mario Bunge. Dictionnaire philosophique. (French) [Philosophical dictionary]. Éditions Matériologiques, Paris, France, 2020. ???? pp.

Bauer:1991:SCP

[BvL91] Eberhard Bauer and Walter v. Lucadou. A strawman called “psi” — or: What is Professor Bunge afraid of? New Ideas in Psychology, 9(2):157–162, 1991. ISSN 0732-118X (print), 1873-3522

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0732118X9190019I. Special issue on Mario Bunge

and Martin Mahner.

Bunge:2004:HCM

[BWM^{+}04] Mario Bunge, Gregorio Weinberg, Tomás Eloy Martinez, Guillermo Jaim Etcheverry, and Pablo Jacovkis. Honoris Causa: Manuel Sadosky en Sus Noventa Años. (Spanish) [Honoris Causa: Manuel Sadosky in His Ninety Years]. Libros del Zorzal, Buenos Aires, Argentina, 2004. ISBN 987-1081-47-2. 73 pp. LCCN MLCS 2007/46947 (Q).

Causey:1984:BRM

[Cau84] Robert L. Causey. Book review: The Mind-Body Problem: A Psychobiological Approach by Mario Bunge. Synthese, 60(3):459– 466, September 1984. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL https://www.jstor.org/stable/20116041.

Cauchy:1988:PCA

[Cau88] Venant Cauchy, editor. Philosophie et culture: actes du XVIIe congrès mondial de philosophie, [Montréal 21–27.8.1983]: Philosophy and culture, volume 2. Beffroi, Montréal, Québec, Canada, 1988. ISBN 2-920449-10-9 (vol. 1), 2-9801018-1-8 (vol. 2), 2-9801018-2-6 (vol. 3), 2-9801018-3-4 (vol. 4), 2-9801018-4-2 (vol. 5). LCCN B20 .W63 1983; HM101 .W88 1983.

Cavallo:2012:MBD

[Cav12] Andrew M. Cavallo. On Mario Bunge’s definition of system and system boundary. Science & Education (Springer), 21(10):1595– 1599, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Chrisomalis:2013:HEI

[CC13] Stephen Chrisomalis and André Costopoulos, editors. Human Expeditions: Inspired by Bruce Trigger. University of Toronto Press, Toronto, ON, Canada, 2013. ISBN 1-4426-6455-X, 1-4426-6455-X (e-book). xx + 295 pp. LCCN CC115.T74 .H863

2013. URL http://www.degruyter.com/doi/book/10.3138/
9781442664555;   http://www.jstor.org/stable/10.3138/
j.ctt5hjxk8.

Crespo:1990:MCS

[CCB^{+}90] Ricardo F. Crespo, Manuel Comesaña, Mario Augusto Bunge, Eleonora Baringoltz, and Eduardo R. Scarano. Metodología de las ciencias sociales : lógica, lenguaje y racionalidad. (Spanish)

[Methodology of the social sciences: logic, language and rationality]. Macchi, Buenos Aires, Argentina, 1990. ???? pp.

Cellerier:1990:PCR

[CD90] G. Cellerier and J.-J. Ducret. Psychology and computation: a response to Bunge. New Ideas in Psychology, 8(2):159– 175, 1990. ISSN 0732-118x (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0732118X9090006N. See [Bun90g] and reply [Bun90e].

Curilaf:2019:RTE

[CD19] Carolina I. García Curilaf and Guillermo M. Denegri. A reconstruction of the theory of ecology based on Mario Bunge’s mechanistic epistemology and systemic ontology. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 30, pages 537–553. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

CelaConde:2014:NDH

[CFF^{+}14] Camilo J. Cela Conde, Atahualpa Fernández Neto, Marly Fernández, Manuella María Fernández, and Lucrecia Burges. Naturaleza y dignidad humana. (Spanish) [Nature and human dignity]. In Denegri [Den14a], pages 121–142. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Cherkaoui:2009:RBL

[CH09] Mohamed Cherkaoui and Peter Hamilton, editors. Raymond Boudon: a life in sociology: essays in honour of Raymond Boudon. Bardwell Press, Oxford, UK, 2009. ISBN 1-905622-18-X (hardcover 4-volume set). ???? pp. LCCN HM585 .R395 2009.

Chapoulie:1969:TES

[Cha69] J.-M. Chapoulie. Un type d’explication en sociologie: les systèmes de variables en relations causales. (French) [A type of explanation in sociology: systems of variables in causal relationships]. Revue Française de Sociologie, 10(3):333–??, ???? 1969. ISSN 2271-7641.

Chavarriage:2006:PDC

[Cha06] John Alexánder Giraldo Chavarriage. 8. la polémica distinción entre ciencias sociales y ciencias naturales: La cuestión del estatuto científico, las bases methodológicas y la naturleza del objeto de estudio . (Spanish) [8. The controversial distinction between social sciences and natural sciences: The question of scientific status, the methodological bases and the nature of the object of study]. Cuadernos de epistemología, ??

(1):29–40, 2006. URL https://issuu.com/juanaguirre/docs/
cuadernos_de_epistemolog_a__n_mero_1.

Cohn:1981:TPE

[Coh81] Priscilla Cohn, editor. Transparencies: philosophical essays in honor of J. Ferrater Mora. Humanities Press, Atlantic Highlands, NJ, USA, 1981. ISBN 0-391-02361-6. xxiii + 235 pp. LCCN B29 .T75.

Collins:1999:BRS

[Col99] Randall Collins. Book review: Social Science under Debate: A Philosophical Perspective by Mario Bunge. Contemporary Sociology — a Journal of Reviews, 28(3):282–284, May 1999. ISSN 0094-3061 (print), 1939-8638 (electronic). URL https://www.

jstor.org/stable/2654142.

Collin:2001:BHC

[Col01] Finn Collin. Bunge and Hacking on constructivism. Philosophy of the Social Sciences = Philosophie des sciences sociales, 31(3): 424–453, 2001. ISSN 0048-3931 (print), 1552-7441 (electronic).

Cortes:1986:BRE

[Cor86] Fernando Cortés. Book review: Economía y filosofía (2 a. edición) by Mario Bunge, Raúl Prebisch. El Trimestre Económico, 53(211):637–641, July 1986. URL https://www.jstor.org/

stable/23396659.

Cordero:2012:MBS

[Cor12] Alberto Cordero. Mario Bunge’s scientific realism. Science & Education (Springer), 21(10):1419–1435, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Corcho:2014:EMB

[Cor14] Roger Corcho. Entrevista a Mario Bunge. (Spanish) [Interview with Mario Bunge]. Mètode: Revista de difusió de la investigació de la Universitat de Valencia, ??(84):80–84, ???? 2014.

Cordero:2019:MBS

[Cor19] Alberto Cordero. Mario Bunge’s scientific approach to realism. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 5, pages 83–100. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Crivos:2014:MBE

[Cri14] Marta Crivos. Mario Bunge y la etnografía. (Spanish) [Mario Bunge and ethnography]. In Denegri [Den14a], pages 143– 148. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL

http://core.cambeiro.com.ar/0-100933-7.pdf;   http://
www.loc.gov/catdir/toc/gc01/9789502323398.pdf.

Crivos:2019:BSA

[Cri19] Marta Crivos. Bunge and scientific anthropology. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 23, pages 389– 396. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Cupani:2019:BCA

[Cup19] Alberto Cupani. Bunge is correct about positivism, but less so about Marxism and hermeneutics. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 16, pages 273–285. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Dahlstrom:1991:NSM

[Dah91] Daniel O. Dahlstrom, editor. Nature and scientific method, volume 22 of Studies in philosophy and the history of philosophy. Catholic University of America Press, Washington, DC, USA, 1991. ISBN 0-8132-0723-1. xii + 328 pp. LCCN B21 .S78 vol.

22 Q175.3. URL http://www.jstor.org/stable/10.2307/j.
ctv176d3.

Delgado:1993:ISS

[DB93] Rafael Rodríguez Delgado and Bela H. Banathy, editors. International Systems Science Handbook. Systematic Publications, Madrid, Spain, 1993. ISBN 84-604-6236-6 (paperback). 512 pp. LCCN Q295 .I58 1993; TA168 .I57x 1993.

Domenech:2014:FSN

[DB14] Antoni Domenech and María Julia Bertomeu. Fiat scientia nec pereat mundus. In Denegri [Den14a], pages 149–160. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Morales:1994:IPH

[dCPM94] Patricia (Maria del Carmen Patricia) Morales, editor. Indigenous Peoples, Human Rights and Global Interdependence, volume 1 of Human rights, development and environment: series A. International Centre for Human and Public Affairs, Tilburg, The Netherlands, 1994. ISBN 90-802139-1-8. 191 pp. LCCN K3240.4 .I52 1994; K3247 I54.1994.

Dechaux:2010:BRV

[Déc10] Jean-Hugues Déchaux. Book review: Voir la société. Le micro et le macro by Pierre Moessinger, Mario Bunge. Revue française de sociologie, 51(4):760–763, October/December 2010. URL https:

//www.jstor.org/stable/40985539.

Deleporte:2012:SEV

[Del12] Pierre Deleporte. The systemist emergentist view of Mahner and Bunge on ‘species as individuals’: What use for science and education? Science & Education (Springer), 21(10):1535–1544, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Deleporte:2013:MSM

[Del13] P. Deleporte. Le matérialisme scientifique de Mario Bunge. (French) [The scientific materialism of Mario Bunge]. In Matériaux philosophiques et scientifiques pour un matérialisme contemporain: Volume 1 [Sil13d], pages 93–110. ISBN 2-919694-51-0 (print), 2-919694-20-0 (e-book). LCCN B825 .M384 2016.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=5336199.

Deleporte:2019:CME

[Del19] Pierre Deleporte. Contrasting materialisms: Engelsian dialectical and Bunge’s emergentist realism. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 6, pages 101–116. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Demeulenaere:1997:BRF

[Dem97] Pierre Demeulenaere. Book review: Finding Philosophy in Social Sciences by Mario Bunge. L’Année sociologique (1940/1948–), 47

(2):250–255, ???? 1997. URL https://www.jstor.org/stable/
27889532.

Denis:2003:MPP

[Den03] Andy Denis. Methodology and policy prescription in economic thought: a response to Mario Bunge. Journal of Socio-Economics, 32(2):219–226, 2003. CODEN JSECFK. ISSN 1053-5357 (print), 1879-1239 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1053535703000209.

Denegri:2014:EHM

[Den14a] Guillermo M. Denegri, editor. Ensayos en homenaje a Mario Bunge en su 95^{∘} ániversario: elogio de la sabiduría. (Spanish) [Essays in homage to Mario Bunge on his 95th anniversary: in praise of wisdom]. Homenaje. Eudeba, Buenos Aires, Argentina, 2014. ISBN 950-23-2339-4. 389 pp. LCCN

BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Denegri:2014:MBE

[Den14b] Guillermo M. Denegri. Mario Bunge o el elogio de la sabiduría (y la longevidad). (Spanish) [Mario Bunge or the praise of wisdom (and longevity)]. In Ensayos en homenaje a Mario Bunge en su 95^{∘} ániversario: elogio de la sabiduría. (Spanish) [Essays in homage to Mario Bunge on his 95th anniversary: in praise of wisdom] [Den14a], pages 15–26. ISBN 950-23-2339-4.

LCCN BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Denegri:2019:MBA

[Den19] Guillermo M. Denegri. Mario Bunge: Argentine’s universal thinker. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 2, pages 31–35. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

deFlorida:1978:EYS

[dF78] Francisco Alonso de Florida, editor. Ensayos de yatrofilosofia. (Spanish) [Yatrophilosophy essays]. Academia Nacional de Medicina y la Asociación Mexicana de Epistemologia, Cuidad de México, DF, México, 1978. xi + 144 pp.

deFelipe:2019:SMM

[dF19] Íñigo Ongay de Felipe. Is simplicity a myth? Mach and Bunge on the principle of parsimony. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 9, pages 163–180. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Diemer:1978:WPS

[Die78] Alwin Diemer, editor. 16. Weltkongress für Philosophie 1978: Sektions-Vorträge = 16ème Congrès mondial de philosophie 1978: résumés des sections = 16th World Congress of philosophy: section papers. P. Lang, Frankfurt am Main, 1978. ISBN 3-8204-5813-1.

Dingle:1960:BRC

[Din60] Herbert Dingle. Book review: Causality — The Place of the Causal Principle in Modern Science, By Mario Bunge. Massachusetts: Harvard University Press, London; Oxford University Press. Pp. xx + 380. Price 60s. Physics Bulletin, 11(5):145–146, May 1960. CODEN PHSBB4. ISSN 0031-9112.

Denegri:2000:TAF

[DM00] Guillermo M. Denegri and Gladys E. Martínez, editors. Tópicos actuales en filosofía de la ciencia: homenaje a Mario Bunge en su 80. aniversario. Universidad Nacional de Mar del Plata, Facultad de Ciencias Exactas y Naturales, Mar del Plata, Argentina, 2000. ISBN 950-9635-96-0. 265 pp. LCCN Q175.3 .T66 2000.

dePaz:2016:BRC

[dP16] María de Paz. Book review: Crítica de la nueva sociología de la ciencia by Mario Bunge. Teorema: Revista Internacional de Filosofía, 35(2):209–214, ???? 2016. CODEN ???? ISSN 0210-

1602. URL https://www.jstor.org/stable/44077451.

Drimer:1987:BRE

[Dri87] Roberto L. Drimer. Book review: Economía y Filosofía by Mario Bunge. Desarrollo Económico, 26(104):674–676, January 1987.

URL https://www.jstor.org/stable/3466938.

Droste:2015:TTG

[Dro15] Heinz W. Droste. Turn of the Tide — Gezeitenwechsel: Einführung in Mario Bunges exakte Philosophie. (German) [Introduction to Mario Bunge’s exact philosophy]. Alibri, Aschaffenburg, Germany, 2015. ISBN 3-86569-189-7 (paperback). 196

pp. LCCN ???? URL http://d-nb.info/107195878X/04.

Droste:2019:MBP

[Dro19] Heinz W. Droste. Mario Bunge as a public intellectual. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 4, pages 63– 80. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

delSolar:2019:MAE

[dSMdC19] Rafael González del Solar, Luis Marone, and Javier Lopez de Casenave. Mechanismic approaches to explanation in ecology. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 31, pages 555–573. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Ducasse:1960:BRB

[Duc60a] C. J. Ducasse. Book review: Causality. The Place of the Causal Principle in Modern Science by Mario Bunge. Isis, 51(1):88–90, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227610.

Ducasse:1960:BRC

[Duc60b] C. J. Ducasse. Book review: Causality. The Place of the Causal Principle in Modern Science by Mario Bunge. Isis, 51(1):88–90, March 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL https://www.jstor.org/stable/227610.

Eberlein:1988:TDE

[EB88] Gerald L. Eberlein and Hal Berghel, editors. Theory and Decision: Essays in Honor of Werner Leinfellner, volume 50 of Theory and decision library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1988. ISBN 90-277-2519-5 (hardcover). xix + 322 pp. LCCN H61 .T466 1988.

Echeverria:1985:BRE

[Ech85] Javier Echeverria. Book review: Elementos para una teoría formal del sistema social by Mario Bunge, F. Parra Luna. Theoria: An International Journal for Theory, History and Foundations of Science, 1(1):315–319, ???? 1985. ISSN 0495-4548 (print),

2171-679X (electronic). URL https://www.jstor.org/stable/
23912586.

Edgerton:1969:BRS

[Edg69] Stephenie G. Edgerton. Book review: Scientific Research II: The Search for Truth by Mario Bunge. History of Education Quarterly, 9(4):492–496, Winter 1969. URL https://www.jstor.

org/stable/367085.

Engels:1947:DNS

[Eng47] Federico Engels. Dialéctica de la naturaleza. (Spanish) [Dialectics of Nature]. Problemas, Buenos Aires, Argentina, 1947.

???? pp. URL http://catalogosuba.sisbi.uba.ar/vufind/
Record/201603170442212069.

Erlichson:1967:CMB

[Erl67] Herman Erlichson. Comment on Mario Bunge’s article. American Journal of Physics, 35(4):359–360, April 1967. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [Bun66b].

Espinoza:1985:BRS

[Esp85] Miguel Espinoza. Book review: Scientific Materialism by Mario Bunge. Revue philosophique de la France et de l’Étranger, ?? (4):546–548, October/December 1985. ISSN 0035-3833 (print),

2104-385X (electronic). URL https://www.jstor.org/stable/
41093943.

Espinoza:1986:BRT

[Esp86] Miguel Espinoza. Book review: Treatise on Basic Philosophy, vol. 7 : Epistemology and Methodology, III : ^{≪} Philosophy of Science and Technology ^{≫}, Part I : ^{≪} Formal and Physical Sciences ^{≫}, Part II : ^{≪} Life Science, Social Science and Technology ^{≫} by Mario Bunge. Revue philosophique de la France et de l’Étranger, ??(3):389–393, July/September 1986. ISSN 0035-3833 (print),

2104-385X (electronic). URL https://www.jstor.org/stable/
41094073.

Espinoza:1992:QCB

[Esp92] Miguel Espinoza. Les quatre causes: De Bunge à Aristote. (French) [The four causes: From Bunge to Aristotle]. Revue Philosophique de la France et de l’Étranger, 182(3):297–316, July/September 1992. ISSN 0035-3833 (print), 2104-385X (elec-

tronic). URL https://www.jstor.org/stable/41096717.

Espinoza:1994:RSM

[Esp94] M. Espinoza. Le réalisme scientifique: une métaphysique tronquée. (French) [Scientific realism: a truncated metaphysics]. Archives de Philosophie, 57(2):325–340, 1994.

Feather:1964:BRI

[Fea64] Norman Feather. Book review: Intuition and Science, by Mario Bunge. 142 pp. Prentice-Hall, Englewood Cliffs, NJ, 1962. Physics Today, 17(2):65–66, February 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/1964PhT....17b..65B.

Feyerabend:1991:EEG

[Fey91] Paul Feyerabend. It’s not easy to exorcize ghosts. New Ideas in Psychology, 9(2):181–186, 1991. ISSN 0732-118X (print),

1873-3522 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0732118X9190022E. Special issue on

Mario Bunge and Martin Mahner.

Fine:1970:BRQ

[Fin70] Arthur Fine. Book review: Quantum Theory and Reality by Mario Bunge. The Philosophical Review, 79(2):295–298, April

1970. URL https://www.jstor.org/stable/2183962.

Fleury:1988:LLF

[FJMT88] N. Fleury, S. Joffily, J. A. Martinho Simoes, and A. Troper, editors. Leite Lopes Festschrift: a Pioneer Physicist in the Third World. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1988. ISBN 9971-5-0693-9. xxvii + 643 pp. LCCN QC170 .L447 1988.

Floistad:1986:CPN

[Flø86] Guttorm Fløistad, editor. Contemporary Philosophy: a New Survey = La philosophie contemporaine: chroniques nouvelles, volume 2 of Philosophy of science = Philosophie des sciences. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1986. ISBN 90-247-2518-6 (hardcover), 90-247-3298-0 (paperback). ix + 550 pp. LCCN ????

Foster:1966:PCS

[FM66] Marguerite H. (Marguerite Herzik) Foster and Michael L. Martin, editors. Probability, Confirmation, and Simplicity: Readings in the Philosophy of Inductive Logic. Odyssey Press, New York, NY, USA, 1966. ix + 470 pp.

Bunge:2009:AMB

[Fonxx] Fondo Bunge. Archivo de Mario Bunge en la Facultad de Ciencias Exactas, Físicas y Naturales de la Universidad Nacional de Buenos Aires, Biblioteca, Ciudad Universitaria, Pabellón 2, Intendente Güiraldes 2160, Ciudad Autónoma de Buenos Aires, C1428EGA. Document archive, 20xx. URL https:/

/digital.bl.fcen.uba.ar/gsdl-282/cgi-bin/library.cgi?
p=FondoBunge.

Forman:1989:MAB

[For89a] Frank Forman. Mario Augusto Bunge and scientific metaphysics. In The Metaphysics of Liberty [For89b], pages 55–82. ISBN 0-7923-0080-7 (U.S.), 94-009-0901-2 (e-book), 94-010-6894-1. LCCN B824.4 .F67 1989.

Forman:1989:ML

[For89b] Frank Forman. The Metaphysics of Liberty, volume 6 of Theory and decision library. Series A, Philosophy and methodology of the social sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 0-7923-0080-7 (U.S.), 94-009-0901-2 (e-book), 94-010-6894-1. xi + 202 pp. LCCN B824.4 .F67 1989.

Forbes:2007:BRC

[For07] Daniel Forbes. Book review: Chasing Reality: Strife over Realism by Mario Bunge. Review of Metaphysics, 61(1):125–127, September 2007. CODEN ???? ISSN 1549-4853 (print), 0034-6632 (elec-

tronic). URL https://www.jstor.org/stable/20130898.

Franco:2018:BHA

[Fra18] Giuseppe Franco. Begegnungen Mit Hans Albert: Eine Hommage. (German) [Encounters With Hans Albert: a Homage]. Vieweg, Wiesbaden, Germany, 2018. ISBN 3-658-22689-7 (hardcover), 3-658-22690-0. xxvi + 385 pp. LCCN B833 .B44 2019. URL http:

//www.springerlink.com/content/978-3-658-22690-9.

Floistad:2003:CPN

[FvWK03] Guttorm Fløistad, G. H. von Wright, and Raymond Klibansky, editors. Contemporary philosophy: a new survey: Philosophy of Latin America, volume 8. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 2003. ISBN 1-4020-1476-7 (hardcover), 90-247-2436-8 (set). LCCN B804 .C573 vol. 8.

G:1973:BRT

[G.73] E. G. Book review: Teoría y realidad by Mario BUNGE, J. L. García Molina, J. Sempere. Teorema: Revista Internacional de Filosofía, 3(2–3):407–410, ???? 1973. CODEN ???? ISSN 0210-

1602. URL https://www.jstor.org/stable/43047680.

Gama:2014:SKT

[Gam14] Cipta Bakti Gama. Studi kritis terhadap teori identitas pikiran– otak Mario Bunge: Perspektif neo-Sadrian. (Indonesian) [Critical study of Mario Bunge’s theory of mind–brain identity: Neo-Sadrian perspective]. Kanz Philosophia: a Journal for Islamic Philosophy and Mysticism, 4(2):139–154, December 1, 2014.

Gillath:2005:ASD

[GBS^{+}05] Omri Gillath, Silvia A. Bunge, Phillip R. Shaver, Carter Wendelken, and Mario Mikulincer. Attachment-style differences in the ability to suppress negative thoughts: Exploring the neural correlates. NeuroImage, 28(4):835–847, 2005. CO-DEN NEIMEF. ISSN 1053-8119 (print), 1095-9572 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S1053811905004556. Special Section: Social Cognitive Neuroscience.

Gopnik:1986:MMS

[GG86] Irwin Gopnik and Myrna Gopnik, editors. From Models to Modules: Studies in Cognitive Science, Theoretical issues in cognitive science. Ablex, Norwood, NJ, USA, 1986. ISBN 0-89391-355-3. LCCN BF311 .F75 1986.

Gil:2019:BRN

[Gil19] José María Gil. Bunge’s requirement of neurological plausibility for a linguistic theory. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 38, pages 717–733. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

GarzonValdes:1997:NSL

[GKvWZ97] Ernesto Garzón Valdés, Werner Krawietz, George Henrik von Wright, and Ruth Zimmerling, editors. Normative Systems in Legal and Moral Theory: Festschrift for Carlos E. Alchourrón and Eugenio Bulygin. Duncker and Humblot, Berlin, Germany, 1997. ISBN 3-428-08691-0 (hardcover). 528 pp. LCCN K235 .G27 1997.

Gross:1996:FSR

[GLL96] Paul R. Gross, Norman Levitt, and Martin W. Lewis, editors. The Flight from Science and Reason, volume 775 of Annals of the New York Academy of Sciences. The New York Academy of Sciences, New York, NY, USA, June 24, 1996. ISBN 1-57331-002-6. xi + 593 pp. LCCN

Q175 .F58 1996. URL http://onlinelibrary.wiley.com/
doi/10.1111/nyas.1995.775.issue-1/issuetoc;   https:/
/nyaspubs.onlinelibrary.wiley.com/toc/17496632/775/1.

This volume represents the proceedings of a conference entitled The Flight from Science and Reason, which was sponsored by the New York Academy of Sciences and held in New York, New York on May 31–June 2, 1995.

GonzalezdelSolar:2014:EMS

[GML14] Rafael González del Solar, Luis Marone, and Javier López de Casenave. Dos enfoques mecanísmicos de la explicación en ecología. (Spanish) [Two mechanistic approaches to explanation in ecology]. In Denegri [Den14a], pages 185–208. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Gomez:2014:MBM

[Góm14] Ricardo Gómez. Mario Bunge y el mito de la interpretación de Copenhague. (Spanish) [Mario Bunge and the myth of the Copenhagen interpretation]. In Denegri [Den14a], pages 161–184. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Gotschl:1995:RCU

[Göt95] Johann Götschl, editor. Revolutionary Changes in Understanding Man and Society: Scopes and Limits, volume 21 of Theory and decision library. Series A, Philosophy and methodology of the social sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-3627-5 (hardcover). ISSN 0921-3384. 300 pp. LCCN H61 .R474 1995.

Grofier:1991:NTH

[GPA91] Ethel Grofier, Michel Paradis, and Konstantinos I. Arvanitakis, editors. The Notion of Tolerance and Human Rights: Essays in Honour of Raymond Klibansky. Carleton University Press, Ottawa, ON, Canada, 1991. ISBN 0-88629-164-X (hardcover), 0-88629-163-1 (paperback), 0-7735-8220-7 (e-book). 193 pp. LCCN

HM1271 .N685 1991. URL http://www.jstor.org/stable/10.
2307/j.ctt7zt3sb.

Aletheia:2005:CHM

[Gru05] Grupo Aletheia. Congreso — Homenaxe a Mario Bunge. Grupo Aletheia, Vigo, Spain, 2005. ???? pp.

Garcia-Sucre:1981:GQS

[GSB81] Máximo García-Sucre and Mario Bunge. Geometry of a quantal system. International Journal of Quantum Chemistry, 19(1):83– 93, January 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hajduk:1965:BRM

[Haj65] Zygmunt Hajduk. Book review: The Myth of Simplicity by Mario Bunge. Roczniki Filozoficzne / Annales de Philosophie / Annals of Philosophy, 13(3):114–115, ???? 1965. CODEN RFLZBF. ISSN 0035-7685. URL https://www.jstor.org/

stable/43406787.

Hajduk:1976:BRF

[Haj76] Zygmunt Hajduk. Book review: Foundations of Physics by Mario Bunge. Roczniki Filozoficzne / Annales de Philosophie / Annals of Philosophy, 24(3):122–123, ???? 1976. CODEN RFLZBF. ISSN 0035-7685. URL https://www.jstor.org/

stable/43407359.

Halloun:2019:CEB

[Hal19] Ibrahim A. Halloun. Cognition and education: a Bungean systemic perspective. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 37, pages 683–714. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Harre:1966:BRM

[Har66] R. Harré. Book review: The Myth of Simplicity by Mario Bunge. The Philosophical Quarterly (1950-), 16(62):85–86, Jan-

uary 1966. URL https://www.jstor.org/stable/2217899.

Harre:1975:BRM

[Har75] R. Harré. Book review: Method, Model and Matter by Mario Bunge. Isis, 66(2):264–266, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229426.

Hartnett:1977:SAT

[Har77] William E. Hartnett, editor. Systems: approaches, theories, applications: including the proceedings of the eighth George Hudson Symposium held at Plattsburgh, New York, April 11-012, 1975, volume 3 of Episteme. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1977. ISBN 90-277-0822-3. LCCN Q295 .G46 1975.

Harman:1991:EFS

[Har91] Willis Harman. The epistemological foundations of science reconsidered. New Ideas in Psychology, 9(2):187– 195, 1991. ISSN 0732-118X (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0732118X9190023F. Special issue on Mario Bunge and Martin Mahner.

Hagan:1994:CW

[HBB^{+}94] William J. Hagan, Jr., Henry H. Bauer, Mario Bunge, A. F. W. Coulson, and Bennett M. Berger. “Culture wars”. Science, 265 (5174):853–855, August 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/2884485.

Hinke-Dornemann:1988:PMW

[HD88] Ulrike Hinke-Dörnemann, editor. Die Philosophie in der modernen Welt: Gedenkschr. für Prof. Dr. med. Dr. phil. Alwin Diemer: 16. April 1920–25. Dez. 1986. Lang, Franfurt a.M., West Germany, 1988. ISBN 3-8204-7349-1 (set).

Hesse:1960:BRC

[Hes60] Mary B. Hesse. Book review: Causality–The Place of the Causal Principle in Modern Science by Mario Bunge. British Journal for the Philosophy of Science, 11(43):252–255, November 1960. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

URL https://www.jstor.org/stable/685315.

Haller:1975:PP

[HG75] Rudolf Haller and Johann Götschl. Philosophie und Physik. (German) [Philosophy and physics], volume 10 of Wissenschaftstheorie Wissenschaft und Philosophie. Vieweg+Teubner Verlag, Wiesbaden, Germany, 1975. ISBN 3-322-86073-6, 3-528-08335-2

(print). LCCN B1-5802. URL http://www.springerlink.com/
content/978-3-322-86073-6.

Harper:1976:FPT

[HH76] William L. Harper and C. A. (Clifford Alan) Hooker, editors. Foundations of Probability Theory, Statistical Inference, and Statistical Theories of Science: Volume III Foundations and Philosophy of Statistical Theories in the Physical Sciences, volume 6c of The University of Western Ontario Series in Philosophy of Science, A Series of Books on Philosophy of Science, Methodology, and Epistemology Published in Connection with The University of Western Ontario Philosophy of Science Programme. Springer Netherlands, Dordrecht, 1976. ISBN 90-277-0621-2, 94-010-1438-8 (e-book). ISSN 1566-659X. 253 pp. LCCN B67. URL http:/

/www.springerlink.com/content/978-94-010-1438-0.

Hinshaw:1961:BRB

[Hin61a] Virgil Hinshaw, Jr. Book review: Causality: The Place of the Causal Principle in Modern Science by Mario Bunge. Philosophy of Science, 28(2):218–222, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185814.

Hinshaw:1961:BRC

[Hin61b] Virgil Hinshaw, Jr. Book review: Causality: The Place of the Causal Principle in Modern Science by Mario Bunge. Philosophy of Science, 28(2):218–222, April 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

jstor.org/stable/185814.

Hinshaw:1964:BRBa

[Hin64] Virgil Hinshaw, Jr. Book review: Intuition and Science by Mario Bunge. Philosophy of Science, 31(2):183–184, April 1964. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185997.

Hindmarsh:1977:BKP

[Hin77] J. L. Hindmarsh. On the Bunge–Kálnay position operator for the Dirac electron. International Journal of Theoretical Physics, 16:443–445, 1977. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Hobson:2019:RAS

[Hob19] Art Hobson. A realist analysis of six controversial quantum issues. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 19, pages 329–348. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Hoffman:1984:BRS

[Hof84] John Hoffman. Book review: Scientific Materialism by Mario Bunge. Science & Society, 47(4):485–487, Winter, 1983/ 1984. ISSN 0036-8237 (print), 1943-2801 (electronic). URL https://

www.jstor.org/stable/40402524.

Ingthorsson:2019:MBC

[Ing19] Rögnvaldur D. Ingthorsson. Mario Bunge and the current revival of causal realism. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 12, pages 205–217. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Ivarola:2019:SML

[Iva19] Leonardo Ivarola. Social mechanisms and the logic of possibility trees. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 24, pages 397–414. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Jarvie:1995:CRM

[JL95] I. C. Jarvie and Nathaniel Laor. Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I, volume 161 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-4211-X, 94-011-0471-9 (ebook). ISSN 0068-0346. xxii + 266 pp. LCCN ???? URL https:

//link.springer.com/book/10.1007/978-94-011-0471-5.

Jodoin:2010:LIM

[Jod10] Laurent Jodoin. L’héritage intellectuel de Mario Bunge: entre science et philosophie. (French) [The intellectual legacy of Mario Bunge: between science and philosophy]. Philosophiques, 37(2): 439–455, ???? 2010.

Jodoin:2013:LRT

[Jod13] Laurent Jodoin. L’émergence et la réalité des états compatibles inobservables: le cas de l’entropie. (French) [the emergence and reality of unobservable compatible states: the case of entropy]. In Matériaux philosophiques et scientifiques pour un matérialisme contemporain. (French) [Philosophical and scientific materials for contemporary materialism], page ?? Éditions Matériologiques, Paris, France, 2013.

Johansson:2019:MMN

[Joh19] Ingvar Johansson. Mechanism models as necessary truths. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 14, pages 241–262. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Junior:2019:LNQ

[Jun19] Olival Freire Junior. On the legacy of a notable quantum dissident: David Bohm (1917–1992). In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 20, pages 349–360. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Kahle:2019:DA

[Kah19] Reinhard Kahle. Dual axiomatics. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 35, pages 633–642. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Kaidesoja:2009:BBS

[Kai09] Tuukka Kaidesoja. Bhaskar and Bunge on social emergence. Journal for the Theory of Social Behaviour, 39(3):300–322, September 2009. CODEN JTSBBM. ISSN 0021-8308 (print),

1468-5914 (electronic). URL https://onlinelibrary.wiley.
com/doi/10.1111/j.1468-5914.2009.00409.x.

Kaldis:2013:EPS

[Kal13] Byron Kaldis, editor. Encyclopedia of Philosophy and the Social Sciences. SAGE Publications, Inc., Thousand Oaks, CA, 2013. ISBN 1-4129-8689-3 (hardcover: set), 1-4522-7604-8 (ebook), 1-4522-7605-6 (e-book), 1-78402-465-1. xxxiii + 1145 (two volumes) pp. LCCN B63 .E53 2013. URL http://sk.

sagepub.com/reference/encyclopedia-of-philosophy-and-
the-social-sciences.

Kaldis:2019:EST

[Kal19] Byron Kaldis. Emergence, systems and technophilosophy. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 40, pages 751–772. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Kantor:2015:MMB

[Kan15] Isabel N. Kantor. Las memorias de Mario Bunge y su filosofía de la ciencia. (Spanish) [Mario Bunge’s memories and his philosophy of science]. Medicina, 75(2):104–108, April 1, 2015.

Kary:2019:EPPa

[Kar19a] Michael Kary. Ethical politics and political ethics I: Agathonism. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 27, pages 467–485. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Kary:2019:EPPb

[Kar19b] Michael Kary. Ethical politics and political ethics II: On socialism through integral democracy. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 29, pages 513–534. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Kary:2019:LLC

[Kar19c] Michael Kary. On leaving as little to chance as possible. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 34, pages 609– 631. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Kurtz:1989:BWC

[KFT89] Paul Kurtz, Levi Fragell, and Rob Tielman, editors. Building a World Community: Humanism in the 21st Century. Prometheus Books, Amherst, NY, USA, 1989. ISBN 0-87975-538-5. LCCN BL2747.6 .B85 1989.

Kilov:2004:REC

[Kil04] Haim Kilov. Review of “Emergence and Convergence: Qualitative Novelty and the Unity of Knowledge” by Mario Bunge, University of Toronto Press, 2004, ISBN 0-8020-8860-0. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33(4):88–90, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kirschenmann:1975:BRB

[Kir75a] Peter Kirschenmann. Book review: Philosophy of Physics by Mario Bunge. Philosophy of Science, 42(1):94–96, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187300.

Kirschenmann:1975:BRP

[Kir75b] Peter Kirschenmann. Book review: Philosophy of Physics by Mario Bunge. Philosophy of Science, 42(1):94–96, March 1975. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL https://www.jstor.org/stable/187300.

Kiwelekar:2010:OOM

[KJ10] Arvind W. Kiwelekar and Rushikesh K. Joshi. An object-oriented metamodel for Bunge–Wand–Weber ontology. arXiv.org, ??(??):

1–8, April 21, 2010. URL https://arxiv.org/abs/1004.3640.

Kurtz:2007:SEC

[KK07] Paul Kurtz and David Koepsell, editors. Science and ethics: can science help us make wise moral judgments? Prometheus Books, Amherst, NY, USA, 2007. ISBN 1-59102-537-0. 359 pp. LCCN Q175.35 .S346 2007.

Klassen:2003:SAN

[Kla03a] Michael Klassen. Systemtheoretische Ansätze von Niklas Luhmann und Mario Bunge im Vergleich und deren Rezeption in der Sozialen Arbeit. (German) [System Theoretical Approaches of Niklas Luhmann and Mario Bunge in Comparison and their Reception in Social Work]. Hochschulschrift, Technische Universität Berlin, Berlin, Germany, 2003. 299 pp.

URL https://search.proquest.com/pqdtglobal/docview/
305256785/;   https://tu-berlin.hosted.exlibrisgroup.
com/primo-explore/fulldisplay?docid=TUB_ALMA_DS21516727720002884.

Klassen:2003:WLS

[Kla03b] Michael Klassen. Was leisten Systemtheorien in der sozialen Arbeit ein Vergleich der systemischen Ansätze von Niklas Luhmann und Mario Bunge. (German) [How do system theories in social work compare the systemic approaches of Niklas Luhmann and Mario Bunge]. Haupt, Bern, Switzerland, 2003. ISBN 3-258-06658-2 (paperback). 298 pp. LCCN ????

Klassen:2004:WLS

[Kla04] Michael Klassen. Was leisten Systemtheorien in der Sozialen Arbeit? Ein Vergleich der systemischen Ansätze von Niklas Luhmann und Mario Bunge. (German) [What do systems theories do in social work? A comparison of the systemic approaches of Niklas Luhmann and Mario Bunge]. Haupt, Bern, Switzerland, 2004. ISBN 3-258-06658-2. 298 pp. LCCN HV41 .K53 2004.

Klibansky:1968:CPS

[Kli68] Raymond Klibansky, editor. Contemporary philosophy, a survey. La philosophie contemporaine, chroniques, volume II. La Nuova Italia Editrice, Florence, Italy, 1968. ???? pp.

Kiefer:1970:MSH

[KM70] Howard Evans Kiefer and Milton Karl Munitz, editors. Mind, Science and History, volume 2 of Contemporary philosophie thought: The International Philosophy Year conferences at Brockport. State University of New York Press, Albany, NY, USA, 1970. ISBN 0-87395-052-6. LCCN B29 .M556 1970. URL

http://catalogue.bnf.fr/ark:/12148/cb353806733.

Kondo:1995:ALS

[Kon95] Michiro Kondo. Approximation logic and strong Bunge algebra. Notre Dame journal of formal logic, 36(4):595–605, 1995. CO-DEN NDJFAM. ISSN 0029-4527 (print), 1939-0726 (electronic).

Kamitz:1974:R

[KPSP74] Reinhard Kamitz, Veit Pittioni, Peter Schmidt, and Peter Posch. Reviews: Mario Bunge (ed.), Exact Philosophy, Problems, Tools and Goals, D. Reidel Publishing Company, Dordrecht-Holland / Boston-U.S.A., 1973, x + 214 pp., Dr. 52,-. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 5(1):99–113, June 1974. CODEN THD-CBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Kranzberg:1980:EAP

[Kra80] Melvin Kranzberg, editor. Ethics in an Age of Pervasive Technology. Westview Press, Boulder, CO, USA, 1980. ISBN 0-89158-686-5. LCCN ????

Kreweras:1991:ST

[Kre91] Germain Kreweras. Skepticism and truth. New Ideas in Psychology, 9(2):197–201, 1991. ISSN 0732-118X (print), 1873-3522

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0732118X9190024G. Special issue on Mario Bunge

and Martin Mahner.

LabordaMolteni:2014:MPE

[Lab14] Jorge Laborda Molteni. Médicos y pacientes, la evolución de la relación. (Spanish) [Doctors and patients, the evolution of the relationship]. In Denegri [Den14a], pages 209–220. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Lacey:1996:RBS

[Lac96] Hugh Lacey. On relations between science and religion. Science & Education (Springer), 5(2):143–153, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Landry:1989:EAM

[Lan89] Cecile Landry. Entretien avec Mario Bunge. (French) [Interview with Mario Bunge]. Philosopher: revue pour tous, 8(??):25–33,

???? 1989. URL https://philpapers.org/rec/LANEAM-4.

Laucken:1991:ONS

[Lau91] Uwe Laucken. The ontology of the natural sciences as a truncheon. New Ideas in Psychology, 9(2):203–213, 1991. ISSN 0732-118X (print), 1873-3522 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0732118X9190025H.

Special issue on Mario Bunge and Martin Mahner.

Llinas:1978:MBP

[LB78a] Rodolfo Llinás and Mario Bunge. The mind–body problem in the light of contemporary biology. In Diemer [Die78], pages 131–133. ISBN 3-8204-5813-1.

Llinas:1978:RAC

[LB78b] Rodolfo Llinás and Mario Bunge. Restricted applicability of the concept of command in the neurosciences: dangers of metaphors. Behavioral and Brain Sciences, 1(1):30–31, March 1978. CODEN BBSCDH. ISSN 0140-525X (print), 1469-1825 (electronic).

Levin:1983:BRM

[Lev83] Michael E. Levin. Book review: The Mind-Body Problem. by Mario Bunge. Noûs, 17(2):316–321, May 1983. ISSN 0029-4624 (print), 1468-0068 (electronic). URL https://www.jstor.org/

stable/2215151.

Lindsay:1969:BRS

[Lin69] R. Bruce Lindsay. Book review: Scientific research I: the search for system: Mario Bunge. 536 pages, 6×9 in. New York, Springer-Verlag, 1967. Price, $21.00. Journal of The Franklin Institute, 287(1):79–80, 1969. CODEN JFINAB. ISSN 0016-0032 (print),

1879-2693 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0016003269900349.

Leinfellner:1974:DMS

[LK74] Werner Leinfellner and Eckehart Köhler, editors. Developments in the methodology of social science, volume 6 of Theory and decision library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0493-7 (hardcover), 90-277-0539-9 (paperback). x

+ 444 pp. LCCN H61 .D45. URL http://swbplus.bsz-bw.de/
bsz011135883cov.htm.

Lakatos:1968:PIC

[LM68] Imre Lakatos and Alan Musgrave, editors. Proceedings of the International Colloquium in the Philosophy of Science, London 1965, volume III. Problems in the Philosophy of Science, volume 49 of Studies in Logic and the Foundations of Mathematics.

Elsevier, Amsterdam, The Netherlands, 1968. ISBN 0-444-53413-X. ISSN 0049-237X. LCCN Q175 .I518 1965. URL http://www.

sciencedirect.com/science/book/978044453413.

Long:1962:BRI

[Lon62] Wilbur Long. Book review: Intuition and Science by Mario Bunge. Philosophy East and West, 12(1):73–75, April 1962. URL

https://www.jstor.org/stable/1397248.

Lemarchand:1988:SPD

[LP88] Guillermo A. Lemarchand and Alberto R. Pedace, editors. Scientists, Peace and Disarmament. Wold Scientific, Teaneck, NJ, USA, 1988. ISBN 9971-5-0761-7. LCCN ????

Lukyanenko:2022:FIT

[LSP22] Roman Lukyanenko, Veda C. Storey, and Oscar Pastor. Foundations of information technology based on Bunge’s systemist philosophy of reality. Metascience: Scientific General Discourse, 2(??):140–181, 2022. CODEN ???? ISSN ???? URL https://

philpapers.org/rec/LUKFOI-2.

Luna:2000:BRS

[Lun00] Francisco Parra Luna. Book review: Social Science under Debate a Philosophical Perspective by Mario Bunge. Reis, ??(91):

191–196, July 2000. URL https://www.jstor.org/stable/
40039190.

MacGaugh:1985:CPB

[Mac85] James L. MacGaugh, editor. Contemporary Psychology: Biological Processes and Theoretical Issues, volume 1 of Proceedings of the XXIII [twenty-third] International Congress of Psychology of the International Union of Psychological Science (IUPsyS). North-Holland Publishing Co., Amsterdam, The Netherlands, 1985. ISBN 0-444-87781-9 (vol. 1), 0-444-87782-7 (vol. 2), 0-444-87783-5 (vol. 3), 0-444-87784-3 (vol. 4), 0-444-87785-1 (vol. 5), 0-444-87786-X (vol. 6), 0-444-87787-8 (vol. 7), 0-444-87788-6 (vol. 8), 0-444-87789-4 (vol. 9), 0-444-87790-8 (set). LCCN BF121 .I564 1984 vol. 1.

MachadoMoretto:2011:AMC

[Mac11] Luís Augusto Machado Moretto. Una arquitetura multiagentes para o compartilhamento do conhecimento em sistemas sociotecnológicos de engenharia de requisitos. (Portuguese) [A multiagent architecture for the sharing of knowledge in sociotechnological systems of requirements engineering]. Dissertação (mestrado), Centro Tecnológico, Programa de Pós-Graduação em Engenharia e Gestão do Conhecimento, Universidade Federal de Santa Catarina UFSC, Florianópolis, Brazil, 2011. 138 pp. URL http://

repositorio.ufsc.br/xmlui/handle/123456789/94810.

Kurtz:1994:CED

[Mad94] Paul Kurtzand Timothy J. Madigan, editor. Challenges to the Enlightenment: in Defense of Reason and Science. Prometheus Books, Amherst, NY, USA, 1994. ISBN 0-87975-869-4. 319 pp. LCCN ????

Mahner:2001:SRS

[Mah01] Martin Mahner, editor. Scientific Realism: Selected Essays of Mario Bunge. Prometheus Books, Amherst, NY, USA, 2001. ISBN 1-57392-892-5. 438 pp. LCCN Q175.32 .R42.

Mahner:2013:RDN

[Mah13] M. Mahner. Le rôle du naturalisme métaphysique en science. (French) [The role of metaphysical naturalism in science]. In Matériaux philosophiques et scientifiques pour un matérialisme contemporain: Volume 1 [Sil13d], page ?? ISBN 2-919694-51-0 (print), 2-919694-20-0 (e-book). LCCN B825 .M384 2016.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=5336199.

Mahner:2014:OFM

[Mah14] Martín Mahner. La ontología y la filosofía de la mente de Mario Bunge. (Spanish) [The ontology and philosophy of the mind of Mario Bunge]. In Denegri [Den14a], pages 221–234. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Mahner:2021:MBC

[Mah21] Martin Mahner. Mario Bunge (1919–2020): Conjoining philosophy of science and scientific philosophy. Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie, 52(1):3–23, March 2021. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL https://link.springer.

com/article/10.1007/s10838-021-09553-7.

Manzano:1985:FTT

[Man85] María Manzano. Formalization in type theory of the existence predicate of Mario Bunge. Theoria: An International Journal for Theory, History and Foundations of Science, 1:513–534, 1985. ISSN 0495-4548 (print), 2171-679X (electronic).

Margain:1972:PDB

[Mar72] Hugo Margáin. La paradoja del Dr. Bunge. (Spanish) [The paradox of Dr. Bunge]. Crítica: Revista Hispanoamericana de Filosofía, 6(18):113–116, September 1972. CODEN ???? ISSN

0011-1503. URL https://www.jstor.org/stable/40103996.

See response [Bun75d].

Margolis:1981:NCP

[Mar81] Joseph Margolis. Nature, culture, and persons. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 13(4):311–329, December 1981. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Marquis:1996:CNB

[Mar96] Jean-Pierre Marquis. A critical note on Bunge’s ‘system boundary’ and a new proposal. International Journal of General Systems, 24(3):245–255, 1996. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic).

Marquis:2012:MBP

[Mar12] Jean-Pierre Marquis. Mario Bunge’s philosophy of mathematics: An appraisal. Science & Education (Springer), 21(10):1567–1594, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Marquis:2019:BMS

[Mar19] Jean-Pierre Marquis. Bunge’s mathematical structuralism is not a fiction. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 33, pages 587–608. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Mattessich:1982:ARS

[Mat82] Richard Mattessich. Axiomatic representation of the systems framework: Similarities and differences between Mario Bunge’s world of systems and my own systems methodology. Cybernetics and Systems, 13(1):51–75, 1982. CODEN CYSYDH. ISSN 0196-9722 (print), 1087-6553 (electronic).

Matthews:1996:E

[Mat96] Michael Matthews. Editorial. Science & Education, 5(2):??, ???? 1996.

Matthews:2003:MBP

[Mat03] Michael R. Matthews. Mario Bunge: Physicist and philosopher. Science & Education (Springer), 12(5–6):431–444, August 2003. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901

(electronic). URL https://link.springer.com/article/10.
1023/A:1025364722916.

Matthews:2012:MBS

[Mat12] Michael R. Matthews. Mario Bunge, systematic philosophy and science education: an introduction. Science & Education (Springer), 21(10):1393–1403, October 2012. CO-DEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11191-012-9530-0.pdf.

Matthews:2016:BRM

[Mat16] Michael R. Matthews. Book review: Mario Bunge: a Long and Productive Life: Between Two Worlds: Memoirs of a Philosopher–Scientist, Springer, Dordrecht. ISBN 978-3-319-29250-2, xviii + 496 pp., EUR 100 (hardback). Science & Education (Springer), 25(9–10):1135–1148, December 2016. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Matthews:2019:MBC

[Mat19a] Michael Matthews. Mario Bunge: a Centenary Festschrift. Springer Nature Switzerland AG, Cham, Switzerland, 2019. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). xxx + 827 pp. LCCN Q175.55.

Matthews:2019:MBI

[Mat19b] Michael R. Matthews. Mario Bunge: An introduction to his life, work and achievements. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 1, pages 1–28. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Matthews:2019:MBE

[Mat19c] Michael R. Matthews. Mario Bunge and the enlightenment project in science education. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 36, pages 645–682. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Matthews:2019:PA

[Mat19d] Michael R. Matthews. Preface and acknowledgements. In Mario Bunge: a Centenary Festschrift [Mat19a], pages ix–xviii. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Maurice:2017:TS

[Mau17] François Maurice. Une triade scientifique? In Marc Silberstein, editor, Qu’est-ce que la science … pour vous?. (French) [What is science … for you?], pages 169–173. Éditions Matériologiques, Paris, France, 2017.

Maurice:2020:E

[Mau20] François Maurice. Editorial. Métascience, 1(??):??, ???? 2020.

Maurice:2022:BMN

[Mau22a] François Maurice. Bunge’s metascience and the naturalization of the general discourse. Metascience: Scientific General Discourse, 2(??):74–93, 2022. CODEN ???? ISSN ???? URL https://

philpapers.org/rec/MAUBMA.

Maurice:2022:PMO

[Mau22b] François Maurice. Presentation: Metascientific ontology. Metascience: Scientific General Discourse, 2(??):10–18, 2022. CO-

DEN ???? ISSN ???? URL https://philpapers.org/rec/
MAUPMO.

Maurice:2022:WMO

[Mau22c] François Maurice. What is metascientific ontology? Metascience: Scientific General Discourse, 2(??):22–44, 2022. CODEN ????

ISSN ???? URL https://philpapers.org/rec/MAUWIM.

Mayer:1990:SPR

[May90] Richard E. Mayer. In search of psychology: a response to Bunge. New Ideas in Psychology, 8(2):151–154, 1990. ISSN 0732-118x (print), 1873-3522 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0732118X9090004L.

See [Bun90g] and reply [Bun90e].

Mahner:1996:ISR

[MB96a] Martin Mahner and Mario Bunge. The incompatibility of science and religion sustained: A reply to our critics. Science & Education (Springer), 5(2):189–199, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Mahner:1996:REC

[MB96b] Martin Mahner and Mario Bunge. Is religious education compatible with science education? Science & Education (Springer), 5(2):101–123, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge. See [WL96] and reply [MB96a].

Mahner:1997:FB

[MB97] Martin Mahner and Mario Bunge. Foundations of Biophilosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61838-4 (hardcover), 3-642-08276-9, 3-662-03368-2 (e-book). xviii + 423 pp. LCCN QH331

.M26 1997. URL http://www.loc.gov/catdir/enhancements/
fy0815/97011448-d.html;   http://www.loc.gov/catdir/
enhancements/fy0815/97011448-t.html;   http://www.springerlink.
com/content/978-3-662-03368-5.

Mahner:2000:PGB

[MB00] Martin Mahner and Mario Bunge. Philosophische Grundlagen der Biologie. (German) [Philosophical Foundations of Biology]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67649-X. xviii + 402 pp. LCCN QH331 .M2615 2000. URL http://www.loc.gov/

catdir/enhancements/fy1109/00435801-d.html;   http://
www.loc.gov/catdir/enhancements/fy1109/00435801-t.html.

Mahner:2001:FFS

[MB01] Martin Mahner and Mario Bunge. Function and functionalism: a synthetic perspective. Philosophy of Science, 68(1):75–94, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3081025.

Machado:2016:CHS

[MB16] Juliana Machado and Marco António Barbosa Braga. Can the history of science contribute to modelling in physics teaching? The case of Galilean studies and Mario Bunge’s epistemology. Science & Education (Springer), 25(7–8):823–836, October 2016. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Morgado-Bernal:2019:CE

[MB19] Ignacio Morgado-Bernal. Is consciousness an epiphenomenon? In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 22, pages 377–385. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Marion:1995:QSP

[MC95] Mathieu Marion and Robert S. Cohen. Québec Studies in the Philosophy of Science: Part I: Logic, Mathematics, Physics and History of Science Essays in Honor of Hugues Leblanc, volume 177 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-7204-3, 94-009-1575-6 (e-book). ISSN 0068-0346. xi + 331 pp. LCCN ???? URL https://link.

springer.com/book/10.1007/978-94-009-1575-6.

McFetridge:1978:BRT

[McF78] I. G. McFetridge. Book review: Treatise on Basic Philosophy. Volume 2, Semantics II: Interpretation and Truth. by Mario Bunge. MIND: A Quarterly Review of Pyschology and Philosophy, 87(345):144–146, January 1978. CODEN MNDDAB. ISSN 0026-4423 (print), 1460-2113 (electronic). URL https://www.

jstor.org/stable/2253375.

McFalls:2007:MWO

[McF07] Laurence McFalls, editor. Max Weber’s Objectivity Reconsidered, volume 5 of German and European studies. University of Toronto Press, Toronto, ON, Canada, 2007. ISBN 0-8020-9224-1. vi + 389 pp. LCCN H61 .M4294 2007. URL http:

//www.degruyter.com/doi/book/10.3138/9781442684553;
http://www.jstor.org/stable/10.3138/j.ctt2tv3d6.

McGowan:1994:WWJ

[McG94] Don McGowan. What is wrong with Jung. Prometheus Books, Amherst, NY, USA, 1994. ISBN 0-87975-859-7 (hardcover). 219 pp. LCCN BF109.J8 M35 1994. US$26.95.

Metzger:1973:BRB

[Met73a] Robert S. Metzger. Book review: Problems in the Foundations of Physics by Mario Bunge. Philosophy of Science, 40(3):464–466, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/
186214.

Metzger:1973:BRP

[Met73b] Robert S. Metzger. Book review: Problems in the Foundations of Physics by Mario Bunge. Philosophy of Science, 40(3):464–466, September 1973. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL https://www.jstor.org/stable/
186214.

Miller:1977:BTP

[Mil77] David Miller. Bunge’s theory of partial truth is no such thing. Philosophical Studies, 31(2):147–150, February 1977. CODEN PLSDA3. ISSN 0031-8116 (print), 1573-0883 (electronic). URL

https://www.jstor.org/stable/4319123.

Miller:1993:MRC

[Mil93] John Laurence Miller. Mentalism revisited: The case of learning disabilities. New Ideas in Psychology, 11(1):39– 59, 1993. ISSN 0732-118X (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0732118X9390019A.

Manzano:2019:QCE

[MM19] María Manzano and Manuel Crescencio Moreno. Quantifiers and conceptual existence. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 7, pages 117–138. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Moessinger:1987:LPP

[Moe87] Pierre Moessinger. Let’s be as precise as possible: an interview with Mario Bunge. New Ideas in Psychology, 5(3):385– 397, 1987. ISSN 0732-118X (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0732118X87900079.

Moessinger:1991:E

[Moe91] Pierre Moessinger. Editorial. New Ideas in Psychology, 9(2): ??, ???? 1991. Special issue dedicated to pseudoscience around Bunge.

Moessinger:2008:VSM

[Moe08] Pierre Moessinger. Voir la société: le micro et le macro. (French) [See society: the micro and the macro]. Société et pensées. Hermann, Paris, France, 2008. ISBN 2-7056-6674-5. 258 pp. LCCN HM588 .M64 2008.

Montada:1992:BKD

[Mon92] Leo Montada, editor. Bericht über den 38. Kongress der Deutschen Gesellschaft für Psychologie in Trier 1992. (German) [Report on the 38th Congress of the German Psychological Society in Trier 1992], volume 2. Verlag für Psychologie Hogrefe, Göttingen, Germany, 1992. ISBN 3-8017-0667-2. LCCN ????

Morgenbesser:1961:BRC

[Mor61] Sidney Morgenbesser. Book review: Causality: The Place of the Causal Principle in Modern Science by Mario Bunge. Scientific American, 204(2):175–178, February 1961. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL https://

www.jstor.org/stable/24937360.

Morscher:1981:PW

[Mor81] Edgar Morscher. Philosophie als Wissenschaft: Paul Weingartner gewidmet = Essays in scientific philosophy: dedicated to Paul Weigartner, volume 1 of Reihe phi. Philosophische Forschung. Comes Verlag, Bad Reichenhall, Germany, 1981. ISBN 3-88820-000-8. 470 pp. LCCN B790 .P48 1981.

Moreno:2016:SPI

[Mor16] Paco Moreno. El “seny” perdido: el independentismo catalán no tiene razón. (Spanish) [The lost sanity: the Catalan independence movement has no reason], volume 20 of Colección Libros abiertos. Laetoli, Pamplona, Spain, 2016. ISBN 84-92422-91-2. 202 pp. LCCN DP302.C68 M67 2016.

Moscovici:1991:RRC

[Mos91] Serge Moscovici. Reflections and reactions to the credo of a true believer. New Ideas in Psychology, 9(2):215– 225, 1991. ISSN 0732-118X (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0732118X9190026I. Special issue on Mario Bunge and Martin Mahner.

Mosterin:2014:OMB

[Mos14] Jesús Mosterín. La ontología de Mario Bunge. (Spanish) [The ontology of Mario Bunge]. In Denegri [Den14a], pages 235– 250. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL

http://core.cambeiro.com.ar/0-100933-7.pdf;   http://
www.loc.gov/catdir/toc/gc01/9789502323398.pdf.

MoyaRojas:2015:NCF

[Moy15] Nicanor Moya Rojas. Naturaleza de los cuestionamientos filosóficos y sociológicos de Mario Bunge a la filosofía materialista dialéctica. (Spanish) [Nature of philosophical and sociological questions of Mario Bunge’s dialectical materialist philosophy]. Horizonte de la Ciencia, 5(8):39–46, ???? 2015.

Mueller:1968:DM

[Mue68] Marvin M. Mueller. In defense of Mach. American Journal of Physics, 36(2):166–167, February 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [Bun68i].

Mullin:1963:BRM

[Mul63] Albert A. Mullin. Book review: The Myth of Simplicity, by Mario Bunge. Pp. 239. Prentice-Hall, Inc., Englewood Cliffs, New Jersey, 1963. American Journal of Physics, 31(11):891, November 1963. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Munier:1993:EF

[Mun93] Bertrand Munier. Editorial foreword. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 35(3):203–205, November 1993. CODEN THD-CBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Michalos:1973:R

[MVR^{+}73] Alex C. Michalos, Viktor Vanberg, Dennis Rohatyn, Mario Bunge, and Karl-Dieter Opp. Reviews. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 3(4):388–402, June 1973. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Narsky:1975:BBV

[Nar75] I. Narsky. Bemerkungen über den Vortrag von Prof. Bunge. (German) [Comments on the lecture of Prof. Bunge]. In Perelman [Per75], pages 78–86. ISBN 90-247-1720-5. LCCN B105.D48 I57 1975; B809.7 .E611 1973. See [Bun75a].

Nuccetelli:2010:CLA

[NSB10] Susana Nuccetelli, Ofelia Schutte, and Otávio Bueno, editors. A Companion to Latin American Philosophy. Blackwell companions to philosophy. Wiley-Blackwell, Malden, MA, USA, 2010. ISBN 1-4051-7979-1 (hardcover). xiv + 555 pp. LCCN B1001 .C65 2010.

Ochoa:1984:CEE

[OB84] García-Bermejo Ochoa and Mario Bunge. Contra la economía escolástica. (Spanish) [Against the scholastic economy]. Teorema: Revista Internacional de Filosofía, 14(3–4):585–595, ???? 1984. CODEN ???? ISSN 0210-1602. URL http://www.jstor.org/

stable/43047293.

Obiedat:2011:MBW

[Obi11] Ahmad Zahaadden Obiedat. Mario Bunge’s worldview and its implications for the modernization of Arabic–Islamic philosophy. Ph.D. thesis, Faculty of Religious Studies, McGill University, Montréal, QC, Canada, May 2011. vii + 220 pp. URL https:/

/search.proquest.com/pqdtglobal/docview/1040759490/.

Obiedat:2019:HCB

[Obi19] A. Z. Obiedat. How can Bunge’s scientific–humanistic ethics engage Islamic moral law? In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 28, pages 487–512. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Obiedat:2022:MIA

[Obi22] A. Z. Obiedat. Modernity and the Ideals of Arab–Islamic and Western-Scientific Philosophy: The Worldviews of Mario Bunge and Taha Abd al-Rahman. Springer International Publishing, Cham, Switzerland, 2022. ISBN 3-030-94264-3 (hardcover), 3-030-94267-8 (paperback), 3-030-94265-1 (e-book), 3-030-94266-X. xix + 411 + 13 pp. LCCN BP166-166.94.

Obrecht:2005:ESM

[Obr05] Werner Obrecht. Der emergentistische Systemismus Mario Bunges und das Systemtheoretische Paradigma der Sozialarbeitswissenschaft und der Sozialen Arbeit (SPSA). (German) [Mario Bunge’s emergentist systemism and the systems theory paradigm of social work science and social work (SPSA)], 2005.

URL http://www.sozialarbeit.ch/kurzinterviews/werner_
obrecht.htm.

Opdahl:2002:OEU

[OHS02] Andreas L. Opdahl and Brian Henderson-Sellers. Ontological evaluation of the UML using the Bunge–Wand–Weber model. Software and Systems Modeling, 1(1):43–67, September 2002. CODEN ???? ISSN 1619-1366 (print), 1619-1374 (electronic).

Orensanz:2022:BHG

[Ore22a] Martín Orensanz. Bunge and Harman on the general theory of objects. Metascience: Scientific General Discourse, 2(??):46–64, 2022. CODEN ???? ISSN ???? URL https://philpapers.

org/rec/OREBAH.

Orensanz:2022:CAM

[Ore22b] Martín Orensanz. Causation according to Mario Bunge and Graham Harman. Metascience: Scientific General Discourse, 2 (??):66–73, 2022. CODEN ???? ISSN ???? URL https://

philpapers.org/rec/ORECAT-2.

Ortiz:2019:MBC

[Ort19] Eduardo L. Ortiz. Mario Bunge in the complex Argentina of the 1940s–1960s. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 3, pages 37–62. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Olive:1996:FME

[OSV96] León Olivé, Fernando Salmerón, and Luis Villoro, editors. Filosofía moral, educación e historia: homenaje a Fernando Salmerón. (Spanish) [Moral Philosophy, Education and History: a Tribute to Fernando Salmerón]. Facultad de Filosofía y Letras, Instituto de Investigaciones Filosóficas, Universidad Nacional Autónoma de México, México, DF, México, 1996. ISBN 968-36-5425-8. 776 pp. LCCN BJ1142.

OConnor:1979:BMS

[OW79] Maeve O’Connor and G. E. W. (Gordon Ethelbert Ward) Wolstenholme, editors. Brain and Mind: [Symposium on Brain and Mind Held at the Ciba Foundation, London, 5th–7th December 1978], volume 69 (new series) of Ciba Foundation symposium. Excerpta Medica, Amsterdam, The Netherlands, 1979. ISBN 0-444-90095-0 (Elsevier/NorthHolland), 0-470-71820-X (e-book), 90-219-4075-2. ix + 425 pp. LCCN QP376 .S878 1978. URL http:

//onlinelibrary.wiley.com/book/10.1002/9780470720523.

Panksepp:1990:CMB

[Pan90] Jaak Panksepp. Can mind and behavior be understood without understanding the brain?: a response to Bunge. New Ideas in Psychology, 8(2):139–149, 1990. ISSN 0732-118x (print),

1873-3522 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0732118X9090003K. See [Bun90g] and

reply [Bun90e].

Paradis:1981:MBP

[Par81] Michel Paradis. Book review: The mind–body problem. A psychobiological approach: Mario Bunge. Pergamon Press, Oxford, 1980, 250 pp. $13.50. Neuropsychologia, 19(4):625– 626, 1981. ISSN 0028-3932 (print), 1873-3514 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0028393281900324.

ParraLuna:1983:EPT

[Par83] Francisco Parra Luna. Elementos para una teoría formal del sistema social: una orientación crítica. (Spanish) [Elements for a Formal Theory of the Social System: a Critical Orientation]. Editorial de la Universidad Complutense, Madrid, Spain, 1983. ISBN 84-7491-068-4. 556 pp. LCCN HM131 .P318 1983.

PardosPerez:2014:P

[Par14] José Luis Pardos Pérez. Prólogo. (Spanish) [Foreword]. In Denegri [Den14a], pages 7–14. ISBN 950-23-2339-4. LCCN

BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Paty:1990:RPM

[Pat90] Michel Paty. Reality and probability in Mario Bunge’s treatise. In Weingartner and Dorn [WD90], pages 301–322. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990.

Poppel:1989:GB

[PB89] Ernst Pöppel and Ansgar Beckermann, editors. Gehirn und Bewusstsein. (German) [Brain and consciousness]. VCH, Weinheim, Germany, 1989. ISBN 3-527-27901-6. LCCN BF311 .G42

1989. URL http://digitool.hbz-nrw.de:1801/webclient/
DeliveryManager?pid=3277305.

Perelman:1975:D

[Per75] Chaïm Perelman, editor. Dialectiques. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1975. ISBN 90-247-1720-5. LCCN B105.D48 I57 1975; B809.7 .E611 1973.

Perrez:1991:DBE

[Per91] Meinrad Perrez. The difference between everyday knowledge, ideology, and scientific knowledge. New Ideas in Psychology, 9(2): 227–231, 1991. ISSN 0732-118X (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0732118X9190027J. Special issue on Mario Bunge and Martin Mahner.

Pfeiffer:1969:AHW

[Pfe69] Heinrich Pfeiffer, editor. Alexander von Humboldt. Werk und Weltgeltung. (German) [Alexander von Humboldt. Work and international standing]. Piper, München, Germany, 1969. 505 pp. LCCN Q143.H9 A68.

Pfeiffer:1977:DUW

[Pfe77] Heinrich Pfeiffer, editor. Denken und Umdenken: zu Werk und Wirkung von Werner Heisenberg. (German) [Thinking and rethinking: the work and influence of Werner Heisenberg]. Piper, München, West Germany, 1977. ISBN 3-492-02275-8. 279 + 10 pp. LCCN QC16.H35 D46. DM28.00.

Pickel:2001:MBP

[Pic01] Andreas Pickel. Mario Bunge’s philosophy of social science. Society, 38(4):71–74, May/June 2001.

Pickel:2004:SMS

[Pic04] Andreas Pickel. Systems and mechanisms: a symposium on Mario Bunge’s philosophy of social science. Philosophy of the Social Sciences = Philosophie des sciences sociales, 34(2):169–181, 2004. ISSN 0048-3931 (print), 1552-7441 (electronic).

Pickel:2007:RBM

[Pic07] Andreas Pickel. Review: Bunge, M. (2003). Emergence and Convergence: Qualitative Novelty and the Unity of Knowledge. Toronto, ON: University of Toronto Press. Philosophy of the Social Sciences = Philosophie des sciences sociales, 37(2):248–251, 2007. ISSN 0048-3931 (print), 1552-7441 (electronic).

Pickel:2019:CSS

[Pic19] Andreas Pickel. Cultures as semiotic systems: Reconceptualizing culture in a systemic perspective. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 25, pages 415–438. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Pietrocola:1999:SRM

[Pie99] Maurício Pietrocola. In search of a reality: Models and modelisation in science teaching according to the epistemology of Mario Bunge. In ????, editor, Proceedings of the 5th International History, Philosophy and Science Teaching Conference, Pavia and Lake Como, Italy, page ?? ????, ????, 1999. ISSN 0926-7220 (print), 1573-1901 (electronic). LCCN ????

Pinilla:2013:MBE

[Pin13] Jorge Ortiz Pinilla. Mario Bunge y la estadística bayesiana. (Spanish) [Mario Bunge and Bayesian statistics]. Comunicaciones en Estadística, 6(2):225, October 12, 2013. CODEN ???? ISSN 2027-3355 (print), 2339-3076 (electronic).

Poole:1996:MBR

[Poo96] Michael Poole. “… for more and better religious education?”. Science & Education (Springer), 5(2):165–174, April 1996. CO-DEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Potok:1970:C

[Pot70] Chaïm Potok. The Chosen. Penguin. Fiction. Penguin Books, Harmondsworth, UK, 1970. ISBN 0-14-003094-8 (paperback). 280 pp. LCCN PS3566.O86 C5 1970.

Potok:1992:CN

[Pot92] Chaim Potok. The Chosen: a Novel. Alfred A. Knopf, New York, NY, USA, 25th anniversary edition, 1992. ISBN 0-679-40222-5. xv + 295 pp. LCCN PS3566.O69 C54 1992.

Parkinson:1994:RHP

[PS94] G. H. R. (George Henry Radcliffe) Parkinson and Stuart Shanker, editors. Philosophy of Science, Logic and Mathematics in the Twentieth Century, volume 9 of Routledge history of philosophy. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-05776-0. xxxviii + 461 pp. LCCN Q174.8.P55 1996.

Puccetti:1981:BRM

[Puc81] Roland Puccetti. Book review: The Mind-Body Problem. A Psychobiological Approach by Mario Bunge. British Journal for the Philosophy of Science, 32(3):282–286, September 1981. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL

https://www.jstor.org/stable/687592.

Pinsker:1980:IPN

[PW80] Harold M. Pinsker and William Darren Willis, editors. Information Processing in the Nervous System. Raven Press, New York, NY, USA, 1980. ISBN 0-89004-422-8. LCCN QP364.5 I53.

Quintanilla:1979:TEM

[QB79] Miguel Angel Quintanilla and Mario Bunge. Teorema entrevista a Mario Bunge. (Spanish) [Theorema interview with Mario Bunge]. Teorema: Revista Internacional de Filosofía, 9(3–4): 353–386, ???? 1979. CODEN ???? ISSN 0210-1602. URL http:

//www.jstor.org/stable/43046024.

QueirozMagalhaes:2013:DFM

[QdAY13] Marcos Thadeu Queiroz Magalhães, Joaquim José Guilherme de Aragão, and Yaeko Yamashita. Definições formais de mobilidade e acessibilidade apoiadas na teoria de sistemas de Mario Bunge. (Portuguese) [Formal definitions of mobility and accessibility supported by the systems theory of Mario Bunge]. Paranoá: Cadernos de Arquitetura e Urbanismo, ??(9):??, December 1, 2013.

Quine:1958:SNL

[Qui58] Willard Van Orman Quine. El sentido de nuestra lógica. (Spanish) [The meaning of our logic], volume 4 of Colección Interciencia. Nueva visión, Buenos Aires, Argentina, 1958. 190 pp. Translation from English to Spanish by Mario Augusto Bunge.

Quintanilla:1974:BRM

[Qui74] Miguel A. Quintanilla. Book review: Method, Model and Matter by Mario Bunge. Teorema: Revista Internacional de Filosofía, 4 (1):147–151, ???? 1974. CODEN ???? ISSN 0210-1602. URL

https://www.jstor.org/stable/43047249.

Quintanilla:1979:TIM

[Qui79] Miguel A. Quintanilla. Teorema interviews Mario Bunge. Teorema: Revista Internacional de Filosofía, 9(3–4):353–386, 1979. ISSN 0210-1602.

Quintanilla:1980:BRO

[Qui80] Miguel A. Quintanilla. Book review: Ontology II: a World of Systems. Treatise on Basic Philosophy Vol. IV by Mario BUNGE. Dialectica: International Review of Philosophy of Knowledge, 34 (4):295–297, ???? 1980. ISSN 0012-2017 (print), 1746-8361 (elec-

tronic). URL https://www.jstor.org/stable/42970458.

QuintanillaFisac:2014:ARF

[Qui14] Miguel Ángel Quintanilla Fisac. Algunos retos filosóficos de la política científica. (Spanish) [Some philosophical challenges of scientific policy]. In Denegri [Den14a], pages 251–268. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL http://core.

cambeiro.com.ar/0-100933-7.pdf;   http://www.loc.gov/
catdir/toc/gc01/9789502323398.pdf.

Raynaud:2015:SCS

[Ray15] Dominique Raynaud. Scientific Controversies: a Socio-historical Perspective on the Advancement of Science. Transaction Publishers, New Brunswick, NJ, USA, 2015. ISBN 1-4128-5571-3 (hardcover), 1-315-12903-5 (e-book), 1-351-49178-4, 1-351-49179-2, 1-351-49180-6. xxii + 298 pp. LCCN Q175.5 .R3813 2015. URL

http://www.tandfebooks.com/isbn/9781315129037.

Raynaud:2019:TPT

[Ray19] Dominique Raynaud. Truth in the post-truth era: Evaluating the theories of truth with a table of contingency. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 8, pages 139–162. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Rescher:1987:SIP

[Res87] Nicholas Rescher, editor. Scientific Inquiry in Philosophical Perspective. University Press of America, Lanham, MD, USA, 1987. ISBN 0-8191-5798-8, 0-8191-5799-6 (paperback). viii + 300 pp. LCCN Q175 .S4235 1987.

Rescher:2019:QE

[Res19] Nicholas Rescher. Quantitative epistemology. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 10, pages 181–184. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Rosemann:2002:DMM

[RG02] Michael Rosemann and Peter Green. Developing a meta model for the Bunge–Wand–Weber ontological constructs. Information systems, 27(2):75–91, 2002. CODEN INSYD6. ISSN 0306-4379 (print), 1873-6076 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0306437901000485.

Richmond:2008:BRB

[Ric08] Sheldon Richmond. Book review: Bunge, Mario. 2006. Chasing Reality: Strife Over Realism. Toronto: University of Toronto Press. Philosophy of the Social Sciences = Philosophie des sciences sociales, 38(4):545–551, 2008. ISSN 0048-3931 (print), 1552-7441 (electronic).

Rivadulla:2019:CET

[Riv19] Andrés Rivadulla. Causal explanations: Are they possible in physics? In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 18, pages 303–328. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Robinson:1974:BRP

[Rob74] M. C. Robinson. Book review: Philosophy of Physics, by M. Bunge. Pp. 248. Reidel, Boston, MA, 1973. American Journal of Physics, 42:797–798, September 1974. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/1974AmJPh..42..797B.

RojasOsorio:2002:MBF

[Roj02] Carlos Rojas Osorio. Mario Bunge: filosofía científica. In Latinoamérica. Cien años de filosofía. (Spanish) [Latin America. One hundred years of philosophy], pages 228–238. Isla Negra Editores, San Juan, Puerto Rico, 2002. ISBN 1-932271-02-3. LCCN B1001 .R648 2002.

Rojas:2013:ADR

[Roj13] Andrés Gutiérrez Rojas. Acerca de la defensa de la racionalidad bayesiana y la obra de Mario Bunge. (Spanish) [On the defense of Bayesian rationality and the work of Mario Bunge]. Comunicaciones en Estadística, 6(2):213, October 12, 2013. CODEN ???? ISSN 2027-3355 (print), 2339-3076 (electronic).

Romero:2017:BGW

[Rom17] Gustavo E. Romero. Bunge on gravitational waves. arXiv.org, ??

(??):1–6, August 6, 2017. URL https://arxiv.org/abs/1708.
04665.

Romero:2018:MBG

[Rom18] Gustavo E. Romero. Mario Bunge on gravitational waves and the reality of spacetime. Foundations of Science, 23(2):405–409, June 2018. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL http://link.springer.com/article/
10.1007/s10699-017-9529-8. See [Bun18b].

Romero:2019:PPP

[Rom19] Gustavo E. Romero. Physics and philosophy of physics in the work of Mario Bunge. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 17, pages 289–301. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Royce:1990:QPD

[Roy90] Joseph R. Royce. Questions on psychology as a discipline: a response to Bunge. New Ideas in Psychology, 8(2):155– 157, 1990. ISSN 0732-118x (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0732118X9090005M. See [Bun90g] and reply [Bun90e].

Ruse:1972:BT

[Rus72] Michael Ruse. Bunge on time. Philosophy of Science, 39(1): 82, March 1972. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/

186596. See [Bun68k].

Russo:1973:LMB

[Rus73] F. Russo. L’épistémologie de Mario Bunge. (French) [the epistemology of Mario Bunge]. Archives de Philosophie, 36 (3):373–393, ???? 1973. ISSN 0003-9632 (print), 1769-

681X (electronic). URL https://www.archivesdephilo.com/
rechercher/auteurs-r/.

Russo:1980:BRT

[Rus80] F. Russo. Book review: Treatise on basic philosophy. Vol. 3, Ontology 1 : The furniture of the World by Mario Bunge. Archives de Philosophie, 43(1):151–152, January/March 1980. ISSN 0003-9632 (print), 1769-681X (electronic). URL https://www.jstor.

org/stable/43034290.

Russo:1986:BRT

[Rus86] François Russo. Book review: Treatise on Basic Philosophy, vol. 5 and 6 : Epistemology and Methodology by Mario Bunge. Archives de Philosophie, 49(3):510–511, July/September 1986. ISSN 0003-9632 (print), 1769-681X (electronic). URL https://

www.jstor.org/stable/43035120.

Rosemann:2005:EEB

[RW05] Michael Rosemann and Boris Wyssusek. Enhancing the expressiveness of the Bunge–Wand–Weber ontology. In ????, editor, Proceedings of the Eleventh Americas Conference on Information Systems, Omaha, NE, USA, 11–14 August, 2005, page ?? ????, ????, 2005. ISBN ???? LCCN ???? URL https://aisel.

aisnet.org/amcis2005/438/.

Suppe:1976:PPB

[SA76] Frederick Suppe and Peter D. Asquith, editors. PSA 1976: Proceedings of the Biennial Meeting of the Philosophy of Science Association 1976, volume 1. Philosophy of Science Association, East Lansing, MI, USA, 1976. ISBN 0-917586-02-6 (hardcover), 0-917586-01-8 (paperback). LCCN Q174.B67.

SanchezSorondo:1993:PCN

[SA93] Marcello Sánchez Sorondo and Dario Antiseri, editors. Physica, cosmologia, Naturphilosophie, nuovi approcci. (Italian) [Physics, cosmology, natural philosophy, new approaches]. Herder/ Università Lateranense, Roma, Italy, 1993. xxvi + 480 pp.

Speroni:2001:PFR

[SA01] José Luis Speroni and Jorge Victoriano Alonso, editors. El pensamiento de Francisco Romero: retrato de un filósofo argentino del siglo XX. Edivérn, Buenos Aires, 2001. ISBN 987-97861-2-2. 387 pp. LCCN B1034.R64 S64 2001; D107 .L4.

Sadovnikov:2004:SSL

[Sad04] Slava Sadovnikov. Systemism, social laws, and the limits of social theory: Themes out of Mario Bunge’s: The sociology–philosophy connection. Philosophy of the Social Sciences = Philosophie des sciences sociales, 34(4):536–587, 2004. ISSN 0048-3931 (print), 1552-7441 (electronic).

Salt:1971:PAF

[Sal71] David Salt. Physical axiomatics: Freudenthal vs. Bunge. Foundations of Physics, 1(4):307–313, ???? 1971. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00708580.

Sobrevilla:1992:LRH

[SB92] D. Sobrevilla and D. García Balaúnde, editors. Lógica, razón y humanismo: La obra filosófica de Francisco Miró Quesada C.. (Spanish) [Logic, reason and humanism: The philosophical work of Francisco Miro Quesada C.]. Universidad de Lima, Lima, Peru, 1992. 443 pp.

Smelser:2001:IES

[SB01] Neil J. Smelser and Paul B. Baltes, editors. International Encyclopedia of the Social & Behavioral Sciences, volume 19 (R– RET). Elsevier, Amsterdam, The Netherlands, 2001. ISBN 0-08-043076-7. xii + 16695 pp.

Stirling:2008:GTM

[SBC^{+}08] James R. Stirling, Ole Bang, Peter L. Christiansen, Maria S. Zakynthinaki, and Steffen Kjær Johansen. Geometry and transport in a model of two coupled quadratic nonlinear waveguides. Chaos: an Interdisciplinary Journal of Nonlinear Science, 18(1): 013116, March 2008.

Serroni-Copello:1989:ECM

[SC89] Raúl Serroni-Copello. Encuentros con Mario Bunge. (Spanish) [Encounters with Mario Bunge]. Colección Entrevistas. ADIP, Buenos Aires, Argentina, 1989. ISBN 950-9272-02-7. 212 pp. LCCN MLCS 99/00424 (B).

Scarano:2014:PEM

[Sca14] Eduardo Scarano. Propuestas epistemológicas de Mario Bunge para comprender la tecnología: reglas, fines, acciones racionales, diseños. (Spanish) [Epistemological proposals of Mario Bunge for understanding technology: rules, goals, rational actions, designs]. In Denegri [Den14a], pages 269–290. ISBN 950-23-2339-4.

LCCN BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Schlegel:1961:MBC

[Sch61] Richard Schlegel. Mario Bunge on causality. Philosophy of Science, 28(1):72–82, January 1961. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/185467.

Scriven:1963:RPI

[Scr63] Michael Scriven. Recent publications: Intuition and Science, by Mario Bunge. American Mathematical Monthly, 70(2):232–233, February 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schurz:1991:ASP

[SD91] Gerhard Schurz and Georg J. W. Dorn, editors. Advances in Scientific Philosophy: Essays in Honour of Paul Weingartner on the Occasion of the 60th Anniversary of His Birthday, volume 24 of Poznań Studies in the Philosophy of the Sciences and the Humanities. Rodopi, Amsterdam, The Netherlands, 1991. ISBN 90-5183-317-2. xii + 644 pp. LCCN B29 .A28 1991; Q175.3A48 1991.

Seising:2006:VMT

[Sei06] Rudolf Seising. From vagueness in medical thought to the foundations of fuzzy reasoning in medical diagnosis. Artificial Intelligence in Medicine, 38(3):237–256, 2006. ISSN 0933-3657 (print),

1873-2860 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0933365706001072.

Seni:1993:ETP

[Sen93] Dan Alexander Seni. Elements of a theory of plans. Ph.D. thesis in City and Regional Planning, Graduate School of Arts and Sciences, University of Pennsylvania, Philadelphia, PA, USA, 1993. xvi + 905 (two volumes) pp. URL https://search.proquest.

com/pqdtglobal/docview/304087462/.

Settle:1996:ASO

[Set96] Tom Settle. Applying scientific openmindedness to religion and science education. Science & Education (Springer), 5(2):125– 141, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Shea:1976:BIP

[She76] William R. Shea, editor. Basic Issues in the Philosophy of Science. Canadian contemporary philosophy series. Science History Publications, New York, NY, USA, 1976. ISBN 0-88202-160-5. 215 pp. LCCN Q175.3 .B37. URL http://catalog.

hathitrust.org/api/volumes/oclc/2331476.html.

Silberstein:2013:BMB

[Sil13a] Marc Silberstein. Bibliographie de Mario Bunge. (French) [Bibliography of Mario Bunge]. Matériaux philosophiques et scientifiques pour un matérialisme contemporain, 2(??):983–1003, December 2, 2013.

Silberstein:2013:FAD

[Sil13b] Marc Silberstein. La fonction architectonique du matérialisme. (French) [The architectural function of materialism]. In Matériaux philosophiques et scientifiques pour un matérialisme contemporain: Volume 1 [Sil13d], page ?? ISBN 2-919694-51-0 (print), 2-919694-20-0 (e-book). LCCN B825 .M384 2016. URL

http://public.eblib.com/choice/publicfullrecord.aspx?
p=5336199.

Silberstein:2013:MSS

[Sil13c] Marc Silberstein. Le matérialisme systémiste selon Mario Bunge. (French) [Systemic materialism according to Mario Bunge]. In Matériaux philosophiques et scientifiques pour un matérialisme contemporain: Volume 1 [Sil13d], chapter 3, page ?? ISBN 2-919694-51-0 (print), 2-919694-20-0 (e-book).

LCCN B825 .M384 2016. URL http://public.eblib.com/
choice/publicfullrecord.aspx?p=5336199.

Silberstein:2013:MPS

[Sil13d] Marc Silberstein. Matériaux philosophiques et scientifiques pour un matérialisme contemporain: Volume 1. Editions Matériologiques, Paris (51, rue de la Fontaine au Roi 75011), France, second edition, 2013. ISBN 2-919694-51-0 (print), 2-919694-20-0 (e-book). 574 pp. LCCN B825 .M384 2016.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=5336199.

Silberstein:2013:SME

[Sil13e] Marc Silberstein. Science(s) et matérialisme(s): examen des conditions d’une synonymie. (French) [science(s) and materialism(s): Examining the conditions of synonymy]. In Matériaux philosophiques et scientifiques pour un matérialisme contemporain: Volume 1 [Sil13d], page ?? ISBN 2-919694-51-0 (print), 2-919694-20-0 (e-book). LCCN B825 .M384 2016.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=5336199.

Silva:2013:DRB

[Sil13f] Luis Carlos Silva. En defensa de la racionalidad bayesiana: a propósito de Mario Bunge y su “Filosofía para médicos”. (Spanish) [In defense of Bayesian rationality: on the subject of Mario Bunge and his “Philosophy for Physicians”]. Comunicaciones en Estadística, 6(2):197, October 12, 2013. CODEN ???? ISSN 2027-3355 (print), 2339-3076 (electronic).

Silberstein:2019:MBP

[Sil19] Marc Silberstein. Mario Bunge publications (all languages). In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 41, pages 775–807. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Sinclair:1981:BRR

[Sin81] S. B. Sinclair. Book review: Rutherford and Physics at the Turn of the Century. Edited by Mario Bunge and William R. Shea. New York: Science History Publications; Folkestone: Dawson, 1979. Pp. viii + 184.£12.50. British Journal for the History of Science, 14(1):96–97, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4026083.

Sinaceur:1991:PAA

[Sin91] Allal Sinaceur, editor. Penser avec Aristote. Erès, Toulouse, France, 1991. ISBN 2-86586-191-0 (paperback). LCCN B485 .P39 1991.

Shook:2009:FN

[SK09] John R. Shook and Paul Kurtz, editors. The Future of Naturalism. Humanity Books, Amherst, NY, USA, 2009. ISBN 1-59102-731-4 (hardcover). LCCN B828.2 .F88 2009.

Sassower:2019:ICR

[SL19] Raphael Sassower and Nathaniel Laor, editors. The impact of critical rationalism: expanding the Popperian legacy through the works of Ian C. Jarvie. Palgrave Macmillan, Basingstoke, Hampshire, UK, 2019. ISBN 3-319-90825-1 (hardcover), 3-319-90826-X (PDF e-book), 3-030-08108-7. xi + 348 pp. LCCN B833 .I473

2019. URL https://link.springer.com/book/10.1007/978-
3-319-90826-7.

Slezak:1996:BRP

[Sle96] Peter Slezak. Book review: Paul Davies, The Mind of God: Science and the Search for Ultimate Meaning, Simon & Schuster (cloth), New York, 1992; Penguin, London (paper), 1993. pp. 254, USD10. Science & Education (Springer), 5(2):201–212, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Slezak:2012:BRMb

[Sle12a] Peter Slezak. Book review: Mario Bunge: Matter and Mind: A Philosophical Inquiry. Science & Education (Springer), 21(8):

1213–1221, August 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Slezak:2012:MBM

[Sle12b] Peter Slezak. Mario Bunge’s materialist theory of mind and contemporary cognitive science. Science & Education (Springer), 21 (10):1475–1484, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Slezak:2019:MBC

[Sle19] Peter Slezak. Mario Bunge and contemporary cognitive science. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 21, pages 363–375. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Slottje:1999:AEI

[Slo99] Daniel J. Slottje, editor. Advances in Econometrics, Income Distribution and Scientific Methodology: Essays in Honor of Camilo Dagum. Physica-Verlag HD, Heidelberg, Germany, 1999. ISBN 3-7908-1226-9 (hardcover), 3-642-93643-1 (print), 3-642-93641-5 (e-book). xii + 388 pp. LCCN HB523 .A3 19994. URL http://

www.springerlink.com/content/978-3-642-93641-8.

Sneed:1968:BRD

[Sne68a] Joseph D. Sneed. Book review: Delaware Seminar in the Foundations of Physics by Mario Bunge. Science, 159(3818):966, March 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL https://www.jstor.org/stable/1723565.

Sneed:1968:DSF

[Sne68b] Joseph D. Sneed. Book review: Delaware Seminar in the Foundations of Physics. Mario Bunge, Ed. Springer-Verlag, New York, 1967. xii + 193 pp. $9.50. Studies in the Foundations, Methodology and Philosophy of Science, vol. 1. Science, 159(3818):966, March 1, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Sneed:1968:BRQ

[Sne68c] Joseph D. Sneed. Book review: Quantum Theory and Reality by Mario Bunge. Synthese, 18(4):464–467, October 1968. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://www.

jstor.org/stable/20114620.

Solano:2022:CCM

[Sol22] David Martín Solano. A constructive critique of Mario Bunge’s theory of truth. Metascience: Scientific General Discourse, 2 (??):130–137, 2022. CODEN ???? ISSN ???? URL https://

philpapers.org/rec/MARACC-8.

Symons:2011:ONU

[SPT11] John Symons, Olga Pombo, and Juan Manuel Torres, editors. Otto Neurath and the Unity of Science, volume 18 of Logic, Epistemology, and the Unity of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-0142-X (hardcover), 94-007-0143-8 (e-book). ISSN 4020-5034. viii + 264 pp. LCCN B3309.N394 O886 2011; Q175 .O88

2011. URL http://link.springer.com/10.1007/978-94-007-
0143-4.

Strauss:1969:CBF

[Str69] M. Strauss. Corrections to Bunge’s Foundations of Physics (1967). Synthese, 19(3–4):433–442, April 1969. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485660;
 http://link.springer.com/content/pdf/10.1007/BF00485660.
pdf.

Schurz:1989:PNS

[SW89] Gerhard Schurz and Paul Weingartner, editors. Philosophy of the natural sciences = Philosophie der naturwissenschaften: Akten des 13. Internationalen Wittgenstein Symposiums: 14. Bis 21. August 1988, Kirchberg am Wechsel, Österreich: ausgewählte beiträge: proceedings of the 13th International Wittgenstein Symposium: 14th to 21st August 1988, Kirchberg am Wechsel, Austria: selected papers, volume 17 of Schriftenreihe der Wittgenstein-Gesellschaft. Hölder-Pichler-Tempsky, Vienna, Austria, 1989. ISBN 3-209-00861-2 (paperback). LCCN B3376.W564; Q174 .I586 1988.

Sampson:2009:ECC

[SW09] Robert J. Sampson and Per-Olof H. Wikström, editors. The Explanation of Crime: Context, Mechanisms and Development. Pathways in crime. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-11905-7 (paperback). vii + 304 pp. LCCN

HV6117 .E96 2009. URL http://www.gbv.de/dms/bowker/
toc/9780521119054.pdf.

Szathmry:1998:US

[Sza98] Eörs Szathmáry. Useful stuff. Trends in Ecology and Evolution, 13(6):251–252, 1998. CODEN TREEEQ. ISSN 0169-5347 (print),

1872-8383 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0169534798013469.

Taylor:1985:SDC

[TBG^{+}85] Mark Kline Taylor, Mario Bunge, Atwood D. Gaines, Yvon Gauthier, Marvin Harris, H. Dieter Heinen, I. C. Jarvie, Robert Layton, Bob Scholte, Anthony Stellato, and Steven Webster. Symbolic dimensions in cultural anthropology [and comments and reply]. Current Anthropology, 26(2):167–185, April 1985. URL

http://www.jstor.org/stable/2743131.

Thom:1991:DI

[Tho91] René Thom. A dangerous illusion. New Ideas in Psychology, 9(2): 233–234, 1991. ISSN 0732-118X (print), 1873-3522 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0732118X9190028K. Special issue on Mario Bunge and Martin Mahner.

Torretti:2014:NHN

[Tor14] Roberto Torretti. Nota histórica sobre la noción de causa. (Spanish) [Historic note on the notion of cause]. In Denegri [Den14a], pages 291–308. ISBN 950-23-2339-4. LCCN

BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Trevino:2001:BRS

[Tre01] A. Javier Treviño. Book review: The Sociology-Philosophy Connection by Mario Bunge. Contemporary Sociology — a Journal of Reviews, 30(5):536–537, September 2001. ISSN 0094-3061 (print), 1939-8638 (electronic). URL https://www.jstor.org/

stable/3089374.

Trigger:2006:BRC

[Tri06] Bruce G. Trigger. Book review: Chasing Reality: Strife over Realism by Mario Bunge. Canadian Journal of Archaeology / Journal Canadien d’Archéologie, 30(2):346–349, ???? 2006. CODEN

???? ISSN 0705-2006. URL https://www.jstor.org/stable/
41103585.

Tuldava:1995:CBC

[Tul95] Juhan Tuldava. A comment on Bunge’s causality and probability in linguistics. Journal of Quantitative Linguistics, 2(1):17–18, January 1995. ISSN 0929-6174 (print), 1744-5035 (electronic).

Tuomela:1974:BRB

[Tuo74a] Raimo Tuomela. Book review: Method, Model and Matter by Mario Bunge. Philosophy of Science, 41(4):429–430, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187016.

Tuomela:1974:BRM

[Tuo74b] Raimo Tuomela. Book review: Method, Model and Matter by Mario Bunge. Philosophy of Science, 41(4):429–430, December 1974. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL https://www.jstor.org/stable/187016.

Turner:1996:RIS

[Tur96] Harold Turner. Religion: impediment or saviour of science? Science & Education (Springer), 5(2):155–164, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Vacher:1993:EAM

[Vac93] Laurent-Michel Vacher. Entretiens avec Mario Bunge: une philosophie pour l’âge de la science. (French) [Interviews with Mario Bunge: a Philosophy for the Age of Science]. De vive voix. Liber, Montréal, QC, Canada, 1993. ISBN 2-921569-07-8. 139 pp. LCCN B995 B86B86 1993.

VanRillaer:1991:SDP

[Van91] Jacques Van Rillaer. Strategies of dissimulation in the pseudosciences. New Ideas in Psychology, 9(2):235–244, 1991. ISSN 0732-118X (print), 1873-3522 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0732118X9190029L.

Special issue on Mario Bunge and Martin Mahner.

VanDenBerg:2001:SSA

[Van01] Axel Van den Berg. The social sciences according to Bunge. Philosophy of the Social Sciences = Philosophie des sciences sociales, 31(1):83–103, March 2001. ISSN 0048-3931 (print), 1552-7441

(electronic). URL https://journals.sagepub.com/doi/pdf/
10.1177/004839310103100105.

VandenBerg:2004:SSR

[Van04] Axel Van den Berg, editor. The Social Sciences and Rationality: Promise, Limits, and Problems [papers presented at a conference held at McGill University in March 1999]. Transaction Publishers, New Brunswick, NJ, USA, 2004. ISBN 0-7658-0232-5. LCCN HM495 .S63 2004.

Vargas:2017:EPS

[Var17] M. L. Vargas. Exact psychiatry: Six axioms. European Psychiatry: the journal of the Association of European Psychiatrists, 41:S330, 2017. ISSN 0924-9338 (print), 1778-3585 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0924933817325403. Abstract of the 25th European Congress of Psychiatry.

Vetter:1975:R

[VBH^{+}75] Hermann Vetter, Radu J. Bogdan, F. Gregory Hayden, Robert H. Puckett, and Mario Bunge. Reviews. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 6(1):103–120, February 1975. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Velazquez:2022:SFS

[Vel22] Matias Velázquez. On some features of the scientific hylorealistic background of crystal chemistry. Metascience: Scientific General Discourse, 2(??):96–128, 2022. CODEN ???? ISSN ???? URL

https://philpapers.org/rec/VELOSF.

Virues-Ortega:2012:PNI

[VOHPMJ12] Javier Virues-Ortega, Camilo Hurtado-Parrado, Toby L. Martin, and Flávia Julio. Psycho-neural identity as the basis for empirical research and theorization in psychology: an interview with Mario A. Bunge. Science & Education (Springer), 21(10):1527–1534, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Vollmer:1973:BRM

[Vol73a] Gerhard Vollmer. Book review: Method, Model and Matter by Mario Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):380–383, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/
25170292.

Vollmer:1973:BRP

[Vol73b] Gerhard Vollmer. Book review: Philosophy of Physics by Mario Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):407–409, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170298.

Vollmer:1979:BRF

[Vol79] Gerhard Vollmer. Book review: The Furniture of the World by M. Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):405–407, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170522.

Vollmer:1981:BRW

[Vol81] Gerhard Vollmer. Book review: A World of Systems by M. Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):178–179, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170581.

Vollmer:1993:WIE

[Vol93] Gerhard Vollmer. Wissenschaftstheorie im Einsatz: Beiträge zu einer selbstkritischen Wissenschaftsphilosophie. (German) [Science Theory in Action: Contributions to a Self-Critical Philosophy of Science]. S. Hirzel, Stuttgart, Germany, 1993. ISBN 3-7776-0499-2. xiv + 226 pp. LCCN BD163 .V653 1993.

vanRootselaar:1977:BTT

[vR77] B. van Rootselaar. On Baunge’s theory of things. International Journal of General Systems, 3(3):175–180, 1977. CODEN IJGSAX. ISSN 0308-1079 (print), 1563-5104 (electronic). See reply [Bun77j].

Vucetich:2014:PCU

[Vuc14] Héctor Vucetich. ‘?Partículas u ondas? Problemas ontológicos en la mecánica cuántica. (Spanish) [Particles or waves? Ontological problems in quantum mechanics]. In Denegri [Den14a], pages 309–322. ISBN 950-23-2339-4. LCCN BD165 .E57 2014. URL

http://core.cambeiro.com.ar/0-100933-7.pdf;   http://
www.loc.gov/catdir/toc/gc01/9789502323398.pdf.

Weingartner:1990:SMB

[WD90] Paul Weingartner and George J. W. Dorn, editors. Studies on Mario Bunge’s Treatise, volume 18 of Poznań studies in the philosophy of the sciences and the humanities. Rodopi, Amsterdam, The Netherlands, 1990. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). 720 pp. LCCN B72.B863 S78 1990.

Wettersten:1982:PMB

[Wet82] John Wettersten. The place of Mario Bunge. In Agassi and Cohen [AC82], chapter 26, pages 465–486. ISBN 90-277-1262-X, 90-277-1263-8 (paperback), 94-009-8462-6 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 67. URL https://link.springer.

com/chapter/10.1007/978-94-009-8462-2_26.

Wheaton:1980:BRB

[Whe80] Bruce R. Wheaton. Book review: Rutherford and Physics at the Turn of the Century by Mario Bunge; William R. Shea. Isis, 71(2):317–318, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230202.

Wiegand:1999:SSW

[Wie99] Herbert Ernst Wiegand, editor. Sprache und Sprachen in den Wissenschaften: Geschichte und Gegenwart: Festschrift für Walter de Gruyter and Co. anlässlich einer 250jährigen Verlagstradition. (German) [Language and Languages in the Sciences: History and the Present: Festschrift for Walter de Gruyter and Co. on the occasion of a 250-year publishing tradition]. Walter de Gruyter, New York, NY, USA, 1999. ISBN 3-11-015579-6 (hardcover), 3-11-080132-9, 3-11-174876-6. LCCN Q226. URL http:

//www.degruyter.com/doi/book/10.1515/9783110801323.

Wren-Lewis:1996:BBN

[WL96] John Wren-Lewis. On babies and bathwater: a non-ideological alternative to the Mahner/Bunge proposals for relating science and religion in education. Science & Education (Springer), 5(2): 185–188, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge. See reply [MB96a].

Woolnough:1996:FCR

[Woo96] Brian E. Woolnough. On the fruitful compatibility of religious education and science. Science & Education (Springer), 5(2): 175–183, April 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic). Special issue on Mario Bunge.

Weinberger:1974:R

[WPG^{+}74] Ota Weinberger, Philip Pettit, Hans-W. Gottinger, Hugh Lehman, Carol Babner Barry, Oliva Blanchette, Michael Ruse, John McMurtry, F. Michael Walsh, J. E. White, and Scott A. Kleiner. Reviews. Theory and Decision: an International Journal for Philosophy and Methodology of the Social Sciences, 4(3-4):373–426, 1974. CODEN THDCBA. ISSN 0040-5833 (print), 1573-7187 (electronic).

Wikstrom:2006:ECC

[WS06] Per-Olof H. Wikström and Robert J. Sampson, editors. The Explanation of Crime: Context, Mechanisms, and Development. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85707-4 (hardcover). xvii + 305 pp. LCCN HV6025 .E86 2006.

Wand:1990:MBO

[WW90] Yair Wand and Ron Weber. Mario Bunge’s ontology as a formal foundation for information systems concepts. In Weingartner and Dorn [WD90], pages 123–150. ISBN 90-5183-187-0 (hardcover), 90-5183-188-9 (paperback). ISSN 0303-8157 (print), 1875-7243 (electronic). LCCN B72.B863 S78 1990.

Weingartner:1970:IPE

[WZ70] Paul Weingartner and Gerhard Zecha, editors. Induction, physics, and ethics: proceedings and discussions of the 1968 Salzburg colloquium in the philosophy of science, volume 31 of Synthese library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1970. ISBN 90-277-0158-X. LCCN Q175 .S235 1968.

Yannarella:2014:BCP

[Yan14] Francisco Gerardo Yannarella. Breves consideraciones sobre parasitismo animal. (Spanish) [Brief considerations on animal parasitism]. In Denegri [Den14a], pages 323–336. ISBN 950-23-2339-4.

LCCN BD165 .E57 2014. URL http://core.cambeiro.com.ar/
0-100933-7.pdf;   http://www.loc.gov/catdir/toc/gc01/
9789502323398.pdf.

Yannarella:2019:BSO

[YCVP^{+}19] Francisco Yannarella, Mauro A. E. Chaparro, José Geiser Villavicencio-Pulido, Martín Orensanz, and Guillermo M. Denegri. Bungean systemic ontology and its application to research in animal parasitism. In Mario Bunge: a Centenary Festschrift [Mat19a], chapter 32, pages 575–583. ISBN 3-030-16672-4 (hardcover), 3-030-16673-2 (e-book), 3-030-16674-0 (print), 3-030-16675-9 (print). LCCN Q175.55.

Zielinska:2016:LRE

[Zie16] Dorota Zielinska. Linguistic research in the empirical paradigm as outlined by Mario Bunge. SpringerPlus, 5(1):1–9, ???? 2016. CODEN SPRIFB. ISSN 2193-1801.

Zielinska:2022:LRE

[Zie22] Dorota Zielińska. Linguistic research in the empirical paradigm as outlined by Mario Bunge. Metascience: Scientific General Discourse, 2(??):182–202, 2022. CODEN ???? ISSN ???? URL

https://philpapers.org/rec/ZIELRI.

Wan:2011:RPL

[zW11a] Poe Yu ze Wan. (Re-)Problematizing the Luhmannian constructivist systems approach: a Bungean intervention. Current Sociology, 59(5):696–716, November 2011. ISSN 0011-3921 (print),

1461-7064 (electronic). URL https://journals.sagepub.com/
home/csi.

Wan:2011:RSE

[zW11b] Poe Yu ze Wan. Reframing the Social: Emergentist Systemism and Social Theory. Ashgate Publishing, Farnham, Surrey GU9 7PT, UK, 2011. ISBN 1-4094-1152-4 (hardcover), 1-283-11536-0 (e-book), 1-315-60433-7 (e-book PDF), 1-317-06933-1 (e-book), 1-4094-1153-2 (e-book). x + 246 pp. LCCN HM585 .W358 2011.

URL http://catalogue.bnf.fr/ark:/12148/cb42430918m.

Wan:2012:ASB

[zW12] Poe Yu ze Wan. Analytical sociology: a Bungean appreciation. Science & Education (Springer), 21(10):1545–1565, October 2012. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).